A Bibliography of Papers in Lecture Notes in Computer Science (1996), Part 1 of 2

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

11 March 2017 Version 1.28

Title word cross-reference

2 [388, 1104, 416, 895]. 3 [2, 255, 1416, 931, 1022, 11, 944, 17, 962, 20, 1023, 24, 904, 1607, 227, 1016, 251, 40, 252, 428, 905, 250, 1308]. i, k [187]. K [1902, 552, 1975, 1606, 1607]. λ(?) [2030, 2029]. Z/nZ [1238]. Z_{n} [1239]. N [384, 919]. O(dSd × dT d) [1884]. o(n03) [1604]. O(n log n) [1822, 529]. q [1366]. r [537]. t [1258].

-Algorithm [1884]. -approximation [1053]. -Average-Time [1822]. -body [384]. -Calculus [2030, 2029]. -Cuts [1607]. -D [11, 388, 1104]. -Dimensional [2, 919]. -Relaxation [1603]. -Resilient [1258]. -rule [784]. -Samples [1366]. -Set [1606].

-Space [1416]. -Trees [1975]. -Tuples [537]. -wise [552].

01U [1155]. 02Log [841]. 0SC0*ECS [1888].

1 [747, 1980]. 1-Coloring [1980]. 12 [1590]. 160 [337].

2 [1370, 1342, 1191, 351, 34]. 2-Layer [34]. 2-Paths [1342]. 2/3 [1370]. 2/ 3-Approximation [1370].

3 [2044]. 3-free [2044]. 368 [1915].

4 [151, 29]. 4-Connected [29]. 4.0 [1182].

5 [386, 488, 140].

83 [1811].

90 [1183, 399, 176]. 95 [164, 144, 1806, 1813, 1799, 138, 1784, 1790, 150, 1793, 1809, 1785, 1819, 1812, 1808, 149, 1796, 1789, 1814, 1807, 1800, 1781].

= [819].

A0* [819]. Ab-initio [412]. Abductive [1650]. ABDUCTOR [51]. ABR [485, 471]. Absence [799]. Abstract [821, 1202, 1905, 605, 560, 567, 768, 1913, 1209, 1033, 678]. Abstraction [1865, 806, 629, 1568, 1540, 591, 195, 784, 1362]. Abstractions [2021]. Acceptance [484]. Access [668, 81, 753, 481, 482]. ACCESS.bus [643]. Accesses [2066]. Accessibility [1635]. Accommodating [750]. ACCOPI [664, 498]. Account [1217]. Accumulated [338]. Accuracy [979, 184, 1079]. Achieve [471]. Achievements [1083]. Achieving [165, 61]. acquire [379]. Acquiring [371]. Acquisition [1577, 1408, 1402, 1410, 1405, 1399, 130, 1398, 1407, 1022, 366, 378, 353]. Across [1893]. Acting [2039, 291]. Action [1548, 537, 624, 1689, 1672, 491, 623, 1501]. Actions [528, 605, 1547, 1909, 1671, 1682]. Activation [1721]. Active [1957, 231, 1943, 229, 1516, 718, 79, 1521, 230, 1115, 87, 232, 2017, 234, 916, 993, 1012, 998]. Activities [835, 836, 1521, 2021]. Actor [1697]. Actor-Based [1697]. Actors [1865]. acts [363]. Acyclic [1627, 1823]. Ada [131, 1795, 143, 164, 162, 163, 1806, 160, 1813, 1799, 135, 138, 1784, 1815, 1790, 150, 1793, 134, 145, 1809, 1785, 1819, 148, 1812, 1808, 149, 1786, 1787, 141, 1789, 157, 1811, 1798, 1814, 1807, 1800, 133, 1794, 139, 1818, 1804, 1781, 136, 1816]. Ada-Object [143]. Ada/O2 [1811]. Ada95 [165, 137, 132, 153]. Adapt [452]. Adaptation [836, 323, 1756, 439]. Adapting [600]. Adaptive [748, 1435, 1232, 1752, 1147, 1753, 402, 1724, 447, 586, 777, 1154, 1436, 1011]. Adaptivity [665]. Add [416]. Adding [1953, 1292]. Addition [1332, 1807]. Additional [1806]. Address [178]. Addressing [501]. Adjunctions [1315]. Adjustment [50]. Admission [502, 666, 667]. Admit [544]. Adocs [4]. Advanced [1690, 1538, 59]. Adversary [2074]. Advice [1551, 1731]. Advice-Giving [1551, 1731]. Advisor [1730]. Advisors [1760]. aerial [909]. Aerodynamic [1192]. Aeronautic [1085]. Aeronautical [1087]. Aerospace [1118]. Aerospatiale [1089]. Aesthetics [44]. Affine [1160, 958, 944]. Affine/ photometric [958]. Affordable [1101]. Against [347, 1251]. AGD [1732]. Agency [447, 301]. Agent [304, 325, 302, 1225, 1232, 440, 1769, 441, 1766, 322, 1229, 1558, 1684, 1778, 1222, 319, 1720, 1230, 442, 327, 1223, 495, 1771, 1699, 308, 1770, 310, 1224, 309, 1780, 1775, 1533, 1233, 1776, 1590, 1767, 290, 311, 297, 298, 439, 312, 1559, 2073, 320, 2081, 1483, 1504, 311]. Agent-Based [1720, 1533]. Agent-Oriented [1223, 1504]. Agents [314, 1779, 306, 299, 1691, 859, 1640, 1642, 307, 296, 118, 317, 328, 287, 329, 694, 890, 305, 1671, 316, 1580, 1774, 297, 1218, 291, 288, 315, 200]. AgentSpeak [316]. Aggregate [727, 182]. Aggregates [824]. Aggregation [271, 751]. Aggressive [812]. Agreement [62, 1972, 1262]. Agreements [58]. AGVs [128]. AHS [1036]. AI [1922]. aide [1727]. Aided [79, 108, 1514, 122]. Air [158, 901, 120]. Aircrafts [1089]. Akratic [1689]. Alarms [301]. Alexandria [1594]. Algebra [1898, 205, 390, 393, 392, 842, 1905, 714, 684, 1361]. Algebraic [1312, 1350, 687, 155, 2027, 1021, 589]. Algebras [1311, 1351, 1324, 694, 1917]. Algorithm [1577, 381, 1546, 896, 473, 1370, 1983, 877, 1363, 1602, 82, 1133, 1191, 1157, 886, 319, 689, 2078, 402, 767, 1948, 882, 282, 1724, 177, 35, 640, 1882, 196, 1884, 895, 426, 40, 323, 429, 1982, 1599, 1620, 902, 471, 884, 480, 1822, 372, 1024]. algorithmic [2080]. Algorithms [2065, 1979, 1387, 1970, 795, 1621, 1841, 8, 9, 12, 879, 2076, 2067, 1600, 1289, 396, 397, 1373, 1162, 1994, 1365, 408, 1608, 462, 899, 564, 1383, 224, 1248, 1626, 1293, 239, 881, 34, 36, 1309, 2042, 901, 1386, 1237, 1236, 1963, 2046, 883, 1837, 1831, 460, 433, 1623, 867, 1567, 1842, 1573, 241, 1379, 1119, 1388, 1464, 674, 979, 675, 681, 357]. Alias [2009]. Alias-Free [2009]. Aliases [825]. Alignment [1374, 1379, 989, 939]. Alignments [1993]. Allow [1429]. Allowing [1741]. Almaden [1593]. Almost [1272, 415]. Alphabet [1385, 1258]. ALPSTONE [1174]. Alternation [2054]. Alternative [877, 764, 1931, 1289, 168]. ALWAN [857]. Amalgamating [733]. ambiguities [964]. Ambiguity [210]. ambiguous [974]. America [1581]. Among [1779, 328, 1917, 450]. Amount [1201]. Anaesthesia [625]. Analogical [1653, 1659]. Analogous [1663]. Analogy [1660, 375]. Analyses [2065, 2067, 169]. Analysing [1887, 1667]. Analysis [1129, 266, 383, 1332, 816, 1525, 687, 621, 469, 166, 643, 1998, 821, 1077, 766, 1526, 1402, 171, 1127, 575, 1644, 567, 198, 547, 1295, 1835, 170, 714, 414, 1624, 1854, 771, 1878, 1879, 233, 713, 1883, 1859, 217, 240, 187, 222, 1386, 711, 1794, 121, 153, 190, 154, 487, 817, 574, 1449, 1893, 1041, 935, 987, 1459, 1038, 920, 904, 963, 1284, 2082, 354]. Analysis/Design [154]. Analyst [1710]. analytic [1276]. Analytical [1793]. Analyze [1435]. Analyzer [686]. Analyzing [321, 704, 120]. Anaphora [1553, 280, 369]. Ancestors [1979]. Ancestral [1993]. ANDF [138]. Angle [230]. angular [966]. Animated [1419]. Animation [1881, 894]. Annealing [17, 40]. Annotation [215]. Annotations [197, 1390]. Anomaly [1090, 2011]. ANSI [1305]. Answering [1674]. Anti [810, 1612]. Anti-Monge [1612]. Anticipatory [1778]. Ants [1220]. Aperiodic [516, 518]. APIs [86]. apparent [911]. APPEAL [1224]. Appearance [236, 913]. Applicability [1542]. Application [158, 1342, 1378, 325, 2005, 621, 1228, 135, 988, 665, 2006, 1642, 208, 537, 494, 899, 660, 1520, 141, 620, 464, 203, 1842, 661, 491, 1765, 1459, 1284, 1491]. Applications [2065, 1211, 1086, 389, 1957, 1111, 1137, 218, 2067, 1289, 658, 850, 1419, 1076, 756, 474, 1638, 723, 1812, 411, 1223, 669, 489, 871, 495, 1192, 1838, 1246, 174, 61, 2040, 286, 1798, 1540, 1563, 1430, 125, 1181, 2046, 133, 1794, 1120, 2017, 136, 1934, 241, 1204, 1468, 996, 925]. Applied [1081, 467, 617]. Applying [158, 369, 267, 1733, 1817, 622, 372]. Apprentice [1734]. apprentissage [1703]. Approach [1296, 1311, 649, 741, 1313, 320, 295, 651, 726, 830, 1153, 1232, 124, 1758, 150, 231, 1289, 1793, 219, 1447, 1734, 1710, 1629, 1365, 1374, 1226, 409, 1875, 91, 542, 1786, 269, 1883, 1564, 1699, 308, 423, 132, 1625, 1224, 890, 465, 1303, 1585, 1233, 1926, 777, 1377, 1148, 637, 97, 1403, 1443, 613, 1306, 1308, 565, 615, 458, 574, 1523, 1063, 1479, 375, 2079, 362, 913, 1490, 1501, 682, 1025, 1481, 1069, 940, 1068, 352, 994, 957, 1061]. Approaches [1554, 1441, 1828, 1746, 220, 1480, 356, 348]. Approche [1717]. appropriate [922]. appropriateness [984]. approximability [1855]. Approximants [784]. Approximate [1363, 1451, 1364, 1389, 1965, 1369, 1386, 1368, 1366]. Approximating [1851, 1606, 1971, 1607, 1850, 1986, 1062]. Approximation [2069, 533, 1387, 1378, 1970, 2044, 1608, 1257, 1835, 1108, 1969, 1968, 73, 1623, 1349, 1379, 673, 1030, 1053, 1370]. Approximations [1253, 1054]. AQ17 [1459]. AQ17-DCI [1459]. Arbitrary [1976, 1599]. Arc [1880, 2043]. Arc-Typed [1880]. Architectural [649, 1447, 1817, 1422, 1020]. Architecture [1737, 304, 302, 837, 1228, 299, 1778, 1134, 1951, 300, 1697, 1698, 1441, 1078, 505, 303, 68, 104, 755, 1699, 1425, 252, 497, 486, 1138, 1776, 1673, 487, 1437, 590, 1498, 1696, 1505, 355, 1496, 1506]. Architectures [166, 1728, 268, 102, 1144, 1789, 194, 743, 1079, 297, 1187, 312]. ARCHON [1219]. Area [790, 1831]. Area-Efficient [1831]. Arguing [1762]. Argumentation [1644, 1648, 1649, 1660, 1686, 1666, 1667, 1670, 1676, 565]. Argumentation-Theoretic [565]. Argumentative [1673]. Arguments [1685, 1658, 1667, 1487]. Aria [144, 1796]. Ariadne [844]. Ariane [140]. Aristotle [126, 1655]. Arithmetic [788, 552, 2026]. Arrangements [1828]. Array [166, 171, 169, 179, 874, 1107, 866]. Arrays [831]. Arrow [2035]. ART [1719]. articular [998]. Articulated [1233, 932]. Articulation [1716]. Artificial [869, 868, 1925, 887, 890]. ASIC [1228]. ASIS [142, 1803]. Asking [1826]. Aspect [1910]. Aspects [1320, 102, 1890]. Assembly [280, 1438]. Assertional [708]. Assertions [1357]. Assessment [160, 1752, 1753, 131, 721, 1715, 1794, 1682]. Assignment [1612, 823, 898, 824]. Associated [1853]. Associativity [1919]. Assumption [782, 1500]. Assumption/ Commitment [782]. Assumptions [1391, 1392]. Astrophysical [428]. Asymmetric [1237, 1241, 1380]. Asymptotic [879]. Asynchronous [2074, 1100, 327, 1875, 2028, 1857, 1891]. Asynchronous-Object [327]. Asynchronously [638]. Asynchrony [759]. ATAC [162]. ATEMPT [1180]. ATM [1131, 484, 483, 459, 477, 478, 664, 486, 460, 476, 1893, 475]. Atmospheric [1097, 1098, 438]. Atom [419]. attachment [361]. Attacks [1260, 343, 77, 79, 330, 344, 87]. Attempt [1435]. Attitudes [1766, 322, 288]. Attraction [99]. Attribute [1576, 1340, 1493]. Attribute-Oriented [1576]. Attributed [1322]. attributes [1476]. Audiences [1676]. audio [996]. audio-visual [996]. Audit [1633]. Augmentation [1838]. Augmenting [1806]. Australian [64, 65]. Authentication [84, 538]. Authenticators [1862]. Author [1268, 2022]. Authoring [489, 1701]. authority [1240]. Auto [1259]. Auto-correlations [1259]. autoepistemic [1277]. Automata [528, 163, 788, 1333, 388, 869, 530, 2049, 456, 2048, 787, 425, 704, 1963, 791, 458, 1066, 1028, 1064, 1060, 1062, 1059, 1071]. Automata-Theoretic [458]. Automated [320, 455, 1502, 60, 1959, 1303, 1731, 1034, 165, 1060, 1031]. Automatic [192, 909, 1176, 387, 1133, 945, 1301, 2006, 693, 1549, 947, 963, 1302, 196, 698, 363, 1300, 675, 568, 1298, 112, 191, 43, 641, 1042, 366, 993]. Automatically [710]. Automating [1577, 1469, 580]. Automation [1223, 620, 1502]. Automaton [75]. Automotive [1086, 1085]. Autonomous [1589, 1921, 1778, 890, 875, 450, 1767, 1774, 258, 1934]. Autonomous-Agent [1767]. Availability [745]. Average [2065, 1852, 534, 1822]. AVL [387]. Avoiding [588]. Awareness [1708]. Away [1571]. Axiomatisation [767]. Axiomatising [792]. Axiomatization [2033]. aye [1071]. aye-automata [1071].

B [395, 463, 622, 481, 467, 623]. B-ISDN [463, 481, 467]. B-Tool [623]. Babel [609]. Back [1191]. Back-Propagation [1191]. Backbone [1755]. Backend [180]. Backtrack [2072]. Bacterial [323]. Bake [1550]. Baker [125]. Balanced [512, 1147, 1724]. Balancing [1821, 1146, 405, 413, 1148, 1145]. Ballot [1240]. Bandwidth [333, 2024]. Bank [1954]. Banking [2004]. Base [1411, 1530, 2042, 278, 480, 1453, 1457]. Based [1552, 527, 304, 795, 1073, 536, 1554, 668, 1932, 135, 138, 1558, 602, 1733, 842, 1301, 182, 1305, 1344, 1749, 1720, 1646, 1697, 1710, 207, 1698, 2023, 1207, 1155, 1653, 1295, 75, 1399, 1231, 463, 282, 714, 1532, 861, 1724, 35, 1879, 1771, 1705, 264, 570, 1519, 1541, 262, 57, 1699, 1693, 1238, 1750, 276, 1780, 849, 252, 1303, 1117, 1575, 1533, 222, 486, 743, 344, 253, 1213, 883, 2012, 1120, 1941, 156, 1748, 113, 1239, 48, 1424, 1110, 263, 1294, 581, 574, 1515, 884, 315, 234, 476, 1757, 1955, 1751, 932, 1479, 1480, 988, 1505, 911]. based [989, 947, 1010, 951, 953, 1585, 1481, 1061, 1024, 1482, 1491, 1461, 1287, 1489, 371, 373, 389, 659, 118, 1121, 1619, 1188, 167]. basée [1745]. Basic [101, 393, 604, 1288]. Basins [99]. Basis [1713, 1126]. Bayesian [1548, 1752, 1935, 1940, 1559]. BDI [300, 317, 316]. Be [468, 1955, 1334]. Bear [341]. Beauty [18]. Becomes [1643]. Beginners [1758]. Behavior [321, 1705, 2021, 865, 1891, 1504]. Behavioral [1869, 445, 1889]. behaviorist [1481]. Behaviors [876]. Behaviour [1947, 1010]. Behavioural [802, 803]. Belief [1683, 1684, 1656, 1646, 313, 1652, 310, 1927, 612, 289, 1755, 1633]. Belief-Desire-Intention [289]. Beliefs [1713]. Benchmark [747, 387, 1309, 1124]. Benchmarking [1535, 1193, 407, 1174]. Bend [26]. Best [525, 679]. best-first [679]. BETA [2019, 1878]. Between [2016, 593, 1902, 1946, 1324, 270, 2035, 2018, 1330]. Beyond [1807]. Bi [1838]. Bi-Level [1838]. bias [1508]. Biased [1858]. Biases [1534]. Bibliography [439, 312]. Bidirectional [255, 967]. bifurcation [920]. Bijections [540]. Billion [869]. Binary [546, 1544, 18, 1997]. Binding [2003]. Bindings [666]. binocular [939]. binocularly [917]. Binomial [2042].

Bioelectromagnetic [1080]. Biological [1107, 1377, 1722]. Biology [1367]. Bipartite [1622]. Bipartitioning [1138]. Birth [873]. Birthday [1260]. BISDN [462]. Bisimulation [2062]. Bit [502, 2059]. Bits [1249]. Bivariate [1249]. Black [515]. Black-Box [515]. BLAS [404]. Block [341, 732, 395, 1155, 224, 415, 347, 342]. Block-Diagonal [415]. blood [990]. Blowfish [332]. BLOWmolding [1128]. BLOWPAC [1128]. blur [969]. Board [140, 162]. Body [227, 384, 1008]. Boltzmann [1134, 1109]. bond [1046, 1169]. Book [1434]. Books [1837]. Boolean [533, 1998, 1835, 1259]. Botanical [18]. Both [1730, 149]. Bottleneck [159]. bottom [1503]. bottom-up [1503]. Bouclettes [1177]. Bound [397, 1157, 1611, 680, 679, 678]. Boundaries [409]. Boundary [243]. Bounded [512, 597, 1342, 508, 1605, 543, 520, 1914, 555, 1916, 1820, 2026, 1325]. Bounded-Degree [1605]. Boundedness [541]. Bounds [1833, 557, 533, 1821, 2057, 5, 2060, 1827, 558, 463, 2049, 1978, 549, 2059, 2026, 1619, 1259, 980]. Box [515, 37]. Boyer [1364]. BPP [2052]. Bracketed [1347]. Brain [869, 869]. Branch [185, 397, 1157, 680, 679, 678]. branch-and-bound [679]. Branching [2051, 1662]. Breaking [159, 692, 1404, 2073]. Bridging [2018]. Broadband [469, 474, 479, 481, 482]. Broadcasting [851]. Browser [3]. Bubble [1572]. Buechi [2047]. Build [585]. Building [302, 680, 1901, 600, 1596, 894, 1926, 1380, 1581, 1764]. Buildings [1079]. BURS [819]. Burstiness [463]. Bus [1172]. Business [741, 756, 1750, 865, 1523, 1454]. Byzantine [1857].

C [584, 2007, 1186, 1197, 628, 198, 165, 2018, 781, 581, 813]. C.N.E.S [141]. CABRI [13]. CABRI-Graph [13]. Cache [171, 407, 595]. cal [173]. Calculations [1088, 419]. Calculi [1217, 805, 1272, 1277]. Calculus [1930, 690, 806, 545, 2030, 801, 840, 173, 1642, 2029, 1662, 849, 1918, 2047, 1360, 700, 807, 1271, 1469, 1906, 1281, 1285, 861]. Calibration [235, 903, 999]. Call [1805, 1197, 465, 657]. Calls [2007, 1182, 1438]. CAM [869]. CAM-BRAIN [869]. CAME [1513]. camera [978, 984, 985]. Can [1997, 1847, 376, 468, 1334]. Cannot [1220]. canonical [1464]. cap [1980, 514]. Capital [2058]. Capturing [670, 1517]. CARABAS [1116]. Card [85]. cardiac [987]. Cards [83, 84, 1261]. Care [1738, 60]. Carlo [1103, 435, 241]. Carried [467]. cartilage [998]. Cartographic [325]. cas [1696]. Cascade [351]. Cascade-2 [351]. Case [2065, 1130, 2001, 1710, 625, 1384, 860, 531, 1824, 1394, 1103, 619, 1751, 1066, 1513, 721]. Case-Based [1710, 1751]. Case-study [1394]. Cases [1612, 1624, 1777]. CASS [88]. Cast [124, 123, 126, 127, 111, 107, 98]. catalogue [990]. Categorical [610, 637]. Categories [802, 1315, 359]. categorization [372]. Category [1313, 1315]. Category-Theoretical [1313]. Causal [1648, 804, 1928, 923]. Causality [605]. Causally [1213]. Causally-Coherent [1213]. Causation [126]. cc [1028]. CCE [849]. CCSDS [157]. CDNF [1839]. CE [1707]. Cell [119, 1346]. Cellular [388, 869, 1141, 530, 1345, 868, 425]. Cellworks [1346]. Census [267]. Center [141]. Centered [1742]. Centralized [465]. Centuries [1692]. Century [1427]. CERN [1105]. Certain [1620]. CFD [1086, 1087, 1088, 1074, 1120, 1085]. Chain [695, 1338]. Chain-Code [1338]. Chained [413]. Chains [799, 1611, 1122, 1860]. Challenge [1121]. CHAM [862]. Change [1647, 1501]. Changes [58]. Changing [1538, 1189, 1669, 1504]. Chaos [885]. Characterisation [1791]. Characteristic [561]. Characterization [1439, 544]. Characterizations [2027]. Characterizing [543, 2031, 251]. Characters [555]. Chart [282, 1550]. Chart-Based [282]. Charts [686]. chattering [1030]. Cheating [1251]. Check [1825, 139]. Check-Out [139]. checkerboards [1269]. Checking [690, 643, 513, 697, 767, 654, 775, 778, 211, 1479, 1070]. Checks [344]. Chemical [832, 1123, 696, 1300, 1103, 1045]. Chemistry [411, 1098]. chi [568]. Children [1534]. Chimera [719]. CHIP [203]. Choice [292]. Choices [63]. Choosing [269]. CICS [622]. Cid [1208]. CIN [1683]. Cipher [73, 340, 342]. Ciphers [341, 346, 347, 344]. Circle [2043]. Circuit [1833, 546, 531]. Circuits [1133, 698, 866, 1844, 1284]. Circular [2043]. Circular-Arc [2043]. Circulation [1097]. Circumscription [1639, 1566]. Circumvent [2011]. circumventing [377]. CL [484]. Clan [48]. Clan-Based [48]. Class [2069, 2037, 1902, 911, 1882, 582, 1199, 1213, 1888, 585, 1438, 906, 1044]. Class-Based [1213]. Classes [534, 1630, 2015]. classical [550, 1283, 1286, 616]. Classification [592, 137, 262, 921, 363, 367, 1460, 994]. Classifications [1565]. Classifier [391, 717, 453]. Classifiers [1849]. Classroom [506]. Clause [598]. Clauses [572]. CLAX [46]. Cliches [1322]. Client [1813, 165, 743]. Client-Based [743]. Client-Server [1813, 165]. Clique [244]. Clock [76]. Clock-Controlled [76]. clocked [1026]. clocks [1061]. Cloning [201]. Closed [1334, 1358, 1500]. closed-world [1500]. Closure [1978, 174, 1499, 937]. CLP [568]. Cluster [241]. Clustering [2016, 1398, 1511, 371]. Clusters [1125]. Clutters [1617]. CM [386]. CM-5 [386]. Co [1427, 1700, 1049]. Co-operative [1427, 1700]. co-simulation [1049]. Coached [1695]. Coalition [328, 1767]. Coalition-Formation [1767]. Coarse [874, 1210]. Coarse-Grained [874]. Code [1077, 1826, 1338, 405, 1087, 628, 744, 1096, 1078, 165, 1094, 1074, 1081, 819, 186, 2020]. Coded [139]. Codes [384, 1170, 2025, 2055, 344, 1084]. Coding [214, 921]. Coevolution [885]. Coevolutionary [1483]. Cofree [2010]. cognitifs [1703]. Cognition [304, 1723]. Cognitive [1534]. COGNITUS [1117]. Cohabitate [1218]. Coherence [171, 269]. Coherency [1178]. Coherent [275, 1213, 929]. Collaboration [1708, 1227, 1424]. Collaborative [294, 844, 1710, 1730, 1425, 1761]. Collages [1338]. Collapsing [1351]. Collected [1692]. Collecting [547]. Collections [585]. Collective [2070, 1170, 2021]. Collusion [87]. Color [966, 236]. Colored [1869, 1873, 1878, 1886]. Coloring [1622, 1606, 1206, 1965, 1980]. Colour [965]. Coloured [1417]. Colourful [1597]. Comb [1631]. combinational [1284]. Combinations [1123]. Combinatorial [4, 672, 542, 1828, 1377, 677]. Combinatory [1351]. Combined [1306]. Combining [1273, 1897, 221, 628, 991, 1715, 157, 1668, 1722, 635, 1007]. COMENIUS [506]. Commands [1651]. Commerce [55, 1775, 59]. Commercial [64, 1078, 61]. Commit [757]. Commitments [314, 288]. Common [1979, 119]. CommonKADS [1412]. commonsense [1451]. Communicating [1578, 638]. Communication [557, 559, 2070, 1170, 1229, 558, 1173, 182, 1160, 1981, 1835, 2081, 499, 1795, 309, 556, 450, 1963, 180, 434, 1187]. Communications [1873, 1426, 656, 663]. Community [1590]. Commutativity [1854, 684]. Compact [549, 829]. Companion [1761]. Comparability [521]. Comparative [1634, 897, 630, 373, 1490].

Comparing [1101]. Comparison

[1102, 9, 343, 1456, 1290, 485, 1124, 470, 1081, 1395, 809, 1079, 433, 1281, 350]. Comparisons [14]. Compatible [544]. Competitive [554, 451]. Compilation [821, 2001, 173, 868]. Compiler [1898, 197, 830, 386, 138, 812, 183, 1183, 1951, 846, 773, 194, 172, 820, 46, 818]. Compilers [765, 830, 1176, 177, 180]. Compiling [761, 179, 777]. Complementary [1606]. Complementation [2048]. Complete [765, 1352, 2033, 123, 576, 1471, 1494]. Completeness [782, 616, 650, 1473, 1859]. Completion [1758, 1624, 576, 1619]. Complex [114, 1543, 1130, 1549, 100, 103, 1667, 90, 116, 934, 1487]. Complex-Evolutive [103]. Complexity [566, 557, 1995, 10, 1962, 513, 534, 558, 1981, 1835, 543, 1847, 606, 1854, 531, 509, 556, 2077, 2061, 960, 1853]. Component [630, 285, 1786, 590, 1489]. Components [1436]. Composite [719, 1455]. Composition [1865, 2005, 1869, 1328, 1875, 491]. Compositional [1027, 839, 693, 768, 774, 535]. Compositionality [1867]. Compound [1653, 43, 51]. Compound-Reaction [1653]. Comprehensive [1765]. Compressed [1365]. Compressibility [508]. Compressible [1152]. Compression [2065]. Computability [1163, 1390]. Computable [316]. Computation [1161, 390, 823, 459, 1114, 79, 792, 1820, 2024, 431, 585, 1844, 916, 1467]. Computational [1845, 1130, 1728, 411, 606, 1938, 1009, 1166, 1154, 1853, 1001]. Computations [786, 1329, 1099, 1150, 2054, 438]. Compute [1324]. Computer [1228, 218, 1698, 1381, 1724, 214, 755, 113, 1125, 1514, 1151, 122, 915]. Computer-Based [1724]. Computerizing [1718]. Computers [1113, 215, 1153, 388, 392, 1323, 1121, 1085, 1218, 677]. Computing [1128, 1105, 1086, 529, 1586, 507, 1089, 584, 1141, 199, 127, 937, 1023, 401, 523, 417, 1972, 1367, 797, 1140, 1625, 1716, 1208, 344, 99, 108, 1857, 1190, 120, 888, 1188, 1063]. Computon [324]. Concatenation [1335]. Concentrated [401]. concentration [1045]. Concentric [1107]. Concept [533, 126, 1320, 1405, 466, 881, 1434, 1696, 1465, 353]. conception [1712]. Concepts [117, 1455]. Conceptual [1396, 1400, 1527, 1427, 2017, 49, 1695, 1456, 1511, 1487]. Conclusions [603, 1394]. Concrete [802]. Concurrency [454, 705, 1907, 414, 1198, 2011, 195, 121, 156]. Concurrent [1865, 835, 1784, 1212, 846, 564, 1529, 847, 793, 710, 156, 791, 1203, 655, 1682, 2079, 770]. Condition [696]. Condition/Event [696]. Conditional [2034, 668, 1650, 933]. Conditions [185, 144, 695, 684]. CondLP [1688]. Conducting [1124]. Cone [1609]. Cone-LP’s [1609]. Conferencing [495, 1425]. Confidence [524]. Confidentiality [1946]. Configurable [1513]. Configurations [27]. Configuring [1412]. Conflict [294, 901, 1777]. Conflicting [1667]. Confluence [800, 1337]. confluent [1276]. Conformance [1789]. Conformationally [1382]. Confusion [588]. Congestion [479, 485]. Congress [1579]. Congruence [1896]. Congruences [803]. Conjecture [556]. Conjugate [402, 1162, 196]. Conjugate-Gradients [1162]. Connected [1172, 29]. Connection [662, 658, 461, 198, 570, 1893, 1281]. Connection-Oriented [461]. connectionist [355, 354]. connectionist/ symbolic [355]. Connectionless [476]. Connections [463, 1274]. Connectivity [1439, 1838]. Connectors [1438]. Consensus [2074, 1380, 525]. Consequences [533, 1807, 872].

Consideration [1422]. Considerations [1933, 290, 1457]. Consistency [634, 1643, 310, 1479]. consistency-checking [1479]. Consistent [1971, 1885, 632]. Constable [1902]. constancy [965]. Constant [813]. Constrained [1614, 1097, 1964, 1967, 1674, 834]. Constraint [204, 1543, 597, 205, 206, 1544, 207, 1226, 1945, 35, 1859, 1968, 833, 1585, 711, 981, 712, 211, 236, 1287, 1458]. Constraint-Based [207, 35, 1585]. Constraint-Directed [1543]. Constraint-Oriented [712]. Constraints [2033, 1855, 1944, 1923, 1569, 208, 1615, 209, 596, 1542, 561, 751, 276, 1664, 750, 1292, 780, 972, 1469, 1472, 1006, 1007, 1072, 978]. Constructed [1334]. Constructing [1841, 1381, 1927, 1444, 238]. Construction [1397, 695, 1258, 2068, 714, 709, 91, 773, 818, 1362, 373]. Constructions [2057, 2010]. Constructive [1562, 611, 1459]. Constructs [1225, 1198]. Consulting [1733]. Contact [31]. containing [1061]. Containment [1354]. Content [1724, 1541, 1475]. Content-Balanced [1724]. Content-Based [1541]. Contention [500]. Contest [1310, 23]. Context [2037, 14, 1530, 2032, 310, 2039]. Context-Driven [1530]. Context-Free [2037, 2039]. Context-Sensitive [2032]. Contexts [1684]. Contextual [1871]. Continuations [1199]. Continued [1599]. Continuous [1614]. Contour [933, 973, 937, 993, 908]. contours [936, 911, 1012]. Contract [318, 321]. Contraction [190]. Contractum [1312]. contraintes [1727]. Control [158, 1929, 314, 1228, 1815, 757, 231, 502, 118, 479, 130, 485, 696, 1556, 141, 666, 568, 202, 465, 1563, 503, 883, 667, 1891, 484, 635, 492, 1464, 1067, 1038, 969, 1035, 1036, 1047, 1052, 1481, 1045, 1069, 1068, 1061]. Controled [1868]. Controlers [1870].

Controlled [1029, 76, 811]. Controllers [108, 1043, 1062]. Controlling [1521, 595, 1051]. Controls [67, 1030]. Controversial [1809, 1762]. Convergence [806]. Conversion [1293, 1304]. Convertible [1265]. Converting [1812, 1814]. Convex [1603]. Convexity [1616]. Convolution [386]. Convolutions [1167]. cooperatifs [1712]. Cooperating [306, 329, 1888]. Cooperation [320, 1779, 886, 1770, 887, 2081, 1712]. Cooperative [1709, 859, 20, 326, 1452, 275, 1221, 1559, 1474, 1714]. cooperatives [1714]. Coordinated [850, 847]. Coordinating [839, 854, 855]. Coordination [835, 832, 836, 837, 1225, 856, 838, 858, 841, 319, 844, 843, 845, 848, 863, 849, 1768, 852, 453, 834]. Coprocessor [162]. Copyright [668]. Corner [1290]. corpora [377]. Corpus [1552, 361, 373]. Corpus-Based [1552, 373]. Correct [644, 652, 722]. Correcting [344]. Correction [467]. Correctness [846, 1661, 1356, 613]. CORRELATE [1101]. Correlation [343, 77, 344]. correlations [1259]. Correspondence [510]. correspondences [936]. COSPAN [1041]. Cost [1949, 567, 450, 1982, 711, 1603, 1757]. Cost-Sensitive [1982]. Counterfactual [1547]. Counting [1833, 1828]. Coupling [626, 1811, 1321]. Course [1732, 1748]. Covariant [2001]. Cover [1855, 1606]. Covergence [1852]. Covering [897, 1602, 1892]. Coverings [1895]. Crafting [830]. Craftsmanship [818]. CRAY [382, 1162]. CRAY-T3D [1162]. crease [924]. Creating [1430]. Creation [712]. Creative [1432]. Crisis [132]. Criteria [2076, 1493, 1501]. criterion [963]. Critical [1793, 800, 620, 133, 1794, 1782, 1358, 635]. Cross [137]. Cross-Classification [137]. Crossing [1909, 34, 47]. Crowd [257].

Cryptanalysis [336, 76, 1253, 80, 347].

Crypto [67, 66]. Cryptographic [1257, 86, 125]. Cryptography [2057, 62, 537, 83, 85]. Cryptosystem [1252, 75, 69, 330, 1238]. Cryptosystems [1863]. Crystal [1383]. Crystalline [383]. CS [1191]. CS-2 [1191]. CS2 [1114]. CSCL [1708, 1713]. CSCW [660]. cTc [709]. cTLA [709]. cTLA-Specifications [709]. Cube [434]. Cubes [518]. Cubic [1959]. Cubicscope [232]. Cultural [1429]. curriculum [1736]. curvature [925, 951]. Curve [1863, 920, 936, 926]. curved [995]. Curves [31, 1290, 1959, 1238, 920, 1023, 924]. curvilinear [957]. Customizable [3]. Customized [52]. Cut [1607, 614]. Cuts [1607]. Cutting [1610]. Cycle [154]. Cyclic [511, 1274].

D [1016, 428, 559, 255, 559, 931, 1022, 11, 388, 1104, 944, 17, 962, 20, 1023, 24, 904, 416, 227, 251, 895, 40, 252, 905, 250, 51, 1308]. D-ABDUCTOR [51]. d-D [559]. DA-Soc [1223]. Dags [30]. Danish [398]. dans [1707, 1736, 1714]. d’apprentissage [1712]. DART [1887]. Dashed [1296, 1310, 1309]. Dashed-Line [1309]. DASME [107]. Data [731, 2065, 1194, 192, 1958, 128, 1163, 738, 6, 166, 1586, 1102, 1589, 171, 189, 629, 173, 1160, 1951, 2006, 1092, 725, 209, 1532, 754, 137, 130, 269, 1541, 57, 1565, 717, 753, 1210, 1522, 427, 191, 188, 1454, 181, 1213, 503, 1404, 2066, 49, 178, 54, 1955, 1063, 931, 1459, 1488, 1509, 1016, 1492, 1460, 1491, 1462]. Data-Driven [269, 1459]. Data-Flow [171, 1210]. data-intensive [1063]. Data-Parallel [1102, 427, 188, 178]. Database [731, 720, 747, 206, 1949, 744, 1953, 1528, 1945, 750, 2017, 1956, 211, 1196, 1091, 1093, 1485, 1488, 1475, 1496, 960]. Databases [1090, 1952, 205, 737, 1589, 1538, 2016, 735, 1574, 1406, 1944, 1092, 1948, 2018, 1521, 755, 596, 1664, 1588, 729, 730, 736, 1524, 734, 733, 1509, 1454, 1461, 1497, 1512]. Dataflow [2050, 782]. Datagram [484]. Datalog [569, 1494, 541]. Datasets [1168]. Dates [1618]. DBMS [1942, 740]. DBMSs [721]. DBS3 [748]. DCI [1459]. deadline [1472]. Deadlocks [1790]. Dealing [759]. deblurring [1013]. Debug [653]. Debugger [831]. Debugging [1201, 1180]. Decade [1818]. Decidability [456, 1337, 1918]. Decidable [2037, 2038, 1889]. Deciding [1055, 2062]. Decimation [71]. Decision [546, 1901, 530, 326, 273, 1912, 1925, 1570, 1940, 1575, 1686, 289, 1148, 1093, 1490, 1512]. Decision-Making [273]. Decision-Theoretic [326]. Decisions [1660, 1667]. Declarative [769, 610, 854, 614, 1508]. Declustering [755, 753]. Decoding [1255]. Decompiled [814]. Decomposing [430]. Decomposition [1073, 1146, 1159, 413, 1329, 521, 926, 1394, 48, 1840, 1145, 1156]. Decompositions [1343]. Decorated [2027]. Decoupling [962]. Dedicated [1730]. Deduction [1897, 633, 561, 1959, 1920]. Deductive [701, 575, 571, 1502, 1496]. Deep [1680]. Default [1634, 1683, 603, 1672, 1674, 1679, 1277, 1470, 1492]. Defaults [1677]. Defeasible [1670]. Defense [442]. Definability [1907, 1916]. Defined [1630, 1334, 1484]. Defining [1666]. Definition [1186, 662, 1343, 1876, 209, 1885, 1765]. Definitional [615]. Definitions [1908, 1561]. Deflection [1824]. Deforestation [780]. Deformable [243, 931, 1011, 908]. deformations [953]. DEGAS [1517]. Degree [1605, 26, 1325]. Delays [559]. Delegating [830]. Deliberative [118]. Delivering [1750]. delta [1980]. deltaP2 [514]. Demand [175, 470, 503, 1580, 493]. Demand-Driven [175]. Demand-Priority [470]. Demoralization [1680]. Denotational [1913, 773, 804]. Dense [1973, 1000, 2028, 940]. Density [1855, 257, 933]. Deontic [1472]. Department [442]. Dependable [134]. Dependence [1773, 187]. Dependences [810]. Dependency [799, 1477]. Dependent [1621, 2066]. Deployment [1818]. Depth [1844, 940]. Depth-Efficient [1844]. Derandomization [1961]. Derandomize [2052]. Derivation [564]. Derivations [1315]. Derived [731, 251]. Deriving [1542]. Describe [1447]. Describing [103]. Description [1163, 94, 1065, 362, 2080, 1455]. Deserve [1438]. Design [649, 483, 720, 1539, 651, 160, 593, 1089, 392, 1600, 119, 1139, 1034, 1723, 1873, 494, 1744, 1446, 1528, 224, 1711, 499, 542, 105, 1760, 582, 1742, 1964, 1693, 1537, 1780, 305, 1701, 750, 1181, 1118, 108, 1444, 156, 116, 342, 1702, 2017, 1304, 1443, 460, 1891, 1196, 635, 1449, 1751, 1485, 1498, 1043, 1025, 1045, 1061, 928]. Design-Centered [1742]. Design-Level [1446]. Design/CPN [1873]. Designated [1246]. Designing [320, 1419, 1873, 1729, 587, 2015, 1849, 669, 848, 1886, 1774]. Designs [1785, 1447, 311, 581]. Desire [289]. Desktop [1424]. Desktop-Based [1424]. Despeckling [1095]. Detail [14]. Detected [243]. Detecting [1290, 949]. Detection [1296, 1090, 1307, 1790, 1291, 294, 1310, 767, 683, 1201, 227, 243, 1309, 895, 698, 247, 969, 977, 970, 926]. detectors [1003]. Determined [790]. determiners [352]. Determinism [804]. Deterministic [1995, 502, 798, 2072, 774, 667, 1837]. Deutsch [127]. d’évaluation [1717, 1725]. Developing [158, 1813, 134, 229, 1771, 1556, 1520, 1746]. Development [701, 2004, 386, 1137, 1411, 705, 1535, 1420, 739, 850, 1749, 1788, 410, 868, 1711, 876, 1532, 1945, 826, 644, 1445, 570, 1787, 140, 632, 1694, 154, 203, 619, 1818, 1358, 650, 1498]. Developments [620]. d’évenement [1696].

Device [85, 1891]. Devices [1744, 1164]. Devil [1762]. DEVS [118, 112]. DEVS-Based [118]. DFA [713]. diabetic [988]. diagnosis [1491]. Diagnostic [1577, 1396, 1729, 1070]. Diagonal [415]. Diagram [1297, 20, 1653]. Diagram-Recognition [1297]. Diagrammatic [1402]. Diagrams [546, 1447, 36, 1300, 2017]. Dialect [584]. Dialectical [1644, 1658, 1660]. dialog [363]. Dialogue [1651, 1549, 1713, 1761, 1673, 1718]. Dialogues [1728, 276]. Diatom [419]. Dictionaries [2059]. Dictionary [1385, 1368]. did [618]. Dienst [1596]. Difference [1078, 1080, 1151]. Different [160, 322, 2035, 1359, 356]. Differential [385, 101, 347, 924, 1053, 971, 957, 1491]. Differentials [331]. Difficult [1809]. Difficulties [1937, 1442]. Difficulty [1849, 1254]. Diffie [1262]. diffusions [1029]. Digital [1579, 536, 1586, 1593, 1133, 1594, 1592, 1290, 213, 1583, 1587, 1582, 1267]. Digraphs [43]. Dimension [1599]. Dimensional [2, 1416, 1387, 1988, 386, 398, 208, 900, 1386, 125, 791, 919, 910]. Dimensioning [1305]. Dimensions [22]. Dining [1799]. Diophantine [788]. Diploid [884]. Dipole [1108]. Direct [1153, 999, 222, 971, 433]. Directed [1543, 9, 318, 1389, 1741, 1823, 33, 201, 48, 1019]. direction [923]. Directions [974, 1467]. Directive [1393]. DIS [147]. Disagreements [1680]. Disambiguation [2008, 184, 358]. DISCO [206]. disconnection [1276]. Discontinuities [219, 940]. Discourse [274]. Discoveries [1367]. Discovering [1377, 1497]. Discovery [1589, 1406, 1948, 1592, 729, 1509, 1454, 1512]. Discreet [1245]. Discrete [89, 1931, 77, 1108, 696, 96, 109, 1051, 679, 93, 1059]. Discretization [1948]. Discrimination [260]. Disjoint [1829]. Disjunctive [768].

Disks [10]. disparities [939]. Dispatch [829]. Dispersion [1966, 394]. displacements [982]. Dissemination [1955]. Dissemination-Based [1955]. Distance [2049, 1856]. Distributed [1779, 2063, 387, 1104, 392, 150, 1593, 699, 659, 1202, 1147, 658, 1215, 1195, 1173, 831, 2015, 151, 1401, 1226, 444, 1783, 411, 1231, 1711, 505, 489, 1532, 1945, 1878, 416, 179, 1980, 176, 149, 1795, 417, 644, 2014, 1445, 329, 666, 246, 1135, 742, 1140, 1165, 1798, 833, 1533, 147, 324, 486, 191, 743, 1279, 1181, 1982, 1213, 585, 1857, 854, 1818, 1417, 460, 1331, 433, 855, 663, 436, 1559, 1765, 1188, 1145, 1777]. Distributed-Memory [387, 179, 176, 191, 433]. Distributing [645]. Distribution [192, 442, 1223, 1168, 253, 190, 1261, 1848, 967]. Distribution-Based [253]. distributions [1488]. Division [482]. DLAB [1508]. DMEROON [1213]. DMM [2071]. DNA [1122]. Do [1220]. Docking [1382]. Document [266, 1289, 1415, 1585, 353]. Documentation [286]. Documents [1591, 739, 1398, 733, 1416]. Does [1810, 891]. Doing [741, 1760]. Dollars [67]. Domain [1929, 1073, 1413, 1639, 1078, 413, 276, 1747, 115, 451, 1764, 1722, 1080, 1151, 1145, 368]. Domain-Specific [1747, 1722, 368]. Domains [1543, 1397, 1729, 1762, 1510]. DOOD [723]. Double [1377]. Doubling [1260]. Dovetailing [808]. Down [1308, 650]. DQDB [484]. DQDB-ATM-DQDB [484]. Draft [1186]. Drago [1798]. Draw [38]. Drawing [4, 7, 9, 12, 17, 22, 23, 26, 27, 33, 42, 43, 44, 48, 1340]. Drawings [5, 1989, 1307, 28, 1298, 1830, 1292, 1831, 1306]. Driven [762, 175, 1530, 1698, 282, 269, 1395, 1459, 375]. Driver [647, 1010]. DSS [1264]. Dual [1608, 1138, 121]. Duality [983]. d’un [1712, 1717]. d’une [1717]. duplication [1272]. Duration [704]. Durations [770].

During [750, 1695, 954]. Dutch [285]. DynaDAG [41]. Dynamic [1979, 2016, 1232, 739, 831, 413, 1978, 472, 753, 222, 191, 2046, 1816, 1947, 2073, 375, 676, 1044, 929, 1071]. Dynamical [1929, 125, 1038, 1056, 1069, 1033]. Dynamically [1504, 829]. Dynamics [551, 1102, 388, 1104, 1100, 412, 417, 422, 1094, 427, 1125, 1084, 1517, 1154, 437, 1009, 934].

E-Unification [1904]. each [1462]. Earley [827]. Early [683]. Ears [1954]. Easy [1612, 2048, 880, 116]. EaSy-Sim [116]. EC [1186]. ECEPP [1121]. ECEPP-Based [1121]. ECLIPS [157]. Ecological [561]. economics [1459]. Edge [1332, 885, 243, 969, 970]. Edges [1, 1976, 1857, 977]. EDI [57, 56]. Editing [39, 1388]. Editors [36]. Education [1739, 1720, 1710, 1431]. Educational [1429, 1430, 1590]. Effect [6, 407]. Effective [2030, 2016, 1784, 825, 1760, 1705, 1915, 2036, 805, 1757]. Effectively [298]. Effectiveness [591]. Effectors [1930]. Effects [322, 815, 243, 1199, 1079]. Efficiency [1102, 1392, 184, 1393, 979]. Efficient [2074, 2070, 1170, 690, 1243, 601, 688, 1195, 1364, 1255, 1994, 1365, 408, 1384, 579, 1198, 1206, 179, 1863, 824, 1811, 472, 776, 1991, 2024, 1561, 585, 1837, 190, 1831, 728, 1567, 1844, 1388, 926, 960]. Efficiently [731]. EFSMs [703]. Eigen [932]. Eigenfaces [906]. Eigenvalues [1848]. Elastic [436]. Elastically [1011]. Elections [1240]. Electrically [1124]. Electro [1126]. Electro-Magnetic [1126]. Electromagnetic [382, 1078]. Electronic [55, 869, 57, 1775, 1583, 59, 1595]. Electrostatic [1109]. Element [1127, 1083, 1840, 1150]. Elements [1123, 949]. ElGamal [1235, 80]. Elicitation [1409, 1402]. Eliminating [2007]. Elimination [1017]. ELINOR [1595]. Elliptic [1863, 1238]. ELM [1719].

ELM-ART [1719]. elsewhere [1280]. Embed [8]. Embedded [653, 598, 370, 1158, 642, 1802, 864, 291]. Embedding [29, 520, 1415, 1837]. Embryonics [874, 873]. Emergence [1691]. Emerging [1188]. Emotions [314]. Emphasis [782]. empirical [1493]. Enabling [143]. eNCE [1336, 1325]. Enclosed [544]. Encoded [1291]. Encoding [281, 764, 1994]. Encounters [1227]. Encryption [333, 74, 65, 1239]. End [666, 1393, 464, 484, 1046]. End-to-End [666, 464, 484]. Endomorphic [118]. Energy [1105, 1106, 930]. Enforcement [62]. Engagement [1770]. Engineering [152, 1307, 1733, 869, 403, 1414, 1529, 860, 1881, 1519, 1140, 626, 1298, 1732, 113, 1514, 1764, 1306, 818, 1510, 1452, 1764]. English [374, 352]. Enhance [808]. Enhanceability [1536]. Enhanced [80, 667, 574, 907]. Enhancements [461]. Enhancing [838, 1838, 1486, 1573]. enseignant [1727]. enseignement [1745]. Entailment [1679]. Enterprises [60]. Entropy [1614]. Enumeration [1915]. Environment [1737, 1738, 1219, 1708, 2004, 1779, 2063, 856, 705, 1141, 1729, 442, 327, 1231, 1711, 489, 1741, 488, 847, 633, 1122, 1513, 418, 447, 1713, 1425, 1750, 1780, 109, 1118, 116, 1514, 1169, 1763, 712, 59, 63, 53, 261]. Environmental [1923, 976]. Environments [1428, 1740, 1229, 2015, 744, 1744, 860, 721, 1564, 1742, 797, 1224, 1165, 223, 1767, 1424, 1436]. ENVISAT [145]. ENVISAT-l [145]. Epidemic [730]. Epidemiological [729]. Epiphyte [1730]. Epipolar [236, 237, 972]. epipole [978]. Epistemological [1716]. EpiTalk [1730]. Equality [604, 98]. Equation [1247, 1249, 1134, 431]. Equational [1959, 1355]. Equations [385, 788, 101, 1127, 1237, 1126, 1507, 924]. equipped [162]. Equitable [668]. Equivalence [802, 776, 1889]. Equivalences [2027, 803, 1917]. Equivalent [1238]. Equivocable [1244]. Erlang [1193]. Error [698, 344, 467, 926]. Error-Correcting [344]. Errors [2034, 767, 245]. ESAUNA [1087]. Escrow [64, 69, 1254]. ESM [1329]. ESSAI [474]. estimates [991]. Estimation [221, 219, 224, 220, 99, 242, 257, 969, 977, 913, 971]. étrangère [1725]. Euclidean [250, 975, 905]. Euler [1155, 1152]. Euler/ Navier [1155, 1152]. Eulerian [398]. EUROPA [1186]. Europe [66, 2000]. European [1131, 475]. Evaluating [2075, 591, 2082]. Evaluation [762, 162, 764, 469, 1176, 1142, 1289, 1160, 859, 1644, 1949, 1087, 1198, 660, 1309, 626, 309, 259, 822, 1748, 460, 484, 734, 1726, 562, 1034, 1503, 1458]. Evaluator [1570]. Event [1327, 96, 109, 804, 491, 1051, 93, 994]. Events [719, 1545]. ever [2000, 974]. Evidence [1633]. Evidential [1926]. Evolution [877, 1130, 1942, 1691, 886, 893, 876, 580, 887, 875, 323, 872, 1515]. Evolutionary [877, 1769, 869, 1466, 878, 899, 111, 894, 1467, 867, 1465]. Evolutive [103]. Evolvable [871, 870]. Evolve [891]. Evolving [1898, 1905, 445, 1463]. Exact [167, 1839, 34, 1267]. Example [653, 1740, 884, 959]. Examples [1037, 1998, 526, 525, 122, 365, 379]. Exchange [1616]. Exchanges [272]. Executables [831]. Executing [688]. Execution [748, 1923, 1516, 1878, 179, 104, 1135, 188, 1776]. Executive [133]. Exercice [1751]. Exercise [1356, 1702]. Existence [532, 1256]. Existential [2047]. existing [372]. Expansion [1554]. Expect [1955]. Expected [1874, 531, 1604]. Experience [1211, 2004, 1089, 1556, 1519, 1214]. Experiences [1086, 299, 1813, 1111, 1100, 1817, 474, 404, 642, 1131, 1083, 284, 1203]. Experiment [1105, 592, 1909, 157]. Experimental [167, 9, 481, 1281]. Experiments [1778, 654]. Expert [1556, 1747, 109, 1757, 1777, 1461]. experts [1456]. Explaining [1499]. Explanation [1740, 1729, 1474, 1499]. Explanations [1750]. Explicit [2008, 832, 1400, 629, 485, 582]. Exploiting [2016, 1406, 744]. Exploration [1188, 1449, 1466]. Exploring [1744, 311, 1433]. explosion [2079]. Exponent [1234]. Exponential [1834]. Exponentiation [1248, 79]. Exponents [1262]. Export [67]. expressed [967]. Expression [279]. Expressions [2034, 2050, 728]. Expressive [2047, 783]. Expressiveness [204, 1918]. Expressivity [545]. Extended [1952, 1600, 1202, 1032, 1209, 792, 616, 1022]. Extending [1551, 144, 14, 1808, 611, 702, 1947]. Extensible [720, 850]. Extension [600, 1734, 256, 1798]. Extensional [1906]. Extensions [726, 607, 107, 1850, 1918]. Extracting [1400, 351, 957, 525]. Extraction [1303, 791, 1294, 122, 909, 365, 978, 993, 367, 368].

Face [227, 253, 945]. Faces

[1421, 253, 1437]. Facets [1627]. Facilities [1952]. Facility [1966, 1600, 1179]. Factoring [1249, 1238]. Factorization [2041, 519, 1024, 429]. Factors [1429, 1422]. Factory [705]. Failure [291]. Fair [462, 472]. Fairly [70]. Fairness [473, 2031, 471]. FALCON [183]. Fallacies [1655]. Families [1237]. Family [266, 874]. Fast [1376, 11, 343, 77, 1257, 2072, 1108, 1248, 79, 2028, 1986, 916, 717, 1117, 338, 45, 1840, 178, 471, 1842, 734, 339, 1007, 1291]. Faster [1363, 1990, 346, 827, 1963]. Fatal [2034]. Fault [1577, 522, 134, 1323, 1783, 871, 683, 523, 1798, 1210, 637, 888, 1044]. Fault-Tolerance [1783]. Fault-Tolerant [522, 871, 523, 1798, 637, 888]. Faulty [2071]. Faulty-Memory [2071]. Fc2TOOLS [706]. FDDI [470]. FDR [692]. Feasibility [1983]. Feasible [1107]. Feature [414, 1303, 236, 1003, 921, 1012]. Features [1806, 863, 498, 1540, 1484]. Feedback [1551, 1629, 1608, 1256, 1756, 1037]. Feistel [342]. FEM [383, 1149]. FEM-Tool [1149]. Few [537]. Fibrational [608]. Fibred [1270]. Field [418, 1520, 874, 866, 997, 955]. Fields [1207, 220, 1237, 974]. FIFO [1869]. Figures [1832]. File [595]. Files [732]. Fill [1976, 2043]. Fill-In [2043]. Filled [1976]. Filling [1936]. Filter [1931, 345, 1010, 928]. Filtering [490]. Filters [501, 970]. Filtration [1366]. Final [630, 126, 793]. Financial [1957]. Finder [212]. Finding [1979, 1973, 1836, 1826, 1247, 1249, 1605, 1015, 1604, 244, 253, 1842]. Fine [534, 1201, 743, 590, 1465, 1707]. Fine-Grain [1201, 590]. fine-grained [1465]. Fine-granularity [743]. Finger [1365]. Finger-Print [1365]. Fingerprinting [1241]. Finite [516, 1899, 1390, 788, 1258, 1127, 1908, 75, 1078, 125, 1083, 2061, 1840, 1080, 1151, 1150, 1494]. Finite-Difference [1080]. Finite-Element [1840]. Finite-State [1390]. Finiteness [2062]. fire [1495, 387]. First [204, 1894, 1086, 1903, 1639, 1547, 141, 1662, 1199, 2047, 1438, 736, 679, 1455, 1511]. First-Class [1199, 1438]. First-Order [204, 1894, 1639, 1547, 1662, 736]. Fischer [1072]. Fisherfaces [906]. Fisheye [25]. FISSURE [212]. fixation [962]. Fixed [1903, 2045, 1916, 1269]. Fixed-Point [1916]. Fixing [692]. Fixpoint [815]. Fixpoints [2048]. Fledged [1416]. Flexible [844, 1698, 1231, 1382, 56]. Flip [548]. Flip-Flop [548]. FLO [1477]. Flop [548]. Flow [1601, 166, 1232, 171, 119, 219, 1605, 405, 1087, 771, 130, 485, 220, 1210, 1603, 1099, 1322, 976, 977, 997, 979, 925]. Flows [1153, 925, 431, 1531]. Flowshop [1624]. Fluid [431, 1084, 1154]. Fluids [394]. Focal [250, 905]. Focus [276]. Focus-Shifting [276]. focused [952]. Focusing [1742, 1107]. Foiling [1260, 87]. Folding [14, 604, 1383, 2077]. fondée [1696]. Forbidden [551]. Force [9, 33, 1286]. Force-Directed [9, 33]. Forecast [410]. Forest [1495]. FORGE [1119]. fork [1199]. Form [823, 825, 395, 1294]. Formal [649, 653, 1396, 738, 621, 546, 630, 697, 2075, 1773, 1792, 1441, 1648, 625, 1876, 1448, 1047, 105, 640, 1330, 1661, 421, 626, 619, 613, 1726, 1065, 1472, 2080, 1046]. Formalising [318, 1410, 288]. Formalism [849, 1508]. Formalization [1672]. Formalizing [1525, 627, 625]. Formation [1405, 328, 1927, 1767, 1725]. Formatting [277]. Formed [530]. Forms [529, 267, 1117]. formula [1468]. Formulas [1839]. Formulations [1600]. Formulator [281]. Forrealization [1373]. Fortran [1177, 1183, 399, 1812, 176]. Forward [485, 467]. Foundation [314]. Foundations [1921, 1637, 308, 1216]. Foveated [230]. FPGA [866]. Fraction [1599]. Fragments [690]. Frame [1548, 233, 980, 983]. Frame-Rate [233]. Framework [2005, 1900, 1766, 2075, 1637, 1733, 850, 1923, 756, 1792, 1638, 567, 1648, 845, 444, 1448, 1938, 696, 610, 1318, 864, 794, 464, 657, 1515, 1757, 1765, 1276, 370, 679, 1492, 1059]. Frameworks [560, 2035, 1917, 1442]. FRD [1490]. Free [2037, 21, 2009, 431, 1272, 2039, 2044]. Freedom [62]. Frenchmen [541]. Frequency [898, 480]. Friendly [652]. Front [409, 1046]. front-end [1046]. Full [806, 1588, 1803]. Full-Text [1588]. Fully [1416, 411, 1913, 1513, 698]. Fully-Fledged [1416]. Function [2007, 897, 550, 335, 1633, 1844, 339, 967]. Functional [762, 1952, 1307, 1553, 579, 100, 1356, 202, 611, 2003, 779, 615, 1497]. Functional-logic [611]. Functionals [1902]. Functions [533, 78, 1258, 629, 1257, 1835, 73, 98, 611, 540, 880, 1259]. Fundamental [758, 916, 983]. Funds [1813]. Further [2020]. Fusion [1525, 1423, 1934]. Future [620]. Fuzzy [1866, 1588, 1941, 1471, 1481, 1496, 1482]. Fuzzy-Timing [1866].

G0n [1353]. G0n-Logics [1353]. Gabidulin [1252]. gain [1067]. Galaxies [1130]. game [1025, 1483]. game-theoretic [1025]. Games [295, 695, 1915, 700, 1853]. Gamma [839, 602]. GAMMA-like [602]. Gap [2018, 501]. Garoña [129]. Gas [130, 881]. Gate [874, 866]. Gateway [1873]. Gaussian [253]. GD [12]. GD-Workbench [12]. Gene [1374]. General [1832, 527, 1105, 1297, 1038, 193, 1139, 1639, 1568, 303, 416, 96, 1767, 94, 1722, 370, 1023]. General-Purpose [1105]. Generalised [972, 1393]. Generalising [2019]. Generalization [325, 1576, 573]. Generalizations [716]. Generalize [1435]. Generalized [1552, 205, 25, 1908, 510, 107, 538, 953, 125, 1620, 1603, 813, 1271]. Generalizing [968]. Generated [1338]. Generating [1235, 513, 283, 1759, 1336, 280, 1713, 40, 352, 52]. Generation [1551, 1313, 1133, 271, 268, 477, 285, 2006, 1004, 410, 273, 1549, 282, 214, 269, 1863, 286, 819, 277, 123, 186, 1862, 278, 284, 178, 691, 1278, 165]. Generator [1255, 279, 1550]. Generators [345, 2055]. Generic [701, 4, 466, 1699, 1746, 51, 909]. Genericity [512]. Genes [1384]. Genetic [896, 879, 1133, 892, 946, 882, 889, 898, 881, 900, 895, 901, 883, 902, 880, 884, 1464, 357]. Gentzen [2033]. Gentzen-style [2033]. Genus [1836]. Geographic [726]. Geometric [936, 1828, 895, 1292, 121, 1007, 907, 957]. Geometrical [1636]. Geometry [1739, 771, 1609, 249, 236, 237, 1021, 904, 915, 981, 983]. geons [908]. Get [62, 618]. Gibbs [242]. GIS [726]. Given [1836, 1822]. Giving [1551, 1731]. GLASS [659, 489]. GLASS-Studio [489]. Global [766, 989, 1969, 65, 681]. GMRES [1155].

GNARL [1801]. GNAT [142, 1803]. Go [1997, 1810, 1571]. Goal [318, 1019]. Goal-Directed [318, 1019]. Goals [1649, 142, 288]. Gödel [1894]. GOLDRUSH [1091]. Good [1849, 1143]. Gossiping [1981]. GOTO [1810]. GOVE [53]. GRACE [1317, 1319]. GRADE [36]. Graded [1665]. Gradient [402, 196]. Gradients [1162]. Grain [1201, 1210, 590]. Grained [874, 1465]. Gram [1159]. Grammar [1328, 1323, 264, 1321, 53, 362, 351]. Grammar-Oriented [53]. Grammars [1315, 1327, 1336, 1330, 1339, 1325, 1804, 1340, 360, 359]. Grammatical [281, 280, 350]. Granularity [1752, 567, 202, 743]. Graph [1311, 816, 1952, 1312, 3, 5, 1976, 1314, 6, 9, 1333, 11, 12, 13, 14, 1315, 1327, 1328, 17, 1323, 1324, 23, 1344, 1320, 1384, 1606, 1838, 1206, 35, 36, 264, 1348, 1330, 2045, 39, 1318, 833, 521, 1321, 42, 1339, 1316, 44, 2046, 1319, 190, 1325, 50, 51, 1331, 95, 1892, 1322, 1340, 1341, 13]. Graph-Drawing [23]. Graph-Like [36]. Graphic [1299, 1828, 36]. Graphical [587, 105, 1298, 99, 632]. Graphically [1657]. Graphs [1, 2, 1342, 559, 1313, 1989, 1352, 1841, 786, 10, 2038, 15, 2044, 814, 16, 1343, 1334, 22, 1344, 1335, 24, 1631, 26, 25, 1389, 1608, 543, 29, 520, 31, 1823, 32, 33, 174, 1980, 2043, 37, 247, 40, 1138, 1830, 2039, 1837, 48, 1848, 1340, 1456, 1046]. Greedily [1973]. Grid [559, 1987, 410, 29, 1939, 37, 1830, 1120, 1099]. Grid-Based [1120]. Gröbner [2042]. GROMOS87 [1094]. Ground [141, 985]. Grounding [382]. Group [515, 1409, 1344, 1207, 656]. Group-Based [1207]. Grouping [483, 935, 936, 949]. Groupoids [516]. Groups [440, 2039]. Groupware [1426]. Grows [869]. Grows/ Evolves [869]. Guarantees [464, 1619]. Guarding [1861]. Guidance [1715, 1713, 258]. Guide [1715]. Guidelines [1525]. Guiding [570].

H [1067]. Handling [1527, 604, 1222, 1585, 503, 1804, 291, 1488]. Handwriting [267, 254]. Handwritten [1117]. Hard [160, 1612, 532, 1544, 2040, 872, 880, 1006]. hardly [974]. Hardness [533, 1971, 1969]. Hardware [1872, 876, 871, 1362]. Harmonic [1835]. Harness [1812]. Harnesses [434]. Hash [78, 335, 742, 339]. Hashing [552, 1373, 338, 907]. Haskell [761]. Having [2038, 1701]. HCSD [1788]. Head [282, 999]. Health [60]. Heap [198]. heart [986]. Heidi [486]. Heidi-II [486]. Height [517]. HELENA [1430]. Helix [588]. Hellman [1262]. Help [1419, 1398, 1750, 376]. Helping [1741]. Hessenberg [395]. Heterogeneity [501]. Heterogeneous [1589, 1507, 401, 137, 194]. Heuristic [1546, 897, 128, 1602, 2075, 34, 1986, 45, 682]. Heuristics [1160, 1984, 1619]. Hexahedral [544]. HFE [1237]. Hidden [1826, 1237, 902, 1497]. Hierarchic [1172]. Hierarchical [920, 1229, 1558, 1785, 1447, 1344, 1759, 1935, 1876, 1045, 1888, 45, 585, 1511, 989, 914, 371]. Hierarchically [1331]. Hierarchies [1752, 578, 1542, 779]. Hierarchy [1833, 547, 2029, 172]. High [381, 114, 1194, 1105, 1106, 1113, 1870, 333, 1950, 726, 747, 584, 1249, 1141, 26, 411, 1875, 745, 754, 106, 1866, 1887, 195, 213, 464, 1148, 646, 1190, 120, 1306, 228, 909]. High-Bandwidth [333]. High-Energy [1105]. High-Level [106, 1866, 1887, 195, 213, 646]. High-Performance [1194, 1113, 120, 228]. High-Speed [464]. High-Throughput [747]. Higher [840, 609, 793, 1355, 1356, 852, 778, 1361, 791]. Higher-Order [840, 1355, 1356, 852, 778, 1361]. Highly [401, 1698]. highway [1034]. histograms [955]. History [1090, 1549]. Hitting [2052]. HKM [330]. HKM/HFX [330]. Hoisting [185]. Homomorphisms [1242]. HOOD [161, 1785, 151, 165, 157]. Hooking [1416]. HOPLa [844]. Horizontal [1291]. Horn [566, 598, 572]. Hot [1158]. Hough [221, 926]. house [909]. HPC [1189, 1076, 1124, 1107]. HPCN [1188]. HPCN3S [1088]. HPF [1176, 1184, 1111, 177]. HRNCE [1336]. HRT [161, 1785]. HRT-HOOD [161, 1785]. HSIIP [1734]. Human [1705, 1008, 1706, 1716, 253, 1422, 1677, 379]. human-made [379]. Humans [1218]. Hybercube [1115]. Hybrid [1299, 665, 1030, 1035, 1028, 111, 1064, 1150, 1037, 1063, 1027, 1067, 1066, 1038, 1065, 1043, 1029, 1036, 1048, 1054, 1056, 1044, 1047, 1057, 1025, 1062, 1058, 1069, 1068, 1049, 1061, 1046, 354, 1071]. Hybridization [1378]. Hydroelectric [641]. hyper [1278]. hyper-resolution [1278]. Hypercubes [520]. Hyperedge [1334, 1338, 1337]. Hypergraph [1965, 1336]. Hypergraphs [1622, 1337]. Hyperincursion [126]. Hypermedia [1416, 1428, 1435, 1215, 739, 1401, 1414, 1415, 1430, 1587, 1436]. hyperplanes [1484]. hypotheses [1050]. Hypothesis [555, 954]. Hypothetical [1637, 1638, 1572].

I/O [1064]. IBM [403, 1192, 1121]. ICC [584]. ICL [1091]. Idas [286]. Ideal [313, 1652, 1617, 569]. Ideals [2042]. Identification [639, 82, 90, 1263]. IDP [274]. if [1884]. IFS [892]. II [486, 2064]. illumination [965]. illusory [938]. ILP [167]. ILP-based [167]. ILP0 [1155]. ILP0-01U [1155]. Image [225, 266, 668, 1167, 214, 1924, 1418, 233, 264, 243, 950, 894, 240, 259, 222, 1116, 927, 433, 959, 242, 257, 903, 1013, 1022, 988, 976, 1004, 1007, 947, 1016, 949, 925, 953, 905, 934, 1024, 929]. image-based [947]. Image-Space [433]. imagery [909]. Images [1095, 391, 1291, 227, 895, 1023, 989, 912, 1018, 998]. Immediate [1909]. Immersive [494]. Immune [888]. Impact [1704, 1173, 202]. imperfect [1488, 1460]. Impersonation [87]. Implement [1796, 581]. Implementability [2031]. Implementation [821, 593, 1212, 1139, 394, 1197, 499, 2014, 609, 1356, 301, 1121, 1801, 1688, 1687, 647, 1116, 1214, 1118, 142, 1803, 1620]. Implementations [652]. Implemented [1193]. Implementing [1195, 329, 1802]. Implication [598, 1471, 1270]. Implications [1383, 366, 606, 464, 94]. Implicit [1908, 1152]. Imposing [1006]. Imprecise [1567]. Improve [896]. Improved [2069, 1821, 2076, 1855, 1827, 2046, 646, 1623, 1379]. Improvement [1788, 689]. Improvements [716]. Improving [185, 582, 184, 432]. Improvisation [1543]. Impulse [1874]. ImpUnity [845]. IN-Rational [517]. inclusions [1053]. incompetence [1493]. Incomplete [1998]. Incompleteness [1894]. Incompressible [1088]. Increasing [1786]. Incremental [281, 1275, 823, 644, 41, 1820]. Incursion [126]. Independence [2026]. Independent [1385, 552, 859, 419, 1764, 1502]. Indeterminacy [1953]. Index [1268, 2022]. Indexing [208, 671, 262, 734, 966, 907, 960]. Indication [485]. Indirect [825]. Individual [1228, 1569, 1730, 1424]. Induced [2044]. Induced-Subgraph [2044]. Induction [1562, 1914, 112, 572, 883, 1561, 1460, 1461, 1459, 1465]. Induction-Based [883]. Inductive [527, 1560, 1485, 1502, 374, 368, 373]. Industrial [1170, 1076, 1088, 1192, 1556, 1746, 1074, 1081, 620, 1084, 884, 1032]. Industry [1086, 1085]. Inequalities [1630, 1631, 1598]. Inexpressible [736].

Inference [527, 6, 1659, 817, 362, 350].

Inferring [1555]. Infinite

[799, 1313, 695, 2038, 174, 519, 805, 1889]. Influence [760]. Influences [1413]. Informal [630]. Information [1163, 1539, 1554, 1813, 101, 1557, 1530, 1535, 1215, 490, 1195, 20, 1432, 1222, 279, 1401, 107, 1201, 1521, 328, 1520, 1164, 1541, 61, 1536, 1522, 1533, 65, 1761, 1580, 1534, 1524, 1433, 1955, 1498, 365, 1452, 963, 1476, 1486, 1474, 1504]. Information-Replay [1201]. Informative [942]. Informix [741]. Informix/Illustra [741]. Infrastructure [65]. Infrastructures [113]. ingénierie [1717]. Inheritance [1569, 2010, 2011, 1665, 2012]. Initial [1551, 144]. Initio [437, 412]. Inproved [1912]. Input [807]. Input-Output [807]. Inputs [691]. Inquiry [1731, 1722, 1718]. Insertion [168]. Insider [87]. Inspection [124]. Instantaneous [1909]. Instantiation [1541]. Instruction [1698, 184]. Instruction-Level [184]. Instructional [280, 1693]. Integer [2069, 1601, 1897, 1249, 552, 775, 186, 780, 178, 1598]. Integrated [469, 2014, 1776]. Integrating [1872, 1923, 636, 1345, 84, 2002, 356, 624, 863, 594, 277, 1633, 650]. Integration [1147, 1134, 1949, 725, 860, 1445, 1537, 1754, 1321, 1533, 1276, 1507]. Integrity [1944, 1945, 1664, 750, 210, 1469]. Intelligence [1925, 1463]. Intelligent [1464, 299, 1719, 1411, 1193, 1642, 1720, 1729, 1697, 1715, 1164, 1746, 1706, 1750, 1580, 1590, 1756, 1700, 1436, 1717, 1475, 1504]. intensity [958]. Intensive [1738, 1063]. Intention [289]. Intentions [1772]. Inter [843]. Inter-Object [843]. Interaction [785, 322, 796, 1345, 771, 2031, 447, 324, 834, 1014]. Interactions [1709]. Interactive [1090, 1591, 11, 183, 893, 274, 669, 1529, 262, 1224, 147, 42, 503, 492]. Interchange [57, 1856, 58]. Interconnected [1059]. Interconnecting [1893]. Interconnection [1438]. Interested [1555]. Interface [1591, 1184, 639, 1954, 399, 447, 106, 591, 147]. Interfaces [1815, 2006, 409, 1425, 854]. Interferometric [85]. Interleaving [1317]. Intermediate [744]. Internal [409, 1748]. Internet [496, 1781]. Interoperability [746, 150, 862]. Interoperable [1165, 1522, 1524, 1523]. Interpolation [1271]. Interpretation [1301, 567, 256, 1939, 1300, 1559, 988]. Interpretations [768]. Interpreting [1306]. Interprocedural [1805, 171, 169, 198, 170, 172]. Interrupted [272]. Intersection [10, 37]. Intersensorial [94]. Intra [843]. Intra- [843]. Introducing [1577, 840]. Introduction [1409, 126, 878, 154, 676]. Introductions [682]. intuitionistic [1273, 1284, 1286, 1287]. Intuitions [313]. Invalidate [1178]. Invariance [226, 710]. Invariant [1383, 1061]. invariant-based [1061]. Invariants [251, 252, 1060, 958]. Inverse [892, 1846]. inverted [1071]. Investigating [859]. Investigation [1107, 435]. Investigative [1448]. Investment [2058]. Invited [610, 612, 614]. Invoked [148]. IP [1237]. IRISS [1196]. Irregular [1205, 1804, 1099, 1204]. Isa [1542]. ISAAC [334]. ISDN [463, 481, 467]. Islands [1433]. ISO [1305]. isolation [929]. Isomorphism [543, 247]. Isomorphisms [1237]. Issues [1591, 1314, 1402, 392, 305, 42, 368, 1718]. Item [1294]. Items [1299]. Iterated [1647]. Iteration [1136]. Iterations [815]. Iterative [382, 1749, 404, 415, 1882, 1787, 787, 698]. ITS-Engineering [1764]. its.Application [406]. ITSs [1716]. IWIM [835]. IWU [1893].

J [1679]. Jacobi [408]. Japanese [266]. Java [1720]. Job [1625]. Job-Shop [1625]. Join [1847]. Joint [1766]. Journey [1584]. Judicial [1660, 1667]. Jumping [558]. junctions [917]. JVTOS [660].

K3 [2044]. K5 [2044]. K5-free [2044]. Kalman [1931]. KBSE [143]. Kernel [1767]. Kernel-Oriented [1767]. Key [1235, 64, 74, 1252, 75, 69, 1254, 692, 1238, 1239, 1261, 1262, 83]. Keyed [78]. Keys [1851, 332]. KGB [3]. Kinds [1659]. Kit [713]. Kleene [684]. KLIC [1212]. Knapsack [1610]. Knowing [1235]. Knowledge [1577, 1554, 1400, 1933, 1558, 1402, 1411, 1406, 1733, 1478, 1301, 1753, 1410, 1646, 1948, 1935, 1405, 1399, 1715, 1509, 1398, 1693, 1750, 276, 1665, 1768, 729, 1748, 1407, 278, 1482, 315, 353, 1681, 292, 1456, 1505, 1454, 1453, 371, 1457]. Knowledge-and [315]. Knowledge-Based [1554, 1733, 1301, 1748, 1482, 1505]. Knowledge-Level [1406, 1768]. Known [1249]. KOALA [1404]. Korean [355]. KQML [309]. Kronecker [786]. KRONOS [1040].

L [316, 1347, 145]. L-Systems [1347]. Labeling [1302, 246, 244]. LabFlow [747]. LabFlow-1 [747]. Lagrangian [394]. Lambda [806, 1360, 807, 1906, 805]. lambda-Calculi [805]. Lambda-Calculus [807, 1906]. Lambert [968]. LAN [470]. Lane [243]. Language [1952, 598, 1225, 856, 857, 1784, 838, 639, 593, 1212, 602, 708, 271, 268, 199, 1943, 1951, 1723, 1516, 273, 1953, 861, 1721, 1881, 609, 793, 670, 309, 286, 1356, 594, 1395, 103, 852, 1205, 316, 711, 779, 1438, 180, 136, 783, 370, 1508, 364, 356, 350, 357, 2082, 348, 1487]. Languages [204, 511, 516, 764, 1333, 173, 1907, 567, 846, 789, 1916, 824, 774, 611, 195, 829, 312]. langue [1725]. LANs [1893]. LAPACK [1183, 399]. Larch [639]. Larch/C [639]. Large [1090, 483, 2004, 2037, 1589, 1836, 1815, 1127, 1819, 1092, 1948, 466, 1192, 745, 1787, 1124, 1083, 1125, 438, 960]. Large-Scale [466]. LARNet [482]. Last [1692]. Late [2003, 1404]. Late-Breaking [1404]. Latency [1422]. Latin [1850]. Lattice [1109, 418, 790]. Lattices [1383, 178]. LAURA [855]. Law [66, 62, 968]. Laws [1991]. Lax [1272]. Layer [661, 1505, 34]. Layered [1302]. Layout [6, 20, 21, 35, 36, 39, 41, 45, 50]. Layouts [40, 52]. Lazy [1197, 579, 1356, 611]. LCL [639]. LCS [1217]. Lean [1898]. leanEA [1898]. Learn [446, 449]. learnable [960]. learned [367]. Learner [1708, 1741, 379]. Learner-Directed [1741]. Learning [1737, 1738, 1552, 1510, 524, 1691, 1428, 1740, 1758, 1557, 1704, 1191, 1733, 1420, 526, 1839, 1723, 360, 361, 1710, 1730, 1744, 443, 444, 380, 1937, 1711, 293, 1741, 365, 1490, 1760, 1938, 1427, 1570, 1564, 1455, 1742, 1565, 1224, 448, 1298, 374, 450, 1563, 1761, 934, 451, 453, 1763, 357, 1695, 1700, 439, 348, 1726, 1436, 369, 370, 364, 356, 350, 372, 953, 376, 352, 349, 368, 354, 1500]. Learning-by-Doing [1760]. LearnMedia [1700]. learns [358]. Least [1931, 1955]. Lecture [1431]. Left [563, 217]. Leftist [1858]. Legacy [1812, 1519]. Legal [1644]. Legged [889]. Lemma [789]. Lempel [1994]. LENA [1687]. Length [1260, 2025, 1291, 1827, 1914, 362]. Length-Doubling [1260]. Lengths [250, 905]. Lenses [230]. Lessons [1638, 1214]. Level [114, 1870, 1299, 1406, 1569, 406, 117, 1446, 1875, 1838, 523, 179, 1564, 106, 1866, 1887, 184, 1288, 1768, 195, 213, 1181, 464, 646, 1306, 655, 2017]. Leveled [30]. Leveled-Planar [30]. Lexical [280, 278, 366]. LH [754]. LH0* [754]. Libraries [1785, 1592, 1168, 582, 1394, 213, 1583, 1587]. Library [1412, 392, 1593, 182, 403, 424, 1580, 1120, 585, 1582, 1595, 680, 1579, 1594]. Life [873, 890, 154]. Life-Cycle [154]. Lifting [1630]. Light [1108, 1017]. Lightning [1079]. Like [2027, 36, 539, 602, 1281]. Likelihood [242, 994]. Limited [1938]. Limitedness [2049]. Linda [836, 842, 1796, 853, 1797]. Line [1129, 554, 1310, 1419, 1626, 1309, 248, 1292, 1831, 1294, 1821, 1044]. Lineage [1345]. Linear [1, 1983, 1597, 2030, 1333, 390, 601, 393, 392, 1240, 1903, 530, 560, 30, 1054, 415, 1828, 1984, 347, 430, 289, 779, 1620, 458, 906, 1014, 1281, 1253, 1058, 1059]. Linear-Time [1984, 289]. Linearly [527, 1778]. Linearly-Ordered [527]. Lines [1296, 1291]. Link [1541, 460, 360]. Linking [1415]. Links [270]. Lion [341]. lip [996]. Lisp [1215]. List [2065, 168]. Lists [1858, 427]. Literal [1573]. Literature [1723]. Liveness [1868, 1869, 631]. liver [921]. LK010q [1282]. Load [1821, 166, 1147, 1146, 1949, 405, 413, 1148, 1145]. Load-Store [166]. Local [473, 987, 690, 786, 638, 969, 1969, 922, 1859, 482, 1436, 930]. Localisation [260]. Localization [1941]. localizing [949]. Locally [1494]. Locating [1299, 1003]. Location [1978, 776]. Locations [1941]. Locking [743]. Locomotion [889]. Log [1201, 556]. Logarithmic [2068]. Logging [743]. Logic [762, 1634, 597, 577, 1894, 1636, 598, 1560, 737, 599, 600, 601, 1212, 602, 1343, 1903, 604, 127, 1547, 573, 567, 547, 1650, 1651, 769, 579, 1911, 606, 607, 792, 608, 571, 1662, 1959, 1355, 1575, 1572, 572, 1668, 612, 576, 778, 1674, 613, 1676, 736, 565, 615, 291, 1203, 458, 574, 1515, 539, 617, 1468, 1471, 1470, 1500, 1280, 374, 611, 1287, 1458, 373]. Logic-Based [1575, 1515]. Logic/ Constraint [1203]. Logical [763, 306, 1900, 1766, 189, 1907, 1910, 91, 308, 2035, 1917, 794, 316, 637, 315]. Logically [1643, 1640]. logiciel [1696]. logicielle [1696]. Logics [1894, 1635, 1066, 1646, 1353, 1912, 1671, 289, 457, 1682, 1272, 1277, 1270, 1455]. logique [1727]. Long [413, 1116, 891]. Long-Chained [413]. Long-Wave [1116]. Longest [805]. Lookahead [554]. Looking [1307]. Loop [1177, 812, 177, 186, 1864]. Loops [187, 808]. Lopsided [2067]. Lot [1623, 753]. Lot-Sizing [1623]. Lotka [885]. Low [515, 333, 1299, 1234, 1997, 1605, 26, 1257, 73, 344, 213, 1306]. Low-Bandwidth [333]. Low-Exponent [1234]. Low-Level [213]. Low-Weight [1605, 344]. Lower [1833, 5, 2060, 1827, 1847, 1978, 549, 2059, 2026, 1619]. Lowness [514]. LP [1619]. LP-Based [1619]. LPMEME [1560]. LP’s [1609]. LU [1156]. LU-Decomposition [1156]. Luby [346]. Luby-Rackoff [346]. Lyndon [519].

M [658]. M-Connection [658]. MAC [473, 1236]. Machine [869, 1905, 605, 1078, 1624, 1626, 1172, 57, 1550, 779, 1196, 369, 370, 678, 350, 372, 371]. Machine-Negotiated [57]. Machines [1996, 390, 179]. Macro [789, 436]. Made [1107, 87, 379]. Magic [211]. Magnetic [1126]. Maintainability [822]. Maintainable [725]. Maintained [1957]. Maintaining [731, 1962]. Maintenance [410]. Make [2048]. Makes [67]. Making [1976, 1392, 273, 1686, 1490]. Management [720, 1958, 747, 1813, 601, 735, 860, 1532, 745, 2014, 1164, 329, 595, 310, 1165, 1811, 1750, 1960, 49, 92, 1091, 1093, 1893, 1063, 1035, 1495]. Managing [685, 1531, 1497]. Mandatory [1748]. Manhattan [45]. Manipulation [1951, 1446]. Manufacturing [1231]. Many [1437, 1273, 1467]. many-valued [1273]. Map [1301, 1937, 1302, 1926, 1304, 940]. Map-Building [1926]. Mapping [1191, 1785, 1146, 1160, 677, 96, 426, 1138]. Mappings [179]. Maps [803, 1303, 1436]. Maptool [826]. Markers [1400]. Market [305]. Market-Oriented [305]. Marketplace [61]. Markets [66]. Markov [1874, 1925, 1611, 1938, 902]. Markovian [468]. masks [971]. massive [1450, 1462]. Massively [1323, 1201, 1180, 423, 585, 1151]. Matching [225, 1385, 1387, 1363, 1364, 1389, 2002, 245, 224, 1978, 997, 1386, 1316, 265, 1360, 1366, 236, 1341, 932, 944, 946, 953, 927]. Material [1232, 1582]. Mathematical [1921]. Mathematics [1728, 1756]. MATISSE [740]. MATLAB [183]. Matrices [1613, 1825, 983, 1512]. Matrix [381, 1612, 395, 416, 432, 916]. Matroids [1384, 1828]. Max [471]. Max-Min [471]. MAXCUT [542]. Maximal [2025, 1842]. Maximally [1631]. Maximum [1614, 1387, 1970, 2044, 1972, 1573, 242, 994]. Maximum-Entropy [1614]. maximum-likelihood [994]. Maxwell [1126]. May [1955]. Maybe [1701]. MBone [495]. McCulloch [127]. McMillan [689]. MD [1101]. MD4 [336]. Mean [1918, 925]. Mean-Value [1918]. Meaning [2064]. Means [383, 101, 688, 1948, 1886, 2050, 95]. Measure [508, 1679]. Measurement [1949, 1999, 1293]. Measurements [248, 987]. Measures [995, 72, 514]. Measuring [1403]. Mechanical [1356]. Mechanics [1420, 578, 1084]. Mechanism [1163, 1942, 1532, 184]. Mechanisms [838, 1173, 1775, 467, 500]. Mechanizing [1684]. MEDEA [1142]. Media [499, 1427, 1694, 590, 663, 261]. Median [1380]. Mediators [1195]. Medical [1749, 1638, 624, 931]. Medicine [1747, 1748]. Meeting [1472, 213, 1424]. Meiko [1191, 1114]. Membership [513, 1839]. Meme [590]. Memoing [562]. Memory [522, 387, 2071, 392, 825, 1195, 407, 179, 176, 2014, 172, 191, 433, 1816, 1145, 238]. meningoenchephalitis [1491]. Mental [1685, 1776]. Mental-States [1776]. MENTOR [1514]. Merging [1115]. Mesh [416, 1824, 544]. Meshed [460]. Meshes [1152, 544, 47, 1840]. Mesoscale [1098]. Message [686, 1102]. Message-Passing [1102]. Messages [1234, 56].

MESSENGERS [843]. Meta [588, 587, 1695]. Meta-Helix [588]. Metacircularity [588]. Metaclasses [582]. Metacomputing [1129, 1130]. Metadata [1594]. MetaEdit [1513]. METAFrame [713]. Metaobject [595]. Metaphoric [2020]. metareasoning [380]. MetaRing [473]. Metasystem [578, 1443]. Method [382, 167, 383, 1560, 1409, 839, 1291, 459, 126, 1817, 1646, 256, 2014, 1939, 1445, 2040, 423, 541, 1124, 1537, 1572, 1395, 1609, 431, 154, 1603, 435, 1151, 884, 437, 1276, 921]. méthodologie [1717]. Methodologies [1297, 1449]. Methodology [152, 1872, 1175, 1647, 1771, 317, 571, 626, 1393, 1498]. Methods [653, 1391, 621, 1147, 1392, 208, 117, 1448, 1290, 1140, 1796, 1394, 123, 1080, 1867, 2008, 1275, 999, 988, 976, 2079, 948, 350, 1283, 1461]. métiers [1745]. Metric [591]. Metrics [1617, 154]. metrology [1019]. MHD [428]. MHEG [659, 488]. MHEG-Based [659]. Microcells [423]. Micron [1108]. Micron-Sized [1108]. MiCroscopic [1113]. Microsoft [1956]. Microsystems [106]. Migratable [2014]. Migration [1815, 1082, 1097]. Mill [1158]. Million [422]. MIMD [1079]. Min [471, 925]. min/max [925]. Mind [1654]. Mines [1404]. Minima [2060]. Minimal [1895, 551, 1976, 1278, 1864, 910, 1285]. Minimization [34, 776, 940]. Minimize [1619]. Minimizing [1372, 1624]. Minimum [1851, 1962, 74, 1974, 1971, 1607, 2043, 38, 1986, 1984, 1982, 1967, 362]. Mining [1194, 1092, 717, 716, 1462, 1509, 1492, 1454, 1491]. Minor [1334]. Minor-Closed [1334]. Minsum [2076]. MIS [1519]. missing [1016]. MIX [303]. Mixed [892, 136]. Mixing [1611]. Mixture [253]. ML [781]. ML-Style [781]. Mobil [1775]. Mobile [2063, 1932, 1921, 1936, 875, 450, 1991, 1934]. Mobility [477, 496]. Modal [1636, 1635, 598, 690, 545, 1911, 1912, 1668, 1671, 700, 457, 1072, 1279]. Mode [778, 1063]. Model [1577, 835, 128, 1396, 738, 690, 802, 643, 724, 1408, 1901, 600, 1769, 1229, 1558, 697, 1527, 394, 1544, 398, 326, 1516, 1295, 327, 1399, 654, 209, 1532, 1913, 105, 35, 1097, 1705, 848, 1541, 1706, 1716, 247, 1098, 1393, 222, 1735, 1768, 1058, 1380, 253, 1213, 779, 121, 1767, 646, 1676, 1308, 1891, 315, 931, 1022, 1479, 1278, 988, 989, 939, 1070, 1010, 993, 1285, 1039]. Model-Based [1295, 222, 1479]. Model-Checking [690, 1070]. modèle [1725, 1714]. Modeled [468]. Modèles [1703]. Modeling [1865, 114, 205, 469, 155, 441, 1728, 699, 1709, 1648, 1056, 881, 1878, 104, 1713, 1060, 425, 1887, 1668, 115, 1755, 2012, 464, 1782, 223, 1534, 110, 1524, 1757, 1027, 1038, 910, 1456, 1504]. modelisation [1743, 1712, 1707]. Modelled [661, 1002]. Modelling [320, 101, 1642, 117, 1223, 862, 317, 1770, 833, 1575, 1671, 1417, 657, 1080, 491, 1128, 1012]. Models [1416, 720, 1708, 1525, 160, 1413, 1872, 2006, 1729, 410, 1685, 239, 1938, 696, 2036, 1522, 1927, 1444, 153, 902, 865, 298, 1523, 438, 973, 1016, 997, 1011, 908, 998, 1020, 984, 985]. Modernization [1813]. modes [967]. Modified [1858, 1159, 1489]. Moding [1203]. Modular [255, 1350, 1247, 1343, 1447, 1951, 889, 1248, 79, 826, 696, 773, 556, 521, 1888, 650]. Modularity [838]. Modulated [1107]. Module [2005, 1320, 1745, 1736]. Modules [1661]. Modulo [168, 809, 1919, 1920]. Molecular [1102, 388, 1104, 1100, 412, 1122, 417, 1367, 422, 1094, 427, 437]. Molecules [413, 1125]. Moment [382]. Moments [1124]. MONACOT [1714]. Monadic [2038, 1343, 539]. Monet [1950, 726]. Monge [1612]. Monitoring [389, 1158, 860, 1044, 992]. monocular [1023, 1008, 982]. Monoids [1907]. Monotone [1612, 815]. Monotonic [1659]. monotonicity [1036]. Monte [1103, 435, 241]. Moore [1364]. Morals [1654]. Morphisms [2056]. Morphogenesis [870, 1346]. Morphological [1296, 922]. MORSE [252]. Motifs [525]. Motion [225, 1983, 1013, 221, 1932, 224, 226, 217, 222, 1233, 237, 235, 986, 987, 974, 962, 1023, 977, 991, 963, 1016, 978, 980, 1024, 964, 985]. motions [934]. Motivated [1770, 301]. Motivation [1706, 1775]. Motivational [288]. moving [999]. MPC [1209]. MPEG [469]. MPI [1185, 1110]. MPP [435]. MPPs [381, 422]. MR [989, 998]. MRF [221, 1016]. Mu [545, 700]. Mu-Calculus [545, 700]. Multi [2008, 304, 314, 325, 320, 1232, 1769, 441, 1766, 322, 1229, 732, 1684, 1240, 1222, 319, 1230, 1759, 208, 1405, 442, 1711, 1513, 755, 1699, 422, 310, 1224, 448, 1864, 1780, 849, 1522, 1233, 1694, 2017, 297, 439, 1545, 1559, 1002, 1055, 1270, 1505, 1483, 1024, 983, 1497]. Multi-Agent [304, 325, 1232, 1769, 441, 1766, 322, 1229, 1684, 1222, 319, 1230, 442, 1699, 310, 1224, 1780, 1233, 297, 439, 1559, 320, 1483]. Multi-Agents [314]. Multi-authority [1240]. Multi-Block [732]. Multi-Computer [755]. Multi-Data [1522]. Multi-Dimensional [208]. multi-frame [983]. multi-image [1024]. multi-implication [1270]. multi-layer [1505]. Multi-level [2017]. Multi-loop [1864]. Multi-Media [1694]. Multi-methods [2008]. Multi-Million [422]. Multi-Object [849]. multi-ocular [1002]. Multi-Perspective [1405]. Multi-Point [1545, 983]. multi-polynomial [1055]. Multi-Robot [448]. multi-step [1497]. Multi-Tool [1513]. Multi-Turn [1759]. Multi-User [1711, 1513]. Multiagent [443, 303, 451, 453]. Multicast [658, 466, 657]. Multicommodity [1601]. Multicomputer [1216, 432]. Multicomputers [408, 754, 176, 1115, 421, 191, 433]. Multicriteria [884]. Multidimensional [1152, 955]. Multifacetted [115]. Multiflows [1617]. Multigrid [1154]. Multilevel [1127]. Multilingual [285, 278, 284, 356]. Multimedia [1591, 662, 659, 658, 1419, 474, 740, 1426, 504, 489, 495, 1418, 670, 470, 664, 498, 656, 1425, 1429, 1431, 465, 486, 496, 1417, 487, 663, 661, 491, 1700, 500, 475]. Multinational [1954]. Multiparty [557, 558, 2031]. Multiphysics [1153]. Multiple [1985, 720, 763, 74, 1656, 831, 76, 1993, 1565, 270, 1369, 450, 853, 249, 480, 1456, 991, 1021, 912, 1506, 981]. Multiple-Valued [763]. Multiple-View [249]. Multiplexed [482]. Multiplexing [463]. Multiplication [381, 1248, 2059, 1844]. Multiplications [1613]. Multiplicative [1243]. Multiply [416, 1559]. Multipoint [475]. Multiprocessor [426]. Multiprocessors [1104]. Multiresolution [221]. Multisearch [1995]. Multiset [832, 602, 840]. Multisets [838]. Multitape [787]. Multiversion [1944]. MuRaLi [1506]. MUSIC [1074]. mutation [1464]. Mutually [443].

NAG [424]. naked [1015]. Names [1885]. Narrowing [762]. Narrowing-Driven [762]. NASA [627]. NASA’s [1578]. National [1579]. Natural [1557, 271, 268, 893, 273, 256, 350, 286, 820, 1112, 370, 364, 993, 2082, 348]. NAture [1714]. Nautical [1814, 146]. Navier [431]. Navigation [1931, 1932, 1921, 627, 1923, 1938, 961]. NC [1842]. Near [2078]. Near-Perfect [2078]. Nearest [1979, 1856]. Necessity [1937]. Need [724, 1658, 1534]. Needham [692]. Needham-Schroeder [692]. Negated [603]. Negation [611, 612]. Negative [525, 2064, 1677]. Negoplan [1575, 1495].

Negotiated [57]. Negotiation [1231, 450]. Negotiation-Based [1231]. Neighbor [427]. Neighbor-Lists [427]. Neighborhood [1980]. Neighbour [1856]. Neighbourhood [1753]. Neighbours [1987]. Nepi [861]. Nested [1352, 757, 606, 188, 808]. Net [318, 707, 321, 1884, 1755, 891, 1891]. Nets [1865, 1868, 1869, 687, 155, 1871, 1752, 796, 1873, 1874, 1875, 1876, 2062, 1878, 1880, 1882, 1883, 1330, 1885, 1886, 1866, 1665, 1887, 548, 646, 1417, 95, 110]. Network [385, 6, 1548, 1600, 199, 1193, 1605, 474, 405, 861, 416, 1721, 1964, 486, 87, 663, 1603, 1934, 484, 480, 1777, 482, 352, 1131]. Network-Flow [1605]. Networked [388, 1104, 1572]. Networks [255, 1113, 2070, 469, 1773, 396, 749, 1935, 461, 889, 478, 130, 1864, 244, 2050, 464, 342, 1857, 782, 1190, 460, 1559, 1822, 1893, 990, 351, 350, 1482, 353]. Neural [391, 396, 889, 891, 1934, 1777, 350, 352, 1482, 353]. Neuron [869]. neuropathy [988]. News [1404, 1583]. Newtonian [1420]. Next [1692, 214, 1427]. Nice [7]. NII [63]. NKRL [1487]. No [2059]. Node [1630, 1143, 811, 1337]. Node-Lifting [1630]. Nodes [753]. Noise [1095, 1510]. Noisy [1930, 1291]. Non [1895, 528, 1871, 1855, 550, 815, 721, 1659, 1253, 774, 123, 1325, 616, 1777, 986, 1283]. Non-approximability [1855]. Non-classical [550, 616, 1283]. Non-conflict [1777]. Non-Deterministic [774]. Non-linear [1253]. Non-monotone [815]. Non-Monotonic [1659]. Non-observable [528]. Non-Obstructing [1325]. Non-Orthogonal [123]. non-rigid [986]. Non-Standard [721]. Non-terminating [1895]. Nondeterminism [638]. Nondeterministic [1547, 2054]. Nonemptiness [1963]. Nonlinear [919, 345, 75, 108, 1126, 1054]. Nonlinearity [1259]. Nonmonotonic [1656, 1641, 617]. Nonstationary [1564]. Nonuniform [71]. Normal [529, 576, 258]. Normalization [798, 956, 1349, 784, 1359, 1282]. Normalized [1542]. Normative [1660, 1666]. Note [1833, 1845, 1902, 335, 71, 1834]. Notes [1856]. Novel [726]. NP [1859]. NP-Completeness [1859]. Nuclear [129]. Nucleotide [734]. null [1488]. Number [1, 555]. Numbering [15]. Numbers [26, 47]. Numeration [550]. Numerical [385, 1153, 410, 1882, 424]. NWChem [411].

O [1064]. OBDD [1971]. Object

[114, 255, 143, 164, 1525, 2004, 1942, 738, 724, 830, 973, 1538, 1813, 2016, 593, 150, 161, 1809, 1477, 1305, 756, 2006, 749, 1792, 843, 587, 327, 209, 1808, 1532, 580, 1924, 948, 264, 596, 1885, 571, 1796, 849, 252, 1442, 591, 1585, 594, 201, 115, 2003, 585, 955, 153, 2017, 154, 657, 1187, 829, 237, 1449, 910, 944, 997, 907, 951]. Object-Oriented [114, 164, 1525, 2004, 1942, 738, 830, 1813, 593, 161, 1809, 2006, 1792, 209, 1532, 580, 596, 571, 1796, 1442, 591, 594, 201, 2003, 585, 153, 2017, 154, 657, 1449]. Object-Orientedness [1808]. Object-Process [1305]. Objects [2000, 830, 7, 2016, 838, 488, 137, 2014, 2011, 1693, 146, 854, 263, 589, 1517, 436, 932, 1002, 1450, 1003, 995, 1018]. Objects-as-Multisets [838]. objets [1745]. objets-règles [1745]. Oblivious [2074, 1244]. Observable [1925, 528]. Observational [1913]. Observations [514]. Observer [229]. Obstacles [1929]. Obstructing [1325]. Obstructions [1334]. occluded [1005]. occlusion [917, 951]. occlusions [950, 954]. Occupancy [1939]. Occurrence [72]. Ocean [1099]. Ocore [594]. OCR [1588]. ocular [1002]. Offs [224, 979]. Oil [420, 1075]. omega [2055]. omega-Generators [2055]. OMG [150].

OMG/CORBA [150]. Omniscient [1640]. On-Board [140]. On-Line [1129, 554, 1419, 1626, 1821, 1044]. once [546]. One [1378, 536, 386, 1600, 75, 828, 2020, 1467]. One-Dimensional [386]. One-Key [75]. One-Pass [828]. One-time [536]. onto [1082, 677]. Ontology [1397, 57]. Ontology-Based [57]. OO [1950]. OODBMS [1957]. Open [803, 1215, 744, 307, 489, 848, 113, 855]. Operational [141, 804, 315, 1468]. Operations [1170, 825, 899, 176, 2059]. operative [1427, 1700]. Operator [614, 1471]. Operators [314, 897, 386, 2013, 900, 1494]. Opponent [441, 449]. Opportunistic [1779, 574]. OPT [713]. OPT-METAFrame [713]. Optical [2070, 976, 219, 220, 977, 997, 979, 971]. Optical-Flow [219]. optics [1001]. Optigraph [33]. Optimal [1851, 1979, 533, 559, 1989, 536, 8, 732, 1043, 2068, 547, 520, 1626, 1838, 2042, 1625, 828, 1982, 1599, 928]. Optimisation [77, 1178, 574, 211]. Optimization [1805, 193, 1160, 1132, 672, 405, 406, 207, 768, 868, 1615, 900, 713, 758, 884, 673, 679, 1503, 677, 681, 956]. Optimizations [166]. Optimize [1664, 1489]. Optimizing [727, 1613, 812, 176, 1156, 902]. Optimum [1251]. Options [278]. Orca [1211]. Order [204, 1952, 1894, 662, 697, 840, 1343, 1903, 1639, 1547, 815, 1257, 609, 793, 1662, 1355, 73, 1356, 2047, 852, 778, 758, 1360, 1361, 736, 539, 2079, 1511]. Ordered [527, 1343]. Orders [1828]. Ordinal [801]. Organic [1653]. Organisations [1960]. organization [938]. Organizational [1536]. Organizations [836]. Organizing [2066, 353]. Oriented [114, 1576, 164, 1525, 2004, 1942, 738, 830, 1813, 593, 150, 161, 1809, 2006, 1792, 461, 1223, 209, 1532, 580, 596, 571, 915, 1796, 305, 1442, 591, 594, 201, 115, 2003, 585, 153, 1767, 2017, 154, 657, 712, 1187, 829, 53, 1449, 1477, 1504]. Orientedness [1808]. Original [1375]. Orthogonal [5, 1989, 42, 123]. Orthographic [1308]. OSI [661]. OSQL [728]. OSSCAR [1187]. Othello [896]. Other [296, 1421, 1120, 297]. Outdoor [256, 259, 260]. Outerplanar [38]. Outline [281]. Output [1829, 789, 807]. Output-Sensitive [1829]. Outputs [691]. Overcome [132]. Overhead [1101]. Overlapping [1531, 1018]. Overview [55, 1209, 1445, 1213].

P [1871]. P/T [1871]. Pablo [389]. Pablo-Based [389]. PAC [1500, 526]. PAC-learning [1500]. Package [1138, 1145]. Packages [1797]. Packet [462, 480]. Packets [1981]. Packing [1988, 1346, 1628]. PadFEM [1149]. Paging [554]. Painless [1407]. Pair [395]. Pairs [799, 800, 787, 1358, 904]. Paper [1685, 1686, 1304, 610, 612, 614]. Paraconsistent [1566]. Paradigm [1102, 821, 119, 1132, 1415, 375, 952]. paradigm-driven [375]. Paradigms [397, 1081]. Parallax [904]. Parallel [1095, 2069, 383, 1105, 1106, 1995, 1211, 832, 1086, 1170, 1870, 1073, 1102, 748, 673, 1153, 1996, 388, 390, 1142, 1186, 1137, 1089, 584, 1212, 393, 1139, 1127, 1133, 1327, 679, 394, 173, 1082, 1323, 1141, 1157, 1149, 182, 199, 1159, 1100, 1185, 1197, 672, 401, 403, 404, 1114, 1092, 831, 1158, 151, 1096, 1088, 882, 800, 411, 1200, 412, 1078, 413, 1108, 415, 1152, 1198, 1172, 2013, 1206, 1201, 1122, 1144, 417, 1180, 418, 419, 420, 682, 246, 1135, 1140, 423, 1124, 1716, 187, 1074, 426, 40, 521, 753, 1081, 427, 1208, 594, 429, 188, 681, 1116]. Parallel [195, 430, 1181, 1079, 181, 2024, 1118, 1120, 431, 432, 585, 1103, 1156, 1169, 180, 1085, 1110, 1084, 120, 178, 1080, 1151, 1099, 1150, 1112, 1196, 1091, 1093, 1154, 1204, 1388, 438, 680, 674, 678, 676, 1465, 677, 675, 1128, 1193].

Parallelisation [1077, 1087, 1075, 1119]. Parallelism [523, 202, 1205]. Parallelization [384, 1176, 387, 396, 397, 170, 1192, 177, 196, 1094, 1098, 1083, 437]. Parallelized [409]. Parallelizer [1177]. Parallelizing [1143, 172]. Parameter [586]. Parameterized [698]. Parameters [1949, 1417]. Parametric [1974, 1910, 931, 908]. Parametricity [772]. paraperspective [975]. Parentheses [1978]. Parents [1534]. PARIS [148]. Parity [1825, 344, 1571]. PARK [718]. PARKBENCH [1175]. Parse [1552]. parsed [377]. Parser [827, 373]. PARSETALK [1553]. Parsing [828, 1322, 377]. Part [920, 270, 1814, 1748]. Partial [1311, 762, 1342, 764, 1351, 634, 662, 697, 1226, 2079, 1975, 1828, 561, 611, 1917, 822, 1458]. Partial-Order [697, 2079]. Partially [1925, 1889]. Particle [394, 422]. Particles [1132, 1108]. Parties [668]. partir [1717]. Partition [742]. Partitioned [732, 148]. Partitioning [1136, 1975, 472, 191]. Partitions [1668]. Parts [1307]. Party [68]. Pass [828]. Passing [1102, 586]. past [374]. Patch [226]. Path [459, 543, 1938, 460, 1063]. Path-Width [543]. Paths [1829, 1604, 790, 805, 1842, 1342]. Patient [329]. Pattern [1387, 1574, 1389, 871, 1386, 1433, 1341]. Patterns [649, 1383, 1518, 1377, 716, 365, 367, 958, 1462]. pcall [1199]. PDE [783]. Pearls [1954]. Pebble [1983]. Pedagogical [1715]. pédagogique [1743]. Penalty [897]. 3-Approximation [1370]. 6000 [1192]. Ada [1801]. Aria [158]. C [639]. Commitment [782]. Constraint [1203]. CORBA [150]. CPN [1873]. Design [154]. Event [696]. Evolves [869]. HFX [330]. Illustra [741]. language [354]. Learning [1709]. Manipulator [51]. max [925]. Mellor [153]. Motif [157]. Navier [1155, 1152]. O2 [1811]. photometric [958].

RSPIN [703]. RT [646]. symbolic [355]. T [1871]. Penetrating [1116]. people [1015]. PEP [707]. Perception [1921, 1923, 1775, 1941, 315, 1009]. Perception-Based [1941, 315]. Perfect [1373, 2078, 555]. Performance [381, 1546, 1194, 473, 185, 1113, 469, 1950, 726, 389, 1142, 322, 584, 392, 1289, 1173, 1141, 1087, 411, 1200, 1143, 1293, 754, 660, 1309, 1174, 470, 626, 1074, 202, 121, 432, 1148, 1619, 1156, 139, 716, 1190, 120, 484, 228, 1893]. Performances [1079]. Performing [1432]. Periodic [160, 71]. Permutable [694]. Permutations [537]. Persistence [1811, 594, 1807]. Persistent [744]. Personal [62, 1432, 496, 1434]. Perspective [64, 1644, 1405, 2002, 1660, 65, 975, 1023, 380]. Perspectives [268, 610, 1190]. Petri [1868, 687, 1873, 1874, 707, 1875, 2062, 1878, 1880, 1882, 1883, 1330, 1885, 1886, 1866, 1887, 646, 1417, 95, 110]. pGVT [640]. phase [976, 1054]. phase-portrait [1054]. Phenomena [1126]. Philosophers [1799]. photographs [1020]. Photometry [249]. Phylogenies [1381]. Phylogeny [2078, 555]. physically [953]. physically-based [953]. Physics [1105, 1106]. Pi [861]. Pi-Calculus [861]. Pictures [1338]. PID [108]. Piecewise [188]. Pilot [1131, 1118, 475]. Pilot-Implementation [1118]. Pipelining [167, 809, 808]. Pipelining-Dovetailing [808]. Pitts [127]. Pixel [225, 1115]. Place [213]. Plan [1551, 1704, 1230]. Planar [1, 1836, 22, 1631, 1608, 30, 1978, 226, 1124, 1055, 958]. Plane [1832, 1989, 29, 1830, 985]. Planes [28, 1610]. planification [1743]. Planner [1932, 274]. Planning [1412, 1557, 1411, 294, 1527, 1413, 575, 859, 319, 1819, 1096, 1231, 118, 1501, 1938, 1940, 1233, 1673, 97, 1718, 1063, 1281]. Plans [1230, 1759, 1772]. Plant [129, 1346, 641]. Plants [696]. Plasma [1109]. Platform [1730, 1529, 481]. Platforms [1082, 1124, 1079]. Play [295]. Playing [1763]. plural [352]. Plus [1911]. Point [8, 1903, 669, 1978, 1916, 1545, 983]. Pointer [1979, 558, 767]. Pointers [2009]. Points [1809, 1290, 817, 1269, 904]. Points-to [817]. Poisson [1378]. polarized [1017]. Policy [66, 63]. Pollution [120]. Polygon [1861]. Polygonal [1832]. polygrams [363]. polyhedra [918]. Polyhedral [1629, 1002, 1020]. Polymers [1122]. Polymorphic [1906, 781, 1360]. Polymorphism [1910, 1885, 2036, 2020]. Polynomial [1630, 1902, 2078, 402, 2041, 247, 449, 1572, 1055]. Polynomial-Time [2078, 1572]. Polynomials [2041, 1237, 1239]. Polytope [1627]. Portability [822]. Portable [1212, 392, 1149, 1135, 39, 1204]. Porting [1078, 1074, 1079]. portrait [976, 1054]. Portraits [1114]. pose [1002]. Position [1685, 1686]. Positive [525, 1920, 1278]. POSIX [1801]. POSIX/Ada [1801]. possibility [1488]. Possible [62, 1643, 1663]. Post [510]. Postprocessing [248]. Postulates [617]. Potentials [467]. pour [1743, 1727, 1703, 1745, 1714, 1725]. Power [2051, 1129, 528, 129, 1121, 43, 2047, 641]. PP [515, 361]. PQ [1632]. PQ-Tree [1632]. Practical [341, 1314, 2001, 135, 2044, 1637, 1817, 1638, 2002, 198, 760, 1746, 1263, 1681]. Practically [1689]. Practice [643, 1741, 794, 1774, 1451]. Pragma [1644, 1660]. Pragma-Dialectical [1644, 1660]. Pragmatic [1320, 300, 2081]. PRAM [1171]. Pre [459]. Pre-Computation [459]. Preconditioners [402]. Preconditioning [1155]. Preconditions [648]. Precongruence [1913]. Predators [445]. Predicate [1957, 1876, 1572, 1487]. Predicate-Logic [1572]. Prediction [1123, 1174]. Predictive [479, 667]. Predictors [1546]. Preelaboration [144]. Preferences [288]. prefixes [1283]. Preliminary [1778, 404, 1079, 1672]. Premises [2064].

Preprocessed [247]. Preprocessor [1562]. Presburger [788]. Presence [1199]. Presentation [488, 1417, 1756]. preserves [940]. Preserving [1869, 219, 561]. press [1032]. Pretexte [279]. Preview [1404]. Prey [445]. Primal [1608]. Primal-Dual [1608]. Primes [552]. Primitives [1538, 895, 290]. Principle [476]. Principles [127, 1394]. Print [1365]. Printed [1295]. Priorities [699, 472, 1670]. Prioritization [1206]. prioritized [1052]. Priority [1996, 470]. priors [1016]. Privacy [63]. Private [1433]. Probabilistic [2053, 1923, 1729, 1993, 1937, 1859, 704, 1926, 248, 360]. Problem [1865, 1370, 1970, 897, 1391, 563, 1548, 1612, 1602, 1413, 892, 1392, 1729, 1629, 1909, 1384, 898, 510, 1615, 1741, 1971, 1354, 1846, 1968, 226, 1742, 1625, 1716, 1750, 1747, 1967, 853, 156, 1325, 1756, 1571, 1695, 884, 1221, 1765, 200, 2073, 2079]. Problem-Solving [1391, 1392, 1741, 1747]. problème [1707]. Problems [2069, 1845, 515, 1130, 1601, 2044, 1791, 1855, 268, 575, 1544, 1157, 1132, 1974, 672, 1226, 1608, 1975, 1823, 1828, 760, 1135, 1964, 259, 1321, 142, 2061, 1620, 1623, 1080, 1603, 1853, 1388, 1777, 673, 1480, 677, 681]. Procedural [295, 1681]. Procedure [1182, 530, 1197, 1650, 576, 1438, 655]. Procedure-Level [655]. Procedures [1901, 1733, 1912, 289]. Process [1737, 1378, 1326, 1217, 1526, 1704, 842, 1134, 1305, 1788, 860, 714, 456, 104, 1556, 694, 849, 1138, 1960, 1443, 865, 1718, 1043, 1497]. Process-Calculus [849]. Process-Sensitive [1737]. Processes [2063, 2037, 119, 638, 844, 831, 1925, 468, 2061, 457, 791, 935, 1029]. Processing [832, 726, 160, 1289, 182, 207, 1924, 1418, 1168, 742, 257, 370, 355, 1486, 925, 2082, 348, 1489]. Processor [185, 416]. Processors [1201, 1112]. processus [1712, 1707]. Produce [2000]. Product [786, 1704]. Production [1596].

Productions [1327, 1328]. ProfGlot [283]. profiles [929]. Program [896, 1332, 816, 701, 1579, 1428, 648, 575, 636, 573, 547, 578, 410, 1446, 1878, 761, 713, 570, 797, 1798, 426, 1735, 572, 1748, 565, 574, 650, 358, 1465, 1028]. Programmable [874, 866]. programmation [1727]. Programmed [1319]. Programming [2069, 1597, 1897, 1560, 1102, 857, 1784, 600, 593, 1758, 1212, 150, 1139, 840, 892, 1809, 604, 1185, 1723, 878, 560, 1650, 769, 1223, 1610, 847, 861, 111, 607, 149, 308, 610, 1536, 775, 2003, 97, 1781, 615, 1203, 617, 1468, 1470, 1477, 676, 374, 1458, 373]. Programs [566, 2051, 2007, 762, 597, 577, 1894, 164, 563, 599, 388, 766, 839, 1790, 2075, 1200, 579, 606, 1144, 608, 1180, 1787, 571, 1789, 541, 568, 1121, 188, 777, 1609, 430, 201, 181, 612, 576, 778, 569, 613, 817, 178, 1598, 1322, 1500, 1494]. Prograph [1202]. PROGRES [1341]. Progress [1216, 1992]. Project [1128, 1683, 1217, 135, 1819, 474, 1721, 1174, 1467, 1395]. projection [906]. Projections [546, 7, 990]. projective [915, 1024]. Projector [435]. Prolog [763, 766, 579, 1561, 614]. Promela [703, 702]. Promela/RSPIN [703]. Promoting [1690, 591]. PRONAOS [141]. Proof [1899, 1352, 601, 1646, 1650, 618, 607, 570, 576, 1287, 1072, 1283]. Proof-theory [607]. Proofs [1245, 513, 1246, 685, 1050, 1266, 2026, 1359, 1469, 1281, 1284]. Propagation [175, 1191, 1521, 423, 1572, 730, 813, 933, 926]. proper [1494, 1487]. Properties [1350, 1123, 693, 874, 775, 576, 710, 95, 918]. Property [1845, 1616]. Proposal [1129, 393, 399, 2011]. Proposed [68]. Propositional [606, 1918, 289, 2026, 1272, 1492]. Propositions [454]. Propositions-as-Types [454]. Prospector [1404]. Prospects [1676]. Protected [1802]. Protection [668, 1031]. Protein [1383, 1122]. Proteins [1382]. Protocol [473, 643, 587, 2015, 321, 654, 692, 595, 1963, 1072]. Protocols [320, 79]. Prototype [859, 1193, 1594, 1803, 261]. Prototyping [12, 708, 1800]. Provable [347]. Provably [341, 1255]. Prove [576]. Proves [541]. Provide [495]. Providing [478, 745, 464]. Proving [801, 2023, 654, 1356, 1919, 710, 95, 1502, 1279]. Proximity [19]. Pruning [599, 1963]. Pseudo [1255]. Pseudo-random [1255]. Pseudonyms [81]. PSR [1872]. Public [1578, 83, 1252, 692, 1238, 1239, 1433, 1680]. Public-Key [692, 83]. Publicly [1250]. PULSAR [1095]. Pulse [1107]. Pumping [789]. punctuation [376]. Pure [1573]. Purpose [1105, 193, 1139, 1568, 303, 1395]. Putting [1450]. PVM [1182, 389, 149, 1179, 863, 424, 1119]. Pyramidal [426].

Q [451]. Q-Learning [451]. Q1 [1167]. QH4 [1134]. QOS [464, 1417, 501]. QR [1159, 429]. QR-Factorization [429]. Quadratic [1612, 1610, 1972]. Quadrics [1167]. Quadries [1134]. Quadrilateral [544]. Qualitative [1666, 1567, 1743, 1717]. Quality [1399, 1403]. Quality-Based [1399]. Quantification [2002, 998]. Quantifiers [597, 2053, 1271]. Quantisation [1117]. Quantitative [935, 621, 1875, 1293, 987]. Quantum [507, 530, 85, 419, 435]. Quarter [1692]. Quasi [1, 1778]. Quasi-Anticipatory [1778]. Quasi-Planar [1]. Queries [205, 1957, 727, 1839, 1949, 728, 1368, 736, 1473]. Query [1554, 748, 749, 207, 262, 742, 1674, 211, 734, 1503, 1486, 1489]. Query-Answering [1674]. Querying [1952, 737]. Question [1816]. Questions [1826]. Queueing [462]. Queues [1996]. Quiescence [691].

R0n [1353]. R0n- [1353]. Rabin [1267, 695, 2048]. Race [1201]. Rackoff [346]. Radar [391, 1116]. Radiation [1107]. Radio [128, 493]. Radio-Taxi [128]. Radiographic [1134]. RADIOSS [1081]. Radiotherapy [1096]. Railroad [1909]. RAISE [158]. RAMs [2059]. Random [18, 552, 2041, 537, 509, 1255]. Randomization [560]. Randomized [2051, 1621, 9, 1962, 674, 1171, 2068, 1365, 1968, 1623]. Randomness [512, 72]. Range [175, 388, 417, 912, 971]. Rangefinding [232]. Rank [556, 981]. Ranking [553, 1679]. Rapid [708, 1594, 907, 323]. Rapide [1789]. RAPT [1096]. Raster [1293, 1303]. Raster-Scanned [1303]. Raster-to-Vector [1293]. Rate [502, 233, 485, 1464]. Rates [1852, 1611, 503]. Rational [1640, 328, 517]. Ray [1112, 990]. rd [853]. Re [1577, 152, 1519, 1214, 1717]. Re-Engineering [152, 1519]. Re-Implementation [1214]. ré-ingénierie [1717]. Re-Use [1577]. Reachability [1879, 1055]. Reaction [832, 1653]. Reactions [1653, 1909, 1103]. Reactive [299, 455, 117, 118, 703]. Reactivity [419]. Reactor [619]. Read [546, 732]. Read-once [546]. Real [144, 1806, 160, 161, 134, 1076, 1819, 636, 1158, 313, 1652, 1835, 327, 224, 996, 227, 233, 1604, 696, 704, 43, 1563, 116, 467, 1085, 291, 702, 655, 1042, 1052, 1070, 1018]. Real-Time [144, 160, 161, 134, 1819, 1158, 327, 224, 227, 233, 696, 704, 43, 1800, 116, 467, 636, 996, 655, 1042, 1052, 1070]. Real-Tune [1806]. Real-Valued [1604]. Real-World [1563]. Reality [1173, 1421, 1422]. Realization [1184]. Reason [1689, 1669]. Reasoner [1396, 1568]. Reasoning [1683, 1930, 1739, 1684, 1637, 1640, 605, 1547, 455, 1656, 1641, 1645, 1729, 1638, 296, 1653, 1399, 60, 1924, 275, 1661, 751, 1663, 1572, 1672, 1761, 954, 1567, 291, 1515, 1677, 1922, 1545, 1679, 1681, 1751, 1275, 1451, 1502, 1285, 1506].

Reasons [1392, 1654]. receptive [955]. RECIBA [481]. Reciprocity [452]. Recognition [266, 255, 1551, 1297, 1002, 267, 254, 490, 1305, 1230, 1374, 871, 1924, 264, 1916, 1298, 521, 252, 1117, 1288, 908, 1776, 90, 232, 237, 942, 906, 944, 945, 351, 948, 996, 907, 950, 951, 955]. Recognize [1322]. Recognized [516]. Recognizing [1342, 10, 1299, 30, 2050]. Reconfigurable [432]. Reconfiguration [1323, 1860]. Reconstruct [70]. Reconstructing [1020]. Reconstruction [990, 1660, 1116, 250, 1308, 920, 975, 1000, 911, 939, 912, 1016, 905, 940]. Recovering [189]. Recovery [1412, 1528, 920]. Rectangular [1830]. Rectangulation [1984]. Rectifying [245]. Rectilinear [1861]. recurrent [350]. Recursion [2036]. Recursions [1357]. Recursive [1931, 541, 1138, 1561, 298]. Redescription [1406]. Redistribution [192]. Redistributions [1184]. Reduce [450]. Reduced [1877, 1883]. Reducing [1476, 1422, 500]. Reduction [1841, 697, 395, 1639, 797, 711, 805, 1039]. Reductions [1852, 532, 553]. Redundant [1146]. Reference [1569, 257, 947]. Referentlal [210]. Refined [1939]. Refinement [634, 440, 739, 636, 977, 96, 569, 1947, 1284]. Refinements [1062]. Refining [623, 1494, 1497]. reflectance [1017]. Reflecting [1827]. reflection [967]. Reflections [1985]. Reflective [307, 1532, 589]. Regenerative [1874]. Region [169, 256, 1302]. regions [1004]. Registers [76, 1256]. Registration [266]. règles [1745]. Regular [395, 553, 1884, 2050, 1099, 1920]. Regularities [1440]. Regularity [2037]. Regularization [970, 940]. Reification [629]. Reinforcement [444, 1563]. Related [1234, 127, 72, 276]. Relation [687, 1400, 1902]. Relation-Algebraic [687]. Relational [205, 724, 1092, 1948, 2018, 2061, 1488, 1496]. Relations [1635, 1175, 269, 270, 1497]. Relationship [1330]. Relationships [2016, 1521]. Relative [1635, 798, 248]. Relaxation [1618, 246, 248, 244, 1603, 943]. Relaxations [1614, 1610]. relaxed [1030]. Relaxing [1541]. Release [1618]. Relevance [1332, 1928]. Reliability [1782]. Reliable [2015, 165, 912, 618, 978, 1117]. relief [939]. RELVIEW [687]. Remarks [510, 439]. Remodeler [1577]. Remote [1161, 1182, 1966, 841, 1197, 492]. Remotely [148]. Removing [764, 1976, 810]. Rendering [433]. Rendezvous [2073]. Reordering [563]. Reorganization [1539]. Repair [897, 1410, 874]. Repairing [810]. repeated [949]. Repeats [1378]. REPERES [1696]. Repetitive [1227]. Replacement [1313, 1334, 1338, 1337]. Replay [1201]. Replicated [730]. Replication [1783, 1210]. Report [1551, 2004, 1596, 23, 761, 1556, 1519, 1216, 1672, 1404]. Reporting [1829]. Repositories [1586]. Representation [825, 16, 272, 24, 1877, 105, 100, 1665, 491, 200, 932, 945, 2080, 914, 1475, 349, 1511, 1745, 1736, 1714]. Representations [2, 216, 1548, 1344, 744, 32, 1879, 961, 947, 352]. representative [995]. Representing [1900]. Reproducing [874]. Reproduction [882, 227]. Requirements [627, 1048, 670, 1514, 487]. resampling [1461]. Rescheduling [645]. Research [13, 1954, 593, 172]. resemblance [995]. Reservoir [1073, 420, 2040, 1075]. Residual [1160, 798]. residue [1039]. Resilient [1258]. Resolution [294, 892, 1646, 901, 369, 1013, 909, 1278, 616]. Resolution-Based [1646]. Resolve [815]. Resource [512, 1480, 508, 601, 1557, 472]. Resource-based [1480]. Resource-Bounded [512]. Resource-Sharing [472]. Resources [806, 1592, 1938, 212]. Response [1551].

Responses [275]. Restoration [460]. restricted [360]. Restructuring [183, 1446]. Result [1859]. Resulting [220]. Results [630, 1855, 1175, 24, 456, 1174, 1079, 2026, 1349, 2054, 373]. Retargetable [812]. Retrieval [1554, 264, 1541, 1587, 261, 927, 1487]. Reusability [1786, 2011, 822]. Reusable [165]. Reuse [1661, 824, 591, 2020, 650]. Reversal [1376]. ReverseNICE [152]. review [2082]. REVISE [1395]. Revising [1759]. Revision [1683, 310, 1679, 617, 1772, 379]. Revisited [1977, 1072]. Rewards [1874]. Rewrite [1332, 816, 1895, 784]. Rewriting [1312, 529, 1314, 602, 801, 1906, 796, 769, 1336, 1348, 2032, 1346, 1318, 833, 1349]. Rhetorical [272, 1676]. Rheumatology [1748]. Rig [235]. Right [1546]. Rights [287]. Rigid [1904, 226, 986]. Rigorous [980]. Ring [1172]. RIPEMD [337]. RIPEMD-160 [337]. RIVUS [670]. RNC [1842]. Road [1113, 258]. Robot [1929, 1931, 1932, 1921, 1923, 1936, 1938, 448, 1233, 121, 223, 1934, 1481]. Robotic [899]. Robotics [1933, 218, 875]. Robots [450, 1991]. Robust [221, 944, 1264, 932, 993, 1068]. ROCK [1943]. Role [1551, 1535, 1215, 1763, 581]. Role-Based [581]. Role-Playing [1763]. Roles [1763]. ROLL [1943]. roofs [909]. Room [1431]. Root [1247, 1249]. rough [1492, 1491, 1461]. Rounding [1621]. Routines [293]. Routing [1824, 549]. Row [2060]. RPC [2014, 1179]. RS [1192]. RS/ 6000 [1192]. RSA [1267, 1234]. r`’solution [1707]. RTL [1032]. Rule [1943, 1518, 541, 122, 1512, 784]. Rules [1552, 1869, 600, 1558, 1946, 718, 1518, 782, 351, 379, 1492, 1460, 368, 1461]. Run [1291, 199]. Run-Time [199]. Runtime [1546].

S.A.M.C.E.F. [1082]. S4 [1912]. SA [646].

SA/RT [646]. Safe [2001, 1991, 133]. SAFER [331]. Safety [135, 620, 133, 1794, 1782, 619, 631, 635]. Safety-Critical [620, 1794, 635]. Salesman [1632, 1157, 1384, 1620]. saliency [936]. SAMMI [1082]. Sample [1961, 364]. sampled [1016]. Sampler [242]. Samples [1366]. Sampling [1614, 2046]. sans [1071]. SAR [1095]. Satellite [478, 479, 976]. Satisfaction [802, 1544, 1226, 1859, 1968, 1287]. Satisfiability [1970, 1846, 1573]. Saturation [2023]. Saturation-Based [2023]. Saving [67]. Scaffolding [1704]. Scalability [461, 470, 1156]. Scalable [1163, 754, 1164, 717, 730, 1581, 657, 315]. ScaLAPACK [1184, 392, 400]. Scalar [185]. Scale [2004, 1589, 466, 1192, 923, 919, 1450, 969, 989, 922, 951, 970, 927, 930, 1462]. Scale-space [923, 919, 922, 970, 927]. Scaled [1385]. Scaling [499, 663, 476, 1044]. Scanned [1303, 1306]. Scattering [1108, 967]. Scenario [1780, 491]. Scenario-Based [1780]. Scenarios [470]. Scene [256, 222, 260, 947, 904, 949, 1009, 993]. Scenes [259, 1112, 965, 1005, 1020]. Sceptical [131]. schedule [1067]. Scheduler [2074]. Schedulers [168]. Schedules [559, 732, 1410, 1625]. Scheduling [128, 1232, 2076, 168, 636, 326, 1618, 462, 899, 1626, 642, 1625, 809, 1619, 500, 1480]. Schema [1942, 752, 1594, 722, 574]. Schema-Based [574]. Schemas [1538, 1537]. Schemata [573]. Scheme [665, 504, 1198, 1201, 80, 1251, 497, 667, 585, 1263, 1893]. Schemes [1243, 536, 592, 1242, 2044, 1191, 1265, 463, 485, 191, 1266, 1239]. Schmidt [1159]. Scholarly [1582]. Schools [1692]. Schroeder [692]. Sciddle [1182]. Science [1578]. Scientific [1586, 1111, 735, 1141, 403, 1144, 1140, 1731]. SCOOP [672]. Scope [757]. SCOTCH [1138]. Screens [1851]. Screws [1307]. sea [1035, 1881]. Search [795, 599, 601, 688, 1997, 898, 2072, 1859, 1369, 946, 1466, 682, 1286, 941, 1287]. Searches [1588]. Searching [1977, 2060, 1574, 1592, 2045, 1484, 263, 1294]. Second [1952, 1343, 1082, 1723, 815, 539]. Second-Order [1343]. Secret [1235, 1242, 1240, 70, 79, 1251, 1250]. Secret-Ballot [1240]. Sectioned [1559]. Secure [341, 630, 1255, 495, 1795, 1256, 1251]. Security [693, 1252, 345, 1886, 347, 497, 1266, 1236, 1263, 59, 63, 1267]. SEdit [1657]. Segment [262]. Segmentation [225, 1301, 1305, 220, 222, 929, 943, 914, 952, 930]. Segmentations [242]. Segments [270, 997]. Seismic [1119]. Selected [259]. Selecting [1546]. Selection [1576, 459, 191, 1992, 364, 947, 1493, 379]. selectionist [943]. Selectivity [1838]. Self [903, 1845, 725, 874, 468, 464, 1941, 2066, 235, 999, 353]. Self-calibration [903, 999]. Self-Localization [1941]. Self-Maintainable [725]. Self-Organizing [2066, 353]. Self-Repair [874]. Self-Reproducing [874]. Self-Similar [468]. Self-Similarity [464]. Self-Witnessing [1845]. Sell [2000]. Semantic [577, 1163, 1897, 746, 127, 1320, 1503, 282, 1945, 1721, 1587, 1275, 1004, 363, 352]. Semantical [290, 1457, 359]. Semantics [785, 597, 763, 1895, 1871, 1327, 547, 718, 760, 105, 608, 1317, 793, 773, 797, 1318, 774, 1663, 1199, 1540, 820, 852, 1675, 614, 589, 1678, 1559, 1468]. Semi [451, 2039]. Semi-Competitive [451]. Semi-Groups [2039]. Semidefinite [1610, 1611, 1609]. Semilinear [2040]. SemNet [1675]. Sense [67]. Sensing [291]. Sensitive [1737, 1829, 2032, 1982]. Sensor [1932, 1940, 1934, 1481]. Sensor-Based [1932, 1481]. Sensors [1930]. Sentences [283, 2047]. Separable [1603]. Separated [119]. Separating [1630, 1631, 1018, 349]. Separation [534, 2054]. Separator [1840]. Sequence [686, 1378, 178, 1013]. Sequences [225, 1133, 1827, 71, 1993, 259, 250, 1377, 1022, 973, 976, 1023, 1004, 1016, 905, 934, 929]. Sequencing [1378]. Sequent [1277, 1271]. Sequential [1372, 1871, 698, 2047, 1156, 716, 942]. Series [517]. serif [1071]. Server [1194, 1813, 1596, 477, 165, 79, 503, 1091]. Server-Aided [79]. Servers [295, 502, 476]. Service [1412, 662, 658, 478, 505, 485, 1964, 664, 498, 712]. Service-Constrained [1964]. Services [749, 148, 68, 472, 1801, 667, 496, 467, 855]. Session [1261]. Set [897, 8, 706, 1602, 2033, 393, 1334, 518, 1629, 1606, 488, 269, 728, 1842, 436, 1492, 1461]. Set-Covering [897]. Sets [527, 2052, 1163, 206, 532, 1656, 2040, 1598, 1269, 1491]. sets-based [1491]. Several [472]. SFDUs [157]. SGI [1121]. SGML [733]. Shake [1550]. Shape [913, 900, 240, 90, 964, 1003, 918, 998, 983]. Shared [522, 1408, 2071, 1195, 70, 1532, 1862, 1797]. Shared-Memory [2071]. Sharing [1243, 1242, 895, 472, 1251, 1250]. Shark [340]. Sharply [1914]. Shell [302, 88, 1430]. Shift [76, 176]. Shifting [276]. Shifts [557]. Shlaer [153]. Shlaer/Mellor [153]. Shop [1625]. Short [1245, 388, 417, 1262]. Short-Range [417]. Shortcomings [639]. Shortest [1370, 1604, 1822]. Shuttle [627]. Side [1615, 1199]. Side-Effects [1199]. sieves [919]. SIGMA [325]. sign [1267]. Signal [1801]. Signals [1229]. Signature [536, 1242, 732, 1265, 80, 1266, 1239]. Signatures [1235, 1264, 1267]. Signed [1468]. Significant [600]. Signs [1229]. Silhouette [951]. Silhouette-based [951]. silhouettes [995]. SIM [311, 116]. SIMD [1134, 1079]. Similar [468]. Similarity [464, 265]. Simple [526, 184, 1853, 1280, 1072]. Simplex [1609]. Simplex-Type [1609]. Simplification [1741, 1478]. Simulated [17, 893, 40]. Simulating [1108, 1079]. Simulation [1106, 1073, 1102, 1153, 1104, 1171, 1134, 1749, 1834, 1122, 1144, 1337, 420, 1355, 1750, 147, 109, 181, 1118, 431, 1103, 1126, 1110, 1075, 1187, 436, 1757, 1751, 1745, 1050, 1049, 1891]. Simulation-Based [1749, 1757, 1751]. Simulations [1113, 522, 388, 2071, 1100, 1109, 413, 1125, 1085, 1060]. Simulator [129, 145, 1158, 106, 1814, 864, 43]. Simulators [1101, 146]. Simultaneous [1904]. Single [823, 1230, 1626, 1604, 824, 541, 480, 2054, 959]. Single-Agent [1230]. Single-Frequency [480]. Single-Machine [1626]. Single-Source [1604]. Single-Tape [2054]. singular [911, 982]. singularity [963]. SINT [1754]. Siphons [1868]. Situation [1930, 1709, 1642, 1469, 1725]. Situational [1281]. Situations [605, 500]. Six [889]. Six-Legged [889]. Sized [1108]. Sizing [1623]. Skeleton [857]. Skeletons [1832]. Skeptical [1674]. sketches [1507]. Skill [1558]. Skills [1747]. Skip [1858]. SLD [616]. SLD-resolution [616]. Slicing [648]. SLIQ [717]. Slotted [1187]. Small [1961, 1841, 1247, 1249, 1201]. Smart [83, 84, 85, 1261]. Smartcards [333]. smooth [920, 968]. smoothing [928]. Snakes [986, 1006]. Snow [425]. Soc [1223]. Social [6, 1228, 1773, 321]. Society [1218]. soft [1006]. Software [167, 653, 2004, 1586, 1691, 119, 705, 134, 1439, 1872, 1792, 1441, 860, 618, 847, 1532, 102, 1786, 1254, 1445, 1787, 1140, 132, 1440, 140, 1321, 591, 486, 103, 1138, 809, 711, 1444, 156, 1694, 1782, 1438, 154, 1304, 1437, 590, 1515, 808, 650]. Software-Development [2004]. Solid [1308]. Solids [383, 1125]. Solution [167, 385, 1127, 477, 1882, 1811, 430]. Solutions [1841, 639, 259, 1126, 212].

Solvable [515]. Solve [1130, 1124]. Solver [1155, 415, 1152, 1716, 1150]. Solvers [404, 783]. Solving [1391, 1413, 575, 1157, 1392, 672, 1729, 1741, 1742, 1716, 1750, 833, 1747, 853, 1695, 1221, 1765, 1777, 924]. Some [1894, 1791, 558, 1778, 1809, 404, 693, 313, 32, 79, 760, 1856, 305, 1918, 1404, 290, 284, 439, 617]. Sonia [836]. sonography [921]. Sorting [1977, 1376]. sound [1471]. Soundness [616]. Source [1604, 1761, 562]. Source-to-Source [562]. Sources [469, 1565]. Sow [1954]. SP [1192]. SP2 [1121]. Space [1416, 1225, 1136, 7, 599, 218, 627, 1877, 227, 220, 433, 1567, 1867, 1433, 919, 962, 922, 923, 951, 970, 927, 930]. Space-and [220]. Spaces [1961, 1877, 1423, 1466, 353]. Spanning [1605, 1974, 1982, 1967]. SPARC [162]. Sparse [197, 532, 1613, 1127, 1159, 404, 1843, 1825, 429, 432]. Sparsification [1974]. Spatial [205, 209, 755, 1524]. Spatiotemporal [961]. Speak [316]. SPEAR [1947]. Special [1624]. Specialization [2001, 592]. Species [881]. Specific [1747, 1722, 906, 368]. Specification [454, 1326, 634, 639, 161, 708, 1641, 117, 625, 640, 624, 421, 626, 1355, 711, 646, 535, 290, 1071, 1032, 1472, 1047]. Specifications [630, 688, 2006, 1792, 628, 709, 644, 1321, 632, 2064, 1048]. Specified [650]. Specify [816]. Specifying [849]. Spectral [1114]. specular [1017]. Speech [490, 996, 354]. speech/language [354]. Speed [464]. Speeding [1948, 2054]. Speeding-Up [2054]. Speeds [869]. Spheres [28]. SPIN [643, 702]. Spinning [1178]. Spinning-on-Coherency [1178]. Spliced [1374]. splines [986]. Splitting [21, 811]. spoken [356, 355]. Spring [35]. Spring-Model [35]. SQL [1953]. Square [1850]. Squares [1931]. SSA [825]. Stability [1129, 1057, 251, 1047, 1058]. stabilization [1037]. Stable [1599]. Stack [480]. Standard [766, 2056, 1305, 587, 721].

Standardisation [56]. Standardized [488]. Star [517]. Star-Height [517]. State [1390, 882, 1877, 2061, 1867, 2079]. state-explosion [2079]. Statecharts [1318, 635]. Statements [179]. States [1753, 1776]. Static [821, 823, 767, 583, 2059, 1138, 1794]. Stations [480]. Statistical [1560, 239, 1012, 1633, 1003, 358, 913, 953]. Statistics [240, 361]. Status [1438]. Steady [882]. Steady-State [882]. Steering [1166]. Steiner [1823]. Steinitz [1616]. Stencil [386]. Step [1082, 1788, 1317, 2020, 1892, 1455, 1497]. Step-Wise [1788]. Steps [76]. Stereo [230, 250, 941, 234, 236, 237, 235, 999, 916]. STI [1736, 1714]. Stochastic [385, 897, 394, 1874, 714, 1882, 110, 1853, 359, 933, 1010]. Stochasticity [512]. Stokes [1155, 1152, 431]. Stop [780]. Storage [1164, 824]. Store [166]. Stored [504]. Straight [1832, 1831]. Straight-Line [1831]. Straightline [34]. Strategic [651, 1722]. Strategies [384, 877, 1576, 1176, 502, 859, 1999, 445, 1198, 1915, 797, 1731, 50, 1573, 1777]. Strategy [1929, 695, 1758, 1364, 570, 742, 1227, 449, 142]. Stream [505, 670, 344]. Streams [502, 504]. Streets [1990]. Streett [695, 1963]. Strength [19]. Strengthened [337]. Stressed [22]. String [1363, 1364, 1369, 2077, 1366, 1283]. Strings [1390, 1373, 1365, 509]. Strip [1158]. STRIPS [1281]. STRIPS-like [1281]. Strokes [1079]. Strong [1282, 1203, 1359, 235, 1497]. Strongest [1627]. Strongly [511, 764]. Strongly-Typed [764]. Structural [1869, 1127, 245, 1883, 1084, 946]. Structure [89, 314, 1327, 272, 1439, 2006, 754, 347, 1300, 1125, 434, 661, 54, 200, 944, 1023, 980, 1024, 964]. Structured [155, 1137, 1185, 1448, 1198, 1879, 1585, 223, 1523, 677, 1033]. Structures [454, 1636, 4, 2027, 189, 1345, 798, 1908, 137, 1124, 112, 804, 2066, 817, 539, 2080, 957]. Structuring [814, 1935, 1401, 1415, 1487]. Structurized [1653]. Student [1698, 1761, 1755, 1757]. Studio [489]. Study [167, 473, 897, 1130, 630, 625, 898, 860, 1824, 1520, 1103, 619, 1443, 813, 1493, 1490, 1394]. Studying [567]. Sturmian [2056]. style [1282, 781, 2033]. Stylistic [273]. Sub [1972]. Sub-quadratic [1972]. Subapical [1347]. Subclasses [1342, 1839]. Subdefinite [212]. Subdivision [433]. Subdomain [1791]. Subgraph [1973, 14, 2044, 1627, 543, 247]. Subgraphs [1836, 1966]. Submodular [1615]. Subobject [2012]. Subobject-Based [2012]. Subprograms [393]. Subroutine [403]. Subsampling [243]. Subscripts [187]. Subset [72]. subspace [948]. Substructure [1851]. Subsumption [795]. Subsystem [1945]. Subtree [1972]. Subtyping [2002, 596, 2036]. Success [61]. Successful [157]. Successor [550]. Suffix [1371, 1372, 2068, 1843]. Suitable [1124]. suite [1042]. suivi [1725]. Super [185, 612, 1013]. super-resolution [1013]. Super-Scalar [185]. Superconnectivity [1864]. Supermodular [1618]. Superscalar [166]. Superstring [1370]. Supervised [443]. Supervisory [130, 1052, 1067, 1062]. Support [731, 701, 1586, 477, 182, 199, 1432, 831, 633, 107, 104, 1427, 1445, 1168, 656, 172, 1732, 1181, 1422, 1205, 822, 1694, 1148, 487, 1443, 1722, 663, 1524, 1515, 1502]. Supporting [2008, 1708, 152, 1529, 709, 826, 1093]. Surface [1077, 1936, 544, 431, 1103, 912, 913, 967, 914, 1017, 928]. surface-reflectance [1017]. Surfaces [37, 910, 938, 968]. Surgical [1738, 378]. surveillance [992]. Survey [1634]. SWAN [1077, 54]. SWIM [261]. Swines [1954]. Switch [483, 471]. Switched [754, 1046].

Switching [1570, 1029]. Syllogisms [1676]. Symbiotic [886]. Symbol [1288]. Symbol-Level [1288]. Symbolic [1211, 551, 175, 1200, 100, 1754, 1208, 99, 108, 378, 1050]. Symbols [1229, 1298]. Symmetric [1612, 1844, 936]. Symmetry [251, 2073]. Symmetry-Breaking [2073]. Synchronization [1869, 592, 504, 505, 833, 1052]. Synchronized [1328, 1157, 182]. Synchronous [1870, 774]. Syndrome [1255]. Synergies [593]. Synphonics [281]. syntactic [358]. Syntax [826]. Synteny [1375]. Synthesis [1870, 571, 890, 572, 1777, 1067, 1505, 675, 1062, 959]. Synthetic [894, 873]. System [266, 1161, 701, 1539, 4, 1899, 1352, 2030, 2019, 1813, 387, 630, 1719, 138, 12, 206, 124, 1299, 322, 1758, 1557, 1139, 659, 1413, 318, 1535, 285, 1173, 1432, 723, 1158, 1401, 1414, 860, 499, 91, 275, 105, 102, 624, 1336, 1881, 130, 1570, 1556, 1520, 230, 1886, 626, 1706, 775, 1575, 1732, 43, 1670, 503, 1118, 121, 632, 1404, 1083, 97, 203, 619, 139, 49, 1304, 781, 92, 95, 2064, 1700, 1091, 888, 54, 1449, 1459, 1471, 1280, 1505, 356, 1036, 366, 1016, 916, 924, 1025, 1045, 1461, 1593, 1679]. System-Theoretic [97]. systematic [1501]. système [1717]. systèmes [1712]. Systemic [285]. Systems [1865, 114, 382, 89, 1129, 1551, 1332, 816, 314, 325, 1895, 1350, 1163, 320, 144, 160, 1769, 441, 1766, 1815, 699, 161, 1793, 134, 1215, 1320, 1222, 455, 796, 1697, 550, 1710, 84, 151, 1730, 117, 443, 1657, 1783, 1414, 871, 415, 1724, 580, 1198, 479, 2028, 745, 624, 1760, 1771, 149, 848, 317, 870, 644, 652, 696, 570, 1164, 100, 1746, 448, 656, 194, 1536, 96, 1440, 276, 833, 704, 1780, 1522, 1887, 1533, 1747, 864, 109, 486, 103, 1666, 620, 125, 1347, 90, 1800, 1888, 1862, 710, 1960, 116, 1148, 113, 637, 453, 535, 1818, 94]. Systems [1889, 434, 855, 1349, 784, 623, 655, 635, 439, 500, 1093, 1726, 1145, 1955, 1777, 1063, 1042, 1027, 1067, 1488, 1038, 1498, 370, 1055, 1065, 1051, 1034, 1043, 1030, 2079, 1035, 2081, 1036, 1048, 1054, 1056, 1044, 1452, 1047, 1026, 1057, 1476, 1052, 1070, 1486, 1060, 1474, 1062, 1058, 1069, 93, 1068, 1483, 1049, 1033, 1059, 2082, 1061, 1046, 1504, 982, 1031, 1489, 1071].

T [1912, 1283]. T-string-unification [1283]. T3D [382, 1162]. T9000 [1171]. Tâches [1714]. Tableau [2030, 1662, 1272, 1280, 1285]. Tableaux [1897, 1273, 1275, 1278, 1270]. Tables [1295, 829]. Tabular [1916]. Tackle [752]. Take [891]. Taking [2020]. TALE [1516]. Talkin’bout [1643]. Tape [2054]. target [952]. Targets [1802, 165]. Task [1408, 1923, 1231, 1814, 450, 1394, 92, 1480]. task-based [1480]. Tasking [162, 1790, 1816]. Tasks [155, 1432, 1014]. Taxi [128]. Taxonomy [1655]. Teach [1747]. Teacher [1734, 1702]. Teacher-Usable [1702]. Teaching [13, 1709, 1762]. Teaching/Learning [1709]. Teamwork [158]. Teamwork/Aria [158]. Technical [1397, 1596, 1657, 286, 1498]. Technique [1342, 697, 1605, 317, 2014, 1124, 1491, 371]. Techniques [1128, 1073, 135, 1901, 1526, 410, 894, 1288, 809, 1422, 99, 758, 263, 257, 911]. Technologies [1690, 1189, 622, 1956, 435, 1188]. Technology [143, 101, 193, 1692, 1720, 1556, 1747, 759]. Telecentric [1001]. Telecommunication [749, 1079]. Teleconferencing [494, 227]. Teleservices [661]. Teleshopping [474]. Teleteaching [475]. Television [490]. Template [670]. Templates [243, 581, 1044]. Tempo [564]. Temporal [1894, 738, 737, 455, 1646, 279, 1516, 1953, 60, 100, 1866, 753, 109, 457, 736, 458, 1472, 1050]. Temporary [1664]. Tenet [497]. tense [374]. Tension [1946, 21]. Tension-Free [21]. tensors [983]. Term [1352, 801, 796, 671, 1348, 1349]. Terminating [563, 1895]. Termination [799, 577, 801, 800, 2032, 778]. Terminological [1399, 380]. Terms [238, 378, 967]. Ternary [541]. Test [1163, 1133, 620, 647, 1598, 691, 963]. Testbed [1223, 310]. Testing [164, 806, 12, 15, 1791, 1792, 1158, 1724, 555, 1789, 1355, 647]. Tests [684]. Text [1400, 285, 1759, 1399, 269, 1398, 270, 1303, 1588, 1735, 277, 284, 372, 367, 368]. Texts [1305, 280, 380]. Textual [270]. Textuality [284]. Texture [921, 930, 943]. Textures [263]. TGEX [1135]. Their [1790, 1246, 696, 241, 1332, 760, 1661, 1018]. Theorem [2023, 654, 2029, 1356, 1919, 1502, 1279]. Theoretic [326, 102, 1675, 97, 565, 458, 1025, 1483]. Theoretical [1313, 231, 1637, 1059]. Theories [1911, 1444, 1920, 312]. Theory [13, 2038, 879, 636, 1647, 645, 456, 107, 418, 1940, 538, 96, 647, 1774, 1403, 1633, 880, 617, 1221, 1275, 1451, 2081, 607, 93, 1461]. thermodynamique [1703]. Thickness [32]. Thickness-Two [32]. Thin [1393]. Thinning [1289]. Third [668, 477, 68, 1360]. Third-Order [1360]. Thorough [435]. Thoughts [1692]. Threads [1201]. Three [1416, 910, 398, 22, 1862, 1308]. Three-Dimensional [1416, 398]. Threshold [1264, 1834, 1844]. thresholding [993]. Throughput [747]. Tier [1658]. Tiger [339]. Tillich [335]. Time [1983, 1852, 529, 144, 160, 2030, 534, 1630, 1902, 161, 134, 199, 2068, 1819, 2078, 1158, 327, 2080, 1078, 224, 30, 1624, 531, 2028, 227, 233, 1604, 696, 419, 1972, 1984, 1662, 247, 449, 704, 220, 1572, 43, 851, 289, 1800, 779, 116, 1619, 1620, 467, 1080, 702, 1151, 292, 1822, 1042, 536, 636, 1044, 996, 1052, 1070, 923, 1059, 655]. Time-Domain [1080]. Time-Variant [220]. Timed [528, 163, 696, 646, 535, 1051, 1060, 1071]. Timing [1041, 1866, 1284]. TIPP [714]. Tissues [1107]. TLM [423]. TM [643, 165]. Toeplitz [1612]. Tolerance [1783, 1210]. Tolerant [522, 134, 1323, 871, 523, 1798, 637, 888]. Tomography [239]. TOMSPIN [110]. Tool [720, 192, 152, 320, 389, 1242, 1142, 13, 14, 1137, 1222, 1723, 1405, 707, 714, 709, 495, 1513, 713, 1519, 582, 1135, 1431, 1533, 1079, 1794, 710, 116, 110, 623, 1042, 1040, 1149]. Tool-Based [1519]. ToolBus [837]. Toolkit [765, 1728, 587, 311]. Tools [2005, 1411, 117, 683, 111, 894, 1165, 1693, 1701, 1694, 1702, 1423, 1034]. Top [488, 1308, 650]. Top-Down [1308, 650]. Topological [1890]. Topologies [195]. Topology [1822]. Torus [1172]. Torus-Ring-Bus [1172]. TOSCANA [49]. Total [1931, 1624, 1619, 613]. Touching [10]. Toupie [711]. Trace [529, 2027]. Traces [1890]. Tracing [924, 1112]. Track [297]. Tracking [931, 932, 1005, 409, 228, 986, 1002, 991, 1007, 933, 996, 1008, 1010]. Trade [224, 979]. Trade-Offs [224, 979]. Tradeoffs [1701]. Traffic [158, 1113, 901, 1991, 468, 464, 260, 258, 471, 1005, 1035]. Train [645]. Training [1690, 129, 1753, 359, 1724, 1746, 1757, 1751]. Transaction [1090, 161]. Transactional [1531]. Transactions [757, 1523]. Transducers [1372, 1324, 789]. Transfer [1244, 1690, 759]. Transfinite [1361]. Transform [221, 926]. Transformation [1311, 816, 563, 575, 1320, 573, 578, 761, 652, 1317, 1319, 1331, 565, 808, 562]. Transformational [765, 651]. Transformations [1260, 1344, 797, 1339, 722, 1319, 94, 574]. Transforms [123]. Transfusion [1747]. Transients [382]. transit [1060]. Transition [2038, 1946, 1819, 1876, 2064, 1026]. Transitions [770]. Transitive [1978, 174, 1628]. Translating [598, 153, 728]. Translation [1552, 579, 1550, 646, 356, 371]. Translational [251]. Translations [1324, 1522]. TransPive [1214]. Transport [1089, 425, 664, 498, 1098]. Transport-Chemistry [1098]. Transportation [326, 115]. Transputer [385, 434]. Transputers [1171]. Traveling [1157, 1620]. Travelling [1632, 1384]. Traversals [1950, 1384]. Treatment [1096, 1936, 538]. Tree [1552, 384, 536, 1372, 2027, 1324, 2068, 1974, 520, 792, 2048, 1570, 789, 1380, 828, 1982, 1967, 1831, 434, 539, 1379, 1388, 677, 1632, 753]. Tree-Based [536]. Tree-Like [2027, 539]. tree-structured [677]. Trees [1979, 1371, 1983, 795, 1899, 1390, 8, 1858, 2067, 18, 1997, 1605, 1975, 1843, 1607, 553, 561, 1380, 828, 363]. Treewidth [1342, 1841]. Trenches [761]. Trends [507]. Triangular [395]. Triangularization [432]. triangulated [914]. Triangulation [1986]. Triangulations [38]. TRICLOPS [228]. Triconnected [1989, 15]. Trigger [95]. Trigger-Graph [95]. trihedral [918]. triplets [903]. TROLL [1520]. Truly [770]. Truly-concurrent [770]. Truncated [331]. Trunk [483]. Trusted [668, 68]. truth [1501]. TSB [753]. TSB-Tree [753]. TSP [1630]. TSQL2 [737]. Tune [1806]. Tuning [139]. Tuples [537]. Turbulence [1109, 428, 1110]. Turbulent [1153, 394]. Turn [1759]. tuteur [1717]. Tutor [1749, 1754, 1747, 1756]. Tutorial [400, 1748]. Tutoring [1719, 1697, 1705, 1706, 1735, 1763, 1700]. Tutors [1705, 1764]. Two [341, 1387, 1988, 1600, 1191, 1147, 1624, 32, 900, 179, 1659, 1564, 1386, 1237, 1236, 1861, 433, 1359, 1101, 980]. Two-Dimensional [1387, 900, 1386]. two-frame [980]. Two-Guarding [1861]. Two-Level [179, 1564]. Two-Machine [1624]. Type [1350, 2053, 2001, 2019, 2002, 1609, 201, 1675, 781, 817, 1745]. Type-2 [2053]. Type-Safe [2001]. Typeable [1349]. Typed [764, 1315, 1880, 2003, 829]. Types [454, 527, 1896, 772, 2035, 1802, 1361, 784, 1269]. Typing [583].

Ubiquitous [1424]. Ultrafast [1126]. UMTS [477]. Unary [540]. Unbalanced [342]. Unbounded [1877]. Unboxing [772]. Uncalibrated [1014, 939, 982]. Uncertain [1222, 1935, 1476]. Uncertainty [1932, 1936, 1924, 1866, 1922]. Unconstrained [872]. Undecidability [541]. Undecidable [1904, 210]. Undeniable [1265]. Underdiagonal [790]. Understanding [1946, 1200, 1205, 918, 865]. Undirected [40]. Undistorted [200]. Unfolding [689]. Unfoldings [1883]. Unification [577, 1900, 1526, 1904, 1316, 1276, 1283]. Unified [1637, 1638, 538, 93]. Uniform [1833, 541, 1585, 1027]. Uniformly [816]. Unifying [237, 1283]. Unions [817]. Uniqueness [238]. Unit [1738, 1788, 488, 1278]. Units [1317, 1319]. Univariate [1247]. Universal [2, 1958, 552, 1580, 476]. Unknown [1547, 1627, 1678]. Unpointed [772]. unpolarized [1017]. Unpredictable [797]. Unranking [553]. Unrolling [812]. Unrooted [1972]. Unstructured [1152, 1135]. Unsupervised [943, 1565, 242]. Untyped [805, 807]. Update [2065, 1530, 1094, 730]. Updates [1679, 1457]. Updating [752, 371]. UPPAAL [715, 1042]. Upper [557, 2049]. Upward [15, 28, 1831]. Upwind [1152]. Urban [423, 120]. Usable [1702]. Use [1577, 1314, 724, 2016, 145, 1324, 1721, 1885, 146, 1794, 1080, 884, 1816, 236, 653, 144]. User [1551, 2006, 1642, 406, 1711, 495, 1555, 1513, 1715, 652, 447, 1536, 106, 1425, 1668, 1118, 1502]. User-Adaptive [447].

User-Enhanceability [1536].

User-Friendly [652]. user-independent [1502]. User-Interface [106]. User-Level [406]. Using [255, 577, 1952, 1113, 128, 1895, 764, 1931, 1400, 1901, 858, 1752, 892, 459, 1171, 575, 1439, 1195, 1157, 283, 1753, 1720, 1185, 1974, 1723, 1792, 1873, 725, 1414, 1223, 1610, 1741, 33, 1878, 377, 1607, 553, 243, 2014, 1520, 1542, 894, 692, 1796, 423, 1664, 1747, 248, 1131, 711, 1800, 367, 121, 1941, 822, 432, 637, 452, 223, 1261, 1443, 178, 663, 435, 581, 574, 298, 982, 480, 354, 1119, 1322, 257, 1751, 943, 906, 932, 643, 1471, 648, 1003, 20, 911, 989, 165, 1044, 948, 1016, 1052, 914, 907, 1072, 978, 1060, 374, 1017, 1585]. using [927, 955, 2017, 971, 357, 998, 353, 930, 373]. utilisation [1743].

V [159, 173]. V-cal [173]. Validating

[1657, 44]. Validation

[128, 155, 1527, 752, 1749, 637, 1049]. Validity [1473, 723]. valuation [1489]. valuations [1489]. Value [2074, 764, 1918]. Value-Oblivious [2074]. Valued [763, 1604, 1273]. Values [764, 1488]. Variable [2025, 502, 1354, 250, 503, 905]. Variable-Length [2025]. Variables [552]. Variant [1197, 220]. Variants [1634]. variation [378]. Variational [219]. varieties [1021]. Various [1917]. varying [965, 1003]. vascular [990]. VAST [594]. VBR [463]. VDM [628]. VDM-SL [628]. Vector [1877, 1293, 1117]. Vectorization [1292]. Vehicle [260, 258, 1433, 991, 1031]. vehicles [1005]. ventes [1745]. Ventricular [217]. verb [379]. verbs [374]. Verifiable [1242, 1250]. Verification [1326, 546, 1411, 697, 693, 1036, 654, 633, 640, 696, 100, 698, 1303, 641, 1867, 631, 655, 1031, 1042, 1032, 2079, 1028, 1060, 954, 1033, 1046, 1071]. Verification-Based [1303]. Verifier [1246]. Verify [1884]. Verifying [699, 1873, 1026]. Version [735, 398, 337]. Versus [1824, 1156]. Vertex [1313, 1855, 21, 1629, 1975]. Very [1092, 745]. vessel [990]. Via [1961, 1507, 910, 1014, 604, 552, 800, 1446, 1124, 775, 1210, 190, 492, 237]. Video [469, 215, 216, 502, 463, 669, 499, 214, 262, 503, 492, 1582, 1019]. View [1942, 1773, 1809, 752, 262, 286, 249, 263, 932, 1507, 1021, 981, 1504]. View-Based [262, 263, 932]. viewed [917]. Viewing [14]. Viewpoint [1540]. Views [1539, 727, 25, 725, 1308, 942, 947]. Violated [1631]. Vipar [1168]. Virtual [2007, 1420, 1173, 1720, 1421, 227, 2014, 447, 472, 1216, 195, 1422, 506, 1169, 1763, 1423, 1018]. Virtual-Space [227]. Virus [1381]. Visibility [2, 1985, 1815, 16, 24]. Visifold [856]. Vision [1228, 218, 231, 214, 233, 230, 213, 223, 234, 236, 1002, 1007, 915, 1008, 916, 1001]. Visual [2057, 1739, 856, 124, 938, 1711, 992, 1536, 1423, 631, 261, 238, 961, 1014, 1007, 996, 1010, 994, 982]. Visualisation [53]. Visualization [11, 1584, 20, 1345, 1168, 54]. Visualized [46]. Visualizer [51]. Visualizer/Manipulator [51]. Visualizing [428]. VLIW [186]. Volterra [885]. Volume [544, 431, 433]. Volumic [914]. vs [906, 1066, 1480, 979, 540]. VSAT [478]. VSE [633]. VSM [406, 1178].

Walking [1990]. Wang [518]. warping [1007]. Watching [215, 992]. Wave [1077, 423, 1116, 929]. Waveguide [1107]. Wavelength [482]. wavelet [930]. Way [1739, 1336, 1489]. Weak [512, 532, 19, 332, 235]. Weakest [648]. Weather [410]. Weaving [858]. Web [741, 1719, 858, 1584, 740, 1529, 492, 1188]. Web-Based [1188]. WebJournal [1584]. Wedge [1393]. Weight [1858, 1605, 38, 1986, 1984, 344]. Weighted [1619]. Weights [1834]. Well [530]. Well-Formed [530]. Whately [1655]. Where [1810, 1955]. Which [869, 1218, 544, 940]. while [219]. Whither [1692]. Wide [1719, 230, 492, 1453, 1529]. Width [543, 520]. Will [2000]. Win [1310]. Wind [425]. Windows [36]. Winning [695]. Wise [1788, 552]. Within [1923, 623]. Without [618, 561, 828, 1539, 1235, 629, 552, 800, 941]. Witnessing [1845]. Won’t [1571]. Word [523]. Word-Level [523]. Words [1371, 551, 519]. Work [1217, 119, 1240, 1744, 1531, 1424, 1450]. Workbench [12]. Workflow [747, 60, 745, 1960, 59, 746]. Workstation [1190]. Workstations [388, 1104]. World [1420, 1663, 1563, 1669, 291, 1459, 1500, 1453, 1719, 1529, 492]. world-wide [1453, 1529]. Worlds [297]. Wormhole [1824]. Worst [531]. Wrap [587]. Write [1178]. Write-Invalidate [1178]. WULPUS [1750].

X [1007, 990, 157]. X-ray [990]. X/Motif [157]. Xnetvideo [499].

Z [634, 624, 635]. Zadeh [1471]. Zemor [335]. Ziv [1994]. zooming [1000]. Zuse [127].

References

Agarwal:1996:QPG

[1] P. K. Agarwal, B. Aronov, J. Pach, and R. Pollack. Quasi-planar graphs have a linear number of edges. Lecture Notes in Computer Science, 1027:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alt:1996:UDV

[2] H. Alt, M. Godau, and S. Whitesides. Universal 3-dimensional visibility representations for graphs. Lecture Notes in Computer Science, 1027:8–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benz:1996:KCG

[3] H. Benz. KGB: A customizable graph browser. Lecture Notes in Computer Science, 1027:20–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertault:1996:ADS

[4] F. Bertault. Adocs: a drawing system for generic combinatorial structures. Lecture Notes in Computer Science, 1027:24–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biedl:1996:NLB

[5] T. C. Biedl. New lower bounds for orthogonal graph drawings. Lecture Notes in Computer Science, 1027:28– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blythe:1996:EGL

[6] J. Blythe, C. McGrath, and D. Krackhardt. The effect of graph layout on inference from social network data. Lecture Notes in Computer Science, 1027: 40–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bose:1996:DNP

[7] P. Bose, F. Gomez, P. Ramos, and G. Toussaint. Drawing nice projections of objects in space. Lecture Notes in Computer Science, 1027:52–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bose:1996:OAE

[8] P. Bose, M. McAllister, and J. Snoeyink. Optimal algorithms to embed trees in a point set. Lecture Notes in Computer Science, 1027:64–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandenburg:1996:ECF

[9] F. J. Brandenburg, M. Himsolt, and C. Rohrer. An experimental comparison of force-directed and randomized graph drawing algorithms. Lecture Notes in Computer Science, 1027: 76–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breu:1996:CRI

[10] H. Breu and D. G. Kirkpatrick. On the complexity of recognizing intersection and touching graphs of disks. Lecture Notes in Computer Science, 1027: 88–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruss:1996:FID

[11] I. Bruss and A. Frick. Fast interactive 3-D graph visualization. Lecture Notes in Computer Science, 1027:99– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buti:1996:GWS

[12] L. Buti, G. Di Battista, G. Liotta, and E. Tassinari. GD-Workbench: A system for prototyping and testing graph drawing algorithms. Lecture Notes in Computer Science, 1027:111–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carbonneaux:1996:CGT

[13] Y. Carbonneaux, J.-M. Laborde, and M. Madani. CABRI-Graph: A tool for research and teaching in graph theory. Lecture Notes in Computer Science, 1027:123–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carpendale:1996:GFE

[14] M. S. T. Carpendale, D. J. Cowperthwaite, F. D. Fracchia, and T. Shermer. Graph folding: Extending detail and context viewing into a tool for subgraph comparisons. Lecture Notes in Computer Science, 1027:127–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chandramouli:1996:UNT

[15] M. Chandramouli and A. A. Diwan. Upward numbering testing for triconnected graphs. Lecture Notes in Computer Science, 1027:140–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cobos:1996:VRG

[16] F. J. Cobos, J. C. Dana, F. Hurtado, and A. Marquez. On a visibility representation of graphs. Lecture Notes in Computer Science, 1027:152–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cruz:1996:GDS

[17] I. F. Cruz and J. P. Twarog. 3D graph drawing with simulated annealing. Lecture Notes in Computer Science, 1027:162–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devroye:1996:BBR

[18] L. Devroye and P. Kruszewski. The botanical beauty of random binary trees. Lecture Notes in Computer Science, 1027:166–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiBattista:1996:SWP

[19] G. Di Battista, G. Liotta, and S. Whitesides. The strength of weak proximity. Lecture Notes in Computer Science, 1027:178–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dodson:1996:IVU

[20] D. Dodson. Information visualization using cooperative 3D diagram layout. Lecture Notes in Computer Science, 1027:190–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1996:VST

[21] P. Eades and C. F. X. De Mendonca. Vertex splitting and tension-free layout. Lecture Notes in Computer Science, 1027:202–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1996:DSP

[22] P. Eades and P. Garvan. Drawing stressed planar graphs in three dimensions. Lecture Notes in Computer Science, 1027:212–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1996:GDC

[23] P. Eades and J. Marks. Graph-drawing contest report. Lecture Notes in Computer Science, 1027:224–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fekete:1996:NRV

[24] S. P. Fekete, M. E. Houle, and S. Whitesides. New results on a visibility representation of graphs in 3D. Lecture Notes in Computer Science, 1027:234–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Formella:1996:GFV

[25] A. Formella and J. Keller. Generalized fisheye views of graphs. Lecture Notes in Computer Science, 1027:242– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foessmeier:1996:DHD

[26] U. Foessmeier and M. Kaufmann. Drawing high degree graphs with low bend numbers. Lecture Notes in Computer Science, 1027:254–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gropp:1996:DC

[27] H. Gropp. The drawing of configurations. Lecture Notes in Computer Science, 1027:267–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hashemi:1996:UDP

[28] S. M. Hashemi, A. Kisielewicz, and I. Rival. Upward drawings on planes and spheres. Lecture Notes in Computer Science, 1027:277–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1996:GEC

[29] X. He. Grid embedding of 4-Connected plane graphs. Lecture Notes in Computer Science, 1027:287–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heath:1996:RLP

[30] L. S. Heath and S. V. Pemmaraju. Recognizing leveled-planar dags in linear time. Lecture Notes in Computer Science, 1027:300–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hlineny:1996:CGC

[31] P. Hlineny. Contact graphs of curves. Lecture Notes in Computer Science, 1027:312–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hutchinson:1996:RST

[32] J. P. Hutchinson, T. Shermer, and A. Vince. On representations of some thickness-two graphs. Lecture Notes in Computer Science, 1027:324–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ignatowicz:1996:DFD

[33] J. Ignatowicz. Drawing force-directed graphs using optigraph. Lecture Notes in Computer Science, 1027:333–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juenger:1996:EHA

[34] M. Juenger and P. Mutzel. Exact and heuristic algorithms for 2-Layer straightline crossing minimization. Lecture Notes in Computer Science, 1027:337–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamps:1996:CBS

[35] T. Kamps, J. Kleinz, and J. Read. Constraint-based spring-model algorithm for graph layout. Lecture Notes in Computer Science, 1027:349–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kikusts:1996:LAG

[36] P. Kikusts and P. Rucevskis. Layout algorithms of graph-like diagrams for GRADE windows graphic editors. Lecture Notes in Computer Science, 1027: 361–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kratochvil:1996:GIB

[37] J. Kratochvil and T. Przytycka. Grid intersection and box intersection graphs on surfaces. Lecture Notes in Computer Science, 1027:365–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenhart:1996:HDO

[38] W. Lenhart and G. Liotta. How to draw outerplanar minimum weight triangulations. Lecture Notes in Computer Science, 1027:373–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madden:1996:PGL

[39] B. Madden, P. Madden, S. Powers, and M. Himsolt. Portable graph layout and editing. Lecture Notes in Computer Science, 1027:385–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monien:1996:PSA

[40] B. Monien, F. Ramme, and H. Salmen. A parallel simulated annealing algorithm for generating 3D layouts of undirected graphs. Lecture Notes in Computer Science, 1027:396–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

North:1996:ILD

[41] S. C. North. Incremental layout in DynaDAG. Lecture Notes in Computer Science, 1027:409–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papakostas:1996:IIO

[42] A. Papakostas and I. G. Tollis. Issues in interactive orthogonal graph drawing. Lecture Notes in Computer Science, 1027:419–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paris:1996:ADC

[43] G. Paris. Automatic drawing of compound digraphs for a real-time power system simulator. Lecture Notes in Computer Science, 1027:431–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Purchase:1996:VGD

[44] H. C. Purchase, R. F. Cohen, and M. James. Validating graph drawing aesthetics. Lecture Notes in Computer Science, 1027:435–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sander:1996:FHH

[45] G. Sander. A fast heuristic for hierarchical Manhattan layout. Lecture Notes in Computer Science, 1027:447– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sander:1996:CVC

[46] G. Sander, M. Alt, C. Ferdinand, and R. Wilhelm. CLAX — A visualized compiler. Lecture Notes in Computer Science, 1027:459–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shahrokhi:1996:CNM

[47] F. Shahrokhi, O. Sykora, L. A. Szekely, and I. Vrt’o. Crossing numbers of meshes. Lecture Notes in Computer Science, 1027:463–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shieh:1996:DGD

[48] F.-S. Shieh and C. L. McCreary. Directed graphs drawing by clan-based decomposition. Lecture Notes in Computer Science, 1027:472–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skorsky:1996:TMS

[49] M. Skorsky. TOSCANA management system for conceptual data. Lecture Notes in Computer Science, 1027:483– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Storey:1996:GLA

[50] M.-A. D. Storey and H. A. Mueller. Graph layout adjustment strategies. Lecture Notes in Computer Science, 1027:487–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sugiyama:1996:GCG

[51] K. Sugiyama and K. Misue. A generic compound graph visualizer/ manipulator: D-ABDUCTOR. Lecture Notes in Computer Science, 1027:500– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:GCL

[52] X. Wang and I. Miyamoto. Generating customized layouts. Lecture Notes in Computer Science, 1027:504–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Webber:1996:GGO

[53] R. Webber and A. Scott. GOVE: Grammar-oriented visualisation environment. Lecture Notes in Computer Science, 1027:516–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1996:SDS

[54] J. Yang, C. A. Shaffer, and L. S. Heath. SWAN: A data structure visualization system. Lecture Notes in Computer Science, 1027:520–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adam:1996:ECO

[55] N. R. Adam and Y. Yesha. Electronic commerce: An overview. Lecture Notes in Computer Science, 1028:5–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steel:1996:SFE

[56] K. Steel. The standardisation of flexible EDI messages. Lecture Notes in Computer Science, 1028:13–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lehmann:1996:MNO

[57] F. Lehmann. Machine-negotiated, ontology-based EDI (electronic data interchange). Lecture Notes in Computer Science, 1028:27–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitrakas:1996:CIA

[58] A. Mitrakas. Changes in the interchange agreements. Lecture Notes in Computer Science, 1028:47–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tewari:1996:AEC

[59] H. Tewari, M. McCourt, and D. O’Mahony. Advanced electronic commerce security in a workflow environment. Lecture Notes in Computer Science, 1028:69– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haimowitz:1996:TRA

[60] I. J. Haimowitz, J. Farley, G. S. Fields, and J. Stillman. Temporal reasoning for automated workflow in health care enterprises. Lecture Notes in Computer Science, 1028:87–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laufmann:1996:IMA

[61] S. Laufmann. The information marketplace:achieving success in commercial applications. Lecture Notes in Computer Science, 1028:115–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caelli:1996:CPF

[62] W. Caelli. Cryptography: Personal freedom and law enforcement — is it possible to get agreement? Lecture Notes in Computer Science, 1029:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ware:1996:PSP

[63] W. Ware. Privacy and security policy choices in an NII environment. Lecture Notes in Computer Science, 1029: 20–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caelli:1996:CKE

[64] W. Caelli. Commercial key escrow: An Australian perspective. Lecture Notes in Computer Science, 1029:40– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orlowski:1996:EGI

[65] S. Orlowski. Encryption and the global information infrastructure: An Australian perspective. Lecture Notes in Computer Science, 1029:65–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1996:CEM

[66] R. Anderson. Crypto in Europe — markets, law and policy. Lecture Notes in Computer Science, 1029:75– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ames:1996:SDM

[67] M. Ames. Saving dollars makes sense of crypto export controls. Lecture Notes in Computer Science, 1029:90– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jefferies:1996:PAT

[68] N. Jefferies, C. Mitchell, and M. Walker. A proposed architecture for trusted third party services. Lecture Notes in Computer Science, 1029:98–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1996:NKE

[69] J. He and E. Dawson. A new key escrow cryptosystem. Lecture Notes in Computer Science, 1029:105–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1996:HFR

[70] J. He and E. Dawson. How to fairly reconstruct a shared secret. Lecture Notes in Computer Science, 1029:115– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golic:1996:NND

[71] J. D. Golic. A note on nonuniform decimation of periodic sequences. Lecture Notes in Computer Science, 1029:125– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gustafson:1996:RMR

[72] H. Gustafson, E. Dawson, and J. D. Golic. Randomness measures related to subset occurrence. Lecture Notes in Computer Science, 1029:132–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Millan:1996:LOA

[73] W. Millan. Low order approximation of cipher functions. Lecture Notes in Computer Science, 1029:144–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damgaard:1996:MEM

[74] I. B. Damgaard and L. R. Knudsen. Multiple encryption with minimum key. Lecture Notes in Computer Science, 1029:156–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gysin:1996:OKC

[75] M. Gysin. A one-key cryptosystem based on a finite nonlinear automaton. Lecture Notes in Computer Science, 1029:165–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golic:1996:CCC

[76] J. D. Golic and L. O’Connor. A cryptanalysis of clock-controlled shift registers with multiple steps. Lecture Notes in Computer Science, 1029:174– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golic:1996:DOF

[77] J. D. Golic, M. Salmasizadeh, A. Clark, and A. Khodkar. Discrete optimisation and fast correlation attacks. Lecture Notes in Computer Science, 1029: 186–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bakhtiari:1996:KHF

[78] S. Bakhtiari, R. Safavi-Naini, and J. Pieprzyk. Keyed hash functions. Lecture Notes in Computer Science, 1029: 201–214, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1996:SAA

[79] S.-J. Hwang, C.-C. Chang, and W.- P. Yang. Some active attacks on fast server-aided secret computation protocols for modular exponentiation. Lecture Notes in Computer Science, 1029: 215–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laih:1996:CEE

[80] C.-S. Laih and W.-C. Kuo. Cryptanalysis of the enhanced ElGamal’s signature scheme. Lecture Notes in Computer Science, 1029:228–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:AP

[81] L. Chen. Access with pseudonyms. Lecture Notes in Computer Science, 1029: 232–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:NIA

[82] K. Chen. A new identification algorithm. Lecture Notes in Computer Science, 1029:244–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchsberger:1996:PKC

[83] A. Fuchsberger, D. Gollmann, P. Lothian, and K. G. Paterson. Public-key cryptography on smart cards. Lecture Notes in Computer Science, 1029:250– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaskell:1996:ISC

[84] G. Gaskell and M. Looi. Integrating smart cards into authentication systems. Lecture Notes in Computer Science, 1029:270–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hruby:1996:SCI

[85] J. Hruby. Smart card with interferometric quantum cryptography device. Lecture Notes in Computer Science, 1029:282–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gollmann:1996:CA

[86] D. Gollmann. Cryptographic APIs. Lecture Notes in Computer Science, 1029:290–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Russell:1996:FAN

[87] S. Russell. Foiling active network impersonation attacks made in collusion with an insider. Lecture Notes in Computer Science, 1029:301–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohay:1996:CS

[88] G. Mohay, H. Morarji, Q. Le-Viet, and L. Munday. The CASS shell. Lecture Notes in Computer Science, 1029:313– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albrecht:1996:SDS

[89] R. F. Albrecht. On the structure of discrete systems. Lecture Notes in Computer Science, 1030:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rattray:1996:IRT

[90] C. Rattray. Identification and recognition through shape in complex systems. Lecture Notes in Computer Science, 1030:19–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hill:1996:LAS

[91] G. Hill. A logical approach to system construction. Lecture Notes in Computer Science, 1030:30–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takahara:1996:TMS

[92] Y. Takahara and X. Chen. Task management system. Lecture Notes in Computer Science, 1030:48–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1996:TUT

[93] R. Sato. Toward a unified theory of discrete event systems. Lecture Notes in Computer Science, 1030:62–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

SuarezAraujo:1996:ITG

[94] C. P. Suarez Araujo and R. Moreno-Diaz. Intersensorial transformations: General systems description and implications. Lecture Notes in Computer Science, 1030:73–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thurner:1996:PSP

[95] E. M. Thurner. Proving system properties by means of trigger-graph and Petri nets. Lecture Notes in Computer Science, 1030:88–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Menezes:1996:RMG

[96] P. B. Menezes, J. F. Costa, and A. Sernadas. Refinement mapping for general (discrete event) systems theory. Lecture Notes in Computer Science, 1030:103–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sierocki:1996:STA

[97] I. Sierocki. A system-theoretic approach to planning and programming. Lecture Notes in Computer Science, 1030:117–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miro:1996:EFC

[98] J. Miro and M. Miro-Julia. Equality of functions in CAST. Lecture Notes in Computer Science, 1030:129–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodriguez-Millan:1996:BAE

[99] J. Rodriguez-Millan. Basins of attraction estimation through symbolic graphical computing techniques. Lecture Notes in Computer Science, 1030: 137–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larnac:1996:TFV

[100] M. Larnac, J. Magnier, E. Vandermeulen, and G. Dray. Temporal and functional verification of a symbolic representation of complex systems. Lecture Notes in Computer Science, 1030:151–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brielmann:1996:MDE

[101] M. Brielmann. Modelling differential equations by basic information technology means. Lecture Notes in Computer Science, 1030:163–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapsner:1996:STA

[102] F. Kapsner. System theoretic aspects of software architectures. Lecture Notes in Computer Science, 1030:175–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parets:1996:LDC

[103] J. Parets and J. C. Torres. A language for describing complex-evolutive software systems. Lecture Notes in Computer Science, 1030:181–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kocourek:1996:APM

[104] C. Kocourek. An architecture for process modeling and execution support. Lecture Notes in Computer Science, 1030:201–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahlert:1996:FSG

[105] S. Kahlert, T. Kruse, and D. Monjau. A formal semantics for a graphical model representation in system design. Lecture Notes in Computer Science, 1030:217–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MeyerzuBexten:1996:UIM

[106] E. Meyer zu Bexten and C. Moraga. A user-interface for a microsystems high-level simulator. Lecture Notes in Computer Science, 1030:228–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joslyn:1996:CED

[107] C. Joslyn and S. Henderson. CAST extensions to DASME to support generalized information theory. Lecture Notes in Computer Science, 1030:237– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodriguez-Millan:1996:SCA

[108] J. Rodriguez-Millan and J. Cardillo. Symbolic computing aided design of nonlinear PID controllers. Lecture Notes in Computer Science, 1030:253– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otero:1996:DES

[109] R. P. Otero, A. Barreiro, P. Cabalar, and D. Lorenzo. Discrete event simulation in an environment for temporal expert systems. Lecture Notes in Computer Science, 1030:271–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thurner:1996:TTM

[110] E. M. Thurner and C. Wincheringer. TOMSPIN — A tool for modeling with stochastic Petri nets. Lecture Notes in Computer Science, 1030:283–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacak:1996:HEP

[111] W. Jacak and S. Dreiseitl. Hybrid evolutionary programming: the tools for CAST. Lecture Notes in Computer Science, 1030:289–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otero:1996:AID

[112] R. P. Otero, D. Lorenzo, and P. Cabalar. Automatic induction of DEVS structures. Lecture Notes in Computer Science, 1030:305–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schweizer:1996:SEI

[113] G. Schweizer and M. Voss. Systems engineering and infrastructures for open computer based systems. Lecture Notes in Computer Science, 1030:317–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agsteiner:1996:OOH

[114] K. Agsteiner, D. Monjau, and S. Schulze. Object-oriented high level modeling of complex systems. Lecture Notes in Computer Science, 1030:332– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Praehofer:1996:MOO

[115] H. Praehofer and F. Pichler. Multifacetted, object oriented modeling in the transportation domain. Lecture Notes in Computer Science, 1030:341– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaffer:1996:EST

[116] C. Schaffer, R. J. Raschhofer, and A. Simma. EaSy-Sim: A tool environment for the design of complex, real-time systems. Lecture Notes in Computer Science, 1030:358–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glaesser:1996:SLS

[117] U. Glaesser. Systems level specification and modelling of reactive systems: Concepts, methods, and tools. Lecture Notes in Computer Science, 1030:375– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamaidi:1996:DBE

[118] L. Hamaidi, P. Bourseau, G. Muratet, and J. M. Couretas. DEVS-Based endomorphic agents: Control through deliberative and reactive planning. Lecture Notes in Computer Science, 1030:386–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chroust:1996:SPW

[119] G. Chroust and W. Jacak. Software processes, work flow and work cell design — separated by a common paradigm? Lecture Notes in Computer Science, 1030:403–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sydow:1996:HPP

[120] A. Sydow, T. Lux, and R.-P. Schaefer. High-performance parallel computing for analyzing urban air pollution. Lecture Notes in Computer Science, 1030:417–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roszkowska:1996:PAD

[121] E. Roszkowska. Performance analysis of a dual robot system using a geometric concurrency model. Lecture Notes in Computer Science, 1030:423– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vysoky:1996:CAR

[122] P. Vysoky and Z. Pechal. Computer aided rule extraction from examples. Lecture Notes in Computer Science, 1030:437–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

QuevedoLosada:1996:CMG

[123] J. C. Quevedo Losada, O. Bolivar Toledo, and R. Moreno Diaz. Cast methods for generation of nonorthogonal complete transforms. Lecture Notes in Computer Science, 1030: 447–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Candela:1996:CSA

[124] S. Candela, C. Garcia, F. Alayon, and J. Muthoz. Cast system approach for visual inspection. Lecture Notes in Computer Science, 1030:459–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pichler:1996:FDG

[125] F. Pichler and J. Scharinger. Finite dimensional generalized Baker dynamical systems for cryptographic applications. Lecture Notes in Computer Science, 1030:465–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1996:IAF

[126] D. M. Dubois. Introduction of the Aristotle’s final causation in CAST: Concept and method of incursion and hyperincursion. Lecture Notes in Computer Science, 1030:477–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1996:SLC

[127] D. M. Dubois. A semantic logic for CAST related to zuse, deutsch and McCulloch and pitts computing principles. Lecture Notes in Computer Science, 1030:494–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BarahonadaFonseca:1996:VMA

[128] J. Barahona da Fonseca. Validation of a model of an AGVs scheduling heuristic using radio-taxi data. Lecture Notes in Computer Science, 1030:511–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bustio:1996:TSG

[129] F. Bustio, P. Corcuera, and E. Mora. Training simulator for Garoña nuclear power plant. Lecture Notes in Computer Science, 1030:523–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klempous:1996:SCD

[130] R. Klempous, B. Lysakowska, and J. Nikodem. Supervisory control and data acquisition system for the gas flow networks. Lecture Notes in Computer Science, 1030:530–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufeler:1996:ASA

[131] J.-F. Kaufeler. Ada: A sceptical assessment. Lecture Notes in Computer Science, 1031:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mangold:1996:AAO

[132] K. Mangold. Ada95 — an approach to overcome the software crisis? Lecture Notes in Computer Science, 1031:4–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Richard-Foy:1996:SAE

[133] M. Richard-Foy. Safe Ada executive: An executive for Ada safety critical applications. Lecture Notes in Computer Science, 1031:11–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

David:1996:DFT

[134] P. David, T. Planche, A. Correge, and J. F. Chane. Developing fault tolerant software in Ada for real-time dependable systems. Lecture Notes in Computer Science, 1031:21–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1996:PAS

[135] R. Brown. The practical application of safety techniques on an Ada based project. Lecture Notes in Computer Science, 1031:51–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taurisano:1996:AML

[136] G. Taurisano. Ada in mixed language applications. Lecture Notes in Computer Science, 1031:58–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kempe:1996:HDS

[137] M. Kempe. Heterogeneous data structures and cross-classification of objects with Ada95. Lecture Notes in Computer Science, 1031:71–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bundgaard:1996:ABA

[138] J. Bundgaard. An ANDF based Ada 95 compiler system. Lecture Notes in Computer Science, 1031:81–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skinderhaug:1996:PTC

[139] P. I. Skinderhaug and B. Rognes. Performance tuning of a check-out system coded in Ada. Lecture Notes in Computer Science, 1031:99–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monfort:1996:ADB

[140] J.-N. Monfort and V. Q. Ribal. Ariane 5: Development of the on-board software. Lecture Notes in Computer Science, 1031:111–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laurens:1996:PGC

[141] A. Laurens. PRONAOS ground control center: First operational Ada application in C.N.E.S. Lecture Notes in Computer Science, 1031:124–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rybin:1996:AGG

[142] S. Rybin, E. Zueff, and A. Strohmeier. ASIS for GNAT: Goals, problems and implementation strategy. Lecture Notes in Computer Science, 1031:139–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bailes:1996:KAO

[143] P. A. Bailes, P. Burnim, M. Chapman, and E. Salzman. KBSE and Adaobject and enabling technology. Lecture Notes in Computer Science, 1031: 152–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birus:1996:EAI

[144] T. Birus, P. Knueven, E. Kuzemchak, and J. Rosenzweig. Extending the Aria 95 initial conditions for preelaboration for use in real-time systems. Lecture Notes in Computer Science, 1031:164–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davies:1996:UAE

[145] B. Davies and D. A. Rothwell. The use of Ada for the ENVISAT-l simulator. Lecture Notes in Computer Science, 1031:170–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Molenmaker:1996:OUN

[146] K. Molenmaker. Objects at use in nautical simulators. Lecture Notes in Computer Science, 1031:178–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Obermayer:1996:DID

[147] P. E. Obermayer, G. Schueer, and R. Landwehr. DIS — an interface to distributed interactive simulation. Lecture Notes in Computer Science, 1031:187–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gargaro:1996:PPA

[148] A. Gargaro, Y. Kermarrec, L. Pautet, and S. Tardieu. PARIS — partitioned Ada for remotely invoked services. Lecture Notes in Computer Science, 1031:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kermarrec:1996:PDS

[149] Y. Kermarrec and L. Pautet. Programming distributed systems with both Ada 95 and PVM. Lecture Notes in Computer Science, 1031:206–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choukair:1996:DOO

[150] Z. Choukair and Y. Kermarrec. Distributed object oriented programming and interoperability for Ada 95: An OMG/CORBA approach. Lecture Notes in Computer Science, 1031:217– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerlich:1996:DPS

[151] R. Gerlich and M. Kerep. Distributed and parallel systems and HOOD 4. Lecture Notes in Computer Science, 1031: 228–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Battaglia:1996:RRE

[152] M. Battaglia and G. Savoia. ReverseNICE: A re-engineering methodology and supporting tool. Lecture Notes in Computer Science, 1031:244– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneeweiss:1996:TSM

[153] H. Schneeweiss, V. Amiot, and O. Vix. Translating shlaer/mellor object-oriented analysis models into Ada95. Lecture Notes in Computer Science, 1031:249–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simoens:1996:IOO

[154] R. Simoens. The introduction of an object-oriented analysis/design method and object-oriented metrics in the software development life-cycle. Lecture Notes in Computer Science, 1031:271–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchs:1996:MVT

[155] D. Buchs, C. Buffard, and P. Racloz. Modeling and validation of tasks with algebraic structured nets. Lecture Notes in Computer Science, 1031:284– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanden:1996:DCS

[156] B. I. Sanden. Design of concurrent software based on problem concurrency. Lecture Notes in Computer Science, 1031:298–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matthewman:1996:ESE

[157] A. Matthewman. ECLIPS — A successful experiment combining CCSDS SFDUs, X/motif, HOOD and Ada. Lecture Notes in Computer Science, 1031:311–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alapide:1996:ATA

[158] A. Alapide, S. Candia, M. Cinnella, and S. Quaranta. Applying teamwork/Aria and RAISE for developing an air traffic control application. Lecture Notes in Computer Science, 1031: 331–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Croxford:1996:BTV

[159] M. Croxford and J. Sutton. Breaking through the V and V bottleneck. Lecture Notes in Computer Science, 1031:344–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bossard:1996:PPH

[160] F. Bossard. Periodic processing in hard real-time systems: Assessment of different design models in Ada. Lecture Notes in Computer Science, 1031:355– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cornwell:1996:TSO

[161] P. Cornwell and A. Wellings. Transaction specification for object-oriented real-time systems in HRT-HOOD. Lecture Notes in Computer Science, 1031: 365–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Battini:1996:ESB

[162] F. Battini, P. L. Mantovani, and M. Mattavelli. Evaluation of a SPARC board equipped with the Ada tasking coprocessor (ATAC). Lecture Notes in Computer Science, 1031:379–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjoernfot:1996:ATA

[163] L. Bjoernfot. Ada and timed automata. Lecture Notes in Computer Science, 1031:389–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbey:1996:TAO

[164] S. Barbey. Testing Ada 95 objectoriented programs. Lecture Notes in Computer Science, 1031:406–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heitz:1996:ARR

[165] M. Heitz. Achieving reusable and reliable client-server code using HOOD[TM] automated code generation for Ada95 and C++ targets. Lecture Notes in Computer Science, 1031: 419–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodik:1996:ADF

[166] R. Bodik and R. Gupta. Array data flow analysis for load-store optimizations in superscalar architectures. Lecture Notes in Computer Science, 1033: 1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Altman:1996:ESI

[167] E. R. Altman, G. R. Gao, and R. Govindarajan. An experimental study of an ILP-based exact solution method for software pipelining. Lecture Notes in Computer Science, 1033: 16–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeDinechin:1996:ISA

[168] B. D. De Dinechin. Insertion scheduling: An alternative to list scheduling for modulo schedulers. Lecture Notes in Computer Science, 1033:31–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Creusillet:1996:IAR

[169] B. Creusillet and F. Irigoin. Interprocedural array region analyses. Lecture Notes in Computer Science, 1033:46– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hall:1996:IAP

[170] M. W. Hall, B. R. Murphy, S. P. Amarasinghe, and S.-W. Liao. Interprocedural analysis for parallelization. Lecture Notes in Computer Science, 1033:61–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1996:IAD

[171] L. Choi and P.-C. Yew. Interprocedural array data-flow analysis for cache coherence. Lecture Notes in Computer Science, 1033:81–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1996:IPC

[172] T. Nguyen, J. Gu, and Z. Li. An interprocedural parallelizing compiler and its support for memory hierarchy research. Lecture Notes in Computer Science, 1033:96–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dechering:1996:VCC

[173] P. F. G. Dechering, J. A. Trescher, J. P. M. De Vreught, and H. J. Sips. V-cal: A calculus for the compilation of data parallel languages. Lecture Notes in Computer Science, 1033:111– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelly:1996:TCI

[174] W. Kelly, W. Pugh, E. Rosser, and T. Shpeisman. Transitive closure of infinite graphs and its applications. Lecture Notes in Computer Science, 1033: 126–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blume:1996:DDS

[175] W. Blume and R. Eigenmann. Demand-driven, symbolic range propagation. Lecture Notes in Computer Science, 1033:141–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kennedy:1996:OFS

[176] K. Kennedy, J. Mellor-Crummey, and G. Roth. Optimizing Fortran 90 shift operations on distributed-memory multicomputers. Lecture Notes in Computer Science, 1033:161–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishizaki:1996:LPA

[177] K. Ishizaki and H. Komatsu. A loop parallelization algorithm for HPF compilers. Lecture Notes in Computer Science, 1033:176–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thirumalai:1996:FAS

[178] A. Thirumalai and J. Ramanujam. Fast address sequence generation for dataparallel programs using integer lattices. Lecture Notes in Computer Science, 1033:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaushik:1996:CAS

[179] S. D. Kaushik, C.-H. Huang, and P. Sadayappan. Compiling array statements for efficient execution on distributedmemory machines: Two-level mappings. Lecture Notes in Computer Science, 1033:209–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stichnoth:1996:CBP

[180] J. M. Stichnoth and T. Gross. A communication backend for parallel language compilers. Lecture Notes in Computer Science, 1033:224–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prakash:1996:PSD

[181] S. Prakash and R. Bagrodia. Parallel simulation of data parallel programs. Lecture Notes in Computer Science, 1033:239–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietz:1996:PPS

[182] H. G. Dietz, T. M. Chung, and T. I. Mattox. A parallel processing support library based on synchronized aggregate communication. Lecture Notes in Computer Science, 1033:254–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeRose:1996:FMI

[183] L. De Rose, K. Gallivan, E. Gallopoulos, and B. Marsolf. FALCON: A MATLAB interactive restructuring compiler. Lecture Notes in Computer Science, 1033:269–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Novack:1996:SMI

[184] S. Novack, J. Hummel, and A. Nicolau. A simple mechanism for improving the accuracy and efficiency of instructionlevel disambiguation. Lecture Notes in Computer Science, 1033:289–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Appelbe:1996:HBC

[185] B. Appelbe, S. Doddapaneni, R. Harmon, and P. May. Hoisting branch conditions — improving super-scalar processor performance. Lecture Notes in Computer Science, 1033:304–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radigan:1996:ILC

[186] J. Radigan, P. Chang, and U. Banerjee. Integer loop code generation for VLIW. Lecture Notes in Computer Science, 1033:318–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Midkiff:1996:DAP

[187] S. P. Midkiff. Dependence analysis in parallel loops with i, k subscripts. Lecture Notes in Computer Science, 1033:331–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palmer:1996:PEN

[188] D. W. Palmer, J. F. Prins, S. Chatterjee, and R. E. Faith. Piecewise execution of nested data-parallel programs. Lecture Notes in Computer Science, 1033:346–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cierniak:1996:RLS

[189] M. Cierniak and W. Li. Recovering logical structures of data. Lecture Notes in Computer Science, 1033:362–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sheffler:1996:EDA

[190] T. J. Sheffler, R. Schreiber, W. Pugh, and J. R. Gilbert. Efficient distribution analysis via graph contraction. Lecture Notes in Computer Science, 1033:377–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palermo:1996:ASD

[191] D. J. Palermo and P. Banerjee. Automatic selection of dynamic data partitioning schemes for distributedmemory multicomputers. Lecture Notes in Computer Science, 1033:392– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayguade:1996:DRA

[192] E. Ayguade, J. Garcia, M. Girones, and M. L. Grande. Data redistribution in an automatic data distribution tool. Lecture Notes in Computer Science, 1033:407–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheatham:1996:GPO

[193] T. Cheatham, A. Fahmy, and D. C. Stefanescu. General purpose optimization technology. Lecture Notes in Computer Science, 1033:422–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McKinley:1996:CAH

[194] K. S. McKinley, S. K. Singhai, G. E. Weaver, and C. C. Weems. Compiler architectures for heterogeneous systems. Lecture Notes in Computer Science, 1033:434–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Philbin:1996:VTN

[195] J. Philbin, S. Jagannathan, and R. Mirani. Virtual topologies: A new concurrency abstraction for high-level parallel languages. Lecture Notes in Computer Science, 1033:450–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kotlyar:1996:APC

[196] V. Kotlyar, K. Pingali, and P. Stodghill. Automatic parallelization of the conjugate gradient algorithm. Lecture Notes in Computer Science, 1033:480–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bik:1996:ASC

[197] A. J. C. Bik and H. A. G. Wijshoff. Annotations for a sparse compiler. Lecture Notes in Computer Science, 1033:500–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghiya:1996:CAP

[198] R. Ghiya and L. J. Hendren. Connection analysis: A practical interprocedural heap analysis for C. Lecture Notes in Computer Science, 1033:515– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dinda:1996:LRT

[199] P. A. Dinda, D. R. O’Hallaron, J. Subhlok, and J. A. Webb. Language and run-time support for network parallel computing. Lecture Notes in Computer Science, 1033:534–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yelon:1996:AUR

[200] J. Yelon and L. V. Kale. Agents: An undistorted representation of problem structure. Lecture Notes in Computer Science, 1033:551–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plevyak:1996:TDC

[201] J. Plevyak and A. A. Chien. Type directed cloning for object-oriented programs. Lecture Notes in Computer Science, 1033:566–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreira:1996:PIG

[202] J. E. Moreira, D. Schouten, and C. Polychronopoulos. The performance impact of granularity control and functional parallelism. Lecture Notes in Computer Science, 1033:581–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simonis:1996:ADC

[203] H. Simonis. Application development with the CHIP system. Lecture Notes in Computer Science, 1034:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Afrati:1996:EFO

[204] F. Afrati, T. Andronikos, and T. G. Kavalieros. On the expressiveness of first-order constraint languages. Lecture Notes in Computer Science, 1034: 22–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belussi:1996:GRA

[205] A. Belussi, E. Bertino, M. Bertolotto, and B. Catania. Generalized relational algebra: Modeling spatial queries in constraint databases. Lecture Notes in Computer Science, 1034:40–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Byon:1996:DCD

[206] J.-H. Byon and P. Z. Revesz. DISCO: A constraint database system with sets. Lecture Notes in Computer Science, 1034:68–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaede:1996:CBQ

[207] V. Gaede and O. Guenther. Constraintbased query optimization and processing. Lecture Notes in Computer Science, 1034:84–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freeston:1996:AMD

[208] M. Freeston. The application of multidimensional indexing methods to constraints. Lecture Notes in Computer Science, 1034:102–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermosilla:1996:TDS

[209] L. Hermosilla and G. M. Kuper. Towards the definition of a spatial objectoriented data model with constraints. Lecture Notes in Computer Science, 1034:120–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reinert:1996:ARI

[210] J. Reinert. Ambiguity for referentlal integrity is undecidable. Lecture Notes in Computer Science, 1034:132–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wallace:1996:MCC

[211] M. Wallace, S. Bressan, and T. Le Provost. Magic checking: Constraint checking for database query optimisation. Lecture Notes in Computer Science, 1034:148–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yakhno:1996:FFS

[212] T. M. Yakhno, G. B. Cheblakov, and V. Z. Zilberfain. FISSURE: Finder of solutions with subdefinite resources. Lecture Notes in Computer Science, 1034:167–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Picard:1996:DLM

[213] R. Picard. Digital libraries: Meeting place for low-level and high-level vision. Lecture Notes in Computer Science, 1035:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1996:CVN

[214] T. S. Huang and R. Lopez. Computer vision in next generation image and video coding. Lecture Notes in Computer Science, 1035:13–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bobick:1996:VAC

[215] A. Bobick. Video annotation: Computers watching video. Lecture Notes in Computer Science, 1035:23–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolle:1996:VR

[216] R. Bolle, Y. Aloimonos, and C. Fermueler. Video representations. Lecture Notes in Computer Science, 1035: 33–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liao:1996:ALV

[217] W.-H. Liao, S. J. Aggarwal, and J. K. Aggarwal. Analysis of left ventricular motion. Lecture Notes in Computer Science, 1035:43–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chien:1996:ACV

[218] C.-H. Chien. Applications of computer vision in space robotics. Lecture Notes in Computer Science, 1035:59– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deriche:1996:OFE

[219] R. Deriche, P. Kornprobst, and G. Aubert. Optical-flow estimation while preserving its discontinuities: A variational approach. Lecture Notes in Computer Science, 1035:71–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagel:1996:STV

[220] H.-H. Nagel, A. Gehrke, M. Haag, and M. Otte. Space-and time-variant estimation approaches and the segmentation of the resulting optical flow fields. Lecture Notes in Computer Science, 1035:81–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bober:1996:CHT

[221] M. Bober and J. Kittler. Combining the Hough transform and multiresolution MRF’s for the robust motion estimation. Lecture Notes in Computer Science, 1035:91–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Odobez:1996:DMB

[222] J.-M. Odobez and P. Bouthemy. Direct model-based image motion segmentation for dynamic scene analysis. Lecture Notes in Computer Science, 1035:101–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shah:1996:MSE

[223] S. Shah and J. K. Aggarwal. Modeling structured environments using robot vision. Lecture Notes in Computer Science, 1035:113–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1996:DTO

[224] Z. He and M. L. Liou. Design trade-offs for real-time block matching motion estimation algorithms. Lecture Notes in Computer Science, 1035:129–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahuja:1996:PMM

[225] N. Ahuja and R. Charan. Pixel matching and motion segmentation in image sequences. Lecture Notes in Computer Science, 1035:139–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:MPR

[226] C.-H. Lee. On the motion problem of a rigid planar patch and its invariance. Lecture Notes in Computer Science, 1035:149–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kishino:1996:VST

[227] F. Kishino, K. Ebihara, and J. Ohya. Virtual-space teleconferencing — realtime detection and reproduction of 3D face and body images. Lecture Notes in Computer Science, 1035:159–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wavering:1996:HPT

[228] A. Wavering, H. Schneiderman, and J. C. Fiala. High-performance tracking with TRICLOPS. Lecture Notes in Computer Science, 1035:171–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eklundh:1996:DAO

[229] J.-O. Eklundh, T. Uhlin, P. Nordlund, and A. Maki. Developing an active observer. Lecture Notes in Computer Science, 1035:181–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuniyoshi:1996:ASV

[230] Y. Kuniyoshi, N. Kita, S. Rougeaux, and T. Suehiro. Active stereo vision system with foveated wide angle lenses. Lecture Notes in Computer Science, 1035:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christensen:1996:CTA

[231] H. I. Christensen, J. Horstmann, and T. Rasmussen. A control theoretical approach to active vision. Lecture Notes in Computer Science, 1035:201– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1996:ARR

[232] Y. Sato. Active rangefinding and recognition with cubicscope. Lecture Notes in Computer Science, 1035:211– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitchen:1996:RTV

[233] L. Kitchen, T.-K. Cheng, and Z.-Q. Liu. Real-time vision is not framerate image analysis. Lecture Notes in Computer Science, 1035:219–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:AVB

[234] C. Wang, Z. Hu, and S. D. Ma. Active vision based stereo vision. Lecture Notes in Computer Science, 1035: 231–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1996:MSR

[235] Z. Zhang. Motion of a stereo rig: Strong weak and self calibration. Lecture Notes in Computer Science, 1035: 241–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xie:1996:CSV

[236] M. Xie and L. Y. Liu. Color stereo vision: Use of appearance constraint and epipolar geometry for feature matching. Lecture Notes in Computer Science, 1035:255–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1996:USM

[237] G. Xu. Unifying stereo, motion and object recognition via epipolar geometry. Lecture Notes in Computer Science, 1035:265–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1996:CVM

[238] J.-Y. Zheng. Constructing a visual memory in terms of uniqueness. Lecture Notes in Computer Science, 1035: 275–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hudson:1996:SMA

[239] M. Hudson. Statistical models and algorithms for tomography. Lecture Notes in Computer Science, 1035:289– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mardia:1996:SSI

[240] K. V. Mardia. Shape statistics and image analysis. Lecture Notes in Computer Science, 1035:297–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:CMC

[241] J.-S. Wang. Cluster Monte Carlo algorithms and their applications. Lecture Notes in Computer Science, 1035: 307–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Won:1996:GSM

[242] C. S. Won. Gibbs sampler and maximum likelihood estimation for unsupervised image segmentations. Lecture Notes in Computer Science, 1035:319– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kluge:1996:LBD

[243] K. Kluge and S. Lakshmanan. Lane boundary detection using deformable templates: Effects of image subsampling on detected lane edge. Lecture Notes in Computer Science, 1035:329– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pelillo:1996:CFR

[244] M. Pelillo. Clique finding relaxation labeling networks. Lecture Notes in Computer Science, 1035:343–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hancock:1996:RSM

[245] E. R. Hancock and R. C. Wilson. Rectifying structural matching errors. Lecture Notes in Computer Science, 1035: 353–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1996:PDR

[246] S. Z. Li. Parallel distributed relaxation labeling. Lecture Notes in Computer Science, 1035:363–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Messmer:1996:SID

[247] B. T. Messmer and H. Bunke. Subgraph isomorphism detection in polynomial time on preprocessed model graphs. Lecture Notes in Computer Science, 1035:373–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papachristou:1996:LPU

[248] P. Papachristou, M. Petrou, and J. Kittler. Line postprocessing using probabilistic relaxation and relative measurements. Lecture Notes in Computer Science, 1035:383–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shashua:1996:MVG

[249] A. Shashua. Multiple-view geometry and photometry. Lecture Notes in Computer Science, 1035:395–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pollefeys:1996:ERS

[250] M. Pollefeys, L. Van Gool, and T. Moons. Euclidean 3D reconstruction from stereo sequences with variable focal lengths. Lecture Notes in Computer Science, 1035:405–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1996:CSI

[251] J. Liu, E. Walker, and J. Mundy. Characterizing the stability of 3D invariants derived from 3D translational symmetry. Lecture Notes in Computer Science, 1035:415–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mundy:1996:MAO

[252] J. Mundy, R. Curwen, J. Liu, and C. Rothwell. MORSE: An architecture for 3D object recognition based on invariants. Lecture Notes in Computer Science, 1035:425–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poggio:1996:FHF

[253] T. Poggio and K.-K. Sung. Finding human faces with a Gaussian mixture distribution-based face model. Lecture Notes in Computer Science, 1035:437– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breuel:1996:HR

[254] T. M. Breuel. Handwriting recognition. Lecture Notes in Computer Science, 1035:447–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ando:1996:ORU

[255] H. Ando. 3D object recognition using bidirectional modular networks. Lecture Notes in Computer Science, 1035: 467–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hild:1996:REM

[256] M. Hild and Y. Shirai. A region extension method for natural outdoor scene interpretation. Lecture Notes in Computer Science, 1035:479–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yin:1996:IPT

[257] J. H. Yin, S. A. Velastin, and A. C. Davies. Image processing techniques for crowd density estimation using a reference image. Lecture Notes in Computer Science, 1035:489–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomanek:1996:ARV

[258] F. Thomanek and E. D. Dickmanns. Autonomous road vehicle guidance in normal traffic. Lecture Notes in Computer Science, 1035:499–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagel:1996:EIS

[259] H.-H. Nagel and H. Kollnig. Evaluation of image sequences from outdoor scenes: Selected problems and solutions. Lecture Notes in Computer Science, 1035:509–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tan:1996:VLD

[260] T. N. Tan, G. D. Sullivan, and K. Baker. Vehicle localisation and discrimination in outdoor traffic scene. Lecture Notes in Computer Science, 1035:519–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1996:SPE

[261] H.-J. Zhang. SWIM: A prototype environment for visual media retrieval. Lecture Notes in Computer Science, 1035: 531–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:VSI

[262] J. C. M. Lee, W. Xiong, D. G. Shen, and R. H. Ma. Video segment indexing through classification and interactive view-based query. Lecture Notes in Computer Science, 1035:541–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swain:1996:VBT

[263] M. J. Swain, C. H. Frankel, and M. Lu. View-based techniques for searching for objects and textures. Lecture Notes in Computer Science, 1035:551–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klauck:1996:GGB

[264] C. Klauck. Graph grammar based object recognition for image retrieval. Lecture Notes in Computer Science, 1035:561–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santini:1996:SM

[265] S. Santini and R. Jain. Similarity matching. Lecture Notes in Computer Science, 1035:571–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amano:1996:DIA

[266] T. Amano, K. Toyokawa, T. Mano, and S. Toriyama. A document image analysis and recognition system for Japanese family registration. Lecture Notes in Computer Science, 1035:583–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breuel:1996:AHR

[267] T. M. Breuel. Applying handwriting recognition to US census forms. Lecture Notes in Computer Science, 1035: 593–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSmedt:1996:ANL

[268] K. De Smedt, H. Horacek, and M. Zock. Architectures for natural language generation: Problems and perspectives. Lecture Notes in Computer Science, 1036:17–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knott:1996:CSC

[269] A. Knott and R. Dale. Choosing a set of coherence relations for text generation: A data-driven approach. Lecture Notes in Computer Science, 1036:47–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maier:1996:TRP

[270] E. Maier. Textual relations as part of multiple links between text segments. Lecture Notes in Computer Science, 1036:68–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dalianis:1996:ANL

[271] H. Dalianis and E. Hovy. Aggregation in natural language generation. Lecture Notes in Computer Science, 1036:88–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daradoumis:1996:TRR

[272] T. Daradoumis. Towards a representation of the rhetorical structure of interrupted exchanges. Lecture Notes in Computer Science, 1036:106–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Green:1996:SDM

[273] S. J. Green and C. DiMarco. Stylistic decision-making in natural language generation. Lecture Notes in Computer Science, 1036:125–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haller:1996:IID

[274] S. M. Haller and S. C. Shapiro. IDP — an interactive discourse planner. Lecture Notes in Computer Science, 1036:144–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jokinen:1996:RAC

[275] K. Jokinen. Reasoning about coherent and cooperative system responses. Lecture Notes in Computer Science, 1036:168–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moeller:1996:DRF

[276] J.-U. Moeller. Domain related focusshifting constraints in dialogues with knowledge based systems. Lecture Notes in Computer Science, 1036:188– ??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pascual:1996:ITF

[277] E. Pascual. Integrating text formatting and text generation. Lecture Notes in Computer Science, 1036:205–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stede:1996:LOM

[278] M. Stede. Lexical options in multilingual generation from a knowledge base. Lecture Notes in Computer Science, 1036:222–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gagnon:1996:PGE

[279] M. Gagnon and G. Lapalme. Pretexte: A generator for the expression of temporal information. Lecture Notes in Computer Science, 1036:238–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kosseim:1996:GGL

[280] L. Kosseim, A. Tutin, R. Kittredge, and G. Lapalme. Generating grammatical and lexical anaphora in assembly instructional texts. Lecture Notes in Computer Science, 1036:260–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abb:1996:IGE

[281] B. Abb, C. Guenther, M. Herweg, and K. Lebeth. Incremental grammatical encoding — an outline of the synphonics formulator. Lecture Notes in Computer Science, 1036:277–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haruno:1996:CBS

[282] M. Haruno, Y. Den, and Y. Matsumoto. A chart-based semantic head driven generation algorithm. Lecture Notes in Computer Science, 1036:300– ??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dik:1996:GSU

[283] S. C. Dik and P. Kahrel. Generating sentences using ProfGlot. Lecture Notes in Computer Science, 1036:314– ??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teich:1996:MTS

[284] E. Teich, L. Degand, and J. Bateman. Multilingual textuality: Some experiences from multilingual text generation. Lecture Notes in Computer Science, 1036:331–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degand:1996:DCM

[285] L. Degand. A Dutch component for a multilingual systemic text generation system. Lecture Notes in Computer Science, 1036:350–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mellish:1996:NLG

[286] C. Mellish, E. Reiter, and J. Levine. Natural language generation applications to technical documentation: A view through Idas. Lecture Notes in Computer Science, 1036:368–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krogh:1996:RA

[287] C. Krogh. The rights of agents. Lecture Notes in Computer Science, 1037:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanLinder:1996:FMA

[288] B. Van Linder, W. Van der Hoek, and J.-J. C. Meyer. Formalising motivational attitudes of agents: On preferences, goals, and commitments. Lecture Notes in Computer Science, 1037: 17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rao:1996:DPP

[289] A. S. Rao. Decision procedures for propositional linear-time belief-desire-intention logics. Lecture Notes in Computer Science, 1037:33–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Singh:1996:SCS

[290] M. P. Singh. Semantical considerations on some primitives for agent specification. Lecture Notes in Computer Science, 1037:49–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Traverso:1996:RAA

[291] P. Traverso, L. Spalazzi, and F. Giunchiglia. Reasoning about acting, sensing, and failure handling: A logic for agents embedded in the real world. Lecture Notes in Computer Science, 1037: 65–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wooldridge:1996:TKC

[292] M. Wooldridge. Time, knowledge, and choice. Lecture Notes in Computer Science, 1037:79–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hexmoor:1996:LR

[293] H. H. Hexmoor. Learning routines. Lecture Notes in Computer Science, 1037: 97–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chu-Carroll:1996:CDR

[294] J. Chu-Carroll and S. Carberry. Conflict detection and resolution in collaborative planning. Lecture Notes in Computer Science, 1037:111–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bicchieri:1996:GSP

[295] C. Bicchieri, E. Ephrati, and A. Antonelli. Games servers play: A procedural approach. Lecture Notes in Computer Science, 1037:127–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gmytrasiewicz:1996:RAO

[296] P. J. Gmytrasiewicz. On reasoning about other agents. Lecture Notes in Computer Science, 1037:143–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tambe:1996:AAT

[297] M. Tambe and P. S. Rosenbloom. Architectures for agents that track other agents in multi-agent worlds. Lecture Notes in Computer Science, 1037:156– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vidal:1996:URA

[298] J. M. Vidal and E. H. Durfee. Using recursive agent models effectively. Lecture Notes in Computer Science, 1037: 171–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonasso:1996:EAI

[299] R. P. Bonasso, D. Kortenkamp, D. P. Miller, and M. Slack. Experiences with an architecture for intelligent, reactive agents. Lecture Notes in Computer Science, 1037:187–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1996:PBA

[300] K. Fischer, J. P. Mueller, and M. Pischel. A pragmatic BDI architecture. Lecture Notes in Computer Science, 1037:203–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Norman:1996:AIM

[301] T. J. Norman and D. Long. Alarms: An implementation of motivated agency. Lecture Notes in Computer Science, 1037:219–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbuceanu:1996:AAB

[302] M. Barbuceanu and M. S. Fox. The architecture of an agent building shell. Lecture Notes in Computer Science, 1037:235–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iglesias:1996:MGP

[303] C. A. Iglesias, J. C. Gonzalez, and J. R. Velasco. MIX: A general purpose multiagent architecture. Lecture Notes in Computer Science, 1037:251–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arcand:1996:CBM

[304] J.-F. Arcand and S.-J. Pelletier. Cognition based multi-agent architecture. Lecture Notes in Computer Science, 1037:267–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mullen:1996:SID

[305] T. Mullen and M. P. Wellman. Some issues in the design of market-oriented agents. Lecture Notes in Computer Science, 1037:283–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beyssade:1996:CLA

[306] C. Beyssade, P. Enjalbert, and C. Lefevre. Cooperating logical agents. Lecture Notes in Computer Science, 1037:299–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giroux:1996:ORA

[307] S. Giroux. Open reflective agents. Lecture Notes in Computer Science, 1037: 315–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lesperance:1996:FLA

[308] Y. Lesperance, H. J. Levesque, F. Lin, and D. Marcu. Foundations of a logical approach to agent programming. Lecture Notes in Computer Science, 1037:331–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayfield:1996:EKA

[309] J. Mayfield, Y. Labrou, and T. Finin. Evaluation of KQML as an agent communication language. Lecture Notes in Computer Science, 1037:347–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malheiro:1996:CCM

[310] B. Malheiro and E. Oliveira. Consistency and context management in a multi-agent belief revision testbed. Lecture Notes in Computer Science, 1037: 361–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sloman:1996:SAT

[311] A. Sloman and R. Poli. SIM AGENT: A toolkit for exploring agent designs. Lecture Notes in Computer Science, 1037:392–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wooldridge:1996:ATA

[312] M. Wooldridge, J. P. Mueller, and M. Tambe. Agent theories, architectures, and languages: A bibliography. Lecture Notes in Computer Science, 1037:408–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giunchiglia:1996:IRBa

[313] F. Giunchiglia and E. Giunchiglia. Ideal and real belief about belief: Some intuitions. Lecture Notes in Computer Science, 1038:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aube:1996:ECO

[314] M. Aube and A. Senteni. Emotions as commitments operators: A foundation for control structure in multi-agents systems. Lecture Notes in Computer Science, 1038:13–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1996:LOM

[315] G. Wagner. A logical and operational model of scalable knowledgeand perception-based agents. Lecture Notes in Computer Science, 1038:26– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rao:1996:ABA

[316] A. S. Rao. AgentSpeak(L): BDI agents speak out in a logical computable language. Lecture Notes in Computer Science, 1038:42–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kinny:1996:MMT

[317] D. Kinny, M. Georgeff, and A. Rao. A methodology and modelling technique for systems of BDI agents. Lecture Notes in Computer Science, 1038:56– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DInverno:1996:FCN

[318] M. D’Inverno and M. Luck. Formalising the contract net as a goal-directed system. Lecture Notes in Computer Science, 1038:72–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ElFallahSeghrouchni:1996:CAM

[319] A. El Fallah Seghrouchni and S. Haddad. A coordination algorithm for multi-agent planning. Lecture Notes in Computer Science, 1038:86–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belakhdar:1996:MAT

[320] O. Belakhdar and J. Ayel. Modelling approach and tool for designing protocols for automated cooperation in multi-agent systems. Lecture Notes in Computer Science, 1038:100–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gu:1996:ASB

[321] C. Gu and T. Ishida. Analyzing the social behavior of contract net protocol. Lecture Notes in Computer Science, 1038:116–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cesta:1996:EDI

[322] A. Cesta, M. Miceli, and P. Rizzo. Effects of different interaction attitudes on a multi-agent system performance. Lecture Notes in Computer Science, 1038:128–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Numaoka:1996:BEA

[323] C. Numaoka. Bacterial evolution algorithm for rapid adaptation. Lecture Notes in Computer Science, 1038:139– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohira:1996:DIC

[324] T. Ohira, R. Sawatari, and M. Tokoro. Distributed interaction with computon. Lecture Notes in Computer Science, 1038:149–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeijs:1996:SAM

[325] C. Baeijs, Y. Demazeau, and L. Alvares. SIGMA: Application of multiagent systems to cartographic generalization. Lecture Notes in Computer Science, 1038:163–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1996:DTM

[326] K. Fischer and J. Mueller. A decisiontheoretic model for cooperative transportation scheduling. Lecture Notes in Computer Science, 1038:177–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guessoum:1996:RTA

[327] Z. Guessoum and M. Dojat. A realtime agent model in an asynchronousobject environment. Lecture Notes in Computer Science, 1038:190–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klusch:1996:CFA

[328] M. Klusch and O. Shehory. Coalition formation among rational information agents. Lecture Notes in Computer Science, 1038:204–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lanzola:1996:CAI

[329] G. Lanzola, S. Falasconi, and M. Stefanelli. Cooperating agents implementing distributed patient management. Lecture Notes in Computer Science, 1038:218–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lai:1996:AHH

[330] X. Lai and R. A. Rueppel. Attacks on the HKM/HFX cryptosystem. Lecture Notes in Computer Science, 1039:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knudsen:1996:TDS

[331] L. R. Knudsen and T. A. Berson. Truncated differentials of SAFER. Lecture Notes in Computer Science, 1039:15– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vaudenay:1996:WKB

[332] S. Vaudenay. On the weak keys of blowfish. Lecture Notes in Computer Science, 1039:27–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blaze:1996:HBE

[333] M. Blaze. High-bandwidth encryption with low-bandwidth smartcards. Lecture Notes in Computer Science, 1039: 33–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jenkins:1996:I

[334] R. J. Jenkins. ISAAC. Lecture Notes in Computer Science, 1039:41–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geiselmann:1996:NHF

[335] W. Geiselmann. A note on the hash function of Tillich and Zemor. Lecture Notes in Computer Science, 1039:51– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobbertin:1996:CM

[336] H. Dobbertin. Cryptanalysis of MD4. Lecture Notes in Computer Science, 1039:53–69, 1996. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobbertin:1996:RSV

[337] H. Dobbertin, A. Bosselaers, and B. Preneel. RIPEMD-160: A strengthened version of RIPEMD. Lecture Notes in Computer Science, 1039:71– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nyberg:1996:FAH

[338] K. Nyberg. Fast accumulated hashing. Lecture Notes in Computer Science, 1039:83–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1996:TFN

[339] Ross Anderson and Eli Biham. Tiger: A fast new hash function. Lecture Notes in Computer Science, 1039:89–97, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.com/link.asp?
id=fx0261047446n136;   http://www.
springerlink.com/openurl.asp?genre=
article&issn=0302-9743&volume=1039&
spage=89;   http://www.springerlink.
com/openurl.asp?genre=article&id=
doi:10.1007/3-540-60865-6_46.

Rijmen:1996:CS

[340] V. Rijmen, J. Daemen, B. Preneel, and A. Bosselaers. The cipher shark. Lecture Notes in Computer Science, 1039: 99–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1996:TPP

[341] R. Anderson and E. Biham. Two practical and provably secure block ciphers: Bear and lion. Lecture Notes in Computer Science, 1039:113–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneier:1996:UFN

[342] B. Schneier and J. Kelsey. Unbalanced Feistel networks and block cipher design. Lecture Notes in Computer Science, 1039:121–144, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

                                        /www.counterpane.com/unbalanced_
                                        feistel.html.

Clark:1996:CFC

[343] A. Clark, J. D. Golic, and E. Dawson. A comparison of fast correlation attacks. Lecture Notes in Computer Science, 1039:145–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Penzhorn:1996:CAS

[344] W. T. Penzhorn. Correlation attacks on stream ciphers: Computing lowweight parity checks based on errorcorrecting codes. Lecture Notes in Computer Science, 1039:159–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golic:1996:SNF

[345] J. D. Golic. On the security of nonlinear filter generators. Lecture Notes in Computer Science, 1039:173–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucks:1996:FLR

[346] S. Lucks. Faster Luby-Rackoff ciphers. Lecture Notes in Computer Science, 1039:189–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsui:1996:NSB

[347] M. Matsui. New structure of block ciphers with provable security against differential and linear cryptanalysis. Lecture Notes in Computer Science, 1039:205–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wermter:1996:LAN

[348] S. Wermter, E. Riloff, and G. Scheler. Learning approaches for natural language processing. Lecture Notes in Computer Science, 1040:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sharkey:1996:SLR

[349] N. E. Sharkey and A. J. C. Sharkey. Separating learning and representation. Lecture Notes in Computer Science, 1040:17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lawrence:1996:NLG

[350] S. Lawrence, S. Fong, and C. L. Giles. Natural language grammatical inference: a comparison of recurrent neural networks and machine learning methods. Lecture Notes in Computer Science, 1040:33–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayward:1996:ERG

[351] R. Hayward, A. Tickle, and J. Diederich. Extracting rules for grammar recognition from Cascade-2 networks. Lecture Notes in Computer Science, 1040:48– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scheler:1996:GEP

[352] G. Scheler. Generating English plural determiners from semantic representations: a neural network learning approach. Lecture Notes in Computer Science, 1040:61–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winiwarter:1996:KAC

[353] W. Winiwarter, E. Schweighofer, and D. Merkl. Knowledge acquisition in concept and document spaces by using self-organizing neural networks. Lecture Notes in Computer Science, 1040: 75–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1996:UHC

[354] V. Weber and S. Wermter. Using hybrid connectionist learning for speech/ language analysis. Lecture Notes in Computer Science, 1040:87–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1996:CSA

[355] G. Lee and J.-H. Lee. A connectionist/symbolic architecture of spoken Korean processing. Lecture Notes in Computer Science, 1040:102–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geutner:1996:IDL

[356] P. Geutner, B. Suhm, F.-D. Buoe, and T. Kemp. Integrating different learning approaches into a multilingual spoken language translation system. Lecture Notes in Computer Science, 1040:117–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1996:LLU

[357] T. C. Smith and I. H. Witten. Learning language using genetic algorithms. Lecture Notes in Computer Science, 1040: 132–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ersan:1996:SSD

[358] M. Ersan and E. Charniak. A statistical syntactic disambiguation program and what it learns. Lecture Notes in Computer Science, 1040:146–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hogenhout:1996:TSG

[359] W. R. Hogenhout and Y. Matsumoto. Training stochastic grammars on semantical categories. Lecture Notes in Computer Science, 1040:160–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fong:1996:LRP

[360] E. W. Fong and D. Wu. Learning restricted probabilistic link grammars. Lecture Notes in Computer Science, 1040:173–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franz:1996:LPA

[361] A. Franz. Learning PP attachment from corpus statistics. Lecture Notes in Computer Science, 1040:188–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruenwald:1996:MDL

[362] P. Gruenwald. A minimum description length approach to grammar inference. Lecture Notes in Computer Science, 1040:203–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mast:1996:ACD

[363] M. Mast, H. Niemann, E. Noeth, and E. G. Schukat-Talamazzini. Automatic classification of dialog acts with semantic classification trees and polygrams. Lecture Notes in Computer Science, 1040:217–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelson:1996:SSN

[364] S. P. Engelson and I. Dagan. Sample selection in natural language learning. Lecture Notes in Computer Science, 1040:230–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huffman:1996:LIE

[365] S. B. Huffman. Learning information extraction patterns from examples. Lecture Notes in Computer Science, 1040:246–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hastings:1996:IAL

[366] P. M. Hastings. Implications of an automatic lexical acquisition system. Lecture Notes in Computer Science, 1040: 261–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riloff:1996:ULE

[367] E. Riloff. Using learned extraction patterns for text classification. Lecture Notes in Computer Science, 1040:275– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soderland:1996:IIL

[368] S. Soderland, D. Fisher, J. Aseltine, and W. Lehnert. Issues in inductive learning of domain-specific text extraction rules. Lecture Notes in Computer Science, 1040:290–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aone:1996:AML

[369] C. Aone and S. W. Bennett. Applying machine learning to anaphora resolution. Lecture Notes in Computer Science, 1040:302–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cardie:1996:EML

[370] C. Cardie. Embedded machine learning systems for natural language processing: a general framework. Lecture Notes in Computer Science, 1040:315– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamazaki:1996:AUH

[371] T. Yamazaki, M. J. Pazzani, and C. Merz. Acquiring and updating hierarchical knowledge for machine translation based on a clustering technique. Lecture Notes in Computer Science, 1040:329–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moulinier:1996:AEM

[372] I. Moulinier and J.-G. Ganascia. Applying an existing machine learning algorithm to text categorization. Lecture Notes in Computer Science, 1040:343– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zelle:1996:CRU

[373] J. M. Zelle and R. J. Mooney. Comparative, results on using inductive logic programming for corpus-based parser construction. Lecture Notes in Computer Science, 1040:355–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mooney:1996:LPT

[374] R. J. Mooney and M. E. Califf. Learning the past tense of English verbs using inductive logic programming. Lecture Notes in Computer Science, 1040: 370–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Federici:1996:DAP

[375] S. Federici, V. Pirrelli, and F. Yvon. A dynamic approach to paradigm-driven analogy. Lecture Notes in Computer Science, 1040:385–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Osborne:1996:CPH

[376] M. Osborne. Can punctuation help learning? Lecture Notes in Computer Science, 1040:399–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joshi:1996:UPC

[377] A. K. Joshi and B. Srinivas. Using parsed corpora for circumventing parsing. Lecture Notes in Computer Science, 1040:413–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacquemin:1996:SSA

[378] C. Jacquemin. A symbolic and surgical acquisition of terms through variation. Lecture Notes in Computer Science, 1040:425–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaneda:1996:RLA

[379] S. Kaneda, H. Almuallim, Y. Akiba, and M. Ishii. A revision learner to acquire verb selection rules from humanmade rules and examples. Lecture Notes in Computer Science, 1040:439– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahn:1996:LTT

[380] U. Hahn, M. Klenner, and K. Schnattinger. Learning from texts — a terminological metareasoning perspective. Lecture Notes in Computer Science, 1040:453–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1996:HPM

[381] R. C. Agarwal, F. G. Gustavson, S. M. Balle, and M. Joshi. A high performance matrix multiplication algorithm for MPPs. Lecture Notes in Computer Science, 1041:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ala:1996:IMM

[382] G. Ala, E. Francomano, and A. Tortorici. Iterative moment method for electromagnetic transients in grounding systems on CRAY T3D. Lecture Notes in Computer Science, 1041:9–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1996:ACS

[383] B. S. Andersen and N. J. Soerensen. Analysis of crystalline solids by means of a parallel FEM method. Lecture Notes in Computer Science, 1041:17– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antonuccio-Delogu:1996:PST

[384] V. Antonuccio-Delogu, U. Becciani, and F. Magugliani. Parallelization strategies for tree N-body codes. Lecture Notes in Computer Science, 1041: 24–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artemiev:1996:NSS

[385] S. S. Artemiev, S. A. Gusev, and O. G. Monakhov. Numerical solution of stochastic differential equations on Transputer network. Lecture Notes in Computer Science, 1041:33–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourbonnais:1996:DSC

[386] R. Bourbonnais. Development of a stencil compiler for one-dimensional convolution operators on the CM-5. Lecture Notes in Computer Science, 1041:38–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brezany:1996:APA

[387] P. Brezany, V. Sipkova, B. Chapman, and R. Greimel. Automatic parallelization of the AVL FIRE benchmark for a distributed-memory system. Lecture Notes in Computer Science, 1041:50–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:1996:DCA

[388] M. Bubak. 2-D cellular automata and short range molecular dynamics programs for simulations on networked workstations and parallel computers. Lecture Notes in Computer Science, 1041:61–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:1996:PBP

[389] M. Bubak, W. Funika, J. Moscinski, and D. Tasak. Pablo-Based performance monitoring tool for PVM applications. Lecture Notes in Computer Science, 1041:69–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:1996:LAC

[390] M. Bubak, W. Owczarz, and J. Wasniewski. Linear algebra computation on parallel machines. Lecture Notes in Computer Science, 1041:79–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casinelli:1996:NCR

[391] T. Casinelli, M. La Manna, and A. Starita. A neural classifier for radar images. Lecture Notes in Computer Science, 1041:88–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1996:SPL

[392] J. Choi, J. Demmel, I. Dhillon, J. Dongarra, S. Ostrouchov, A. Petitet, K. Stanley, D. Walker, and R. C. Whaley. ScaLAPACK: A portable linear algebra library for distributed memory computers — design issues and performance. Lecture Notes in Computer Science, 1041:95–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1996:PSP

[393] Jaeyoung Choi, J. Dongarra, S. Ostrouchov, A. Petitet, D. Walker, and R. C. Whaley. A proposal for a set of parallel basic linear algebra subprograms. Lecture Notes in Computer Science, 1041:107–114, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crone:1996:PIL

[394] L. G. C. Crone, H. Van Dop, and W. Lourens. Parallel implementation of a Lagrangian stochastic particle model for turbulent dispersion in fluids. Lecture Notes in Computer Science, 1041:115–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dackland:1996:RRM

[395] K. Dackland and B. Kaagstroem. Reduction of a regular matrix pair (A, B) to block Hessenberg triangular form. Lecture Notes in Computer Science, 1041:125–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiMartino:1996:PAN

[396] B. Di Martino. Parallelization of algorithms for neural networks. Lecture Notes in Computer Science, 1041:134– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiMartino:1996:PPB

[397] B. Di Martino, N. Mazzocca, and S. Russo. Paradigms for the parallelization of branch and bound algorithms. Lecture Notes in Computer Science, 1041:141–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimov:1996:TDV

[398] I. Dimov, K. Georgiev, J. Wasniewski, and Z. Zlatev. Three-dimensional version of the Danish Eulerian model. Lecture Notes in Computer Science, 1041: 151–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1996:PFI

[399] J. J. Dongarra, J. Du Croz, S. Hammarling, and J. Wasniewski. A proposal for a Fortran 90 interface for LA-PACK. Lecture Notes in Computer Science, 1041:158–165, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1996:ST

[400] J. Dongarra and A. Petitet. ScaLA-PACK tutorial. Lecture Notes in Computer Science, 1041:166–176, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fet:1996:HPC

[401] Y. I. Fet and A. P. Vazhenin. Highly parallel concentrated heterogeneous computing. Lecture Notes in Computer Science, 1041:177–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Field:1996:APP

[402] M. R. Field. Adaptive polynomial preconditioners for the conjugate gradient algorithm. Lecture Notes in Computer Science, 1041:189–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filippone:1996:IPE

[403] S. Filippone. The IBM parallel engineering and scientific subroutine library. Lecture Notes in Computer Science, 1041:199–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filippone:1996:SPE

[404] S. Filippone and C. Vittoli. Some preliminary experiences with sparse BLAS in parallel iterative solvers. Lecture Notes in Computer Science, 1041:207– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fontdecaba:1996:LBN

[405] E. Fontdecaba, A. Gonzalez, and J. Labarta. Load balancing in a network flow optimization code. Lecture Notes in Computer Science, 1041:214– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ford:1996:ULV

[406] R. W. Ford, A. P. Nisbet, and J. M. Bull. User-level VSM optimization and its.Application. Lecture Notes in Computer Science, 1041:223–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Getov:1996:BCM

[407] V. Getov. Benchmarking the cache memory effect. Lecture Notes in Computer Science, 1041:233–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gimenez:1996:EJA

[408] D. Gimenez, V. Hernandez, and A. M. Vidal. Efficient Jacobi algorithms on multicomputers. Lecture Notes in Computer Science, 1041:247–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glimm:1996:FTP

[409] J. Glimm, J. Grove, X. L. Li, and R. Young. Front tracking: A parallelized approach for internal boundaries and interfaces. Lecture Notes in Computer Science, 1041:257–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldman:1996:PGT

[410] V. V. Goldman and G. Cats. Program generation techniques for the development and maintenance of numerical weather forecast grid models. Lecture Notes in Computer Science, 1041:267–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guest:1996:HPC

[411] M. F. Guest, E. Apra, D. E. Bernholdt, and H. A. Fruechtl. High performance computational chemistry: NWChem and fully distributed parallel applications. Lecture Notes in Computer Science, 1041:278–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hammer:1996:PIM

[412] B. Hammer and O. H. Nielsen. Parallel ab-initio molecular dynamics. Lecture Notes in Computer Science, 1041: 295–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hegarty:1996:DDD

[413] D. F. Hegarty, M. T. Kechadi, and K. A. Dawson. Dynamic domain decomposition and load balancing for parallel simulations of long-chained molecules. Lecture Notes in Computer Science, 1041:303–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holz:1996:CFA

[414] H. J. Holz and M. H. Loew. Concurrency in feature analysis. Lecture Notes in Computer Science, 1041:313– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1996:PIS

[415] D. Hu, F. Lou, and A. Sameh. A parallel iterative solver for almost blockdiagonal linear systems. Lecture Notes in Computer Science, 1041:323–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaagstroem:1996:DGM

[416] B. Kaagstroem and M. Raennar. Distributed general matrix multiply and add for a 2D mesh processor network. Lecture Notes in Computer Science, 1041:333–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitowski:1996:DPC

[417] J. Kitowski. Distributed and parallel computing of short-range molecular dynamics. Lecture Notes in Computer Science, 1041:345–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krasnitz:1996:LFT

[418] A. Krasnitz. Lattice field theory in a parallel environment. Lecture Notes in Computer Science, 1041:352–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lagana:1996:PTI

[419] A. Lagana, S. Crocchianti, G. O. De Aspuru, and R. Gargano. Parallel time independent quantum calculations of atom diatom reactivity. Lecture Notes in Computer Science, 1041:361–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1996:POR

[420] J. Larsen, L. Frellesen, J. Jansson, and F. If. Parallel oil reservoir simulation. Lecture Notes in Computer Science, 1041:371–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lazaro:1996:FSM

[421] J. A. G. Lazaro. Formal specification of multicomputers. Lecture Notes in Computer Science, 1041:380–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lomdahl:1996:MMP

[422] P. S. Lomdahl and D. M. Beazley. Multi-million particle molecular dynamics on MPPs. Lecture Notes in Computer Science, 1041:391–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luthi:1996:WPU

[423] P. O. Luthi, B. Chopard, and J. F. Wagen. Wave propagation in urban microcells: a massively parallel approach using the TLM method. Lecture Notes in Computer Science, 1041:408–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McDonald:1996:NNP

[424] K. McDonald. The NAG numerical PVM library. Lecture Notes in Computer Science, 1041:419–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Masselot:1996:CAM

[425] W. A. Masselot and B. Chopard. Cellular automata modeling of snow transport by wind. Lecture Notes in Computer Science, 1041:429–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monakhov:1996:PAM

[426] O. G. Monakhov. Parallel algorithm for mapping parallel program into pyramidal multiprocessor. Lecture Notes in Computer Science, 1041:436–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:1996:DPM

[427] O. H. Nielsen. Data-parallel molecular dynamics with neighbor-lists. Lecture Notes in Computer Science, 1041: 443–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nordlund:1996:VAM

[428] A. Nordlund, K. Galsgaard, and R. F. Stein. Visualizing astrophysical 3D MHD turbulence. Lecture Notes in Computer Science, 1041:450–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ostromsky:1996:PSQ

[429] T. Ostromsky, P. C. Hansen, and Z. Zlatev. A parallel sparse QR-Factorization algorithm. Lecture Notes in Computer Science, 1041:462–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinar:1996:DLP

[430] A. Pinar, U. V. Catalyuerek, C. Aykanat, and M. Pinar. Decomposing linear programs for parallel solution. Lecture Notes in Computer Science, 1041:473–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sabeur:1996:PCN

[431] Z. A. Sabeur. A parallel computation of the Navier–Stokes equation for the simulation of free surface flows with the volume of fluid method. Lecture Notes in Computer Science, 1041:483– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchez:1996:IPP

[432] J. L. Sanchez, J. M. Garcia, and J. Fernandez. Improving the performance of parallel triangularization of a sparse matrix using a reconfigurable multicomputer. Lecture Notes in Computer Science, 1041:493–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanin:1996:CTI

[433] E. Tanin, T. M. Kurc, C. Aykanat, and B. Oezguec. Comparison of two image-space subdivision algorithms for direct volume rendering on distributed-memory multicomputers. Lecture Notes in Computer Science, 1041:503–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tingleff:1996:CHT

[434] O. Tingleff. Communication harnesses for Transputer systems with tree structure and cube structure. Lecture Notes in Computer Science, 1041:513–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valasek:1996:TIP

[435] P. Valasek and R. Hetzel. A thorough investigation of the projector quantum Monte Carlo method using MPP technologies. Lecture Notes in Computer Science, 1041:524–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wcislo:1996:DSS

[436] R. Wcislo, J. Kitowski, and J. Moscinski. Distributed simulation of a set of elastic macro objects. Lecture Notes in Computer Science, 1041:534–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamasaki:1996:PIM

[437] T. Yamasaki. Parallelization of ab initio molecular dynamics method. Lecture Notes in Computer Science, 1041: 540–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zlatev:1996:PCL

[438] Z. Zlatev, I. Dimov, K. Georgiev, and J. Wasniewski. Parallel computations with large atmospheric models. Lecture Notes in Computer Science, 1041:550–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weiss:1996:ALM

[439] G. Weiss. Adaptation and learning in multi-agent systems: Some remarks and a bibliography. Lecture Notes in Computer Science, 1042:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Byrne:1996:RAG

[440] C. Byrne and P. Edwards. Refinement in agent groups. Lecture Notes in Computer Science, 1042:22–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carmel:1996:OMM

[441] D. Carmel and S. Markovitch. Opponent modeling in multi-agent systems. Lecture Notes in Computer Science, 1042:40–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glicoes:1996:MAE

[442] L. Glicoes, R. Staats, and M. Huhns. A multi-agent environment for department of defense distribution. Lecture Notes in Computer Science, 1042:53– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldman:1996:MSL

[443] C. V. Goldman and J. S. Rosenschein. Mutually supervised learning in multiagent systems. Lecture Notes in Computer Science, 1042:85–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gu:1996:FDR

[444] P. Gu and A. B. Maddox. A framework for distributed reinforcement learning. Lecture Notes in Computer Science, 1042:97–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haynes:1996:EBS

[445] T. Haynes and S. Sen. Evolving behavioral strategies in predators and prey. Lecture Notes in Computer Science, 1042:113–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joshi:1996:LL

[446] A. Joshi. To learn or not to learn. Lecture Notes in Computer Science, 1042: 127–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenzmann:1996:UAI

[447] B. Lenzmann and I. Wachsmuth. A user-adaptive interface agency for interaction with a virtual environment. Lecture Notes in Computer Science, 1042:140–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mataric:1996:LMR

[448] M. J. Mataric. Learning in multi-robot systems. Lecture Notes in Computer Science, 1042:152–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mor:1996:LYO

[449] Y. Mor, C. V. Goldman, and J. S. Rosenschein. Learn your opponent’s strategy (in polynomial time)! Lecture Notes in Computer Science, 1042: 164–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohko:1996:LRC

[450] T. Ohko, K. Hiraki, and Y. Anzai. Learning to reduce communication cost on task negotiation among multiple autonomous mobile robots. Lecture Notes in Computer Science, 1042:177– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sandholm:1996:MQL

[451] T. W. Sandholm and R. H. Crites. On multiagent Q-learning in a semicompetitive domain. Lecture Notes in Computer Science, 1042:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sen:1996:URA

[452] S. Sen and M. Sekaran. Using reciprocity to adapt to others. Lecture Notes in Computer Science, 1042:206– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sen:1996:MCL

[453] S. Sen and M. Sekaran. Multiagent coordination with learning classifier systems. Lecture Notes in Computer Science, 1042:218–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abramsky:1996:SSP

[454] S. Abramsky, S. Gay, and R. Nagarajan. Specification structures and propositions-as-types for concurrency. Lecture Notes in Computer Science, 1043:5–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emerson:1996:ATR

[455] E. A. Emerson. Automated temporal reasoning about reactive systems. Lecture Notes in Computer Science, 1043:41–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirshfeld:1996:DRA

[456] Y. Hirshfeld and F. Moller. Decidability results in automata and process theory. Lecture Notes in Computer Science, 1043:102–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stirling:1996:MTL

[457] C. Stirling. Modal and temporal logics for processes. Lecture Notes in Computer Science, 1043:149–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vardi:1996:ATA

[458] M. Y. Vardi. An automata-theoretic approach to linear temporal logic. Lecture Notes in Computer Science, 1043: 238–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crochat:1996:PSM

[459] O. Crochat, J.-Y. Le Boudec, and T. Przygienda. A path selection method in ATM using precomputation. Lecture Notes in Computer Science, 1044:1–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Struyve:1996:DED

[460] K. Struyve, P. Demeester, L. Nederlof, and L. Van Hauwermeiren. Design and evaluation of distributed link and path restoration algorithms for ATM meshed networks. Lecture Notes in Computer Science, 1044:13–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gauthier:1996:SEC

[461] E. Gauthier and J.-Y. Le Boudec. Scalability enhancements for connectionoriented networks. Lecture Notes in Computer Science, 1044:27–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golestani:1996:FQA

[462] S. J. Golestani. Fair queueing algorithms for packet scheduling in BISDN. Lecture Notes in Computer Science, 1044:39–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamdi:1996:BBB

[463] M. Hamdi and J. W. Roberts. Burstiness bounds based multiplexing schemes for VBR video connections in the B-ISDN. Lecture Notes in Computer Science, 1044:53–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ryu:1996:ISS

[464] B. K. Ryu. Implications of selfsimilarity for providing end-to-end QOS guarantees in high-speed networks: A framework of application level traffic modeling. Lecture Notes in Computer Science, 1044:65–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1996:MCC

[465] H. Mueller. Multimedia call control: A centralized approach. Lecture Notes in Computer Science, 1044:81–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmann:1996:GCL

[466] M. Hofmann. A generic concept for large-scale multicast. Lecture Notes in Computer Science, 1044:95–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stock:1996:PFE

[467] T. Stock and X. Garcia. On the potentials of forward error correction mechanisms applied to real-time services carried over B-ISDN. Lecture Notes in Computer Science, 1044:107–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robert:1996:CSS

[468] S. Robert and J.-Y. Le Boudec. Can self-similar traffic be modeled by Markovian processes? Lecture Notes in Computer Science, 1044:119–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blefari-Melazzi:1996:MAM

[469] N. Blefari-Melazzi. Modeling and analysis of MPEG video sources for performance evaluation of broadband integrated networks. Lecture Notes in Computer Science, 1044:131–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martini:1996:SDP

[470] P. Martini and J. Ottensmeyer. On the scalability of the demand-priority LAN — A performance comparison to FDDI for multimedia scenarios. Lecture Notes in Computer Science, 1044: 149–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsang:1996:FSA

[471] D. H. K. Tsang, W. K. F. Wong, S. M. Jiang, and E. Y. S. Liu. A fast switch algorithm for ABR traffic to achieve max-min fairness. Lecture Notes in Computer Science, 1044:161–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitra:1996:VPD

[472] D. Mitra and I. Ziedins. Virtual partitioning by dynamic priorities: Fair and efficient resource-sharing by several services. Lecture Notes in Computer Science, 1044:173–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anastasi:1996:PSL

[473] G. Anastasi, M. La Porta, and L. Lenzini. A performance study of the local fairness algorithm for MetaRing MAC protocol. Lecture Notes in Computer Science, 1044:187–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felix:1996:EMT

[474] D. Felix, E. Paolillo, F. Mercalli, and Y.-S. Wu. Experiences with multimedia teleshopping applications over a broadband network — the project ESSAI. Lecture Notes in Computer Science, 1044:209–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Znaty:1996:MMT

[475] S. Znaty, T. Walter, M. Brunner, and J.-P. Hubaux. Multimedia multipoint teleteaching over the European ATM pilot. Lecture Notes in Computer Science, 1044:225–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winkler:1996:USP

[476] C. M. Winkler. A universal scaling principle for ATM based connectionless servers. Lecture Notes in Computer Science, 1044:239–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeVriendt:1996:UMS

[477] J. De Vriendt, L. Vercauteren, K. Georgokitsos, and A. Saidi. The UMTS mobility server: a solution to support third generation mobility in ATM. Lecture Notes in Computer Science, 1044: 251–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hadjitheodosiou:1996:VSN

[478] M. H. Hadjitheodosiou. VSAT satellite networks providing ATM service. Lecture Notes in Computer Science, 1044: 263–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jang:1996:PCC

[479] Y. M. Jang, A. Ganz, and J. F. Hayes. Predictive congestion control for broadband satellite systems. Lecture Notes in Computer Science, 1044:277–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vvedenskaya:1996:SFP

[480] N. D. Vvedenskaya and J. P. M. G. Linnartz. Single-frequency packet network using stack algorithm and multiple base stations. Lecture Notes in Computer Science, 1044:289–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solana:1996:BAR

[481] J. I. Solana, J. Marino, and R. Caravantes. Broadband access in RECIBA B-ISDN experimental platform. Lecture Notes in Computer Science, 1044: 299–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zirngibl:1996:LWD

[482] M. Zirngibl, C. H. Joyner, C. R. Doerr, and L. W. Stulz. LARNet, a wavelength division multiplexed network for broadband local access. Lecture Notes in Computer Science, 1044: 311–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aryal:1996:DLA

[483] S. Aryal and J. S. Meditch. Design of a large ATM switch with trunk grouping. Lecture Notes in Computer Science, 1044:321–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogt:1996:EEP

[484] R. Vogt and U. Killat. End-to-end performance evaluation of datagram acceptance control in DQDB-ATM-DQDB CL network. Lecture Notes in Computer Science, 1044:335–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolarov:1996:CER

[485] A. Kolarov and G. Ramamurthy. Comparison of explicit rate and explicit forward congestion indication flow control schemes for ABR service. Lecture Notes in Computer Science, 1044:347– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ott:1996:HIS

[486] M. Ott, D. Reininger, G. Michelitsch, and V. Bansal. Heidi-II: A software architecture for ATM network based distributed multimedia systems. Lecture Notes in Computer Science, 1045:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1996:MAS

[487] J. D. Smith and K. Takahashi. Multimedia architecture to support requirements analysis. Lecture Notes in Computer Science, 1045:15–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofrichter:1996:MSP

[488] K. Hofrichter. MHEG 5 — standardized presentation objects for the set top unit environment. Lecture Notes in Computer Science, 1045:33–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heumann:1996:GSO

[489] B. Heumann, T. Leidig, and P. Roesch. GLASS-Studio: An open authoring environment for distributed multimedia applications. Lecture Notes in Computer Science, 1045:45–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeVries:1996:TIF

[490] A. P. De Vries. Television information filtering through speech recognition. Lecture Notes in Computer Science, 1045:59–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vazirgiannis:1996:EAR

[491] M. Vazirgiannis and T. Sellis. Event and action representation and composition for multimedia application scenario modelling. Lecture Notes in Computer Science, 1045:71–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolf:1996:IVR

[492] K. H. Wolf, K. Froitzheim, and M. Weber. Interactive video and remote control via the World Wide Web. Lecture Notes in Computer Science, 1045:91– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reimann:1996:RD

[493] K. Reimann and D. Rueffler. Radio on demand. Lecture Notes in Computer Science, 1045:105–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georgiannakis:1996:DIT

[494] G. Y. Georgiannakis, P. E. Trahanias, C. Nikolaou, and S. C. Orphanoudakis. Design of an immersive teleconferencing application. Lecture Notes in Computer Science, 1045:117–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hinsch:1996:SCU

[495] E. Hinsch, A. Jaegemann, L. Wang, and I. C. Roper. The secure conferencing user agent: A tool to provide secure conferencing with MBone multimedia conferencing applications. Lecture Notes in Computer Science, 1045: 131–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulzrinne:1996:PMM

[496] H. Schulzrinne. Personal mobility for multimedia services in the Internet. Lecture Notes in Computer Science, 1045:143–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oppliger:1996:SAT

[497] R. Oppliger, A. Gupta, M. Moran, and R. Bettati. A security architecture for tenet scheme 2. Lecture Notes in Computer Science, 1045:163–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mathy:1996:FAM

[498] L. Mathy. Features of the ACCOPI multimedia transport service. Lecture Notes in Computer Science, 1045:175– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hess:1996:VCM

[499] R. Hess, T. Hutschenreuther, and A. Schill. Video communication and media scaling system “xnetvideo”: Design and implementation. Lecture Notes in Computer Science, 1045:195– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Werner:1996:SMR

[500] J. Werner and L. C. Wolf. Scheduling mechanisms reducing contention situations in multimedia systems. Lecture Notes in Computer Science, 1045:211– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yeadon:1996:QFA

[501] N. Yeadon, A. Mauthe, D. Hutchison, and F. Garcia. QoS filters: Addressing the heterogeneity gap. Lecture Notes in Computer Science, 1045:227–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dengler:1996:DAC

[502] J. Dengler, C. Bernhardt, and E. Biersack. Deterministic admission control strategies in video servers with variable bit rate streams. Lecture Notes in Computer Science, 1045:245–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rautenberg:1996:CSI

[503] M. Rautenberg and H. Rzehak. A control system for an interactive video on demand server handling variable data rates. Lecture Notes in Computer Science, 1045:265–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geyer:1996:SSS

[504] W. Geyer, C. Bernhardt, and E. Biersack. A synchronization scheme for stored multimedia streams. Lecture Notes in Computer Science, 1045:277– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helbig:1996:ADS

[505] T. Helbig and K. Rothermel. An architecture for a distributed stream synchronization service. Lecture Notes in Computer Science, 1045:297–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seidel:1996:CVC

[506] C. Seidel. COMENIUS — the virtual classroom. Lecture Notes in Computer Science, 1045:315–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brassard:1996:NTQ

[507] G. Brassard. New trends in quantum computing. Lecture Notes in Computer Science, 1046:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buhrman:1996:CRB

[508] H. Buhrman and L. Longpre. Compressibility and resource bounded measure. Lecture Notes in Computer Science, 1046:13–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kummer:1996:CRS

[509] M. Kummer. On the complexity of random strings. Lecture Notes in Computer Science, 1046:25–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harju:1996:RGP

[510] T. Harju, J. Karhumaeki, and D. Krob. Remarks on generalized post correspondence problem. Lecture Notes in Computer Science, 1046:39–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beal:1996:CLS

[511] M.-P. Beal, O. Carton, and C. Reutenauer. Cyclic languages and strongly cyclic languages. Lecture Notes in Computer Science, 1046:49–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambos-Spies:1996:RBB

[512] K. Ambos-Spies, E. Mayordomo, Y. Wang, and X. Zheng. Resourcebounded balanced genericity, stochasticity and weak randomness. Lecture Notes in Computer Science, 1046:63– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buhrman:1996:CGC

[513] H. Buhrman and T. Thierauf. The complexity of generating and checking proofs of membership. Lecture Notes in Computer Science, 1046:75–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lutz:1996:OML

[514] J. H. Lutz. Observations on measures and lowness for cap deltaP2. Lecture Notes in Computer Science, 1046:87– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1996:SBB

[515] V. Arvind and N. V. Vinodchandran. Solvable black-box group problems are low for PP. Lecture Notes in Computer Science, 1046:99–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beaudry:1996:LRF

[516] M. Beaudry. Languages recognized by finite aperiodic groupoids. Lecture Notes in Computer Science, 1046:113– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bassino:1996:SHR

[517] F. Bassino. Star-height of an IN-Rational series. Lecture Notes in Computer Science, 1046:125–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Culik:1996:ASW

[518] K. Culik and J. Kari. An aperiodic set of Wang cubes. Lecture Notes in Computer Science, 1046:137–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melancon:1996:LFI

[519] G. Melancon. Lyndon factorization of infinite words. Lecture Notes in Computer Science, 1046:147–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heun:1996:EGB

[520] V. Heun and E. W. Mayr. Embedding graphs with bounded tree width into optimal hypercubes. Lecture Notes in Computer Science, 1046:157–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morvan:1996:PCG

[521] M. Morvan and L. Viennot. Parallel comparability graph recognition and modular decomposition. Lecture Notes in Computer Science, 1046:169– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berenbrink:1996:FTS

[522] P. Berenbrink, F. Meyer Auf Der Heide, and V. Stemann. Fault-tolerant shared memory simulations. Lecture Notes in Computer Science, 1046:181–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Indyk:1996:WLP

[523] P. Indyk. On word-level parallelism in fault-tolerant computing. Lecture Notes in Computer Science, 1046:193– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barzdins:1996:LC

[524] J. Barzdins, R. Freivalds, and C. H. Smith. Learning with confidence. Lecture Notes in Computer Science, 1046: 207–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tateishi:1996:EBC

[525] E. Tateishi, O. Maruyama, and S. Miyano. Extracting best consensus motifs from positive and negative examples. Lecture Notes in Computer Science, 1046:219–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denis:1996:PLS

[526] F. Denis, C. D’Halluin, and R. Gilleron. PAC learning with simple examples. Lecture Notes in Computer Science, 1046:231–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambainis:1996:GII

[527] A. Ambainis, R. Freivalds, and C. H. Smith. General inductive inference types based on linearly-ordered sets. Lecture Notes in Computer Science, 1046:243–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berard:1996:PNO

[528] B. Berard, P. Gastin, and A. Petit. On the power of non-observable actions in timed automata. Lecture Notes in Computer Science, 1046:257–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertol:1996:TRC

[529] M. Bertol and V. Diekert. Trace rewriting: Computing normal forms in time O(n log n). Lecture Notes in Computer Science, 1046:269–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duerr:1996:DPW

[530] C. Duerr, H. Le Thanh, and M. Santha. A decision procedure for wellformed linear quantum cellular automata. Lecture Notes in Computer Science, 1046:281–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacoby:1996:CWC

[531] A. Jacoby and C. Schindelhauer. On the complexity of worst case and expected time in a circuit. Lecture Notes in Computer Science, 1046:295– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:1996:EHS

[532] J.-Y. Cai, A. V. Naik, and D. Sivakumar. On the existence of hard sparse sets under weak reductions. Lecture Notes in Computer Science, 1046:307– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreev:1996:OBA

[533] A. E. Andreev, A. E. F. Clementi, and J. D. P. Rolim. Optimal bounds on the approximation of Boolean functions with consequences on the concept of hardness. Lecture Notes in Computer Science, 1046:319–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:1996:FSA

[534] J.-Y. Cai and A. L. Selman. Fine separation of average time complexity classes. Lecture Notes in Computer Science, 1046:331–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sifakis:1996:CST

[535] J. Sifakis and S. Yovine. Compositional specification of timed systems. Lecture Notes in Computer Science, 1046:347–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bleichenbacher:1996:OTB

[536] D. Bleichenbacher and U. M. Maurer. Optimal tree-based one-time digital signature schemes. Lecture Notes in Computer Science, 1046:363–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Friedman:1996:AFR

[537] J. Friedman, A. Joux, Y. Roichman, and J. Stern. The action of a few random permutations on r-tuples and an application to cryptography. Lecture Notes in Computer Science, 1046:375– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurer:1996:UGT

[538] U. M. Maurer. A unified and generalized treatment of authentication theory. Lecture Notes in Computer Science, 1046:387–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walukiewicz:1996:MSO

[539] I. Walukiewicz. Monadic second order logic on tree-like structures. Lecture Notes in Computer Science, 1046: 401–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwentick:1996:BVU

[540] T. Schwentick. On bijections vs. unary functions. Lecture Notes in Computer Science, 1046:415–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marcinkowski:1996:FMP

[541] J. Marcinkowski. The 3 Frenchmen method proves undecidability of the uniform boundedness for single recursive rule ternary DATALOG programs. Lecture Notes in Computer Science, 1046:427–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmeister:1996:CDA

[542] T. Hofmeister and H. Lefmann. A combinatorial design approach to MAX-CUT. Lecture Notes in Computer Science, 1046:441–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1996:CCS

[543] A. Gupta and N. Nishimura. Characterizing the complexity of subgraph isomorphism for graphs of bounded pathwidth. Lecture Notes in Computer Science, 1046:453–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitchell:1996:CQM

[544] S. A. Mitchell. A characterization of the quadrilateral meshes of a surface which admit a compatible hexahedral mesh of enclosed volume. Lecture Notes in Computer Science, 1046:

465–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bradfield:1996:EMM

[545] J. C. Bradfield. On the expressivity of the modal mu-calculus. Lecture Notes in Computer Science, 1046:479– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bollig:1996:ROP

[546] B. Bollig and I. Wegener. Readonce projections and formal circuit verification with binary decision diagrams. Lecture Notes in Computer Science, 1046:491–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giacobazzi:1996:OCS

[547] R. Giacobazzi. “optimal” collecting semantics for analysis in a hierarchy of logic program semantics. Lecture Notes in Computer Science, 1046:503– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmitt:1996:FFN

[548] V. Schmitt. Flip-flop nets. Lecture Notes in Computer Science, 1046:517– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kranakis:1996:LBC

[549] E. Kranakis and D. Krizanc. Lower bounds for compact routing. Lecture Notes in Computer Science, 1046:529– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frougny:1996:SFN

[550] C. Frougny. On the successor function in non-classical numeration systems. Lecture Notes in Computer Science, 1046:543–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beal:1996:MFW

[551] M.-P. Beal, F. Mignosi, and A. Restivo. Minimal forbidden words and symbolic dynamics. Lecture Notes in Computer Science, 1046:555–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietzfelbinger:1996:UHW

[552] M. Dietzfelbinger. Universal hashing and k-wise independent random variables via integer arithmetic without primes. Lecture Notes in Computer Science, 1046:569–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelsen:1996:RUT

[553] P. Kelsen. Ranking and unranking trees using regular reductions. Lecture Notes in Computer Science, 1046:581– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breslauer:1996:CLP

[554] D. Breslauer. On competitive online paging with lookahead. Lecture Notes in Computer Science, 1046:593– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lagergren:1996:HTP

[555] J. Lagergren. Hypothesis testing in perfect phylogeny for a bounded number of characters. Lecture Notes in Computer Science, 1046:605–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meinel:1996:LRC

[556] C. Meinel and S. Waack. The “log rank” conjecture for modular communication complexity. Lecture Notes in Computer Science, 1046:619–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambainis:1996:UBM

[557] A. Ambainis. Upper bounds on multiparty communication complexity of shifts. Lecture Notes in Computer Science, 1046:631–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damm:1996:SBM

[558] C. Damm, S. Jukna, and J. Sgall. Some bounds on multiparty communication complexity of pointer jumping. Lecture Notes in Computer Science, 1046:643–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bampis:1996:OSD

[559] E. Bampis, C. Delorme, and J.-C. Koenig. Optimal schedules for d-D grid graphs with communication delays. Lecture Notes in Computer Science, 1046:655–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaertner:1996:LPR

[560] B. Gaertner and E. Welzl. Linear programming — randomization and abstract frameworks. Lecture Notes in Computer Science, 1046:669–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leuschel:1996:EPD

[561] M. Leuschel. Ecological partial deduction: Preserving characteristic trees without constraints. Lecture Notes in Computer Science, 1048:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wunderwald:1996:MES

[562] J. E. Wunderwald. Memoing evaluation by source-to-source transformation. Lecture Notes in Computer Science, 1048:17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bossi:1996:TLT

[563] A. Bossi, N. Cocco, and S. Etalle. Transformation of left terminating programs: The reordering problem. Lecture Notes in Computer Science, 1048: 33–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gregory:1996:DCA

[564] S. Gregory. Derivation of concurrent algorithms in tempo. Lecture Notes in Computer Science, 1048:46–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toni:1996:ATA

[565] F. Toni and R. A. Kowalski. An argumentation-theoretic approach to logic program transformation. Lecture Notes in Computer Science, 1048:61– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aarts:1996:CHP

[566] E. Aarts. Complexity of Horn programs. Lecture Notes in Computer Science, 1048:76–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gallardo:1996:SCL

[567] M. M. Gallardo and J. M. Troya. Studying the cost of logic languages in an abstract interpretation framework for granularity analysis. Lecture Notes in Computer Science, 1048:91– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mesnard:1996:TAC

[568] F. Mesnard. Towards automatic control for CLP(chi) programs. Lecture Notes in Computer Science, 1048:106– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Semeraro:1996:IRD

[569] G. Semeraro, F. Esposito, and D. Malerba. Ideal refinement of datalog programs. Lecture Notes in Computer Science, 1048:120–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kreitz:1996:GPD

[570] C. Kreitz, J. Otten, and S. Schmitt. Guiding program development systems by a connection based proof strategy. Lecture Notes in Computer Science, 1048:137–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lau:1996:TOO

[571] K.-K. Lau and M. Ornaghi. Towards an object-oriented methodology for deductive synthesis of logic programs. Lecture Notes in Computer Science, 1048:152–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parkes:1996:LPS

[572] A. J. Parkes and G. A. Wiggins. Logic program synthesis by induction over Horn clauses. Lecture Notes in Computer Science, 1048:170–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flener:1996:LPT

[573] P. Flener and Y. Deville. Logic program transformation through generalization schemata. Lecture Notes in Computer Science, 1048:171–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vasconcelos:1996:OAL

[574] W. W. Vasconcelos and N. E. Fuchs. An opportunistic approach for logic program analysis and optimisation using enhanced schema-based transformations. Lecture Notes in Computer Science, 1048:174–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeWaal:1996:SDP

[575] D. A. De Waal and M. Thielscher. Solving deductive planning problems using program analysis and transformation. Lecture Notes in Computer Science, 1048:189–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Renault:1996:TCP

[576] S. Renault. Towards a complete proof procedure to prove properties of normal logic programs under the completion. Lecture Notes in Computer Science, 1048:204–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arts:1996:TLP

[577] T. Arts and H. Zantema. Termination of logic programs using semantic unification. Lecture Notes in Computer Science, 1048:219–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glueck:1996:MMH

[578] R. Glueck. On the mechanics of metasystem hierarchies in program transformation. Lecture Notes in Computer Science, 1048:234–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanus:1996:ETL

[579] M. Hanus. Efficient translation of lazy functional logic programs into Prolog. Lecture Notes in Computer Science, 1048:252–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huersch:1996:AEO

[580] W. L. Huersch and L. M. Seiter. Automating the evolution of objectoriented systems. Lecture Notes in Computer Science, 1049:2–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanHilst:1996:UCT

[581] M. VanHilst and D. Notkin. Using C++ templates to implement rolebased designs. Lecture Notes in Computer Science, 1049:22–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ledoux:1996:EMT

[582] T. Ledoux and P. Cointe. Explicit metaclasses as a tool for improving the design of class libraries. Lecture Notes in Computer Science, 1049:38– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1996:ST

[583] B. Meyer. Static typing. Lecture Notes in Computer Science, 1049:57– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chien:1996:ICD

[584] A. Chien, U. Reddy, J. Plevyak, and J. Dolby. ICC++ — A C++ dialect for high performance parallel computing. Lecture Notes in Computer Science, 1049:76–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1996:HCE

[585] N. Sato, S. Matsuoka, and A. Yonezawa. Hierarchical collections: an efficient scheme to build an object-oriented distributed class library for massively parallel computation. Lecture Notes in Computer Science, 1049:96–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopes:1996:APP

[586] C. Lopes. Adaptive parameter passing. Lecture Notes in Computer Science, 1049:118–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gallesio:1996:DMO

[587] E. Gallesio. Designing a meta object protocol to wrap a standard graphical toolkit. Lecture Notes in Computer Science, 1049:137–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiba:1996:ACM

[588] S. Chiba, G. Kiczales, and J. Lamping. Avoiding confusion in metacircularity: The meta-helix. Lecture Notes in Computer Science, 1049:157–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tahara:1996:ASR

[589] Y. Tahara, F. Kumeno, A. Ohsuga, and S. Honiden. An algebraic semantics of reflective objects. Lecture Notes in Computer Science, 1049:173–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanaka:1996:MMA

[590] Y. Tanaka. A meme media architecture for fine-grain component software. Lecture Notes in Computer Science, 1049:190–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakanishi:1996:MEE

[591] K. Nakanishi and T. Arano. A metric for evaluating effectiveness of objectoriented interface abstraction for promoting software reuse. Lecture Notes in Computer Science, 1049:215–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Briot:1996:ECS

[592] J.-P. Briot. An experiment in classification and specialization of synchronization schemes. Lecture Notes in Computer Science, 1049:227–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chambers:1996:SBO

[593] C. Chambers. Synergies between object-oriented programming language design and implementation research. Lecture Notes in Computer Science, 1049:250–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nolte:1996:OVI

[594] J. Nolte, H. Konaka, A. Hori, and Y. Ishikawa. Ocore and VAST — integrating persistence into a parallel object-oriented language. Lecture Notes in Computer Science, 1049:260– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maeda:1996:MPC

[595] C. Maeda. A metaobject protocol for controlling file cache management. Lecture Notes in Computer Science, 1049:275–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kowarschick:1996:SCO

[596] W. Kowarschick, G. Koestler, and W. Kiessling. Subtyping by constraints in object-oriented databases. Lecture Notes in Computer Science, 1049:287– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Argenius:1996:SCL

[597] M. Argenius and A. Voronkov. Semantics of constraint logic programs with bounded quantifiers. Lecture Notes in Computer Science, 1050:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldoni:1996:TML

[598] M. Baldoni, L. Giordano, and A. Martelli. Translating a modal language with embedded implication into Horn clause logic. Lecture Notes in Computer Science, 1050:19–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brogi:1996:PSS

[599] A. Brogi and C. Guarino. Pruning the search space of logic programs. Lecture Notes in Computer Science, 1050:35–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caferra:1996:SEL

[600] R. Caferra and N. Peltier. A significant extension of logic programming by adapting model building rules. Lecture Notes in Computer Science, 1050: 51–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cervesato:1996:ERM

[601] I. Cervesato, J. S. Hodas, and F. Pfenning. Efficient resource management for linear logic proof search. Lecture Notes in Computer Science, 1050:67– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciancarini:1996:LLB

[602] P. Ciancarini, D. Fogli, and M. Gaspari. A logic language based on GAMMA-like multiset rewriting. Lecture Notes in Computer Science, 1050: 83–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damasio:1996:DNC

[603] C. V. Damasio and L. M. Pereira. Default negated conclusions: why not? Lecture Notes in Computer Science, 1050:103–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degtyarev:1996:HEL

[604] A. Degtyarev and A. Voronkov. Handling equality in logic programming via basic folding. Lecture Notes in Computer Science, 1050:119–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eder:1996:AMR

[605] K. Eder, S. Hoelldobler, and M. Thielscher. An abstract machine for reasoning about situations, actions, and causality. Lecture Notes in Computer Science, 1050:137–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hudelmaier:1996:CCP

[606] J. Hudelmaier. On the computational complexity of propositional logic programs with nested implications. Lecture Notes in Computer Science, 1050: 153–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonker:1996:PTE

[607] C. M. Jonker. Proof-theory for extensions of logic programming. Lecture Notes in Computer Science, 1050:161– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kinoshita:1996:FSL

[608] Y. Kinoshita and A. John Power. A fibrational semantics for logic programs. Lecture Notes in Computer Science, 1050:177–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuchen:1996:HOB

[609] H. Kuchen and J. Anastasiadi. Higher order babel: Language and implementation. Lecture Notes in Computer Science, 1050:193–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lipton:1996:NFD

[610] J. Lipton, S. E. Finkelstein, and P. Freyd. A new framework for declarative programming: Categorical perspectives (invited paper). Lecture Notes in Computer Science, 1050:209– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreno-Navarro:1996:ECN

[611] J. J. Moreno-Navarro. Extending constructive negation for partial functions in lazy functional-logic languages. Lecture Notes in Computer Science, 1050: 213–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Przymusinski:1996:SLP

[612] T. C. Przymusinski. Super logic programs and negation as belief (invited paper). Lecture Notes in Computer Science, 1050:229–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Staerk:1996:TCL

[613] R. F. Staerk. Total correctness of logic programs: A formal approach. Lecture Notes in Computer Science, 1050:237– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stroetmann:1996:DSP

[614] K. Stroetmann and T. Glass. A declarative semantics for the Prolog cut operator (invited paper). Lecture Notes in Computer Science, 1050:255–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Torgersson:1996:DAF

[615] O. Torgersson. A definitional approach to functional logic programming. Lecture Notes in Computer Science, 1050:273–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vojtas:1996:SCN

[616] P. Vojtas and L. Paulik. Soundness and completeness of non-classical extended SLD-resolution. Lecture Notes in Computer Science, 1050:289–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Witteveen:1996:SPN

[617] C. Witteveen and W. Van der Hoek. Some postulates for nonmonotonic theory revision applied to logic programming. Lecture Notes in Computer Science, 1050:303–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoare:1996:HDS

[618] C. A. R. Hoare. How did software get so reliable without proof? Lecture Notes in Computer Science, 1051:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sivertsen:1996:CSF

[619] T. Sivertsen. A case study on the formal development of a reactor safety system. Lecture Notes in Computer Science, 1051:18–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peleska:1996:TAS

[620] J. Peleska. Test automation for safetycritical systems: Industrial application and future developments. Lecture Notes in Computer Science, 1051:39– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bicarregui:1996:QAA

[621] J. Bicarregui, J. Dick, and E. Woods. Quantitative analysis of an application of formal methods. Lecture Notes in Computer Science, 1051:60–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoare:1996:ABT

[622] J. Hoare, J. Dick, D. Neilson, and I. Soerensen. Applying the B technologies to CICS. Lecture Notes in Computer Science, 1051:74–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walden:1996:RAS

[623] M. Walden and K. Sere. Refining action systems within B-tool. Lecture Notes in Computer Science, 1051:85–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kasurinen:1996:IAS

[624] V. Kasurinen and K. Sere. Integrating action systems and Z in a medical system specification. Lecture Notes in Computer Science, 1051:105–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groenboom:1996:FAC

[625] R. Groenboom, E. Saaman, E. Rotterdam, and G. Renardel de Lavalette. Formalizing anaesthesia: a case study in formal specification. Lecture Notes in Computer Science, 1051:120–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martins:1996:NSE

[626] J. Martins and J.-P. Hubaux. A new system engineering methodology coupling formal specification and performance evaluation. Lecture Notes in Computer Science, 1051:140–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiVito:1996:FNN

[627] B. L. Di Vito. Formalizing new navigation requirements for NASA’s space shuttle. Lecture Notes in Computer Science, 1051:160–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Froehlich:1996:CVS

[628] B. Froehlich and P. G. Larsen. Combining VDM-SL specifications with C++ code. Lecture Notes in Computer Science, 1051:179–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clement:1996:DRE

[629] T. Clement. Data reification without explicit abstraction functions. Lecture Notes in Computer Science, 1051:195– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brookes:1996:FIS

[630] T. M. Brookes, J. S. Fitzgerald, and P. G. Larsen. Formal and informal specifications of a secure system component: Final results in a comparative study. Lecture Notes in Computer Science, 1051:214–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valmari:1996:VVS

[631] A. Valmari and M. Setaelae. Visual verification of safety and liveness. Lecture Notes in Computer Science, 1051:228–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaetz:1996:GDC

[632] B. Schaetz, H. Hussmann, and M. Broy. Graphical development of consistent system specifications. Lecture Notes in Computer Science, 1051:248–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hutter:1996:DVS

[633] D. Hutter, B. Langenstein, C. Sengler, and J. H. Siekmann. Deduction in the verification support environment (VSE). Lecture Notes in Computer Science, 1051:268–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boiten:1996:CRP

[634] E. Boiten, J. Derrick, H. Bowman, and M. Steen. Consistency and refinement for partial specification in Z. Lecture Notes in Computer Science, 1051:287–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1996:CSZ

[635] M. Weber. Combining statecharts and Z for the design of safety-critical control systems. Lecture Notes in Computer Science, 1051:307–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fidge:1996:IRT

[636] C. Fidge, M. Utting, P. Kearney, and I. Hayes. Integrating real-time scheduling theory and program refinement. Lecture Notes in Computer Science, 1051:327–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seguin:1996:ULC

[637] C. Seguin and V. Wiels. Using a logical and categorical approach for the validation of fault-tolerant systems. Lecture Notes in Computer Science, 1051:347–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBoer:1996:LNA

[638] F. S. De Boer and M. Van Hulst. Local nondeterminism in asynchronously communicating processes. Lecture Notes in Computer Science, 1051:367– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chalin:1996:ISS

[639] P. Chalin, P. Grogono, and T. Radhakrishnan. Identification of and solutions to shortcomings of LCL, a larch/ C interface specification language. Lecture Notes in Computer Science, 1051: 385–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kannikeswaran:1996:FSV

[640] B. Kannikeswaran, R. Radhakrishnan, P. Frey, and P. Alexander. Formal specification and verification of the pGVT algorithm. Lecture Notes in Computer Science, 1051:405–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pugliese:1996:AVH

[641] R. Pugliese and E. Tronci. Automatic verification of a hydroelectric power plant. Lecture Notes in Computer Science, 1051:425–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jackson:1996:EES

[642] D. M. Jackson. Experiences in embedded scheduling. Lecture Notes in Computer Science, 1051:445–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boigelot:1996:MCP

[643] B. Boigelot and P. Godefroid. Model checking in practice: An analysis of the ACCESS.bus[TM] protocol using SPIN. Lecture Notes in Computer Science, 1051:465–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kleuker:1996:IDC

[644] S. Kleuker and H. Tjabben. The incremental development of correct specifications for distributed systems. Lecture Notes in Computer Science, 1051: 479–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

George:1996:TDT

[645] C. George. A theory of distributing train rescheduling. Lecture Notes in Computer Science, 1051:499–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shi:1996:ITS

[646] L. Shi and P. Nixon. An improved translation of SA/RT specification model to high-level timed Petri nets. Lecture Notes in Computer Science, 1051:518–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peleska:1996:TTT

[647] J. Peleska and M. Siegel. From testing theory to test driver implementation. Lecture Notes in Computer Science, 1051:538–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Comuzzi:1996:PSU

[648] J. J. Comuzzi and J. M. Hart. Program slicing using weakest preconditions. Lecture Notes in Computer Science, 1051:557–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alencar:1996:FAA

[649] P. S. C. Alencar, D. D. Cowan, and C. J. P. Lucena. A formal approach to architectural design patterns. Lecture Notes in Computer Science, 1051: 576–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zwiers:1996:MCI

[650] J. Zwiers, U. Hannemann, Y. Lakhneche, and W.-P. De Roever. Modular completeness: Integrating the reuse of specified software in top-down program development. Lecture Notes in Computer Science, 1051:595–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bohn:1996:SAT

[651] J. Bohn and W. Janssen. A strategic approach to transformational design. Lecture Notes in Computer Science, 1051:609–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolyang:1996:CUF

[652] Kolyang, T. Santen, and B. Wolff. Correct and user-friendly implementations of transformation systems. Lecture Notes in Computer Science, 1051:629– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnold:1996:EUF

[653] A. Arnold, D. Begay, and J.-P. Radoux. An example of use of formal methods to debug an embedded software. Lecture Notes in Computer Science, 1051: 649–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Havelund:1996:ETP

[654] K. Havelund and N. Shankar. Experiments in theorem proving and model checking for protocol verification. Lecture Notes in Computer Science, 1051: 662–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:PLV

[655] F. Wang and C.-T. Lo. Procedure-level verification of real-time concurrent systems. Lecture Notes in Computer Science, 1051:682–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mauthe:1996:GSM

[656] A. Mauthe, G. Coulson, D. Hutchison, and S. Namuye. Group support in multimedia communications systems. Lecture Notes in Computer Science, 1052:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smirnov:1996:OOF

[657] M. I. Smirnov. Object-oriented framework for a scalable multicast call modelling. Lecture Notes in Computer Science, 1052:19–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeRezende:1996:MCS

[658] J. F. De Rezende, S. Fdida, A. Mauthe, and D. Hutchison. M-connection service: A multicast service for distributed multimedia applications. Lecture Notes in Computer Science, 1052: 40–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cossmann:1996:GDM

[659] H. Cossmann, C. Griwodz, G. Grassel, and M. Puehlhoefer. GLASS: A distributed MHEG-Based multimedia system. Lecture Notes in Computer Science, 1052:61–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klovning:1996:PEC

[660] E. Klovning and O. Bonaventure. Performance evaluation of the CSCW application JVTOS. Lecture Notes in Computer Science, 1052:79–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanRijssen:1996:MTM

[661] E. Van Rijssen, I. Widya, and E. Michiels. Multimedia teleservices modelled with the OSI application layer structure. Lecture Notes in Computer Science, 1052:99–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chassot:1996:SDM

[662] C. Chassot, M. Fournier, M. Diaz, and A. Lozes. Service definition of a multimedia partial order connection. Lecture Notes in Computer Science, 1052:118–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toutain:1996:NSM

[663] F. Toutain and L. Toutain. Network support for multimedia communications using distributed media scaling. Lecture Notes in Computer Science, 1052:138–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mathy:1996:AMT

[664] L. Mathy and O. Bonaventure. The ACCOPI multimedia transport service over ATM. Lecture Notes in Computer Science, 1052:158–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cho:1996:THS

[665] H. S. Cho, M. R. Fry, A. Seneviratne, and V. Witana. Towards a hybrid scheme for application adaptivity. Lecture Notes in Computer Science, 1052:177–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leboucher:1996:ACE

[666] L. Leboucher and J.-B. Stefani. Admission control for end-to-end distributed bindings. Lecture Notes in Computer Science, 1052:192–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reinhardt:1996:EAC

[667] W. Reinhardt and D. Trossen. An enhanced admission control scheme for deterministic and predictive services. Lecture Notes in Computer Science, 1052:209–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boucqueau:1996:ECA

[668] J.-M. Boucqueau, S. Lacroix, B. Macq, and J.-J. Quisquater. Equitable conditional access and copyright protection for image based on trusted third parties. Lecture Notes in Computer Science, 1052:229–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamdi:1996:DPP

[669] M. Hamdi, H. Afifi, P. Rolin, and L. M. Rojas. Designing point to point interactive video applications. Lecture Notes in Computer Science, 1052:244– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindsey:1996:RST

[670] D. Lindsey and P. F. Linington. RIVUS: A stream template language for capturing multimedia requirements. Lecture Notes in Computer Science, 1052:259–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graf:1996:TI

[671] P. Graf. Term indexing. Lecture Notes in Computer Science, 1053:1–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferreira:1996:SSC

[672] A. Ferreira and P. M. Pardalos. SCOOP: Solving combinatorial optimization problems in parallel. Lecture Notes in Computer Science, 1054:1–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bovet:1996:PAO

[673] D. P. Bovet, A. Clementi, P. Crescenzi, and R. Silvestri. Parallel approximation of optimization problems. Lecture Notes in Computer Science, 1054:7–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clementi:1996:RPA

[674] A. Clementi, J. Rolim, and E. Urland. Randomized parallel algorithms. Lecture Notes in Computer Science, 1054:25–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Megson:1996:ASP

[675] G. M. Megson, L. Rapanotti, and X. Chen. Automatic synthesis of parallel algorithms. Lecture Notes in Computer Science, 1054:51–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gengler:1996:IPD

[676] M. Gengler. An introduction to parallel dynamic programming. Lecture Notes in Computer Science, 1054:87–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lueling:1996:MTS

[677] R. Lueling, B. Monien, A. Reinefeld, and S. Tschoeke. Mapping treestructured combinatorial optimization problems onto parallel computers. Lecture Notes in Computer Science, 1054: 115–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBruin:1996:TAP

[678] A. De Bruin, G. A. P. Kindervater, and H. W. J. M. Trienekens. Towards an abstract parallel branch and bound machine. Lecture Notes in Computer Science, 1054:145–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Correa:1996:PBF

[679] R. Correa and A. Ferreira. Parallel best-first branch-and-bound in discrete optimization: a framework. Lecture Notes in Computer Science, 1054:171– ??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benaichouche:1996:BPB

[680] M. Benaichouche, V.-D. Cung, S. Dowaji, and B. Le Cun. Building a parallel branch and bound library. Lecture Notes in Computer Science, 1054:201– ??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pardalos:1996:PAG

[681] P. M. Pardalos, G. Xue, and P. D. Panagiotopoulos. Parallel algorithms for global optimization problems. Lecture Notes in Computer Science, 1054: 232–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laursen:1996:PHS

[682] P. S. Laursen. Parallel heuristic search — introductions and a new approach. Lecture Notes in Computer Science, 1054:248–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzmann:1996:EFD

[683] G. Holzmann. Early fault detection tools. Lecture Notes in Computer Science, 1055:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kozen:1996:KAT

[684] D. Kozen. Kleene algebra with tests and commutativity conditions. Lecture Notes in Computer Science, 1055:14– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamport:1996:MP

[685] L. Lamport. Managing proofs. Lecture Notes in Computer Science, 1055: 34–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1996:AMS

[686] R. Alur, G. Holzmann, and D. Peled. An analyzer for message sequence charts. Lecture Notes in Computer Science, 1055:35–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berghammer:1996:RAA

[687] R. Berghammer, B. Von Karger, and C. Ulke. Relation-algebraic analysis of Petri nets with RELVIEW. Lecture Notes in Computer Science, 1055:49– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damon:1996:ESM

[688] C. Damon and D. Jackson. Efficient search as a means of executing specifications. Lecture Notes in Computer Science, 1055:70–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esparza:1996:IMU

[689] J. Esparza, S. Roemaer, and W. Vogler. An improvement of McMillan’s unfolding algorithm. Lecture Notes in Computer Science, 1055:87–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhat:1996:ELM

[690] G. Bhat and R. Cleaveland. Efficient local model-checking for fragments of the modal — calculus. Lecture Notes in Computer Science, 1055:107–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tretmans:1996:TGI

[691] J. Tretmans. Test generation with inputs, outputs, and quiescence. Lecture Notes in Computer Science, 1055:127– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lowe:1996:BFN

[692] G. Lowe. Breaking and fixing the Needham-Schroeder public-key protocol using FDR. Lecture Notes in Computer Science, 1055:147–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Focardi:1996:ACV

[693] R. Focardi and R. Gorrieri. Automatic compositional verification of some security properties. Lecture Notes in Computer Science, 1055:167–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michael:1996:PAP

[694] F. Michael, P. Azema, and F. Vernadat. Permutable agents in process algebras. Lecture Notes in Computer Science, 1055:187–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buhrke:1996:SCI

[695] N. Buhrke, H. Lescow, and J. Voege. Strategy construction in infinite games with Streett and Rabin chain winning conditions. Lecture Notes in Computer Science, 1055:207–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kowalewski:1996:TCE

[696] S. Kowalewski and J. Preussig. Timed condition/event systems: A framework for modular discrete models of chemical plants and verification of their realtime discrete control. Lecture Notes in Computer Science, 1055:225–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chou:1996:FVP

[697] C.-T. Chou and D. Peled. Formal verification of a partial-order reduction technique for model checking. Lecture Notes in Computer Science, 1055:241– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Margaria:1996:FAV

[698] T. Margaria. Fully automatic verification and error detection for parameterized iterative sequential circuits. Lecture Notes in Computer Science, 1055:258–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cleaveland:1996:PMV

[699] B. Cleaveland, G. Luettgen, V. Natarajan, and S. Sims. Priorities for modeling and verifying distributed systems. Lecture Notes in Computer Science, 1055:278–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stirling:1996:GMM

[700] C. Stirling. Games and modal mucalculus. Lecture Notes in Computer Science, 1055:298–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayari:1996:GSS

[701] A. Ayari and D. Basin. Generic system support for deductive program development. Lecture Notes in Computer Science, 1055:313–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tripakis:1996:EPS

[702] S. Tripakis and C. Courcoubetis. Extending Promela and Spin for real time. Lecture Notes in Computer Science, 1055:329–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Najm:1996:RER

[703] E. Najm and F. Olsen. Reactive EF-SMs — reactive Promela/RSPIN. Lecture Notes in Computer Science, 1055: 349–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moser:1996:PDA

[704] L. Moser and P. Melliar-Smith. Probabilistic duration automata for analyzing real-time systems. Lecture Notes in Computer Science, 1055:369–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cleaveland:1996:CFS

[705] R. Cleaveland, P. Lewis, S. Smolka, and O. Sokolsky. The concurrency factory software development environment. Lecture Notes in Computer Science, 1055:391–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouali:1996:FS

[706] A. Bouali, A. Ressouche, V. Roy, and R. De Simone. The Fc2TOOLS set. Lecture Notes in Computer Science, 1055:396–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grahlmann:1996:PMT

[707] B. Grahlmann and E. Best. PEP — more than a Petri net tool. Lecture Notes in Computer Science, 1055:397– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuellar:1996:RPA

[708] J. Cuellar, D. Barnard, and M. Huber. Rapid prototyping for an assertional specification language. Lecture Notes in Computer Science, 1055:402–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heyl:1996:CTS

[709] C. Heyl, A. Mester, and H. Krumm. cTc — A tool supporting the construction of cTLA-specifications. Lecture Notes in Computer Science, 1055:407– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saidi:1996:TPI

[710] H. Saidi. A tool for proving invariance properties of concurrent systems automatically. Lecture Notes in Computer Science, 1055:412–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauzy:1996:UCL

[711] A. Rauzy. Using the constraint language toupie for “software cost reduction” specification analysis. Lecture Notes in Computer Science, 1055:417– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steffen:1996:COS

[712] B. Steffen, T. Margaria, A. Classen, and V. Braun. A constraint-oriented service creation environment. Lecture Notes in Computer Science, 1055:418– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klein:1996:DOM

[713] M. Klein, J. Knoop, D. Koschuetzki, and B. Steffen. DFA and OPT-METAFrame: A tool kit for program analysis and optimization. Lecture Notes in Computer Science, 1055:422– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermanns:1996:CAT

[714] H. Hermanns, V. Mertsiotakis, and M. Rettelbach. A construction and analysis tool based on the stochastic process algebra TIPP. Lecture Notes in Computer Science, 1055:427–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bengtsson:1996:U

[715] J. Bengtsson, K. G. Larsen, F. Larsson, and P. Pettersson. UPPAAL in 1995. Lecture Notes in Computer Science, 1055:431–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srikant:1996:MSP

[716] R. Srikant and R. Agrawal. Mining sequential patterns: Generalizations and performance improvements. Lecture Notes in Computer Science, 1057:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehta:1996:SFS

[717] M. Mehta, R. Agrawal, and J. Rissanen. SLIQ: A fast scalable classifier for data mining. Lecture Notes in Computer Science, 1057:18–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gottlob:1996:PSA

[718] G. Gottlob, G. Moerkotte, and V. S. Subrahmanian. The PARK semantics for active rules. Lecture Notes in Computer Science, 1057:35–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meo:1996:CEC

[719] R. Meo, G. Psaila, and S. Ceri. Composite events in chimera. Lecture Notes in Computer Science, 1057:56– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atzeni:1996:MMM

[720] P. Atzeni and R. Torlone. Management of multiple models in an extensible database design tool. Lecture Notes in Computer Science, 1057:79–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelter:1996:ANS

[721] U. Kelter and D. Daeberitz. An assessment of non-standard DBMSs for CASE environments. Lecture Notes in Computer Science, 1057:96–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qian:1996:CST

[722] X. Qian. Correct schema transformations. Lecture Notes in Computer Science, 1057:114–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Friesen:1996:VAD

[723] O. Friesen, A. Lefebvre, and L. Vieille. VALIDITY: Applications of a DOOD system. Lecture Notes in Computer Science, 1057:131–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonjour:1996:NOR

[724] D. Bonjour. The need for an object relational model and its use. Lecture Notes in Computer Science, 1057:135– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1996:DIU

[725] A. Gupta, H. V. Jagadish, and I. S. Mumick. Data integration using selfmaintainable views. Lecture Notes in Computer Science, 1057:140–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boncz:1996:MGE

[726] P. A. Boncz, W. Quak, and M. L. Kersten. Monet and its geographic extensions: A novel approach to high performance GIS processing. Lecture Notes in Computer Science, 1057:147–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaudhuri:1996:OQA

[727] S. Chaudhuri and K. Shim. Optimizing queries with aggregate views. Lecture Notes in Computer Science, 1057: 167–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steenhagen:1996:TOQ

[728] H. J. Steenhagen, R. A. De By, and H. M. Blanken. Translating OSQL queries into efficient set expressions. Lecture Notes in Computer Science, 1057:183–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pavilion:1996:KDE

[729] G. Pavilion. Knowledge discovery from epidemiological databases. Lecture Notes in Computer Science, 1057: 201–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabinovich:1996:SUP

[730] M. Rabinovich, N. Gehani, and A. Kononov. Scalable update propagation in epidemic replicated databases.

Lecture Notes in Computer Science, 1057:207–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adelberg:1996:DSE

[731] B. Adelberg, B. Kao, and H. Garcia-Molina. Database support for efficiently maintaining derived data. Lecture Notes in Computer Science, 1057: 223–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciaccia:1996:OMB

[732] P. Ciaccia. Optimal multi-block read schedules for partitioned signature files. Lecture Notes in Computer Science, 1057:241–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoshikawa:1996:ASD

[733] M. Yoshikawa, O. Ichikawa, and S. Uemura. Amalgamating SGML documents and databases. Lecture Notes in Computer Science, 1057:259–274, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1996:IND

[734] H. Williams and J. Zobel. Indexing nucleotide databases for fast query evaluation. Lecture Notes in Computer Science, 1057:275–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:VMS

[735] I.-M. A. Chen, V. M. Markowitz, S. Letovsky, and P. Li. Version management for scientific databases. Lecture Notes in Computer Science, 1057:

289–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toman:1996:FOQ

[736] D. Toman and D. Niwinski. First-order queries over temporal databases inexpressible in temporal logic. Lecture Notes in Computer Science, 1057:307– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boehlen:1996:QTD

[737] M. H. Boehlen, J. Chomicki, R. T. Snodgrass, and D. Toman. Querying TSQL2 databases with temporal logic. Lecture Notes in Computer Science, 1057:325–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertino:1996:FTO

[738] E. Bertino, E. Ferrari, and G. Guerrini. A formal temporal object-oriented data model. Lecture Notes in Computer Science, 1057:342–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delcambre:1996:DDR

[739] L. Delcambre, C. Hamon, M. Biezunski, and R. Reddy. Dynamic development and refinement of hypermedia documents. Lecture Notes in Computer Science, 1057:359–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finkelstein:1996:MMW

[740] S. Finkelstein, E. Lemoine, and R. Lenaers. MATISSE: A multimedia Web DBMS. Lecture Notes in Computer Science, 1057:363–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baker:1996:DBW

[741] B. Baker. Doing business with the Web: The Informix/Illustra approach. Lecture Notes in Computer Science, 1057: 364–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1996:HPS

[742] C. Liu and H. Chen. A hash partition strategy for distributed query processing. Lecture Notes in Computer Science, 1057:373–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Panagos:1996:FGL

[743] E. Panagos, A. Biliris, H. V. Jagadish, and R. Rastogi. Fine-granularity locking and client-based logging for distributed architectures. Lecture Notes in Computer Science, 1057:388–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gawecki:1996:EPI

[744] A. Gawecki and F. Matthes. Exploiting persistent intermediate code representations in open database environments. Lecture Notes in Computer Science, 1057:403–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamath:1996:PHA

[745] M. Kamath, G. Alonso, R. Guenthoer, and C. Mohan. Providing high availability in very large workflow management systems. Lecture Notes in Computer Science, 1057:427–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casati:1996:SWI

[746] F. Casati, S. Ceri, B. Pernici, and G. Pozzi. Semantic WorkFlow interoperability. Lecture Notes in Computer Science, 1057:443–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonner:1996:LDB

[747] A. Bonner, A. Shrufi, and S. Rozen. LabFlow-1: A database benchmark for high-throughput workflow management. Lecture Notes in Computer Science, 1057:463–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouganim:1996:APQ

[748] L. Bouganim, B. Dageville, and P. Valduriez. Adaptive parallel query execution in DBS3. Lecture Notes in Computer Science, 1057:481–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fessy:1996:OQS

[749] J. Fessy, Y. Lepetit, and P. Pucheral. Object query services for telecommunication networks. Lecture Notes in Computer Science, 1057:490–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plexousakis:1996:AIC

[750] D. Plexousakis and J. Mylopoulos. Accommodating integrity constraints during database design. Lecture Notes in Computer Science, 1057:497–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levy:1996:RAC

[751] A. Y. Levy and I. S. Mumick. Reasoning with aggregation constraints. Lecture Notes in Computer Science, 1057: 514–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Decker:1996:HTS

[752] H. Decker, E. Teniente, and T. Urpi. How to tackle schema validation by view updating. Lecture Notes in Computer Science, 1057:535–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muth:1996:LDD

[753] P. Muth, A. Kraiss, and G. Weikum. LoT: Dynamic declustering of TSB-Tree nodes for parallel access to temporal data. Lecture Notes in Computer Science, 1057:553–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karlsson:1996:LLS

[754] J. S. Karlsson, W. Litwin, and T. Risch. LH0* LH: A scalable high performance data structure for switched multicomputers. Lecture Notes in Computer Science, 1057:573– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koudas:1996:DSD

[755] N. Koudas, C. Faloutsos, and I. Kamel. Declustering spatial databases on a multi-computer architecture. Lecture Notes in Computer Science, 1057:592– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emrich:1996:OFB

[756] M. Emrich. Object framework for business applications. Lecture Notes in Computer Science, 1057:617–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:CSC

[757] Q. Chen and U. Dayal. Commit scope control in nested transactions. Lecture Notes in Computer Science, 1057:621– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simmen:1996:FTO

[758] D. Simmen, E. Shekita, and T. Malkemus. Fundamental techniques for order optimization. Lecture Notes in Computer Science, 1057:625–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiederhold:1996:DAT

[759] G. Wiederhold. Dealing with asynchrony in technology transfer. Lecture Notes in Computer Science, 1057:631– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1996:SPP

[760] C. B. Jones. Some practical problems and their influence on semantics. Lecture Notes in Computer Science, 1058:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1996:CHP

[761] S. L. P. Jones. Compiling Haskell by program transformation: A report from the trenches. Lecture Notes in Computer Science, 1058:18–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alpuente:1996:NDP

[762] M. Alpuente, M. Falaschi, and G. Vidal. Narrowing-driven partial evaluation of functional logic programs. Lecture Notes in Computer Science, 1058: 45–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbuti:1996:MVL

[763] R. Barbuti and P. Mancarella. A multiple-valued logical semantics for Prolog. Lecture Notes in Computer Science, 1058:62–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bechet:1996:RVE

[764] D. Bechet. Removing value encoding using alternative values in partial evaluation of strongly-typed languages. Lecture Notes in Computer Science, 1058:77–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergstra:1996:CTT

[765] J. A. Bergstra, T. B. Dinesh, J. Field, and J. Heering. A complete transformational toolkit for compilers. Lecture Notes in Computer Science, 1058:92– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bueno:1996:GAS

[766] F. Bueno, D. Cabeza, M. Hermenegildo, and G. Puebla. Global analysis of standard Prolog programs. Lecture Notes in Computer Science, 1058:108–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fradet:1996:SDP

[767] P. Fradet, R. Gaugne, and D. Le Metayer. Static detection of pointer errors: An axiomatisation and a checking algorithm. Lecture Notes in Computer Science, 1058:125–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giacobazzi:1996:COD

[768] R. Giacobazzi and F. Ranzato. Compositional optimization of disjunctive abstract interpretations. Lecture Notes in Computer Science, 1058:141–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez-Moreno:1996:RLD

[769] J. C. Gonzalez-Moreno, M. T. Hortala-Gonzalez, F. J. Lopez-Fraguas, and M. Rodriguez-Artalejo. A rewriting logic for declarative programming. Lecture Notes in Computer Science, 1058: 156–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goubault:1996:DTC

[770] E. Goubault. Durations for trulyconcurrent transitions. Lecture Notes in Computer Science, 1058:173–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jensen:1996:FAG

[771] T. P. Jensen and I. Mackie. Flow analysis in the geometry of interaction. Lecture Notes in Computer Science, 1058:188–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Launchbury:1996:PUU

[772] J. Launchbury and R. Paterson. Parametricity and unboxing with unpointed types. Lecture Notes in Computer Science, 1058:204–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liang:1996:MDS

[773] S. Liang and P. Hudak. Modular denotational semantics for compiler construction. Lecture Notes in Computer Science, 1058:219–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maraninchi:1996:CSN

[774] F. Maraninchi and N. Halbwachs. Compositional semantics of nondeterministic synchronous languages. Lecture Notes in Computer Science, 1058:235–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melzer:1996:CSP

[775] S. Melzer and J. Esparza. Checking system properties via integer programming. Lecture Notes in Computer Science, 1058:250–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montanari:1996:EML

[776] U. Montanari, M. Pistore, and D. Yankelevich. Efficient minimization up to location equivalence. Lecture Notes in Computer Science, 1058: 265–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palsberg:1996:NAC

[777] J. Palsberg, B. Patt-Shamir, and K. Lieberherr. A new approach to compiling adaptive programs. Lecture Notes in Computer Science, 1058:280– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rohwedder:1996:MTC

[778] E. Rohwedder and F. Pfenning. Mode and termination checking for higherorder logic programs. Lecture Notes in Computer Science, 1058:296–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rose:1996:LTH

[779] E. Rose. Linear time hierarchies for a functional language machine model. Lecture Notes in Computer Science, 1058:311–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seidl:1996:ICS

[780] H. Seidl. Integer constraints to stop deforestation. Lecture Notes in Computer Science, 1058:326–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1996:TMS

[781] G. Smith and D. Volpano. Towards an ML-Style polymorphic type system for C. Lecture Notes in Computer Science, 1058:341–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoelen:1996:ACR

[782] K. Stoelen. Assumption/commitment rules for dataflow networks — with an emphasis on completeness. Lecture Notes in Computer Science, 1058:356– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thune:1996:TEL

[783] M. Thune and K. Aahlander. Towards an expressive language for PDE solvers. Lecture Notes in Computer Science, 1058:373–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanBakel:1996:RSA

[784] S. Van Bakel, F. Barbanera, and M. Fernandez. Rewrite systems with abstraction and ??-rule: Types, approximants and normalization. Lecture Notes in Computer Science, 1058:387– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abramsky:1996:SI

[785] S. Abramsky. Semantics of interaction. Lecture Notes in Computer Science, 1059:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bottreau:1996:KPL

[786] A. Bottreau and Y. Metivier. The Kronecker product and local computations in graphs. Lecture Notes in Computer Science, 1059:2–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madonia:1996:IPM

[787] M. Madonia and S. Varricchio. Iterative pairs and multitape automata. Lecture Notes in Computer Science, 1059:17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudet:1996:DEP

[788] A. Boudet and H. Comon. Diophantine equations, Presburger arithmetic and finite automata. Lecture Notes in Computer Science, 1059:30–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuehnemann:1996:PLO

[789] A. Kuehnemann. A pumping lemma for output languages of macro tree transducers. Lecture Notes in Computer Science, 1059:44–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merlini:1996:ADU

[790] D. Merlini, R. Sprugnoli, and M. C. Verri. The area determined by underdiagonal lattice paths. Lecture Notes in Computer Science, 1059:59–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takayama:1996:ECP

[791] Y. Takayama. Extraction of concurrent processes from higher dimensional automata. Lecture Notes in Computer Science, 1059:72–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaivola:1996:AEC

[792] R. Kaivola. Axiomatising extended computation tree logic. Lecture Notes in Computer Science, 1059:87– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenisa:1996:FSH

[793] M. Lenisa. Final semantics for a higher order concurrent language. Lecture Notes in Computer Science, 1059:102– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfenning:1996:PLF

[794] F. Pfenning. The practice of logical framework. Lecture Notes in Computer Science, 1059:119–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmair:1996:SAB

[795] L. Bachmair, T. Chen, C. R. Ramakrishnan, and I. V. Ramakrishnan. Subsumption algorithms based on search trees. Lecture Notes in Computer Science, 1059:135–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1996:INT

[796] M. Fernandez and I. Mackie. Interaction nets and term rewriting systems. Lecture Notes in Computer Science, 1059:149–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lisper:1996:CUE

[797] B. Lisper. Computing in unpredictable environments: Semantics, reduction strategies, and program transformations. Lecture Notes in Computer Science, 1059:165–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glauert:1996:RND

[798] J. Glauert and Z. Khasidashvili. Relative normalization in deterministic residual structures. Lecture Notes in Computer Science, 1059:180–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arts:1996:TAI

[799] T. Arts. Termination by absence of infinite chains of dependency pairs. Lecture Notes in Computer Science, 1059: 196–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gramlich:1996:CTP

[800] B. Gramlich. Confluence without termination via parallel critical pairs. Lecture Notes in Computer Science, 1059: 211–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cichon:1996:OCP

[801] E. A. Cichon and H. Touzet. An ordinal calculus for proving termination in term rewriting. Lecture Notes in Computer Science, 1059:226–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1996:BSE

[802] M. Bidoit and A. Tarlecki. Behavioural satisfaction and equivalence in concrete model categories. Lecture Notes in Computer Science, 1059:241–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1996:OMB

[803] A. Cheng and M. Nielsen. Open maps, behavioural equivalences, and congruences. Lecture Notes in Computer Science, 1059:257–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rensink:1996:DCO

[804] A. Rensink. Denotational, causal, and operational determinism in event structures. Lecture Notes in Computer Science, 1059:272–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soerensen:1996:ELI

[805] M. H. Soerensen. Effective longest and infinite reduction paths in untyped lambda-calculi. Lecture Notes in Computer Science, 1059:287–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudol:1996:FAL

[806] G. Boudol and C. Lavatelli. Full abstraction for lambda calculus with resources and convergence testing. Lecture Notes in Computer Science, 1059: 302–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tiuryn:1996:ULC

[807] J. Tiuryn and M. Wand. Untyped lambda-calculus with inputoutput. Lecture Notes in Computer Science, 1059:317–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:PDT

[808] J. Wang and G. R. Gao. Pipeliningdovetailing: A transformation to enhance software pipelining for nested loops. Lecture Notes in Computer Science, 1060:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfahler:1996:CMS

[809] P. Pfahler and G. Piepenbrock. A comparison of modulo scheduling techniques for software pipelining. Lecture Notes in Computer Science, 1060:18– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ertl:1996:RAD

[810] M. A. Ertl and A. Krall. Removing anti dependences by repairing. Lecture Notes in Computer Science, 1060:33– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janssen:1996:CNS

[811] J. Janssen and H. Corporaal. Controlled node splitting. Lecture Notes in Computer Science, 1060:44–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davidson:1996:ALU

[812] J. W. Davidson and S. Jinturkar. Aggressive loop unrolling in a retargetable, optimizing compiler. Lecture Notes in Computer Science, 1060:59– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verbrugge:1996:GCP

[813] C. Verbrugge, P. Co, and L. Hendren. Generalized constant propagation: A study in C. Lecture Notes in Computer Science, 1060:74–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cifuentes:1996:SDG

[814] C. Cifuentes. Structuring decompiled graphs. Lecture Notes in Computer Science, 1060:91–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geser:1996:NMF

[815] A. Geser, J. Knoop, G. Luettgen, and O. Ruething. Non-monotone fixpoint iterations to resolve second order effects. Lecture Notes in Computer Science, 1060:106–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Assmann:1996:HUS

[816] U. Assmann. How to uniformly specify program analysis and transformation with graph rewrite systems. Lecture Notes in Computer Science, 1060: 121–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steensgaard:1996:PAT

[817] B. Steensgaard. Points-to analysis by type inference of programs with structures and unions. Lecture Notes in Computer Science, 1060:136–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waite:1996:CCC

[818] W. M. Waite. Compiler construction: Craftsmanship or engineering? Lecture Notes in Computer Science, 1060: 151–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nymeyer:1996:CGA

[819] A. Nymeyer, J.-P. Katoen, Y. Westra, and H. Alblas. Code generation = A0* + BURS. Lecture Notes in Computer Science, 1060:160–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pettersson:1996:CNS

[820] M. Pettersson. A compiler for natural semantics. Lecture Notes in Computer Science, 1060:177–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boucher:1996:ACN

[821] D. Boucher and M. Feeley. Abstract compilation: A new implementation paradigm for static analysis. Lecture Notes in Computer Science, 1060:192– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salomon:1996:UPE

[822] D. J. Salomon. Using partial evaluation in support of portability, reusability, and maintainability. Lecture Notes in Computer Science, 1060:208–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1996:ICS

[823] J.-D. Choi, V. Sarkar, and E. Schonberg. Incremental computation of static single assignment form. Lecture Notes in Computer Science, 1060:223– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1996:ESR

[824] Z. Li and C. C. Kirkham. Efficient storage reuse of aggregates in single assignment languages. Lecture Notes in Computer Science, 1060:238–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chow:1996:ERA

[825] F. Chow, S. Chan, S.-M. Liu, and R. Lo. Effective representation of aliases and indirect memory operations in SSA form. Lecture Notes in Computer Science, 1060:253–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kadhim:1996:MSM

[826] B. M. Kadhim and W. M. Waite. Maptool — supporting modular syntax development. Lecture Notes in Computer Science, 1060:268–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McLean:1996:FEP

[827] P. McLean and R. N. Horspool. A faster Earley parser. Lecture Notes in Computer Science, 1060:281–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Proebsting:1996:OPO

[828] T. A. Proebsting and B. R. Whaley. One-pass, optimal tree parsing — with or without trees. Lecture Notes in Computer Science, 1060:294–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vitek:1996:CDT

[829] J. Vitek and R. N. Horspool. Compact dispatch tables for dynamically typed object oriented languages. Lecture Notes in Computer Science, 1060: 309–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosch:1996:DCO

[830] J. Bosch. Delegating compiler objects — an object-oriented approach to crafting compilers. Lecture Notes in Computer Science, 1060:326–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fritzson:1996:PDS

[831] P. Fritzson, R. Wismueller, O. Hansen, and J. Sala. A parallel debugger with support for distributed arrays, multiple executables and dynamic processes. Lecture Notes in Computer Science, 1060:341–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banatre:1996:PMP

[832] J.-P. Banâtre. Parallel multiset processing: From explicit coordination to chemical reaction. Lecture Notes in Computer Science, 1061:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montanari:1996:GRC

[833] U. Montanari and F. Rossi. Graph rewriting and constraint solving for modelling distributed systems with synchronization. Lecture Notes in Computer Science, 1061:12–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wegner:1996:CCI

[834] P. Wegner. Coordination as constrained interaction. Lecture Notes in Computer Science, 1061:28–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arbab:1996:IMC

[835] F. Arbab. The IWIM model for coordination of concurrent activities. Lecture Notes in Computer Science, 1061: 34–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banville:1996:SAL

[836] M. Banville. Sonia: an adaptation of Linda for coordination of activities in organizations. Lecture Notes in Computer Science, 1061:57–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergstra:1996:TCA

[837] J. A. Bergstra and P. Klint. The Tool-Bus coordination architecture. Lecture Notes in Computer Science, 1061:75– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castellani:1996:ECM

[838] S. Castellani and P. Ciancarini. Enhancing coordination and modularity mechanisms for a language with objects-as-multisets. Lecture Notes in Computer Science, 1061:89–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaudron:1996:TCM

[839] M. Chaudron and E. De Jong. Towards a compositional method for coordinating gamma programs. Lecture Notes in Computer Science, 1061:107–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1996:ICH

[840] D. Cohen and J. Muylaert-Filho. Introducing a calculus for higher-order multiset programming. Lecture Notes in Computer Science, 1061:124–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBosschere:1996:TRC

[841] K. De Bosschere and J.-M. Jacquet. 02Log: Towards remote coordination. Lecture Notes in Computer Science, 1061:142–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNicola:1996:PAB

[842] R. De Nicola and R. Pugliese. A process algebra based on Linda. Lecture Notes in Computer Science, 1061:160– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fukuda:1996:IIO

[843] M. Fukuda, L. F. Bic, M. B. Dillencourt, and F. Merchant. Intra- and inter-object coordination with MES-SENGERS. Lecture Notes in Computer Science, 1061:179–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Florijn:1996:AHF

[844] G. Florijn, T. Besamusca, and D. Greefhorst. Ariadne and HOPLa: Flexible coordination of collaborative processes. Lecture Notes in Computer Science, 1061:197–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goeman:1996:CIF

[845] H. J. M. Goeman, J. N. Kok, K. Sere, and R. T. Udink. Coordination in the ImpUnity framework. Lecture Notes in Computer Science, 1061:215–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gladstein:1996:CCC

[846] D. S. Gladstein and M. Wand. Compiler correctness for concurrent languages. Lecture Notes in Computer Science, 1061:231–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzbacher:1996:SEC

[847] A. A. Holzbacher. A software environment for concurrent coordinated programming. Lecture Notes in Computer Science, 1061:249–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kielmann:1996:DCM

[848] T. Kielmann. Designing a coordination model for open systems. Lecture Notes in Computer Science, 1061:267– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukherji:1996:CPC

[849] M. Mukherji and D. Kafura. CCE: A process-calculus based formalism for specifying multi-object coordination. Lecture Notes in Computer Science, 1061:285–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denti:1996:EFD

[850] E. Denti, A. Omicini, A. Natali, and M. Venuti. An extensible framework for the development of coordinated applications. Lecture Notes in Computer Science, 1061:305–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prasad:1996:BT

[851] K. V. S. Prasad. Broadcasting in time. Lecture Notes in Computer Science, 1061:321–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radestock:1996:SHO

[852] M. Radestock and S. Eisenbach. Semantics of a higher-order coordination language. Lecture Notes in Computer Science, 1061:339–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rowstron:1996:SLM

[853] A. Rowstron and A. Wood. Solving the Linda multiple rd problem. Lecture Notes in Computer Science, 1061: 357–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Singh:1996:CDO

[854] N. Singh and M. A. Gisi. Coordinating distributed objects with declarative interfaces. Lecture Notes in Computer Science, 1061:368–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tolksdorf:1996:CSO

[855] R. Tolksdorf. Coordinating services in open distributed systems with LAURA. Lecture Notes in Computer Science, 1061:386–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouvry:1996:VVE

[856] P. Bouvry and F. Arbab. Visifold: A visual environment for a coordination language. Lecture Notes in Computer Science, 1061:403–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burkhart:1996:ASP

[857] H. Burkhart, R. Frank, and G. Haechler. ALWAN: A skeleton programming language. Lecture Notes in Computer Science, 1061:407–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciancarini:1996:WWU

[858] P. Ciancarini, R. Tolksdorf, and F. Vitali. Weaving the Web using coordination. Lecture Notes in Computer Science, 1061:411–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doberkat:1996:ISC

[859] E.-E. Doberkat, W. Hasselbring, and C. Pahl. Investigating strategies for cooperative planning of independent agents through prototype evaluation. Lecture Notes in Computer Science, 1061:416–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hazeyama:1996:CSI

[860] A. Hazeyama and S. Komiya. A case study of integration of a software process management system with software engineering environments for process monitoring and management. Lecture Notes in Computer Science, 1061:420– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horita:1996:NNP

[861] E. Horita and K. Mano. Nepi: A network programming language based on the Pi-Calculus. Lecture Notes in Computer Science, 1061:424–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inverardi:1996:MIC

[862] P. Inverardi and D. Compare. Modelling interoperability by CHAM. Lecture Notes in Computer Science, 1061: 428–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krone:1996:ICF

[863] O. Krone, M. Aguilar, B. Hirsbrunner, and V. Sunderam. Integrating coordination features in PVM. Lecture Notes in Computer Science, 1061:432– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olivier:1996:SFE

[864] P. A. Olivier. A simulator framework for embedded systems. Lecture Notes in Computer Science, 1061:436–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Straub:1996:UBB

[865] P. A. Straub and C. L. Hurtado. Understanding behavior of business process models. Lecture Notes in Computer Science, 1061:440–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchez:1996:FPG

[866] E. Sanchez. Field programmable gate array (FPGA) circuits. Lecture Notes in Computer Science, 1062:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomassini:1996:EA

[867] M. Tomassini. Evolutionary algorithms. Lecture Notes in Computer Science, 1062:19–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruau:1996:ACD

[868] F. Gruau. Artificial cellular development in optimization and compilation. Lecture Notes in Computer Science, 1062:48–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeGaris:1996:CBE

[869] H. De Garis. CAM-BRAIN: The evolutionary engineering of a billion neuron artificial brain by 2001 which grows/ evolves at electronic speeds inside a cellular automata machine (CAM). Lecture Notes in Computer Science, 1062: 76–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitano:1996:MES

[870] H. Kitano. Morphogenesis for evolvable systems. Lecture Notes in Computer Science, 1062:99–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Higuchi:1996:EHA

[871] T. Higuchi, M. Iwata, I. Kajitani, and H. Iba. Evolvable hardware and its applications to pattern recognition and fault-tolerant systems. Lecture Notes in Computer Science, 1062:118–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thompson:1996:UEH

[872] A. Thompson, I. Harvey, and P. Husbands. Unconstrained evolution and hard consequences. Lecture Notes in Computer Science, 1062:136–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marchal:1996:EBS

[873] P. Marchal, P. Nussbaum, C. Piguet, and S. Durand. Embryonics: The birth of synthetic life. Lecture Notes in Computer Science, 1062:166–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mange:1996:ENF

[874] D. Mange, M. Goeke, D. Madon, and A. Stauffer. Embryonics: A new family of coarse-grained field programmable gate array with self-repair and selfreproducing properties. Lecture Notes in Computer Science, 1062:197–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mondada:1996:EMA

[875] F. Mondada and D. Floreano. Evolution and mobile autonomous robotics. Lecture Notes in Computer Science, 1062:221–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemmi:1996:DEH

[876] H. Hemmi, J. Mizoguchi, and K. Shimohara. Development and evolution of hardware behaviors. Lecture Notes in Computer Science, 1062:250–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeck:1996:ESA

[877] T. Baeck. Evolution strategies: An alternative evolutionary algorithm. Lecture Notes in Computer Science, 1063: 3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fogel:1996:IEP

[878] D. B. Fogel and L. J. Fogel. An introduction to evolutionary programming. Lecture Notes in Computer Science, 1063:21–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerf:1996:ATG

[879] R. Cerf. An asymptotic theory of genetic algorithms. Lecture Notes in Computer Science, 1063:37–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Venturini:1996:TGT

[880] G. Venturini. Towards a genetic theory of easy and hard functions. Lecture Notes in Computer Science, 1063: 54–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jelasity:1996:GCM

[881] M. Jelasity and J. Dombi. GAs, a concept of modeling species in genetic algorithms. Lecture Notes in Computer Science, 1063:69–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godart:1996:GAP

[882] C. Godart and M. Krueger. A genetic algorithm with parallel steadystate reproduction. Lecture Notes in Computer Science, 1063:86–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravise:1996:IBC

[883] C. Ravise, M. Sebag, and M. Schoenauer. Induction-based control of genetic algorithms. Lecture Notes in Computer Science, 1063:100–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viennet:1996:NMO

[884] R. Viennet, C. Fonteix, and I. Marc. New multicriteria optimization method based on the use of a diploid genetic algorithm: Example of an industrial problem. Lecture Notes in Computer Science, 1063:120–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourgine:1996:LVC

[885] P. Bourgine and D. Snyers. Lotka Volterra coevolution at the edge of chaos. Lecture Notes in Computer Science, 1063:131–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dumeur:1996:ETC

[886] R. Dumeur. Evolution through cooperation: The symbiotic algorithm. Lecture Notes in Computer Science, 1063:

145–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meuleau:1996:AEC

[887] N. Meuleau and C. Lattaud. The artificial evolution of cooperation. Lecture Notes in Computer Science, 1063: 159–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xanthakis:1996:ISF

[888] S. Xanthakis, C. Karapoulios, R. Pajot, and A. Rozz. Immune system and faulttolerant computing. Lecture Notes in Computer Science, 1063:181–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruau:1996:MGN

[889] F. Gruau. Modular genetic neural networks for six-legged locomotion. Lecture Notes in Computer Science, 1063: 201–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michel:1996:ALA

[890] O. Michel. An artificial life approach for the synthesis of autonomous agents. Lecture Notes in Computer Science, 1063:220–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoenauer:1996:HLD

[891] M. Schoenauer and E. Ronald. How long does it take to evolve a neural net? Lecture Notes in Computer Science, 1063:232–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cretin:1996:MIR

[892] G. Cretin, E. Lutton, J. Levy-Vehel, and P. Glevarec. Mixed IFS: Resolution of the inverse problem using genetic programming. Lecture Notes in Computer Science, 1063:247–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graf:1996:IES

[893] J. Graf and W. Banzhaf. Interactive evolution for simulated natural evolution. Lecture Notes in Computer Science, 1063:259–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Louchet:1996:BNT

[894] J. Louchet, M. Boccara, D. Crochemore, and X. Provot. Building new tools for synthetic image animation by using evolutionary techniques. Lecture Notes in Computer Science, 1063:273– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lutton:1996:GAS

[895] E. Lutton and P. Martinez. A genetic algorithm with sharing for the detection of 2D geometric primitives in images. Lecture Notes in Computer Science, 1063:287–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alliot:1996:GAI

[896] J.-M. Alliot and N. Durand. A genetic algorithm to improve an Othello program. Lecture Notes in Computer Science, 1063:307–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeck:1996:CSP

[897] T. Baeck, M. Schuetz, and S. Khuri. A comparative study of a penalty function, a repair heuristic and stochastic operators with the set-covering problem. Lecture Notes in Computer Science, 1063:320–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hao:1996:SGS

[898] J.-K. Hao and R. Dorne. Study of genetic search for the frequency assignment problem. Lecture Notes in Computer Science, 1063:333–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorrini:1996:AEA

[899] V. Gorrini and M. Dorigo. An application of evolutionary algorithms to the scheduling of robotic operations. Lecture Notes in Computer Science, 1063:345–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kane:1996:GOT

[900] C. Kane and M. Schoenauer. Genetic operators for two-dimensional shape optimization. Lecture Notes in Computer Science, 1063:355–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Medioni:1996:ATC

[901] F. Medioni, N. Durand, and J.-M Alliot. Air traffic conflict resolution by genetic algorithms. Lecture Notes in Computer Science, 1063:370–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slimane:1996:OHM

[902] M. Slimane, G. Venturini, J.-P. Asselin de Beauville, and T. Brouard. Optimizing hidden Markov models with a genetic algorithm. Lecture Notes in Computer Science, 1063:384–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armstrong:1996:SCI

[903] M. Armstrong, A. Zisserman, and R. Hartley. Self-calibration from image triplets. Lecture Notes in Computer Science, 1064:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Irani:1996:PGP

[904] M. Irani and P. Anandan. Parallax geometry of pairs of points for 3D scene analysis. Lecture Notes in Computer Science, 1064:17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pollefeys:1996:ERI

[905] M. Pollefeys, L. Van Cool, and M. Proesmans. Euclidean 3D reconstruction from image sequences with variable focal lengths. Lecture Notes in Computer Science, 1064:31–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belhumeur:1996:EVF

[906] P. N. Belhumeur, J. P. Hespanha, and D. J. Kriegman. Eigenfaces vs. fisherfaces: recognition using class specific linear projection. Lecture Notes in Computer Science, 1064:45–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamiroy:1996:ROI

[907] B. Lamiroy and P. Gros. Rapid object indexing and recognition using enhanced geometric hashing. Lecture Notes in Computer Science, 1064:59– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pilu:1996:RGP

[908] M. Pilu and R. B. Fisher. Recognition of geons by parametric deformable contour models. Lecture Notes in Computer Science, 1064:71–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bignone:1996:AEG

[909] F. Bignone, O. Henricsson, P. Fua, and M. Stricker. Automatic extraction of generic house roofs from high resolution aerial imagery. Lecture Notes in Computer Science, 1064:85–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caselles:1996:TDO

[910] V. Caselles, R. Kimmel, G. Sapiro, and C. Sbert. Three dimensional object modeling via minimal surfaces. Lecture Notes in Computer Science, 1064: 97–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fletcher:1996:CBR

[911] G. J. Fletcher and P. J. Giblin. Class based reconstruction techniques using singular apparent contours. Lecture Notes in Computer Science, 1064:107– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hilton:1996:RSR

[912] A. Hilton, A. J. Stoddart, J. Illingworth, and T. Windeatt. Reliable surface reconstruction from multiple range images. Lecture Notes in Computer Science, 1064:117–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hougen:1996:SAS

[913] D. R. Hougen and N. Ahuja. Shape from appearance: a statistical approach to surface shape estimation. Lecture Notes in Computer Science, 1064:127–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lachaud:1996:VSU

[914] J.-O. Lachaud and A. Montanvert. Volumic segmentation using hierarchical representation and triangulated surface. Lecture Notes in Computer Science, 1064:137–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laveau:1996:OPG

[915] S. Laveau and O. Faugeras. Oriented projective geometry for computer vision. Lecture Notes in Computer Science, 1064:147–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1996:FCF

[916] F. Li, J. M. Brady, and C. Wiles. Fast computation of the fundamental matrix for an active stereo vision system. Lecture Notes in Computer Science, 1064:157–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malik:1996:BVO

[917] J. Malik. On binocularly viewed occlusion junctions. Lecture Notes in Computer Science, 1064:167–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rothwell:1996:USP

[918] C. Rothwell and J. Stern. Understanding the shape properties of trihedral polyhedra. Lecture Notes in Computer Science, 1064:175–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bangham:1996:NSS

[919] J. A. Bangham, R. Harvey, P. D. Ling, and R. V. Aldridge. Nonlinear scale-space from n-dimensional sieves. Lecture Notes in Computer Science, 1064:189–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casadei:1996:HCR

[920] S. Casadei and S. Mitter. Hierarchical curve reconstruction. part I: bifurcation analysis and recovery of smooth curves. Lecture Notes in Computer Science, 1064:199–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horng:1996:TFC

[921] M.-H. Horng, Y.-N. Sun, and X.-Z. Lin. Texture feature coding method for classification of liver sonography. Lecture Notes in Computer Science, 1064:209– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koethe:1996:LAS

[922] U. Koethe. Local appropriate scale in morphological scale-space. Lecture Notes in Computer Science, 1064:219– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindeberg:1996:SSC

[923] T. Lindeberg and D. Fagerstroem. Scale-space with causal time direction. Lecture Notes in Computer Science, 1064:229–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez:1996:TCC

[924] A. M. Lopez and J. Serrat. Tracing crease curves by solving a system of differential equations. Lecture Notes in Computer Science, 1064:241–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malladi:1996:FUM

[925] R. Malladi and J. A. Sethian. Flows under min/max curvature flow and mean curvature: applications in image processing. Lecture Notes in Computer Science, 1064:251–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olson:1996:DHT

[926] C. F. Olson. Decomposition of the Hough Transform: curve detection with efficient error propagation. Lecture Notes in Computer Science, 1064: 263–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravela:1996:IRU

[927] S. Ravela, R. Manmatha, and E. M. Riseman. Image retrieval using scalespace matching. Lecture Notes in Computer Science, 1064:273–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taubin:1996:OSS

[928] G. Taubin, T. Zhang, and G. Golub. Optimal surface smoothing as filter design. Lecture Notes in Computer Science, 1064:283–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vernon:1996:SDI

[929] D. Vernon. Segmentation in dynamic image sequences by isolation of coherent wave profiles. Lecture Notes in Computer Science, 1064:293–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xie:1996:TSU

[930] Z.-Y. Xie and J. M. Brady. Texture segmentation using local energy in wavelet scale space. Lecture Notes in Computer Science, 1064:304–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bardinet:1996:TMD

[931] E. Bardinet, L. Cohen, and N. Ayache. Tracking medical 3D data with a deformable parametric model. Lecture Notes in Computer Science, 1064: 317–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Black:1996:ETR

[932] M. J. Black and A. D. Jepson. Eigen tracking: robust matching and tracking of articulated objects using a viewbased representation. Lecture Notes in Computer Science, 1064:329–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Isard:1996:CTS

[933] M. Isard and A. Blake. Contour tracking by stochastic propagation of conditional density. Lecture Notes in Computer Science, 1064:343–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reynard:1996:LDC

[934] D. Reynard, A. Wildenberg, A. Blake, and J. Marchant. Learning dynamics of complex motions from image sequences. Lecture Notes in Computer Science, 1064:357–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amir:1996:QAG

[935] A. Amir and M. Lindenbaum. Quantitative analysis of grouping processes. Lecture Notes in Computer Science, 1064:371–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cham:1996:GSC

[936] T.-J. Cham and R. Cipolla. Geometric saliency of curve correspondences and grouping of symmetric contours. Lecture Notes in Computer Science, 1064:385–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elder:1996:CCC

[937] J. H. Elder and S. W. Zucker. Computing contour closure. Lecture Notes in Computer Science, 1064:399–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geiger:1996:VOI

[938] D. Geiger and K. Kumaran. Visual organization of illusory surfaces. Lecture Notes in Computer Science, 1064:413– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaarding:1996:URR

[939] J. Gaarding, J. Porrill, J. P. Frisby, and J. E. W. Mayhew. Uncalibrated relief reconstruction and model alignment from binocular disparities. Lecture Notes in Computer Science, 1064: 427–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robert:1996:DDM

[940] L. Robert and R. Deriche. Dense depth map reconstruction: a minimization and regularization approach which preserves discontinuities. Lecture Notes in Computer Science, 1064:439–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomasi:1996:SS

[941] C. Tomasi and R. Manduchi. Stereo without search. Lecture Notes in Computer Science, 1064:452–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arbel:1996:IVS

[942] T. Arbel and F. P. Ferrie. Informative views and sequential recognition. Lecture Notes in Computer Science, 1064:469–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrey:1996:UTS

[943] P. Andrey and P. Tarroux. Unsupervised texture segmentation using selectionist relaxation. Lecture Notes in Computer Science, 1064:482–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cass:1996:RAS

[944] T. A. Cass. Robust affine structure matching for 3D object recognition. Lecture Notes in Computer Science, 1064:492–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costen:1996:AFR

[945] N. Costen, L. Craw, G. Robertson, and S. Akamatsu. Automatic face recognition: what representation? Lecture Notes in Computer Science, 1064:504– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cross:1996:GSS

[946] A. D. J. Cross, R. C. Wilson, and E. R. Hancock. Genetic search for structural matching. Lecture Notes in Computer Science, 1064:514–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hlavac:1996:ASR

[947] V. Hlavac, A. Leonardis, and T. Werner. Automatic selection of reference views for image-based scene representations. Lecture Notes in Computer Science, 1064:526–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huttenlocher:1996:ORU

[948] D. P. Huttenlocher, R. H. Lilien, and C. F. Olson. Object recognition using subspace methods. Lecture Notes in Computer Science, 1064:536–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leung:1996:DLG

[949] T. Leung and J. Malik. Detecting, localizing and grouping repeated scene elements from an image. Lecture Notes in Computer Science, 1064:546–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1996:IRO

[950] T.-L. Liu, M. Donahue, D. Geiger, and R. Hummel. Image recognition with occlusions. Lecture Notes in Computer Science, 1064:556–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mokhtarian:1996:SBO

[951] F. Mokhtarian. Silhouette-based object recognition with occlusion through curvature scale space. Lecture Notes in Computer Science, 1064:566–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nair:1996:FTS

[952] D. Nair and J. K. Aggarwal. A focused target segmentation paradigm. Lecture Notes in Computer Science, 1064:579– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nastar:1996:GIM

[953] C. Nastar, B. Moghaddam, and A. Pentland. Generalized image matching: statistical learning of physicallybased deformations. Lecture Notes in Computer Science, 1064:589–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rothwell:1996:RAO

[954] C. Rothwell. Reasoning about occlusions during hypothesis verification. Lecture Notes in Computer Science, 1064:599–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiele:1996:ORU

[955] B. Schiele and J. L. Crowley. Object recognition using multidimensional receptive field histograms. Lecture Notes in Computer Science, 1064:610– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiller:1996:NO

[956] R. Schiller. Normalization by optimization. Lecture Notes in Computer Science, 1064:620–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steger:1996:ECS

[957] C. Steger. Extracting curvilinear structures: a differential geometric approach. Lecture Notes in Computer Science, 1064:630–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanGool:1996:API

[958] L. Van Gool, T. Moons, and D. Ungureanu. Affine/photometric invariants for planar intensity patterns. Lecture Notes in Computer Science, 1064:642– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vetter:1996:ISS

[959] T. Vetter and T. Poggio. Image synthesis from a single example image. Lecture Notes in Computer Science, 1064:652–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Werman:1996:CIE

[960] M. Werman and D. Weinshall. Complexity of indexing: efficient and learnable large database indexing. Lecture Notes in Computer Science, 1064:660– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheong:1996:SRV

[961] L. F. Cheong, C. Fermueller, and Y. Aloimonos. Spatiotemporal representations for visual navigation. Lecture Notes in Computer Science, 1064: 673–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daniilidis:1996:DMS

[962] K. Daniilidis and I. Thomas. Decoupling the 3D motion space by fixation. Lecture Notes in Computer Science, 1064:685–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanatani:1996:AST

[963] K. Kanatani. Automatic singularity test for motion analysis by an information criterion. Lecture Notes in Computer Science, 1064:697–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szeliski:1996:SAS

[964] R. Szeliski and S. B. Kang. Shape ambiguities in structure from motion. Lecture Notes in Computer Science, 1064: 709–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barnard:1996:CCS

[965] K. Barnard, G. Finlayson, and B. Funt. Colour constancy for scenes with varying illumination. Lecture Notes in Computer Science, 1065:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finlayson:1996:CAI

[966] G. D. Finlayson, S. S. Chatterjee, and B. V. Funt. Color angular indexing. Lecture Notes in Computer Science, 1065:16–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koenderink:1996:BRD

[967] J. J. Koenderink, A. J. Van Doorn, and M. Stavridi. Bidirectional reflection distribution function expressed in terms of surface scattering modes. Lecture Notes in Computer Science, 1065: 28–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolff:1996:GLL

[968] L. B. Wolff. Generalizing Lambert’s law for smooth surfaces. Lecture Notes in Computer Science, 1065:40– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elder:1996:LSC

[969] J. H. Elder and S. W. Zucker. Local scale control for edge detection and blur estimation. Lecture Notes in Computer Science, 1065:57–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielsen:1996:RSS

[970] M. Nielsen, L. Florack, and R. Deriche. Regularization, scale-space, and edge detection filters. Lecture Notes in Computer Science, 1065:70–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simoncelli:1996:DDR

[971] E. P. Simoncelli and H. Farid. Direct differential range estimation using optical masks. Lecture Notes in Computer Science, 1065:82–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aastroem:1996:GEC

[972] K. Aastroem, R. Cipolla, and P. J. Giblin. Generalised epipolar constraints. Lecture Notes in Computer Science, 1065:97–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyer:1996:OMC

[973] E. Boyer. Object models from contour sequences. Lecture Notes in Computer Science, 1065:109–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodsky:1996:DMF

[974] T. Brodsky, C. Fermueller, and Y. Aloimonos. Directions of motion fields are hardly ever ambiguous. Lecture Notes in Computer Science, 1065:119– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christy:1996:ERP

[975] S. Christy and R. Horaud. Euclidean reconstruction: from paraperspective to perspective. Lecture Notes in Computer Science, 1065:129–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1996:OFP

[976] I. Cohen and I. Herlin. Optical flow and phase portrait methods for environmental satellite image sequences. Lecture Notes in Computer Science, 1065: 141–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giachetti:1996:ROF

[977] A. Giachetti and V. Torre. Refinement of optical flow estimation and detection of motion edges. Lecture Notes in Computer Science, 1065:151–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lawn:1996:REC

[978] J. M. Lawn and R. Cipolla. Reliable extraction of the camera motion using constraints on the epipole. Lecture Notes in Computer Science, 1065: 161–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1996:AVE

[979] H. Liu, T.-H. Hong, M. Herman, and R. Chellappa. Accuracy vs. efficiency trade-offs in optical flow algorithms. Lecture Notes in Computer Science, 1065:174–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oliensis:1996:RBT

[980] J. Oliensis. Rigorous bounds for twoframe structure from motion. Lecture Notes in Computer Science, 1065:184– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shashua:1996:RCM

[981] A. Shashua and S. Avidan. The rank 4 constraint in multiple (3) view geometry. Lecture Notes in Computer Science, 1065:196–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vieville:1996:USD

[982] T. Vieville and D. Lingrand. Using singular displacements for uncalibrated monocular visual systems. Lecture Notes in Computer Science, 1065: 207–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weinshall:1996:DMP

[983] D. Weinshall, M. Werman, and A. Shashua. Duality of multi-point and multi-frame geometry: fundamental shape matrices and tensors. Lecture Notes in Computer Science, 1065: 217–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiles:1996:ACM

[984] C. Wiles and J. M. Brady. On the appropriateness of camera models. Lecture Notes in Computer Science, 1065:228–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiles:1996:GPM

[985] C. Wiles and J. M. Brady. Ground plane motion camera models. Lecture Notes in Computer Science, 1065:238– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amini:1996:SST

[986] A. A. Amini, R. W. Curwen, and J. C. Gore. Snakes and splines for tracking non-rigid heart motion. Lecture Notes in Computer Science, 1065:251– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benayoun:1996:LQM

[987] S. Benayoun, D. Kharitonsky, A. Zilberman, and S. Peleg. Local quantitative measurements for cardiac motion analysis. Lecture Notes in Computer Science, 1065:262–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Byrne:1996:AMB

[988] M. J. Byrne and J. Graham. Application of model based image interpretation methods to diabetic neuropathy. Lecture Notes in Computer Science, 1065:272–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fletcher:1996:GAM

[989] S. Fletcher, A. Bulpitt, and D. C. Hogg. Global alignment of MR images using a scale based hierarchical model. Lecture Notes in Computer Science, 1065:283–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hall:1996:RBV

[990] P. Hall, M. Ngan, and P. Andreae. Reconstruction of blood vessel networks from X-ray projections and a vascular catalogue. Lecture Notes in Computer Science, 1065:293–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gil:1996:CMM

[991] S. Gil, R. Milanese, and T. Pun. Combining multiple motion estimates for vehicle tracking. Lecture Notes in Computer Science, 1065:307–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Howarth:1996:VSM

[992] R. Howarth and H. Buxton. Visual surveillance monitoring and watching. Lecture Notes in Computer Science, 1065:321–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ngoi:1996:RAC

[993] K. P. Ngoi and J. Jia. A robust active contour model for natural scene contour extraction with automatic thresholding. Lecture Notes in Computer Science, 1065:335–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siskind:1996:MLA

[994] J. Y. Siskind and Q. Morris. A maximum-likelihood approach to visual event classification. Lecture Notes in Computer Science, 1065:347–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gdalyahu:1996:MSR

[995] Y. Gdalyahu and D. Weinshall. Measures for silhouettes resemblance and representative silhouettes of curved objects. Lecture Notes in Computer Science, 1065:363–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaucic:1996:RTL

[996] R. Kaucic, B. Dalton, and A. Blake. Real-time lip tracking for audio-visual speech recognition applications. Lecture Notes in Computer Science, 1065: 376–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kollnig:1996:MOM

[997] H. Kollnig and H.-H. Nagel. Matching object models to segments from an optical flow field. Lecture Notes in Computer Science, 1065:388–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solloway:1996:QAC

[998] S. Solloway, C. J. Taylor, C. E. Hutchinson, and J. C. Waterton. Quantification of articular cartilage from MR images using active shape models. Lecture Notes in Computer Science, 1065:400–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brooks:1996:DMS

[999] M. J. Brooks, L. De Agapito, D. Q. Huynh, and L. Baumela. Direct methods for self-calibration of a moving stereo head. Lecture Notes in Computer Science, 1065:415–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delherm:1996:DRZ

[1000] C. Delherm, J. M. Lavest, M. Dhome, and J. T. Lapreste. Dense reconstruction by zooming. Lecture Notes in Computer Science, 1065:427–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watanabe:1996:TOC

[1001] M. Watanabe and S. K. Nayar. Telecentric optics for computational vision. Lecture Notes in Computer Science, 1065:439–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braud:1996:RPT

[1002] P. Braud, J.-T. Lapreste, and M. Dhome. Recognition, pose and tracking of modelled polyhedral objects by multi-ocular vision. Lecture Notes in Computer Science, 1065:455–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cootes:1996:LOV

[1003] T. F. Cootes and C. J. Taylor. Locating objects of varying shape using statistical feature detectors. Lecture Notes in Computer Science, 1065:465–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernyhough:1996:GSR

[1004] J. H. Fernyhough, A. G. Cohn, and D. C. Hogg. Generation of semantic regions from image sequences. Lecture Notes in Computer Science, 1065:475– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frank:1996:TOV

[1005] T. Frank, M. Haag, H. Kollnig, and H.- H Nagel. Tracking of occluded vehicles in traffic scenes. Lecture Notes in Computer Science, 1065:485–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fua:1996:IHC

[1006] P. Fua and U. Brechbuehler. Imposing hard constraints on soft snakes. Lecture Notes in Computer Science, 1065: 495–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hager:1996:XVC

[1007] G. D. Hager and K. Toyama. X vision: combining image warping and geometric constraints for fast visual tracking. Lecture Notes in Computer Science, 1065:507–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lerasle:1996:HBT

[1008] F. Lerasle, G. Rives, M. Dhome, and A. Yassine. Human body tracking by monocular vision. Lecture Notes in Computer Science, 1065:518–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mann:1996:CPS

[1009] R. Mann, A. D. Jepson, and J. M.

Siskind. Computational perception of scene dynamics. Lecture Notes in Computer Science, 1065:528–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maybank:1996:FVT

[1010] S. J. Maybank, A. D. Worrall, and G. D. Sullivan. A filter for visual tracking based on a stochastic model for driver behaviour. Lecture Notes in Computer Science, 1065:540–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Metaxas:1996:EAD

[1011] D. Metaxas and I. A. Kakadiaris.

Elastically adaptive deformable models. Lecture Notes in Computer Science, 1065:550–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rowe:1996:SFM

[1012] S. Rowe and A. Blake. Statistical feature modelling for active contours. Lecture Notes in Computer Science, 1065:560–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bascle:1996:MDS

[1013] B. Bascle, A. Blake, and A. Zisserman. Motion deblurring and super-resolution from an image sequence. Lecture Notes in Computer Science, 1065:573–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colombo:1996:UVT

[1014] C. Colombo and J. L. Crowley. Uncalibrated visual tasks via linear interaction. Lecture Notes in Computer Science, 1065:583–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fleck:1996:FNP

[1015] M. M. Fleck, D. A. Forsyth, and C. Bregler. Finding naked people. Lecture Notes in Computer Science, 1065: 593–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kokaram:1996:SRM

[1016] A. C. Kokaram and S. J. Godsill. A system for reconstruction of missing data in image sequences using sampled 3D AR models and MRF motion priors. Lecture Notes in Computer Science, 1065:613–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1996:ESS

[1017] V. Mueller. Elimination of specular surface-reflectance using polarized and unpolarized light. Lecture Notes in Computer Science, 1065:625–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohnishi:1996:SRV

[1018] N. Ohnishi, K. Kumaki, T. Yamamura, and T. Tanaka. Separating real and virtual objects from their overlapping images. Lecture Notes in Computer Science, 1065:636–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reid:1996:GDV

[1019] I. Reid and A. Zisserman. Goaldirected video metrology. Lecture Notes in Computer Science, 1065:647–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taylor:1996:RPM

[1020] C. J. Taylor, P. E. Debevec, and J. Malik. Reconstructing polyhedral models of architectural scenes from photographs. Lecture Notes in Computer Science, 1065:659–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heyden:1996:AVM

[1021] A. Heyden and K. Aastroem. Algebraic varieties in multiple view geometry. Lecture Notes in Computer Science, 1065:671–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beardsley:1996:MAE

[1022] P. Beardsley, P. Torr, and A. Zisserman. 3D model acquisition from extended image sequences. Lecture Notes in Computer Science, 1065:683– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faugeras:1996:CSM

[1023] T. Faugeras and O. Papadopoulo. Computing structure and motion of general 3D curves from monocular sequences of perspective images. Lecture Notes in Computer Science, 1065:696– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sturm:1996:FBA

[1024] P. Sturm and B. Triggs. A factorization based algorithm for multiimage projective structure and motion. Lecture Notes in Computer Science, 1065:709–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lygeros:1996:GTA

[1025] J. Lygeros, D. N. Godbole, and S. Sastry. A game-theoretic approach to hybrid system design. Lecture Notes in Computer Science, 1066:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kesten:1996:VCT

[1026] Y. Kesten, Z. Manna, and A. Pnueli. Verifying clocked transition systems. Lecture Notes in Computer Science, 1066:13–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benveniste:1996:CUM

[1027] A. Benveniste. Compositional and uniform modeling of hybrid systems. Lecture Notes in Computer Science, 1066: 41–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1996:HCH

[1028] V. Gupta, R. Jagadeesan, and V. Saraswat. Hybrid cc, hybrid automata, and program verification. Lecture Notes in Computer Science, 1066: 52–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghosh:1996:CSD

[1029] M. K. Ghosh, S. I. Marcus, and A. Arapostathis. Controlled switching diffusions as hybrid processes. Lecture Notes in Computer Science, 1066:64– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ge:1996:HSC

[1030] X. Ge, W. Kohn, A. Nerode, and J. B. Remmel. Hybrid systems: chattering approximation to relaxed controls. Lecture Notes in Computer Science, 1066:76–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weinberg:1996:VAV

[1031] H. B. Weinberg, N. Lynch, and N. Delisle. Verification of automated vehicle protection systems. Lecture Notes in Computer Science, 1066:101– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeLemos:1996:ERS

[1032] R. De Lemos and J. G. Hall. Extended RTL in the specification and verification of an industrial press. Lecture Notes in Computer Science, 1066:114– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sintzoff:1996:AVS

[1033] M. Sintzoff. Abstract verification of structured dynamical systems. Lecture Notes in Computer Science, 1066:126– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deshpande:1996:DET

[1034] A. Deshpande, D. Godbole, A. Goellue, and P. Varaiya. Design and evaluation tools for automated highway systems. Lecture Notes in Computer Science, 1066:138–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godhavn:1996:HCS

[1035] J.-M. Godhavn, T. Lauvdal, and O. Egeland. Hybrid control in sea traffic management systems. Lecture Notes in Computer Science, 1066:149– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haddon:1996:VHS

[1036] J. A. Haddon, D. N. Godbole, A. Deshpande, and J. Lygeros. Verification of hybrid systems: monotonicity in the AHS control system. Lecture Notes in Computer Science, 1066:161–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artstein:1996:ESH

[1037] Z. Artstein. Examples of stabilization with hybrid feedback. Lecture Notes in Computer Science, 1066:173–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branicky:1996:GHD

[1038] M. S. Branicky. General hybrid dynamical systems: modeling, analysis, and control. Lecture Notes in Computer Science, 1066:186–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seidman:1996:RMR

[1039] T. I. Seidman. The residue of model reduction. Lecture Notes in Computer Science, 1066:201–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daws:1996:TK

[1040] C. Daws, A. Olivero, S. Tripakis, and S. Yovine. The tool KRONOS. Lecture Notes in Computer Science, 1066: 208–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1996:TAC

[1041] R. Alur and R. P. Kurshan. Timing analysis in COSPAN. Lecture Notes in Computer Science, 1066:220–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bengtsson:1996:UTS

[1042] J. Bengtsson, K. G. Larsen, F. Larsson, and P. Pettersson. UPPAAL: a tool suite for automatic verification of real-time systems. Lecture Notes in Computer Science, 1066:232–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimitriadis:1996:ODH

[1043] V. D. Dimitriadis, N. Shah, and C. C. Pantelides. Optimal design of hybrid controllers for hybrid process systems. Lecture Notes in Computer Science, 1066:244–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holloway:1996:LFM

[1044] L. E. Holloway. On-line fault monitoring of a class of hybrid systems using templates with dynamic time scaling. Lecture Notes in Computer Science, 1066:258–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qiwen:1996:HDC

[1045] X. Qiwen and H. Weidong. Hierarchical design of a chemical concentration control system. Lecture Notes in Computer Science, 1066:270–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stroemberg:1996:SBG

[1046] J.-E. Stroemberg, S. Nadjm-Tehrani, and J. L. Top. Switched bond graphs as front-end to formal verification of hybrid systems. Lecture Notes in Computer Science, 1066:282–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ji:1996:FSS

[1047] W. Ji and H. Weidong. Formal specification of stability in hybrid control systems. Lecture Notes in Computer Science, 1066:294–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heitmeyer:1996:RSH

[1048] C. Heitmeyer. Requirements specifications for hybrid systems. Lecture Notes in Computer Science, 1066:304– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sinclair:1996:VHS

[1049] D. Sinclair, E. Holz, D. Witaszek, and M. Wasowski. Validation of hybrid systems by co-simulation. Lecture Notes in Computer Science, 1066:315–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narain:1996:PTH

[1050] S. Narain. Proofs from temporal hypotheses by symbolic simulation. Lecture Notes in Computer Science, 1066: 327–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cofer:1996:CTD

[1051] D. D. Cofer and V. K. Garg. On controlling timed discrete event systems. Lecture Notes in Computer Science, 1066:340–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kumar:1996:SCR

[1052] R. Kumar and M. A. Shayman. Supervisory control of real-time systems using prioritized synchronization. Lecture Notes in Computer Science, 1066: 350–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puri:1996:ADI

[1053] A. Puri, V. Borkar, and P. Varaiya. ??-approximation of differential inclusions. Lecture Notes in Computer Science, 1066:362–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1996:LPP

[1054] T. A. Henzinger and H. Wong-Toi. Linear phase-portrait approximations for nonlinear hybrid systems. Lecture Notes in Computer Science, 1066:377– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ceraens:1996:DRP

[1055] K. Ceraens and J. Viksna. Deciding reachability for planar multipolynomial systems. Lecture Notes in Computer Science, 1066:389–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffmann:1996:MHD

[1056] I. Hoffmann and K.-U Klatt. Modeling hybrid dynamical systems. Lecture Notes in Computer Science, 1066:401– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kourjanski:1996:SHS

[1057] M. Kourjanski and P. Varaiya. Stability of hybrid systems. Lecture Notes in Computer Science, 1066:413–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1996:MSH

[1058] H. S. Park, Y. S. Kim, W. H. Kwon, and S. J. Lee. Model and stability of hybrid linear systems. Lecture Notes in Computer Science, 1066:424–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sontag:1996:IAL

[1059] E. D. Sontag. Interconnected automata and linear systems: a theoretical framework in discrete time. Lecture Notes in Computer Science, 1066:436–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lynch:1996:MVA

[1060] N. Lynch. Modeling and verification of automated transit systems, using timed automata, invariants, and simulations. Lecture Notes in Computer Science, 1066:449–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stiver:1996:IBA

[1061] J. A. Stiver, P. J. Antsaklis, and M. D.

Lemmon. An invariant-based approach to the design of hybrid control systems containing clocks. Lecture Notes in Computer Science, 1066:464–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niinomi:1996:RAA

[1062] T. Niinomi, B. H. Krogh, and J. E. R.

Cury. Refinements of approximating automata for synthesis of supervisory controllers for hybrid systems. Lecture Notes in Computer Science, 1066:475– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bailey:1996:DIC

[1063] S. Bailey, R. L. Grossman, L. Gu, and D. Hanley. A data-intensive computing approach to path planning and mode management for hybrid systems. Lecture Notes in Computer Science, 1066:485–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lynch:1996:HA

[1064] N. Lynch, R. Segala, F. Vaandrager, and H. B. Weinberg. Hybrid I/O automata. Lecture Notes in Computer Science, 1066:496–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaochen:1996:FDH

[1065] Z. Chaochen, W. Ji, and A. P. Ravn. A formal description of hybrid systems. Lecture Notes in Computer Science, 1066:511–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouajjani:1996:LVA

[1066] A. Bouajjani and Y. Lakhnech. Logics vs. automata: the hybrid case. Lecture Notes in Computer Science, 1066: 531–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bett:1996:HGS

[1067] C. J. Bett and M. D. Lemmon. H gain schedule synthesis of supervisory hybrid control systems. Lecture Notes in Computer Science, 1066:543–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Savkin:1996:NAR

[1068] A. V. Savkin, R. J. Evans, and I. R. Petersen. A new approach to robust control of hybrid systems. Lecture Notes in Computer Science, 1066:553–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raisch:1996:ACH

[1069] J. Raisch and S. O’Young. A DES approach to control of hybrid dynamical systems. Lecture Notes in Computer Science, 1066:563–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1996:DMC

[1070] K. G. Larsen, P. Pettersson, and W. Yi. Diagnostic model-checking for real-time systems. Lecture Notes in Computer Science, 1066:575–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1996:SVH

[1071] Y. Zhang and A. K. Mackworth. Specification and verification of hybrid dynamic systems with timed inverted sans serif aye-automata. Lecture Notes in Computer Science, 1066:587–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1996:FPR

[1072] K. G. Larsen, B. Steffen, and C. Weise.

Fischer’s protocol revisited: a simple proof using modal constraints. Lecture Notes in Computer Science, 1066:604– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjoerstad:1996:PRS

[1073] P. Bjoerstad, R. Moe, R. Olufsen, and E. Vainikko. Parallel reservoir simulation based on domain decomposition techniques. Lecture Notes in Computer Science, 1067:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moe:1996:PPP

[1074] R. Moe, M. Talal Rahman, O. Saevareid, and R. Teigland. Porting and parallel performance of the industrial CFD code MUSIC. Lecture Notes in Computer Science, 1067:12–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thompson:1996:POR

[1075] A. M. Thompson and G. R. Bowen. Parallelisation of a oil reservoir simulation. Lecture Notes in Computer Science, 1067:20–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elliott:1996:RIH

[1076] J. A. Elliott, S. H. Brindle, A. Colbrook, and D. G. Green. Real industrial HPC applications. Lecture Notes in Computer Science, 1067:29–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braathen:1996:PSS

[1077] A. Braathen, J. Cook, A. C. Damhaug, and M. T. Rahman. Parallelisation of the SWAN surface wave analysis code. Lecture Notes in Computer Science, 1067:36–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hargreaves:1996:PCF

[1078] M. L. Hargreaves, D. C. Spicer, C. C. R. Jones, and S. Whittle. Porting a commercial finite difference time domain electromagnetic code to a parallel architecture machine. Lecture Notes in Computer Science, 1067:43–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pomponi:1996:PPP

[1079] R. Pomponi, P. D’Atanasio, E. Rubino, and M. Busuoli. Porting on parallel platforms of a tool for simulating the effects of lightning strokes on telecommunication buildings: A comparison on preliminary results about performances and accuracy on SIMD and MIMD architectures. Lecture Notes in Computer Science, 1067:51–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tinniswood:1996:UPF

[1080] A. D. Tinniswood and P. S. Excell. On the use of parallel finitedifference time-domain methods in the modelling of bioelectromagnetic problems. Lecture Notes in Computer Science, 1067:60–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicolopoulos:1996:CPP

[1081] D. Nicolopoulos and G. Winkelmuller.

A comparison of parallel paradigms applied to an industrial code: RA-DIOSS. Lecture Notes in Computer Science, 1067:66–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delsemme:1996:SMC

[1082] J.-P. Delsemme and O. Zone. SAMMI: S.A.M.C.E.F. migration onto parallel platforms: Second step. Lecture Notes in Computer Science, 1067:74– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz:1996:EAP

[1083] U. Schulz, M. Ast, J. Labarta, and H. Manz. Experiences and achievements with the parallelization of a large finite element system. Lecture Notes in Computer Science, 1067:82–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stueben:1996:PIF

[1084] K. Stueben, H. Mierendorff, C.-A. Thole, and O. Thomas. Parallel industrial fluid dynamics and structural mechanics codes. Lecture Notes in Computer Science, 1067:90–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoessel:1996:TRC

[1085] A. Stoessel, E. Issman, and M. Loriot. Towards real CFD simulations on parallel computers in the aeronautic and automotive industry. Lecture Notes in Computer Science, 1067:99–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauer:1996:PCC

[1086] W. Bauer, H. Reister, F. Ross, and D. Robinson. Parallel computing for CFD applications in the automotive industry — first experiences. Lecture Notes in Computer Science, 1067:106– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forsey:1996:PPE

[1087] C. R. Forsey, C. S. Ierotheou, U. Block, and M. Leatham. Parallelisation and performance evaluation of the aeronautical CFD flow code ESAUNA. Lecture Notes in Computer Science, 1067: 116–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giraud:1996:PII

[1088] L. Giraud, N. Maman, P. Menegazzi, and A. Micelotta. Parallel industrial incompressible CFD calculations with HPCN3S. Lecture Notes in Computer Science, 1067:122–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaput:1996:EPC

[1089] E. Chaput, C. Gacherieu, and L. Tourrette. Experience with parallel computing for the design of transport aircrafts at aerospatiale. Lecture Notes in Computer Science, 1067:128–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allen:1996:IAD

[1090] P. Allen, R. McKendrick, C. Scott, and M. Buonanno. Interactive anomaly detection in large transaction history databases. Lecture Notes in Computer Science, 1067:143–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watson:1996:SMI

[1091] P. Watson, M. Ward, and K. Hoyle. The system management of the ICL GOLDRUSH parallel database server. Lecture Notes in Computer Science, 1067:150–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freitas:1996:PDM

[1092] A. A. Freitas and S. H. Lavington. Parallel data mining for very large relational databases. Lecture Notes in Computer Science, 1067:158–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1996:DSM

[1093] M. H. Williams, S. Zhou, H. Taylor, and N. Tomov. Decision supporting for management of parallel database systems. Lecture Notes in Computer Science, 1067:164–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meacham:1996:PGM

[1094] K. Meacham and D. Green. Parallelization of the GROMOS87 molecular dynamics code: An update. Lecture Notes in Computer Science, 1067: 170–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Addison:1996:PPN

[1095] C. Addison, E. Appiani, R. Cook, and M. Corvi. PULSAR: Parallel noise despeckling of SAR images. Lecture Notes in Computer Science, 1067:177– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gill:1996:RPR

[1096] A. Gill, M. Surridge, G. Scielzo, and R. Felici. RAPT: A parallel radiotherapy treatment planning code. Lecture Notes in Computer Science, 1067:183– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kandhai:1996:CMA

[1097] B. D. Kandhai, P. M. A. Sloot, and J. P. Huot. Constrained migration of an atmospheric circulation model. Lecture Notes in Computer Science, 1067: 193–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1996:PMA

[1098] A. Mueller. Parallelization of a mesoscale atmospheric transportchemistry model. Lecture Notes in Computer Science, 1067:200–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanGijzen:1996:POF

[1099] M. B. Van Gijzen. Parallel ocean flow computations on a regular and on an irregular grid. Lecture Notes in Computer Science, 1067:207–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dormanns:1996:EAP

[1100] M. Dormanns and W. Sprangers. Experiences with asynchronous parallel molecular dynamics simulations. Lecture Notes in Computer Science, 1067: 213–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joosen:1996:AOC

[1101] W. Joosen, S. Bijnens, J. Van Oeyen, and B. Robben. Affordable overhead in CORRELATE: Comparing two MD simulators. Lecture Notes in Computer Science, 1067:219–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boryczko:1996:ECD

[1102] K. Boryczko, J. Kitowski, and J. Moscinski. Efficiency comparison of data-parallel programming and message-passing paradigm for molecular dynamics simulation. Lecture Notes in Computer Science, 1067:228–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Segers:1996:PMC

[1103] J. Segers, J. Lukkien, and P. Hilbers. Parallel Monte Carlo simulation of chemical surface reactions: A case study. Lecture Notes in Computer Science, 1067:235–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:1996:DDM

[1104] M. Bubak, J. Moscinski, and M. Pogoda. Distributed 2-D molecular dynamics simulation on networked workstations and multiprocessors. Lecture Notes in Computer Science, 1067:243–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apostolakis:1996:GPP

[1105] J. Apostolakis, L. M. Bertolotto, C. E. Bruschini, and P. Calafiura. Generalpurpose parallel computing in a highenergy physics experiment at CERN.

Lecture Notes in Computer Science, 1067:251–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apostolakis:1996:PSH

[1106] J. Apostolakis. Parallel simulation for high energy physics. Lecture Notes in Computer Science, 1067:258–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikita:1996:FPM

[1107] K. S. Nikita and A. J. Marsh. Focusing of the pulse modulated radiation of a concentric waveguide array in biological tissues: An investigation made feasible by HPC. Lecture Notes in Computer Science, 1067:263–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoekstra:1996:SLS

[1108] A. G. Hoekstra, M. D. Grimminck, and P. M. A. Sloot. Simulating light scattering from micron-sized particles: A parallel fast discrete dipole approximation. Lecture Notes in Computer Science, 1067:269–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fogaccia:1996:LBS

[1109] G. Fogaccia, R. Benzi, and F. Romanelli. Lattice Boltzmann simulations of electrostatic plasma turbulence. Lecture Notes in Computer Science, 1067:276–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strietzel:1996:PTS

[1110] M. Strietzel. Parallel turbulence simulation based on MPI. Lecture Notes in Computer Science, 1067:283–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cabitza:1996:EHS

[1111] G. Cabitza, C. Nardone, C. Bagaini, and A. Balzano. Experiences with HPF for scientific applications. Lecture Notes in Computer Science, 1067: 290–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verdu:1996:RTN

[1112] I. Verdu, D. Gimenez, and J. C. Torres. Ray tracing for natural scenes in parallel processors. Lecture Notes in Computer Science, 1067:297–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachem:1996:MTS

[1113] A. Bachem, C. Gawron, C. Moll, and M. Rickert. MiCroscopic traffic simulations of road networks using highperformance computers. Lecture Notes in Computer Science, 1067:306–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraysse:1996:PCS

[1114] V. Fraysse, L. Giraud, and V. Toumazou. Parallel computation of spectral portraits on the Meiko CS2. Lecture Notes in Computer Science, 1067:312– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurc:1996:APM

[1115] T. M. Kurc, C. Aykanat, and B. Oezguec. Active pixel merging on hybercube multicomputers. Lecture Notes in Computer Science, 1067:319– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Persson:1996:PII

[1116] T. Persson and P. Fritzson. Parallel implementation of image reconstruction for the CARABAS long-wave penetrating radar. Lecture Notes in Computer Science, 1067:327–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neschen:1996:CFR

[1117] M. Neschen and F. Nuebel. COG-NITUS — fast and reliable recognition of handwritten forms based on vector quantisation. Lecture Notes in Computer Science, 1067:333–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rieger:1996:PIA

[1118] H. Rieger, L. Fornasier, D. Ockel, and S. Haberhauer. Pilot-implementation of an aerospace design system into a parallel user simulation environment. Lecture Notes in Computer Science, 1067:340–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wedemeijer:1996:PSA

[1119] H. Wedemeijer, H. L. H. Cox, D. J.

Verschuur, and I. L. Ritsema. Parallelisation of seismic algorithms using PVM and FORGE. Lecture Notes in Computer Science, 1067:352–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rudgyard:1996:PLC

[1120] M. Rudgyard, T. Schoenfeld, and I. D’Ast. A parallel library for CFD and other grid-based applications. Lecture Notes in Computer Science, 1067: 358–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ODonnell:1996:IEB

[1121] T. J. O’Donnell, S. G. Hotovy, M. S. Pottle, and D. R. Ripoll. Implementation of ECEPP-Based programs on the IBM SP2 and the SGI power challenge computers. Lecture Notes in Computer Science, 1067:365–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kechadi:1996:PSE

[1122] M. T. Kechadi, P. Kiernan, D. Hegarty, and K. Dawson. Parallel simulation environment for polymers, DNA and protein molecular chains. Lecture Notes in Computer Science, 1067:373–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fet:1996:PPC

[1123] A. I. Fet and L. N. Romanov. Prediction of properties of chemical elements and combinations. Lecture Notes in Computer Science, 1067:379–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marsh:1996:BCS

[1124] A. Marsh and D. I. Kaklamani. A benchmark comparison for suitable HPC platforms to solve electrically large planar conducting structures via a parallel method of moments technique. Lecture Notes in Computer Science, 1067:387–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seifert:1996:CSS

[1125] G. Seifert, T. Heine, O. Knospe, and R. Schmidt. Computer simulations for the structure and dynamics of large molecules, clusters and solids. Lecture Notes in Computer Science, 1067:393– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Serkin:1996:SUN

[1126] V. N. Serkin, T. L. Belyaeva, and E. V.

Samarina. Simulation of ultrafast nonlinear electro-magnetic phenomena on the basis of Maxwell’s equations solutions. Lecture Notes in Computer Science, 1067:402–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cook:1996:MPS

[1127] J. Cook, P. Bjoerstad, and J. Braekhus.

Multilevel parallel solution of large, sparse finite element equations from structural analysis. Lecture Notes in Computer Science, 1067:404–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adorni:1996:BPB

[1128] G. Adorni, P. Ornelli, W. Bellettato, and F. D’Oria. BLOWPAC project: BLOWmolding modelling by PArallel computing techniques. Lecture Notes in Computer Science, 1067:413–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aloisio:1996:MLS

[1129] G. Aloisio, M. A. Bochicchio, and M. La Scala. Metacomputing for online stability analysis in power systems: A proposal. Lecture Notes in Computer Science, 1067:421–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baraglia:1996:MSC

[1130] R. Baraglia, G. Faieta, M. Formica, and D. Laforenza. Metacomputing to solve complex computational problems. A case study: Evolution of galaxies. Lecture Notes in Computer Science, 1067:429–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rankin:1996:EUE

[1131] R. F. Rankin and P. Sage. Experiences using the European ATM Pilot Network. Lecture Notes in Computer Science, 1067:441–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dzwinel:1996:PPO

[1132] D. Dzwinel. Particles paradigm and optimization problems. Lecture Notes in Computer Science, 1067:447–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corno:1996:PGA

[1133] F. Corno, P. Prinetto, M. Rebaudengo, and M. Sonza Reorda. A parallel genetic algorithm for automatic generation of test sequences for digital circuits. Lecture Notes in Computer Science, 1067:454–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DellaRocca:1996:RPS

[1134] A. B. Della Rocca, L. La Porta, and F. Valentinotti. Radiographic process simulation by integration of Boltzmann equation on SIMD architecture (quadries QH4). Lecture Notes in Computer Science, 1067:460–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1996:TTP

[1135] H. X. Lin, A. J. C. Van Gemund, J. Meijdam, and P. Nauta. TGEX: a tool for portable parallel and distributed execution of unstructured problems. Lecture Notes in Computer Science, 1067:467–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bik:1996:ISP

[1136] A. J. C. Bik and H. A. G. Wijshoff.

Iteration space partitioning. Lecture Notes in Computer Science, 1067:475– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ceccolini:1996:TDS

[1137] A. Ceccolini, M. Danelutto, G. Orsini, and S. Pelagatti. A tool for the development of structured parallel applications. Lecture Notes in Computer Science, 1067:485–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pellegrini:1996:SSP

[1138] F. Pellegrini and J. Roman. SCOTCH: A software package for static mapping by dual recursive bipartitioning of process and architecture graphs. Lecture Notes in Computer Science, 1067:493– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chu-Carroll:1996:DIG

[1139] M. Chu-Carroll and L. L. Pollock. Design and implementation of a general purpose parallel programming system. Lecture Notes in Computer Science, 1067:499–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luksch:1996:SEM

[1140] P. Luksch, U. Maier, S. Rathmayer, and M. Weidmann. Software engineering methods for parallel and distributed scientific computing. Lecture Notes in Computer Science, 1067:508– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiGregorio:1996:PCE

[1141] S. Di Gregorio, R. Rongo, W. Spataro, and G. Spezzano. A parallel cellular environment for high performance scientific computing. Lecture Notes in Computer Science, 1067:514–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calzarossa:1996:PPE

[1142] M. Calzarossa, L. Massari, A. Merlo, and D. Tessera. Parallel performance evaluation: The MEDEA tool. Lecture Notes in Computer Science, 1067: 522–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hernandez:1996:PGN

[1143] E. Hernandez. Parallelizing for a good node performance. Lecture Notes in Computer Science, 1067:530–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kempf:1996:SSP

[1144] G. Kempf, A. Van der Steen, C. Caremoli, and W.-Y. Thang. Simulation of scientific programs on parallel architectures. Lecture Notes in Computer Science, 1067:536–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuan:1996:LBP

[1145] X. Yuan, B. He, D. Balsara, and R. Melhem. A load balancing package for domain decomposition on distributed memory systems. Lecture Notes in Computer Science, 1067:547– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeRonde:1996:LBR

[1146] J. F. De Ronde, A. Schoneveld, P. M. A. Sloot, and N. Floros. Load balancing by redundant decomposition and mapping. Lecture Notes in Computer Science, 1067:555–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeDoncker:1996:TML

[1147] E. De Doncker, P. Ealy, and A. Gupta. Two methods for load balanced distributed adaptive integration. Lecture Notes in Computer Science, 1067:562– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiemann:1996:NAL

[1148] B. Schiemann and L. Borrmann. A new approach for load balancing in high performance decision support systems. Lecture Notes in Computer Science, 1067:571–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diekmann:1996:PPP

[1149] R. Diekmann, U. Dralle, F. Neugebauer, and T. Roemke. PadFEM: A portable parallel FEM-Tool. Lecture Notes in Computer Science, 1067:580– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vanderstraeten:1996:HPS

[1150] D. Vanderstraeten, F.-X. Roux, and R. Keunings. A hybrid parallel solver for finite element computations. Lecture Notes in Computer Science, 1067: 586–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanEwijk:1996:FDT

[1151] L. J. Van Ewijk. The finite difference time domain method on a massively parallel computer. Lecture Notes in Computer Science, 1067:593–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Issman:1996:PIC

[1152] E. Issman and G. Degrez. A parallel implicit compressible multidimensional upwind Euler/Navier–Stokes solver on unstructured meshes. Lecture Notes in Computer Science, 1067:599–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brenner:1996:MAD

[1153] G. Brenner, T. Schoenfeld, and N. Mangiavacchi. A multiphysics approach to the direct numerical simulation of turbulent flows on parallel computers. Lecture Notes in Computer Science, 1067:607–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1996:APM

[1154] J. Wu. Adaptive parallel multigrid in computational fluid dynamics. Lecture Notes in Computer Science, 1067:613– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goossens:1996:BIP

[1155] S. Goossens, E. Issman, G. Degrez, and D. Roose. Block ILP0-01U(0) preconditioning for a GMRES based Euler/ Navier–Stokes solver. Lecture Notes in Computer Science, 1067:619–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simon:1996:SPV

[1156] J. Simon and J.-M. Wierum. Sequential performance versus scalability: Optimizing parallel LU-Decomposition. Lecture Notes in Computer Science, 1067:627–??, 1996. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Diderich:1996:STS

[1157] C. G. Diderich and M. Gengler. Solving traveling salesman problems using a parallel synchronized branch and bound algorithm. Lecture Notes in Computer Science, 1067:633–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia:1996:PHS

[1158] D. Garcia, F. Suarez, J. Garcia, and M. Cuervo. A parallel hot strip mill simulator for testing embedded realtime monitoring system. Lecture Notes in Computer Science, 1067:639–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doallo:1996:PSM

[1159] R. Doallo, B. B. Fraguela, J. Touriño, and E. L. Zapata. Parallel sparse modified Gram–Schmidt QR decomposition. Lecture Notes in Computer Science, 1067:646–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dion:1996:ADM

[1160] M. Dion. Affine data mapping with residual communication optimization: Evaluation of heuristics. Lecture Notes in Computer Science, 1067:654–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arbenz:1996:RCS

[1161] P. Arbenz, W. Gander, and M. Oettli. The remote computation system. Lecture Notes in Computer Science, 1067:662–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fornasari:1996:CGA

[1162] N. Fornasari and S. Rovida. Conjugategradients algorithms on a CRAY-T3D. Lecture Notes in Computer Science, 1067:668–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bates:1996:SCM

[1163] C. Bates, I. E. Jelly, I. Forgacs, and E. Szabo. Semantic computability: A mechanism for the description of test data sets for scalable information systems. Lecture Notes in Computer Science, 1067:677–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kukla:1996:ISD

[1164] R. Kukla and J. Kerridge. Intelligent storage devices for scalable information management systems. Lecture Notes in Computer Science, 1067:683–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCann:1996:MTD

[1165] J. A. McCann, A. MacFarlane, and H. M. Liddell. Management tools for distributed interoperable environments. Lecture Notes in Computer Science, 1067:689–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanLiere:1996:CS

[1166] R. Van Liere, J. D. Mulder, and J. J. Van Wijk. Computational steering. Lecture Notes in Computer Science, 1067:696–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DellaRocca:1996:ICQ

[1167] A. B. Della Rocca, S. Ferriani, and L. La Porta. Image convolutions by quadrics Q1. Lecture Notes in Computer Science, 1067:703–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larkin:1996:VLS

[1168] S. Larkin, A. J. Grant, and W. T. Hewitt. ‘Vipar’ libraries to support distribution and processing of visualization datasets. Lecture Notes in Computer Science, 1067:711–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sirbu:1996:BPV

[1169] M. G. Sirbu and D. C. Marinescu. Bond — A parallel virtual environment. Lecture Notes in Computer Science, 1067:722–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernaschi:1996:ECC

[1170] M. Bernaschi, F. Papetti, and G. Iannello. Efficient collective communication operations for parallel industrial codes. Lecture Notes in Computer Science, 1067:729–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czech:1996:RPS

[1171] Z. J. Czech and W. Mikanik. Randomized PRAM simulation using T9000 transputers. Lecture Notes in Computer Science, 1067:736–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Itoh:1996:TRB

[1172] T. Itoh and M. Kohata. Torus-ring-bus connected hierarchic parallel machine. Lecture Notes in Computer Science, 1067:744–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demuynck:1996:ICM

[1173] K. Demuynck, J. Broeckhove, and F. Arickx. The impact of communication mechanisms on the performance in a distributed virtual reality system. Lecture Notes in Computer Science, 1067:755–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuhn:1996:PPB

[1174] W. Kuhn. Performance prediction and benchmarking: Results from the ALP-STONE project. Lecture Notes in Computer Science, 1067:762–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1996:PMR

[1175] J. J. Dongarra, T. Hey, and E. Strohmaier.

PARKBENCH: Methodology, relations and results. Lecture Notes in Computer Science, 1067:770–777, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boulet:1996:EAP

[1176] P. Boulet and T. Brandes. Evaluation of automatic parallelization strategies for HPF compilers. Lecture Notes in Computer Science, 1067:778–791, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boulet:1996:BFL

[1177] P. Boulet. Bouclettes: A Fortran loop parallelizer. Lecture Notes in Computer Science, 1067:784–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nisbet:1996:SCN

[1178] A. P. Nisbet and R. W. Ford. Spinningon-coherency: A new VSM optimisation for write-invalidate. Lecture Notes in Computer Science, 1067:792– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krantz:1996:RFP

[1179] A. T. Krantz, A. Zadroga, S. E. Chodrow, and V. S. Sunderam. An RPC facility for PVM. Lecture Notes in Computer Science, 1067:798–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kranzlmueller:1996:DMP

[1180] D. Kranzlmueller, S. Grabner, and J. Volkert. Debugging massively parallel programs with ATEMPT. Lecture Notes in Computer Science, 1067:806– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Polman:1996:DLS

[1181] M. Polman and M. Van Steen. Design level support for parallel and distributed applications. Lecture Notes in Computer Science, 1067:812–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arbenz:1996:SRP

[1182] P. Arbenz, W. Gander, H. P. Luethi, and U. Von Matt. Sciddle 4.0, or, remote procedure calls in PVM. Lecture Notes in Computer Science, 1067:820– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1996:LFC

[1183] J. J. Dongarra, J. Du Croz, S. Hammarling, and J. Wasniewski. LAPACK for Fortran 90 compiler. Lecture Notes in Computer Science, 1067:826–833, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandes:1996:RHI

[1184] T. Brandes and D. Greco. Realization of an HPF interface to ScaLAPACK with redistributions. Lecture Notes in Computer Science, 1067:834–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fang:1996:SPP

[1185] N. Fang and H. Burkhart. Structured parallel programming using MPI. Lecture Notes in Computer Science, 1067: 840–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caromel:1996:EEP

[1186] D. Caromel, P. Dzwig, R. Kauffman, and H. Liddell. EC++ — EU-ROPA parallel C++. A draft definition. Lecture Notes in Computer Science, 1067:848–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ulrich:1996:OOO

[1187] R. Ulrich and W. Dulz. OSSCAR — object oriented simulation of slotted communication architectures. Lecture Notes in Computer Science, 1067:858– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yau:1996:EEH

[1188] H. W. Yau, A. Leung, W. Furmanski, and G. C. Fox. Exploration of emerging HPCN technologies for Web-Based distributed computing. Lecture Notes in Computer Science, 1067:869–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1996:CTH

[1189] J. J. Dongarra, H. W. Meuer, H. D. Simon, and E. Strohmaier. Changing technologies of HPC. Lecture Notes in Computer Science, 1067:875–879, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strumpen:1996:PHP

[1190] V. Strumpen, B. Ramkumar, T. L. Casavant, and S. M. Reddy. Perspectives for high performance computing in workstation networks. Lecture Notes in Computer Science, 1067:880– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DAcierno:1996:BPL

[1191] A. D’Acierno and S. Palma. The back-propagation learning algorithm on the Meiko CS-2: Two mapping schemes. Lecture Notes in Computer Science, 1067:890–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holthoff:1996:PLS

[1192] H. Holthoff, W. Roensch, H. Bleecke, and B. Eisfeld. Parallelization of large scale industrial aerodynamic applications on the IBM RS/6000 SP. Lecture Notes in Computer Science, 1067: 901–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engstroem:1996:BIN

[1193] K. Engstroem, E. Reitsma, and P. Fritzson. Benchmarking an intelligent network prototype implemented in Parallel Erlang. Lecture Notes in Computer Science, 1067:905–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anand:1996:HPD

[1194] S. S. Anand, D. A. Bell, J. G.

Hughes, and C. M. Shapcott. A highperformance data mining server. Lecture Notes in Computer Science, 1067: 907–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dechamboux:1996:UDS

[1195] P. Dechamboux, D. Hagimont, and M. Lopez. Using a distributed shared memory for implementing efficient information mediators. Lecture Notes in Computer Science, 1067:909–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walter:1996:DIP

[1196] D. Walter and J. Kerridge. The design of the IRISS parallel database machine. Lecture Notes in Computer Science, 1067:913–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feeley:1996:LRP

[1197] M. Feeley. Lazy remote procedure call and its implementation in a parallel variant of C. Lecture Notes in Computer Science, 1068:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ito:1996:EES

[1198] T. Ito. Efficient evaluation strategies for structured concurrency constructs in parallel scheme systems. Lecture Notes in Computer Science, 1068:22– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreau:1996:SPF

[1199] L. Moreau and D. Ribbens. The semantics of pcall and fork in the presence of first-class continuations and sideeffects. Lecture Notes in Computer Science, 1068:53–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halstead:1996:UPP

[1200] R. H. Halstead. Understanding the performance of parallel symbolic programs. Lecture Notes in Computer Science, 1068:81–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamada:1996:DSF

[1201] T. Kamada and A. Yonezawa. A debugging scheme for fine-grain threads on massively parallel processors with a small amount of log information-replay and race detection. Lecture Notes in Computer Science, 1068:108–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cox:1996:DPE

[1202] P. Cox, H. Glaser, and B. Lanaspre.

Distributed prograph: Extended abstract. Lecture Notes in Computer Science, 1068:128–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ueda:1996:ESM

[1203] K. Ueda. Experiences with strong moding in concurrent logic/constraint programming. Lecture Notes in Computer Science, 1068:134–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yelick:1996:PPI

[1204] K. Yelick, C.-P. Wen, S. Chakrabarti, and E. Deprit. Portable parallel irregular applications. Lecture Notes in Computer Science, 1068:157–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raghavachari:1996:ULS

[1205] M. Raghavachari and A. Rogers. Understanding language support for irregular parallelism. Lecture Notes in Computer Science, 1068:174–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kale:1996:EPG

[1206] L. V. Kale, B. H. Richards, and T. D.

Allen. Efficient parallel graph coloring with prioritization. Lecture Notes in Computer Science, 1068:190–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giavitto:1996:GBF

[1207] J. Giavitto, O. Michel, and J. Sansonnet. Group-based fields. Lecture Notes in Computer Science, 1068:209– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikhil:1996:PSC

[1208] R. S. Nikhil. Parallel symbolic computing in Cid. Lecture Notes in Computer Science, 1068:217–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishikawa:1996:OME

[1209] Y. Ishikawa, A. Hori, M. Sato, and M. Matsuda. An overview of MPC++ — extended abstract. Lecture Notes in Computer Science, 1068:243–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen-Tuong:1996:FTR

[1210] A. Nguyen-Tuong, A. G. Grimshaw, and J. F. Karpovich. Fault tolerance via replication in coarse grain data-flow. Lecture Notes in Computer Science, 1068:250–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bal:1996:EPS

[1211] H. Bal, K. Langendoen, and R. Bhoedjang. Experience with parallel symbolic applications in Orca. Lecture Notes in Computer Science, 1068:266–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chikayama:1996:KPP

[1212] T. Chikayama. KLIC: A portable parallel implementation of a concurrent logic programming language. Lecture Notes in Computer Science, 1068:286– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Queinnec:1996:DOD

[1213] C. Queinnec. DMEROON, overview of a distributed class-based causallycoherent data model. Lecture Notes in Computer Science, 1068:297–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piquer:1996:RIT

[1214] J. M. Piquer. A re-implementation of TransPive: Lessons from the experience. Lecture Notes in Computer Science, 1068:310–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeRoure:1996:RDL

[1215] D. De Roure. The role of distributed Lisp in open hypermedia information systems. Lecture Notes in Computer Science, 1068:330–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padget:1996:FVM

[1216] J. Padget. Foundations for a virtual multicomputer — progress report. Lecture Notes in Computer Science, 1068:336–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berthomieu:1996:PCW

[1217] B. Berthomieu. Process calculi at work — an account of the LCS project. Lecture Notes in Computer Science, 1068: 344–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tokoro:1996:ATS

[1218] M. Tokoro. Agents: Towards a society in which humans and computers cohabitate. Lecture Notes in Computer Science, 1069:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arlabosse:1996:AE

[1219] F. Arlabosse. ARCHON and its environment. Lecture Notes in Computer Science, 1069:11–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Werner:1996:WAC

[1220] E. Werner. What ants cannot do. Lecture Notes in Computer Science, 1069: 19–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wooldridge:1996:TTC

[1221] M. Wooldridge and N. Jennings. Towards a theory of cooperative problem solving. Lecture Notes in Computer Science, 1069:40–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ekenberg:1996:THU

[1222] L. Ekenberg, M. Danielson, and M. Boman. A tool for handling uncertain information in multi-agent systems. Lecture Notes in Computer Science, 1069:54–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haegg:1996:STM

[1223] S. Haegg, F. Ygge, R. Gustavsson, and H. Ottosson. DA-Soc: A testbed for modelling distribution automation applications using agent-oriented programming. Lecture Notes in Computer Science, 1069:63–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Masthoff:1996:AMA

[1224] J. Masthoff and R. Van Hoe. AP-PEAL: A multi-agent approach to interactive learning environments. Lecture Notes in Computer Science, 1069: 77–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bijnens:1996:LCC

[1225] S. Bijnens, W. Joosen, and P. Verbaeten. Language constructs for coordination in an agent space. Lecture Notes in Computer Science, 1069:90– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghedira:1996:DAP

[1226] K. Ghedira. A distributed approach to partial constraint satisfaction problems. Lecture Notes in Computer Science, 1069:106–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsubayashi:1996:CSR

[1227] K. Matsubayashi and M. Tokoro. A collaboration strategy for repetitive encounters. Lecture Notes in Computer Science, 1069:123–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boissier:1996:AAS

[1228] O. Boissier and Y. Demazeau. ASIC: An architecture for social and individual control and its application to computer vision. Lecture Notes in Computer Science, 1069:135–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaib-Draa:1996:HMC

[1229] B. Chaib-Draa and P. Levesque. Hierarchical model and communication by signs, signals and symbols in multiagent environments. Lecture Notes in Computer Science, 1069:150–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Falcone:1996:PRS

[1230] R. Falcone and C. Castelfranchi. Plan recognition: from single-agent to multiagent plans. Lecture Notes in Computer Science, 1069:166–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahndel:1996:DNB

[1231] S. Hahndel, F. Fuchs, and P. Levi. Distributed negotiation-based task planning for a flexible manufacturing environment. Lecture Notes in Computer Science, 1069:179–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bussman:1996:MAA

[1232] S. Bussman. A multi-agent approach to dynamic, adaptive scheduling of material flow. Lecture Notes in Computer Science, 1069:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Overgaard:1996:MPA

[1233] L. Overgaard, H. Petersen, and J. Perram. Motion planning for an articulated robot: A multi-agent approach. Lecture Notes in Computer Science, 1069:206–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coppersmith:1996:LER

[1234] Don Coppersmith, Matthew Franklin, Jacques Patarin, and Michael Reiter. Low-exponent RSA with related messages. Lecture Notes in Computer Science, 1070:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700001.
pdf.

Bleichenbacher:1996:GES

[1235] Daniel Bleichenbacher. Generating ElGamal signatures without knowing the secret key. Lecture Notes in Computer Science, 1070:10–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700010.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1070/10700010.
pdf.

Preneel:1996:STM

[1236] Bart Preneel and Paul C. van Oorschot. On the security of two MAC algorithms. Lecture Notes in Computer Science, 1070:19–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700019.
pdf.

Patarin:1996:HFE

[1237] Jacques Patarin. Hidden fields equations (HFE) and isomorphisms of polynomials (IP): Two new families of asymmetric algorithms. Lecture Notes in Computer Science, 1070:33– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700033.
pdf.

Meyer:1996:PKC

[1238] Bernd Meyer and Volker Müller. A public key cryptosystem based on elliptic curves over Z/nZ equivalent to factoring. Lecture Notes in Computer Science, 1070:49–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/1070/10700049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700049.
pdf.

Schwenk:1996:PKE

[1239] Jörg Schwenk and Jörg Eisfeld. Public key encryption and signature schemes based on polynomials over Z_{n}. Lecture Notes in Computer Science, 1070: 60–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700060.
pdf.

Cramer:1996:MAS

[1240] Ronald Cramer, Matthew Franklin, Berry Schoenmakers, and Moti Yung. Multi-authority secret-ballot elections with linear work. Lecture Notes in Computer Science, 1070:72–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700072.
pdf.

Pfitzmann:1996:AF

[1241] Birgit Pfitzmann and Matthias Schunter.

Asymmetric fingerprinting. Lecture Notes in Computer Science, 1070:84– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700084.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1070/10700084.
                                        pdf.

Burmester:1996:HSS

[1242] Mike V. D. Burmester. Homomorphisms of secret sharing schemes: A tool for verifiable signature sharing. Lecture Notes in Computer Science, 1070:96–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1070/10700096.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1070/10700096.
                                        pdf.

Blackburn:1996:EMS

[1243] Simon R. Blackburn, Mike V. D. Burmester, Yvo Desmedt, and Peter R. Wild. Efficient multiplicative sharing schemes. Lecture Notes in Computer Science, 1070:107–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1070/10700107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1070/10700107.
                                        pdf.

Beaver:1996:EOT

[1244] Donald Beaver. Equivocable oblivious transfer. Lecture Notes in Computer Science, 1070:119–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1070/10700119.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
1070/10700119.pdf.

Boyar:1996:SDP

[1245] Joan Boyar and René Peralta. Short discreet proofs. Lecture Notes in Computer Science, 1070:131–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700131.
pdf.

Jakobsson:1996:DVP

[1246] Markus Jakobsson, Kazue Sako, and Russell Impagliazzo. Designated verifier proofs and their applications. Lecture Notes in Computer Science, 1070: 143–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700143.
pdf.

Coppersmith:1996:FSRa

[1247] Don Coppersmith. Finding a small root of a univariate modular equation. Lecture Notes in Computer Science, 1070: 155–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700155.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700155.
pdf.

Hong:1996:NMM

[1248] Seong-Min Hong, Sang-Yeop Oh, and Hyunsoo Yoon. New modular multiplication algorithms for fast modular exponentiation. Lecture Notes in Computer Science, 1070:166–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1070/10700166.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1070/10700166.
                                        pdf.

Coppersmith:1996:FSRb

[1249] Don Coppersmith. Finding a small root of a bivariate integer equation; factoring with high bits known. Lecture Notes in Computer Science, 1070: 178–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1070/10700178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1070/10700178.
                                        pdf.

Stadler:1996:PVS

[1250] Markus A. Stadler. Publicly verifiable secret sharing. Lecture Notes in Computer Science, 1070:190–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1070/10700190.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1070/10700190.
                                        pdf.

Ogata:1996:OSS

[1251] Wakaha Ogata and Kaoru Kurosawa.

Optimum secret sharing scheme secure against cheating. Lecture Notes in Computer Science, 1070:200–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700200.
pdf.

Gibson:1996:SGP

[1252] Keith Gibson. The security of the Gabidulin public key cryptosystem. Lecture Notes in Computer Science, 1070:212–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1070/10700212.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1070/10700212.
pdf.

Knudsen:1996:NLA

[1253] Lars R. Knudsen and Matthew J. B.

Robshaw. Non-linear approximations in linear cryptanalysis. Lecture Notes in Computer Science, 1070:224–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700224.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700224.
pdf.

Knudsen:1996:DSK

[1254] Lars R. Knudsen and Torben P. Pedersen. On the difficulty of software key escrow. Lecture Notes in Computer Science, 1070:237–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700237.
pdf.

Fischer:1996:EPR

[1255] Jean-Bernard Fischer and Jacques Stern. An efficient pseudo-random generator provably as secure as syndrome decoding. Lecture Notes in Computer Science, 1070:245–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700245.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700245.
pdf.

Klapper:1996:ESF

[1256] Andrew Klapper. On the existence of secure feedback registers. Lecture Notes in Computer Science, 1070:256– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700256.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700256.
pdf.

Golic:1996:FLO

[1257] Jovan Dj. Golic. Fast low order approximation of cryptographic functions. Lecture Notes in Computer Science, 1070:268–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1070/10700268.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1070/10700268.
pdf.

Camion:1996:CRF

[1258] Paul Camion and Anne Canteaut.

Construction of t-resilient functions over a finite alphabet. Lecture Notes in Computer Science, 1070:283–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700283.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700283.
pdf.

Zhang:1996:ACN

[1259] Xian-Mo Zhang and Yuliang Zheng.

Auto-correlations and new bounds on the nonlinearity of Boolean functions. Lecture Notes in Computer Science, 1070:294–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1070/10700294.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1070/10700294.
pdf.

Aiello:1996:FBA

[1260] William Aiello and Ramarathnam Venkatesan. Foiling birthday attacks in length-doubling transformations. Lecture Notes in Computer Science, 1070: 307–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1070/10700307.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1070/10700307.
                                        pdf.

Shoup:1996:SKD

[1261] Victor Shoup and Avi Rubin. Session key distribution using smart cards. Lecture Notes in Computer Science, 1070: 321–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1070/10700321.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1070/10700321.
                                        pdf.

vanOorschot:1996:DHK

[1262] Paul C. van Oorschot and Michael J. Wiener. On Diffie–Hellman key agreement with short exponents. Lecture Notes in Computer Science, 1070:332– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1070/10700332.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1070/10700332.
                                        pdf.

Shoup:1996:SPI

[1263] Victor Shoup. On the security of a practical identification scheme. Lecture Notes in Computer Science, 1070: 344–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700344.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700344.
pdf.

Gennaro:1996:RTD

[1264] Rosario Gennaro, Stanislaw Jarecki, Hugo Krawczyk, and Tal Rabin. Robust threshold DSS signatures. Lecture Notes in Computer Science, 1070: 354–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700354.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700354.
pdf.

Damgaard:1996:NCU

[1265] Ivan Bjerre Damgård and Torben Pryds Pedersen. New convertible undeniable signature schemes. Lecture Notes in Computer Science, 1070:372– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1070/10700372.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1070/10700372.
pdf.

Pointcheval:1996:SPS

[1266] David Pointcheval and Jacques Stern. Security proofs for signature schemes. Lecture Notes in Computer Science, 1070:387–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1070/10700387.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1070/10700387.
                                        pdf.

Bellare:1996:ESD

[1267] Mihir Bellare and Phillip Rogaway. The exact security of digital signatures — how to sign with RSA and Rabin. Lecture Notes in Computer Science, 1070:399–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1070/10700399.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1070/10700399.
                                        pdf.

Anonymous:1996:AIa

[1268] Anonymous. Author index. Lecture Notes in Computer Science, 1070:417– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1070/1070indx.pdf.

Andrews:1996:STF

[1269] P. B. Andrews and M. Bishop. On sets, types, fixed points, and checkerboards. Lecture Notes in Computer Science, 1071:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DAgostino:1996:FTM

[1270] M. D’Agostino and D. M. Gabbay.

Fibred tableaux for multi-implication logics. Lecture Notes in Computer Science, 1071:16–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alechina:1996:ISC

[1271] N. Alechina. Interpolation for a sequent calculus of generalized quantifiers. Lecture Notes in Computer Science, 1071:36–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Avellone:1996:ADF

[1272] A. Avellone and M. Ferrari. Almost duplication free tableau calculi for propositional Lax logics. Lecture Notes in Computer Science, 1071:48–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1996:CMV

[1273] M. Baaz and C. G. Fermueller. Combining many-valued and intuitionistic tableaux. Lecture Notes in Computer Science, 1071:65–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becher:1996:CC

[1274] G. Becher. Cyclic connections. Lecture Notes in Computer Science, 1071: 80–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beckert:1996:ITR

[1275] B. Beckert and C. Pape. Incremental theory reasoning methods for semantic tableaux. Lecture Notes in Computer Science, 1071:93–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Billon:1996:DMC

[1276] J.-P Billon. The disconnection method: a confluent integration of unification in the analytic framework. Lecture Notes in Computer Science, 1071:110– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonatti:1996:SCD

[1277] P. A. Bonatti. Sequent calculi for default and autoepistemic logics. Lecture Notes in Computer Science, 1071:127– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bry:1996:MMG

[1278] F. Bry and A. Yahya. Minimal model generation with positive unit hyperresolution tableaux. Lecture Notes in Computer Science, 1071:143–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pith:1996:DMT

[1279] J. Pith and J. Cunningham. Distributed modal theorem proving with KE. Lecture Notes in Computer Science, 1071:160–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demri:1996:STS

[1280] S. Demri. A simple tableau system for the logic of elsewhere. Lecture Notes in Computer Science, 1071:177–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fronhoefer:1996:SCL

[1281] B. Fronhoefer. Situational calculus, linear connection proofs and STRIPSlike planning: an experimental comparison. Lecture Notes in Computer Science, 1071:193–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joinet:1996:SNA

[1282] J.-B. Joinet, H. Schellinx, and L. Tortora De Falco. Strong normalization for all style LK010q. Lecture Notes in Computer Science, 1071:226–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otten:1996:SUU

[1283] J. Otten and C. Kreitz. T-string-unification: unifying prefixes in nonclassical proof methods. Lecture Notes in Computer Science, 1071:244–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mendler:1996:TRI

[1284] M. Mendler. A timing refinement of intuitionistic proofs and its application to the timing analysis of combinational circuits. Lecture Notes in Computer Science, 1071:261–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niemelae:1996:TCM

[1285] I. Niemelae. A tableau calculus for minimal model reasoning. Lecture Notes in Computer Science, 1071:278–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ritter:1996:IFC

[1286] E. Ritter, D. Pym, and L. Wallen. On the intuitionistic force of classical search. Lecture Notes in Computer Science, 1071:295–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voronkov:1996:PSI

[1287] A. Voronkov. Proof search in intuitionistic logic based on constraint satisfaction. Lecture Notes in Computer Science, 1071:312–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OGorman:1996:BTS

[1288] L. O’Gorman. Basic techniques and symbol-level recognition. Lecture Notes in Computer Science, 1072:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cordella:1996:AAP

[1289] L. P. Cordella and A. Marcelli. An alternative approach to the performance evaluation of thinning algorithms for document processing applications. Lecture Notes in Computer Science, 1072:13–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kadonaga:1996:CMD

[1290] T. Kadonaga and K. Abe. Comparison of methods for detecting corner points from digital curves. Lecture Notes in Computer Science, 1072:23–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chhabra:1996:DHL

[1291] A. K. Chhabra, V. Misra, and J. Arias.

Detection of horizontal lines in noisy run length encoded images: The FAST method. Lecture Notes in Computer Science, 1072:35–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roeoesli:1996:AGC

[1292] M. Roeoesli and G. Monagan. Adding geometric constraints to the vectorization of line drawings. Lecture Notes in Computer Science, 1072:49–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hori:1996:QMP

[1293] O. Hori and D. S. Doermann. Quantitative measurement of the performance of raster-to-vector conversion algorithms. Lecture Notes in Computer Science, 1072:57–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turolla:1996:FIE

[1294] E. Turolla, Y. Belaid, and A. Belaid.

Form item extraction based on line searching. Lecture Notes in Computer Science, 1072:69–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Green:1996:MBA

[1295] E. A. Green and M. S. Krishnamoorthy. Model-based analysis of printed tables. Lecture Notes in Computer Science, 1072:80–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agam:1996:MAD

[1296] G. Agam, H. Luo, and I. Dinstein. Morphological approach for dashed lines detection. Lecture Notes in Computer Science, 1072:92–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blostein:1996:GDR

[1297] D. Blostein. General diagramrecognition methodologies. Lecture Notes in Computer Science, 1072:106– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Messmer:1996:ALR

[1298] B. T. Messmer and H. Bunke. Automatic learning and recognition of graphical symbols in engineering drawings. Lecture Notes in Computer Science, 1072:123–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cesarini:1996:HSL

[1299] F. Cesarini, M. Gori, S. Marinai, and G. Soda. A hybrid system for locating and recognizing low level graphic items. Lecture Notes in Computer Science, 1072:135–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McDaniel:1996:AIC

[1300] J. R. McDaniel and J. R. Balmuth.

Automatic interpretation of chemical structure diagrams. Lecture Notes in Computer Science, 1072:148–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DenHartog:1996:KBS

[1301] J. Den Hartog, T. Ten Kate, and J. Gerbrands. Knowledge-based segmentation for automatic map interpretation. Lecture Notes in Computer Science, 1072:159–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1996:ARL

[1302] M.-K. Kim, M.-K. Park, O.-S. Kwon, and Y.-B Kwon. Automatic region labeling of the layered map. Lecture Notes in Computer Science, 1072:179– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myers:1996:VBA

[1303] G. K. Myers, P. G. Mulgaonkar, C.-H. Chen, and J. L. DeCurtins. Verification-based approach for automated text and feature extraction from raster-scanned maps. Lecture Notes in Computer Science, 1072:190–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smeulders:1996:SSD

[1304] A. W. M. Smeulders and T. Ten Kate.

Software system design for paper map conversion. Lecture Notes in Computer Science, 1072:204–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dori:1996:OPB

[1305] D. Dori, Y. Velkovitch, and L. Wenyin. Object-process based segmentation and recognition of ANSI and ISO standard dimensioning texts. Lecture Notes in Computer Science, 1072:212–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1996:CHL

[1306] P. D. Thomas, J. F. Poliakoff, S. M. Razzaq, and R. J. Whitrow. A combined high and low level approach to interpreting scanned engineering drawings. Lecture Notes in Computer Science, 1072:233–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Capellades:1996:FPD

[1307] M. A. Capellades and O. I. Camps. Functional parts detection in engineering drawings: Looking for the screws. Lecture Notes in Computer Science, 1072:246–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomiyama:1996:RSM

[1308] K. Tomiyama and K. Nakaniwa. Reconstruction of 3D solid model from three orthographic views — top-down approach. Lecture Notes in Computer Science, 1072:260–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kong:1996:BPE

[1309] B. Kong, I. T. Phillips, R. M. Haralick, and A. Prasad. A benchmark: Performance evaluation of dashed-line detection algorithms. Lecture Notes in Computer Science, 1072:270–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dori:1996:HWD

[1310] D. Dori, L. Wenyin, and M. Peleg.

How to win a dashed line detection contest. Lecture Notes in Computer Science, 1072:286–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alberich:1996:PAA

[1311] R. Alberich, P. Burmeister, F. Rossello, and G. Valiente. A partial algebras approach to graph transformation. Lecture Notes in Computer Science, 1073:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banach:1996:CAG

[1312] R. Banach. The contractum in algebraic graph rewriting. Lecture Notes in Computer Science, 1073:16–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauderon:1996:CTA

[1313] M. Bauderon. A category-theoretical approach to vertex replacement: The generation of infinite graphs. Lecture Notes in Computer Science, 1073:27– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blostein:1996:IPU

[1314] D. Blostein, H. Fahmy, and A. Grbavec. Issues in the practical use of graph rewriting. Lecture Notes in Computer Science, 1073:38–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1996:CTG

[1315] A. Corradini, H. Ehrig, M. Loewe, and U. Montanari. The category of typed graph grammars and its adjunctions with categories of derivations. Lecture Notes in Computer Science, 1073:56–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plump:1996:GUM

[1316] D. Plump and A. Habel. Graph unification and matching. Lecture Notes in Computer Science, 1073:75–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kreowski:1996:IST

[1317] H.-J. Kreowski and S. Kuske. On the interleaving semantics of transformation units — A step into GRACE. Lecture Notes in Computer Science, 1073: 89–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maggiolo-Schettini:1996:GRF

[1318] A. Maggiolo-Schettini and A. Peron. A graph rewriting framework for statecharts semantics. Lecture Notes in Computer Science, 1073:107–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuerr:1996:PGT

[1319] A. Schuerr. Programmed graph transformations and graph transformation units in GRACE. Lecture Notes in Computer Science, 1073:122–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1996:PSA

[1320] H. Ehrig and G. Engels. Pragmatic and semantic aspects of a module concept for graph transformation systems. Lecture Notes in Computer Science, 1073:137–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagl:1996:SIP

[1321] M. Nagl and A. Schuerr. Software integration problems and coupling of graph grammar specifications. Lecture Notes in Computer Science, 1073:155– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wills:1996:UAF

[1322] L. M. Wills. Using attributed flow graph parsing to recognize cliches in programs. Lecture Notes in Computer Science, 1073:170–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Derk:1996:RGG

[1323] M. D. Derk and L. S. DeBrunner. Reconfiguration graph grammar for massively parallel, fault tolerant computers. Lecture Notes in Computer Science, 1073:185–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drewes:1996:UTT

[1324] F. Drewes. The use of tree transducers to compute translations between graph algebras. Lecture Notes in Computer Science, 1073:196–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skodinis:1996:BDP

[1325] K. Skodinis and E. Wanke. The bounded degree problem for nonobstructing eNCE graph grammars. Lecture Notes in Computer Science, 1073:211–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthelmann:1996:PSV

[1326] K. Barthelmann. Process specification and verification. Lecture Notes in Computer Science, 1073:225–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1996:ESS

[1327] A. Corradini, H. Ehrig, M. Loewe, and U. Montanari. An event structure semantics for graph grammars with parallel productions. Lecture Notes in Computer Science, 1073:240–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1996:SCG

[1328] A. Corradini and F. Rossi. Synchronized composition of graph grammar productions. Lecture Notes in Computer Science, 1073:257–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janssens:1996:DEC

[1329] D. Janssens. The decomposition of ESM computations. Lecture Notes in Computer Science, 1073:271–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korff:1996:FRB

[1330] M. Korff and L. Ribeiro. Formal relationship between graph grammars and Petri nets. Lecture Notes in Computer Science, 1073:288–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taentzer:1996:HDG

[1331] G. Taentzer. Hierarchically distributed graph transformation. Lecture Notes in Computer Science, 1073:304–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Assmann:1996:EAR

[1332] U. Assmann. On edge addition rewrite systems and their relevance to program analysis. Lecture Notes in Computer Science, 1073:321–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandenburg:1996:GAL

[1333] F. J. Brandenburg and K. Skodinis.

Graph automata for linear graph languages. Lecture Notes in Computer Science, 1073:336–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcelle:1996:OMC

[1334] B. Courcelle and G. Senizergues. The obstructions of a minor-closed set of graphs defined by hyperedge replacement can be constructed. Lecture Notes in Computer Science, 1073:351– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1996:CG

[1335] J. Engelfriet and J. J. Vereijken. Concatenation of graphs. Lecture Notes in Computer Science, 1073:368–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1996:HGH

[1336] C. Kim and T. E. Jeong. HRNCE grammars — A hypergraph generating system with an eNCE way of rewriting. Lecture Notes in Computer Science, 1073:383–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klempien-Hinrichs:1996:NRH

[1337] R. Klempien-Hinrichs. Node replacement in hypergraphs: Simulation of hyperedge replacement, and decidability of confluence. Lecture Notes in Computer Science, 1073:397–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dassow:1996:CCP

[1338] J. Dassow, A. Habel, and S. Taubenberger. Chain-code pictures and collages generated by hyperedge replacement. Lecture Notes in Computer Science, 1073:412–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parisi-Presicce:1996:TGG

[1339] F. Parisi-Presicce. Transformations of graph grammars. Lecture Notes in Computer Science, 1073:428–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zinssmeister:1996:DGA

[1340] G. Zinssmeister and C. L. McCreary.

Drawing graphs with attribute graph grammars. Lecture Notes in Computer Science, 1073:443–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zuendorf:1996:GPM

[1341] A. Zuendorf. Graph pattern matching in PROGRES. Lecture Notes in Computer Science, 1073:454–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnborg:1996:TRG

[1342] S. Arnborg and A. Proskurowski. A technique for recognizing graphs of bounded treewidth with application to subclasses of partial 2-Paths. Lecture Notes in Computer Science, 1073:469– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcelle:1996:DMS

[1343] B. Courcelle. The definition in monadic second-order logic of modular decompositions of ordered graphs. Lecture Notes in Computer Science, 1073:487– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrenfeucht:1996:GBG

[1344] A. Ehrenfeucht, T. Harju, and G. Rozenberg. Group based graph transformations and hierarchical representations of graphs. Lecture Notes in Computer Science, 1073:502–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fracchia:1996:ILI

[1345] F. D. Fracchia. Integrating lineage and interaction for the visualization of cellular structures. Lecture Notes in Computer Science, 1073:521–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lueck:1996:CCR

[1346] J. Lueck and H. B. Lueck. Cellworks with cell rewriting and cell packing for plant morphogenesis. Lecture Notes in Computer Science, 1073:536–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prusinkiewicz:1996:SBS

[1347] P. Prusinkiewicz and L. Kari. Subapical bracketed L-systems. Lecture Notes in Computer Science, 1073:550– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klop:1996:TGR

[1348] J. W. Klop. Term graph rewriting. Lecture Notes in Computer Science, 1074: 1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanBakel:1996:ANR

[1349] S. Van Bakel and M. Fernandez. Approximation and normalization results for typeable term rewriting systems. Lecture Notes in Computer Science, 1074:17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1996:MPA

[1350] G. Barthe and H. Geuvers. Modular properties of algebraic type systems. Lecture Notes in Computer Science, 1074:37–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bethke:1996:CPC

[1351] I. Bethke and J. W. Klop. Collapsing partial combinatory algebras. Lecture Notes in Computer Science, 1074: 57–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blom:1996:CPS

[1352] S. Blom. A complete proof system for nested term graphs. Lecture Notes in Computer Science, 1074:74–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hintermeier:1996:RGL

[1353] C. Hintermeier, H. Kirchner, and P. D.

Mosses. R0n- and G0n-logics. Lecture Notes in Computer Science, 1074:90– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahrs:1996:VCP

[1354] S. Kahrs. The variable containment problem. Lecture Notes in Computer Science, 1074:109–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meinke:1996:HOE

[1355] K. Meinke. Higher-order equational logic for specification, simulation and testing. Lecture Notes in Computer Science, 1074:124–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mintchev:1996:CHO

[1356] S. Mintchev and D. Lester. The correctness of a higher-order lazy functional language implementation: An exercise in mechanical theorem proving. Lecture Notes in Computer Science, 1074:144–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moeller:1996:AR

[1357] B. Moeller. Assertions and recursions. Lecture Notes in Computer Science, 1074:163–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanOostrom:1996:DCC

[1358] V. Van Oostrom. Development closed critical pairs. Lecture Notes in Computer Science, 1074:185–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandePol:1996:TDS

[1359] J. Van de Pol. Two different strong normalization proofs? Lecture Notes in Computer Science, 1074:201–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Springintveld:1996:TOM

[1360] J. Springintveld. Third-order matching in the polymorphic lambda calculus. Lecture Notes in Computer Science, 1074:221–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steggles:1996:HOA

[1361] L. J. Steggles. Higher-order algebra with transfinite types. Lecture Notes in Computer Science, 1074:238–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:AHC

[1362] L.-G. Wang and M. Mendler. Abstraction of hardware construction. Lecture Notes in Computer Science, 1074:264– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeza-Yates:1996:FAA

[1363] R. Baeza-Yates and G. Navarro. A faster algorithm for approximate string matching. Lecture Notes in Computer Science, 1075:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

El-Mabrouk:1996:BMS

[1364] N. El-Mabrouk and M. Crochemore.

Boyer–Moore strategy to efficient approximate string matching. Lecture Notes in Computer Science, 1075:24– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasieniec:1996:REA

[1365] L. Gasieniec, M. Karpinski, W. Plandowski, and W. Rytter. Randomized efficient algorithms for compressed strings: The finger-print approach. Lecture Notes in Computer Science, 1075:39–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutinen:1996:FQS

[1366] E. Sutinen and J. Tarhio. Filtration with q-samples in approximate string matching. Lecture Notes in Computer Science, 1075:50–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lipman:1996:CDM

[1367] D. J. Lipman. Computing discoveries in molecular biology. Lecture Notes in Computer Science, 1075:64–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

StoeltingBrodal:1996:ADQ

[1368] G. Stoelting Brodal and L. Gasieniec. Approximate dictionary queries. Lecture Notes in Computer Science, 1075: 65–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muth:1996:AMS

[1369] R. Muth and U. Manber. Approximate multiple string search. Lecture Notes in Computer Science, 1075:75–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armen:1996:AAS

[1370] C. Armen and C. Stein. A 2 2/3-Approximation algorithm for the shortest superstring problem. Lecture Notes in Computer Science, 1075:87– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersson:1996:STW

[1371] A. Andersson, N. J. Larsson, and K. Swanson. Suffix trees on words. Lecture Notes in Computer Science, 1075: 102–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breslauer:1996:STT

[1372] D. Breslauer. The suffix tree of a tree and minimizing sequential transducers. Lecture Notes in Computer Science, 1075:116–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farach:1996:PHS

[1373] M. Farach and S. Muthukrishnan. Perfect hashing for strings: Forrealization and algorithms. Lecture Notes in Computer Science, 1075:130–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gelfand:1996:SAN

[1374] M. S. Gelfand, A. A. Mironov, and P. A. Pevzner. Spliced alignment: A new approach to gene recognition. Lecture Notes in Computer Science, 1075: 141–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferretti:1996:OS

[1375] V. Ferretti, J. H. Nadeau, and D. Sankoff. Original synteny. Lecture Notes in Computer Science, 1075:159– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berman:1996:FSR

[1376] P. Berman and S. Hannenhalli. Fast sorting by reversal. Lecture Notes in Computer Science, 1075:168–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sagot:1996:DCA

[1377] M.-F. Sagot and A. Viari. A double combinatorial approach to discovering patterns in biological sequences. Lecture Notes in Computer Science, 1075: 186–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arratia:1996:PPA

[1378] R. Arratia and G. Reinert. Poisson process approximation for repeats in one sequence and its application to sequencing by hybridization. Lecture Notes in Computer Science, 1075:209– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:IAA

[1379] L. Wang and D. Gusfield. Improved approximation algorithms for tree alignment. Lecture Notes in Computer Science, 1075:220–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Phillips:1996:AMT

[1380] C. A. Phillips and T. J. Warnow. The asymmetric median tree — A new model for building consensus trees. Lecture Notes in Computer Science, 1075: 234–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldberg:1996:CCV

[1381] L. A. Goldberg, P. W. Goldberg, C. A. Phillips, and G. B. Sorkin. Constructing computer virus phylogenies. Lecture Notes in Computer Science, 1075: 253–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sandak:1996:DCF

[1382] B. Sandak, R. Nussinov, and H. J.

Wolfson. Docking of conformationally flexible proteins. Lecture Notes in Computer Science, 1075:271–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hart:1996:IPC

[1383] W. E. Hart and S. Istrail. Invariant patterns in crystal lattices: Implications for protein folding algorithms. Lecture Notes in Computer Science, 1075:288–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gusfield:1996:GTG

[1384] D. Gusfield, R. Karp, L. Wang, and P. Stelling. Graph traversals, genes, and matroids: An efficient case of the travelling salesman problem. Lecture Notes in Computer Science, 1075:304– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amir:1996:AID

[1385] A. Amir and G. Calinescu. Alphabet independent and dictionary scaled matching. Lecture Notes in Computer Science, 1075:320–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1996:ATD

[1386] K. Park. Analysis of two-dimensional approximate pattern matching algorithms. Lecture Notes in Computer Science, 1075:335–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arikati:1996:AAM

[1387] S. R. Arikati, A. Dessmark, A. Lingas, and M. Marathe. Approximation algorithms for maximum two-dimensional pattern matching. Lecture Notes in Computer Science, 1075:348–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1996:EPA

[1388] K. Zhang. Efficient parallel algorithms for tree editing problems. Lecture Notes in Computer Science, 1075:361– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fu:1996:APM

[1389] J. J. Fu. Approximate pattern matching in directed graphs. Lecture Notes in Computer Science, 1075:373–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1996:FSC

[1390] H. L. Bodlaender, P. A. Evans, and M. R. Fellows. Finite-state computability of annotations of strings and trees. Lecture Notes in Computer Science, 1075:384–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benjamins:1996:APS

[1391] R. Benjamins and C. Pierret-Golbreich. Assumptions of problem-solving methods. Lecture Notes in Computer Science, 1076:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fensel:1996:PSM

[1392] D. Fensel and R. Straatman. Problemsolving methods: Making assumptions for efficiency reasons. Lecture Notes in Computer Science, 1076:17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OHara:1996:TEW

[1393] K. O’Hara and N. Shadbolt. The thin end of the wedge: Efficiency and the generalised directive model methodology. Lecture Notes in Computer Science, 1076:33–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orsvaern:1996:PLT

[1394] K. Orsvaern. Principles for libraries of task decomposition methods — conclusions from a case-study. Lecture Notes in Computer Science, 1076:48– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orsvaern:1996:PDM

[1395] K. Orsvaern. A purpose driven method for language comparison: The REVISE Project. Lecture Notes in Computer Science, 1076:66–81, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benjamins:1996:CFM

[1396] R. Benjamins and M. Aben. A conceptual and formal model of a diagnostic reasoner. Lecture Notes in Computer Science, 1076:82–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benjamin:1996:OCT

[1397] J. Benjamin, P. Borst, H. Akkermans, and B. Wielinga. Ontology construction for technical domains. Lecture Notes in Computer Science, 1076:98– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lapalut:1996:TCH

[1398] S. Lapalut. Text clustering to help knowledge acquisition from documents. Lecture Notes in Computer Science, 1076:115–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahn:1996:QBT

[1399] U. Hahn, M. Klenner, and K. Schnattinger. A quality-based terminological reasoning model for text knowledge acquisition. Lecture Notes in Computer Science, 1076:131–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bowden:1996:ECK

[1400] P. R. Bowden, P. Halstead, and T. G. Rose. Extracting conceptual knowledge from text using explicit relation markers. Lecture Notes in Computer Science, 1076:147–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

GhediniRalha:1996:SID

[1401] C. Ghedini Ralha. Structuring information in a distributed hypermedia system. Lecture Notes in Computer Science, 1076:163–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1996:DKA

[1402] P. C.-H. Cheng. Diagrammatic knowledge acquisition: Elicitation, analysis and issues. Lecture Notes in Computer Science, 1076:179–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sommer:1996:AMT

[1403] E. Sommer. An approach to measuring theory quality. Lecture Notes in Computer Science, 1076:195–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmalhofer:1996:SLB

[1404] F. Schmalhofer and C. Kozieja. Some late-breaking news from the data mines and a preview of the KOALA system: A prospector’s report. Lecture Notes in Computer Science, 1076:213–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

FurtadoVasco:1996:KAT

[1405] J. J. Furtado Vasco, C. Faucher, and E. Chouraqui. A knowledge acquisition tool for multi-perspective concept formation. Lecture Notes in Computer Science, 1076:229–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cupit:1996:KDD

[1406] J. Cupit and N. Shadbolt. Knowledge discovery in databases: Exploiting knowledge-level redescription. Lecture Notes in Computer Science, 1076: 245–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sleeman:1996:TPK

[1407] D. Sleeman and F. Mitchell. Towards painless knowledge acquisition. Lecture Notes in Computer Science, 1076:262– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brazier:1996:AST

[1408] F. Brazier, J. Treur, and N. Wijngaards. The acquisition of a shared task model. Lecture Notes in Computer Science, 1076:278–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boy:1996:GEM

[1409] G. Boy. The group elicitation method: An introduction. Lecture Notes in Computer Science, 1076:290–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Efstathiou:1996:FRS

[1410] J. Efstathiou. Formalising the repair of schedules through knowledge acquisition. Lecture Notes in Computer Science, 1076:306–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chien:1996:ITP

[1411] S. A. Chien. Intelligent tools for planning knowledge base development and verification. Lecture Notes in Computer Science, 1076:321–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arlanzon:1996:CSR

[1412] V. Arlanzon, A. Bernaras, and I. Laresgoiti. Configuring service recovery planning with the CommonKADS library. Lecture Notes in Computer Science, 1076:338–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cottam:1996:DSI

[1413] H. Cottam and N. Shadbolt. Domain and system influences in problem solving models for planning. Lecture Notes in Computer Science, 1076: 354–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haake:1996:UHS

[1414] J. M. Haake, A. Bapat, and T. Knopik.

Using a hypermedia system for systems engineering. Lecture Notes in Computer Science, 1077:3–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurer:1996:DLE

[1415] H. Maurer and N. Scherbakov. Document linking and embedding: A new hypermedia structuring paradigm. Lecture Notes in Computer Science, 1077: 17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrews:1996:HST

[1416] K. Andrews and M. Pichler. Hooking up 3-space: Three-dimensional models as fully-fledged hypermedia documents. Lecture Notes in Computer Science, 1077:28–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smeaton:1996:DMQ

[1417] A. F. Smeaton and A. Gregan. Distributed multimedia QoS parameters from presentation modelling by coloured Petri nets. Lecture Notes in Computer Science, 1077:47–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jin:1996:IPM

[1418] J. S. Jin. Image processing in multimedia. Lecture Notes in Computer Science, 1077:61–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dormann:1996:DLA

[1419] C. Dormann. Designing on-line animated help for multimedia applications. Lecture Notes in Computer Science, 1077:73–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dede:1996:DVW

[1420] C. J. Dede, M. Salzman, and R. B. Loftin. The development of a virtual world for learning Newtonian mechanics. Lecture Notes in Computer Science, 1077:87–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hand:1996:OFV

[1421] C. Hand. Other faces of virtual reality. Lecture Notes in Computer Science, 1077:107–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pose:1996:TRV

[1422] R. Pose and M. Regan. Techniques for reducing virtual reality latency with architectural support and consideration of human factors. Lecture Notes in Computer Science, 1077:117–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tegolo:1996:FVT

[1423] D. Tegolo. Fusion of visual tools in virtual spaces. Lecture Notes in Computer Science, 1077:136–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Streitz:1996:IWD

[1424] N. A. Streitz. From individual work and desktop-based collaboration to ubiquitous meeting environments. Lecture Notes in Computer Science, 1077: 149–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miah:1996:MCA

[1425] T. Miah. Multimedia conferencing architecture and user interfaces for collaborative environment. Lecture Notes in Computer Science, 1077:164–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gershman:1996:MCG

[1426] A. V. Gershman and S. H. Sato. Multimedia communications and groupware. Lecture Notes in Computer Science, 1077:178–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kommers:1996:CSN

[1427] P. Kommers. Conceptual support by the new media for co-operative learning in the next century. Lecture Notes in Computer Science, 1077:193–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyle:1996:HEL

[1428] T. Boyle and M. Davies. Hypermedia environments for learning to program. Lecture Notes in Computer Science, 1077:216–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Millwood:1996:EMH

[1429] R. Millwood and G. Mladenova. Educational multimedia: How to allow for cultural factors. Lecture Notes in Computer Science, 1077:230–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petrushin:1996:HSC

[1430] V. A. Petrushin. HELENA: A shell for creating educational hypermedia applications. Lecture Notes in Computer Science, 1077:237–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morozov:1996:MLR

[1431] M. Morozov. Multimedia lecture room: A new tool for education. Lecture Notes in Computer Science, 1077:246– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egov:1996:PIS

[1432] A. S. Egov, E. I. Podolny, A. V.

Polykakhin, and A. V. Polishjuk. A personal information support system for performing creative tasks. Lecture Notes in Computer Science, 1077:257– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waterworth:1996:PIE

[1433] J. A. Waterworth. A pattern of islands:

Exploring public information space in a private vehicle. Lecture Notes in Computer Science, 1077:265–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Subbotin:1996:CPB

[1434] M. Subbotin and D. Subbotin. The concept of a “personal book”. Lecture Notes in Computer Science, 1077:279– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brusilovsky:1996:AHA

[1435] P. Brusilovsky. Adaptive hypermedia: An attempt to analyze and generalize. Lecture Notes in Computer Science, 1077:288–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zyryanov:1996:ALM

[1436] M. I. Zyryanov. Adaptive local maps in hypermedia components of intelligent learning environments. Lecture Notes in Computer Science, 1077:305– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soni:1996:MFS

[1437] D. Soni, R. L. Nord, L. Hsu, and P. J. Drongowski. Many faces of software architecture. Lecture Notes in Computer Science, 1078:6–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shaw:1996:PCA

[1438] M. Shaw. Procedure calls are the assembly language of software interconnection: Connectors deserve first-class status. Lecture Notes in Computer Science, 1078:17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dean:1996:SSC

[1439] T. R. Dean and J. R. Cordy. Software structure characterization using connectivity. Lecture Notes in Computer Science, 1078:33–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minsky:1996:RSS

[1440] N. H. Minsky. Regularities in software systems. Lecture Notes in Computer Science, 1078:49–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garlan:1996:FAS

[1441] D. Garlan. Formal approaches to software architecture. Lecture Notes in Computer Science, 1078:64–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murphy:1996:DOO

[1442] G. C. Murphy and D. Notkin. Difficulties with object-oriented frameworks. Lecture Notes in Computer Science, 1078:77–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sorenson:1996:UMA

[1443] P. G. Sorenson and J.-P. Tremblay. Using a metasystem approach to support and study the design process. Lecture Notes in Computer Science, 1078:88– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ryman:1996:CSD

[1444] A. Ryman. Constructing software design theories and models. Lecture Notes in Computer Science, 1078:103– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kramer:1996:MIS

[1445] J. Kramer, A. Finkelstein, and B. Nuseibeh. Method integration and support for distributed software development: An overview. Lecture Notes in Computer Science, 1078:115–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griswold:1996:PRD

[1446] W. G. Griswold and R. W. Bowdidge.

Program restructuring via design-level manipulation. Lecture Notes in Computer Science, 1078:127–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaz-Herrera:1996:HMD

[1447] J. L. Diaz-Herrera. Hierarchical modular diagrams: An approach to describe architectural designs. Lecture Notes in Computer Science, 1078:140–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hinchey:1996:SFM

[1448] M. G. Hinchey. Structured and formal methods: An investigative framework. Lecture Notes in Computer Science, 1078:151–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuan:1996:EOO

[1449] G. Yuan and N. Patel. An exploration of object-oriented methodologies for system analysis and design. Lecture Notes in Computer Science, 1078:164–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodie:1996:POW

[1450] M. Brodie. Putting objects to work on a massive scale. Lecture Notes in Computer Science, 1079:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1996:ACR

[1451] D. Dubois and H. Prade. Approximate and commonsense reasoning: From theory to practice. Lecture Notes in Computer Science, 1079:19–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jarke:1996:CIS

[1452] M. Jarke, M. Jeusfeld, P. Peters, and P. Szczurko. Cooperative information systems engineering. Lecture Notes in Computer Science, 1079:34–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sandewall:1996:TWW

[1453] E. Sandewall. Towards a world-wide knowledge base. Lecture Notes in Computer Science, 1079:50–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piatetsky-Shapiro:1996:DMK

[1454] G. Piatetsky-Shapiro. Data mining and knowledge discovery in business databases. Lecture Notes in Computer Science, 1079:56–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lambrix:1996:LCC

[1455] P. Lambrix and J. Maleki. Learning composite concepts in description logics: a first step. Lecture Notes in Computer Science, 1079:68–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dieng:1996:CCG

[1456] R. Dieng. Comparison of conceptual graphs for modeling knowledge of multiple experts. Lecture Notes in Computer Science, 1079:78–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1996:SCK

[1457] Y. Zhang. Semantical considerations for knowledge base updates. Lecture Notes in Computer Science, 1079:88– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wrzos-Kaminska:1996:PEC

[1458] A. Wrzos-Kaminska. Partial evaluation in constraint logic programming. Lecture Notes in Computer Science, 1079:98–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloedorn:1996:ADS

[1459] E. Bloedorn and R. Michalski. The AQ17-DCI system for data-driven constructive induction and its application to the analysis of world economics. Lecture Notes in Computer Science, 1079:108–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shan:1996:ICR

[1460] N. Shan, H. Hamilton, and N. Cercone. Induction of classification rules from imperfect data. Lecture Notes in Computer Science, 1079:118–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsumoto:1996:IES

[1461] S. Tsumoto and H. Tanaka. Induction of expert system rules from databases based on rough set theory and resampling methods. Lecture Notes in Computer Science, 1079:128–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zytkow:1996:MPE

[1462] J. Zytkow and R. Zembowicz. Mining patterns at each scale in massive data. Lecture Notes in Computer Science, 1079:139–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeJong:1996:EI

[1463] K. De Jong. On evolving intelligence. Lecture Notes in Computer Science, 1079:149–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeck:1996:IMR

[1464] T. Baeck and M. Schuetz. Intelligent mutation rate control in canonical genetic algorithms. Lecture Notes in Computer Science, 1079:158–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giordana:1996:FGP

[1465] A. Giordana, F. Neri, and L. Saitta. A fine-grained parallel evolutionary program for concept induction. Lecture Notes in Computer Science, 1079:168– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiben:1996:EES

[1466] A. E. Eiben. Evolutionary exploration of search spaces. Lecture Notes in Computer Science, 1079:178–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michalewicz:1996:ECO

[1467] Z. Michalewicz, J. Xiao, and K. Trojanowski. Evolutionary computation: one project, many directions. Lecture Notes in Computer Science, 1079:189– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calmet:1996:SFL

[1468] J. Calmet, I. Lu, M. Rodriguez, and J. Schue. Signed formula logic programming: operational semantics and applications. Lecture Notes in Computer Science, 1079:202–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertossi:1996:API

[1469] L. Bertossi, J. Pinto, P. Saez, and D. Kapur. Automating proofs of integrity constraints in situation calculus. Lecture Notes in Computer Science, 1079:212–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cholewinski:1996:TPD

[1470] P. Cholewinski. Towards programming in default logic. Lecture Notes in Computer Science, 1079:223–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:SCF

[1471] J. Chen and S. Kundu. A sound and complete fuzzy logic system using Zadeh’s implication operator. Lecture Notes in Computer Science, 1079:233– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dignum:1996:MDF

[1472] F. Dignum, H. Weigand, and E. Verharen. Meeting the deadline: on the formal specification of temporal Deontic constraints. Lecture Notes in Computer Science, 1079:243–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demolombe:1996:VQC

[1473] R. Demolombe. Validity queries and completeness queries. Lecture Notes in Computer Science, 1079:253–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minock:1996:ECI

[1474] M. J. Minock and W. W. Chu. Explanation for cooperative information systems. Lecture Notes in Computer Science, 1079:264–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oldham:1996:TIR

[1475] J. Oldham and V. Marek. Toward intelligent representation of database content. Lecture Notes in Computer Science, 1079:274–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kryszkiewicz:1996:RIS

[1476] M. Kryszkiewicz and H. Rybinski. Reducing information systems with uncertain attributes. Lecture Notes in Computer Science, 1079:285–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dery:1996:ODO

[1477] A.-M. Dery, S. Ducasse, and M. Fornarino. Object and dependency oriented programming in FLO. Lecture Notes in Computer Science, 1079:295–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debenham:1996:KS

[1478] J. Debenham. Knowledge simplification. Lecture Notes in Computer Science, 1079:305–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruening:1996:MBA

[1479] S. Bruening and T. Schaub. A model-based approach to consistencychecking. Lecture Notes in Computer Science, 1079:315–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brusoni:1996:RBV

[1480] V. Brusoni, L. Console, E. Lamma, and P. Mello. Resource-based vs taskbased approaches for scheduling problems. Lecture Notes in Computer Science, 1079:325–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pin:1996:FBA

[1481] F. G. Pin. A fuzzy behaviorist approach to sensor-based robot control. Lecture Notes in Computer Science, 1079:335–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sztandera:1996:KBF

[1482] L. Sztandera. Knowledge-based fuzzy neural networks. Lecture Notes in Computer Science, 1079:346–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seredynski:1996:CGT

[1483] F. Seredynski. Coevolutionary game theoretic multi-agent systems. Lecture Notes in Computer Science, 1079:356– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Son:1996:SFD

[1484] N. Son, N. Hoa, and A. Skowron. Searching for features defined by hyperplanes. Lecture Notes in Computer Science, 1079:366–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blockeel:1996:IDD

[1485] H. Blockeel and L. De Raedt. Inductive database design. Lecture Notes in Computer Science, 1079:376–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loo:1996:EQP

[1486] G. S. Loo, T. Dillon, J. Zeleznikow, and K.-H Lee. Enhancing query processing of information systems. Lecture Notes in Computer Science, 1079: 386–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zarri:1996:SRC

[1487] G. P. Zarri and L. Gilardoni. Structuring and retrieval of the complex predicate arguments proper to the NKRL conceptual language. Lecture Notes in Computer Science, 1079:398–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosc:1996:HID

[1488] P. Bosc and O. Pivert. On the handling of imperfect data in relational database systems — from null values to possibility distributions. Lecture Notes in Computer Science, 1079:408–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wierzchon:1996:MCV

[1489] S. Wierzchon and M. Klopotek. Modified component valuations in valuation based systems as a way to optimize query processing. Lecture Notes in Computer Science, 1079:418–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imam:1996:LDM

[1490] I. F. Imam and R. S. Michalski. Learning for decision making: the FRD approach and a comparative study. Lecture Notes in Computer Science, 1079: 428–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsumoto:1996:ARS

[1491] S. Tsumoto and W. Ziarko. The application of rough sets-based data mining technique to differential diagnosis of meningoenchephalitis. Lecture Notes in Computer Science, 1079:438–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mollestad:1996:RSF

[1492] T. Mollestad and A. Skowron. A rough set framework for data mining of propositional default rules. Lecture Notes in Computer Science, 1079:448– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imam:1996:ESI

[1493] I. F. Imam. An empirical study on the incompetence of attribute selection criteria. Lecture Notes in Computer Science, 1079:458–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esposito:1996:LFP

[1494] F. Esposito, A. Laterza, D. Malerba, and G. Semeraro. Locally finite, proper, and complete operators for refining datalog programs. Lecture Notes in Computer Science, 1079:468– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noronha:1996:FFM

[1495] S. J. Noronha and S. Szpakowicz.

Forest fire management with Negoplan. Lecture Notes in Computer Science, 1079:479–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pons:1996:ADF

[1496] O. Pons, J. Medina, J. Cubero, and A. Vila. An architecture for a deductive fuzzy relational database. Lecture Notes in Computer Science, 1079:491– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhong:1996:MSP

[1497] N. Zhong and S. Ohsuga. A multistep process for discovering, managing, and refining strong functional relations hidden in databases. Lecture Notes in Computer Science, 1079:501–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Capobianchi:1996:AMD

[1498] R. Capobianchi, M. Mautref, M. Van Keulen, and H. Balsters. An architecture and methodology for the design and development of technical information systems. Lecture Notes in Computer Science, 1079:511–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doherty:1996:EEC

[1499] P. Doherty, W. Lukaszewics, and A. Szalas. Explaining explanation closure. Lecture Notes in Computer Science, 1079:521–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeRaedt:1996:PLL

[1500] L. De Raedt. PAC-learning logic programs under the closed-world assumption. Lecture Notes in Computer Science, 1079:531–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karlsson:1996:PTC

[1501] L. Karlsson. Planning, truth criteria, and the systematic approach to action and change. Lecture Notes in Computer Science, 1079:541–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franova:1996:AIR

[1502] M. Franova. Automated inductive reasoning as a support of deductive reasoning in a user-independent automation of inductive theorem proving. Lecture Notes in Computer Science, 1079:551–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godfrey:1996:SQO

[1503] P. Godfrey, J. Gryz, and J. Minker. Semantic query optimization for bottomup evaluation. Lecture Notes in Computer Science, 1079:561–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tuerker:1996:DCB

[1504] C. Tuerker, S. Conrad, and G. Saake.

Dynamically changing behavior: an agent-oriented view to modeling intelligent information systems. Lecture Notes in Computer Science, 1079:572– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eusterbrock:1996:MLA

[1505] J. Eusterbrock. A multi-layer architecture for knowledge-based system synthesis. Lecture Notes in Computer Science, 1079:582–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravindranathan:1996:MAM

[1506] M. Ravindranathan and R. Leitch. Mu-RaLi: an architecture for multiple reasoning. Lecture Notes in Computer Science, 1079:593–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cadish:1996:HVI

[1507] B. Cadish and Z. Diskin. Heterogeneous view integration via sketches and equations. Lecture Notes in Computer Science, 1079:603–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dehaspe:1996:DDL

[1508] L. Dehaspe and L. De Raedt. DLAB:

a declarative language bias formalism. Lecture Notes in Computer Science, 1079:613–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloesgen:1996:KDD

[1509] W. Kloesgen. Knowledge discovery in databases and data mining. Lecture Notes in Computer Science, 1079:623– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bala:1996:LNE

[1510] J. Bala, P. Pachowicz, and H. Wechsler. Learning with noise in engineering domains. Lecture Notes in Computer Science, 1079:633–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vrain:1996:HCC

[1511] C. Vrain. Hierarchical conceptual clustering in a first order representation. Lecture Notes in Computer Science, 1079:643–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ziarko:1996:RDD

[1512] W. Ziarko, N. Cercone, and X. Hu. Rule discovery from databases with decision matrices. Lecture Notes in Computer Science, 1079:653–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelly:1996:MFC

[1513] S. Kelly, K. Lyytinen, and M. Rossi. MetaEdit+: A fully configurable multiuser and multi-tool CASE and CAME environment. Lecture Notes in Computer Science, 1080:1–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Si-Said:1996:MCA

[1514] S. Si-Said, C. Rolland, and G. Grosz.

MENTOR: A computer aided requirements engineering environment. Lecture Notes in Computer Science, 1080: 22–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vescoukis:1996:LBF

[1515] V. C. Vescoukis, N. Papaspyrou, and E. Skordalakis. A logic-based framework for reasoning support in software evolution. Lecture Notes in Computer Science, 1080:44–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gal:1996:TTA

[1516] A. Gal, O. Etzion, and A. Segev.

TALE: A temporal active language and execution model. Lecture Notes in Computer Science, 1080:60–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandenAkker:1996:DCD

[1517] J. Van den Akker and A. Siebes.

DEGAS: Capturing dynamics in objects. Lecture Notes in Computer Science, 1080:82–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kappel:1996:RRP

[1518] G. Kappel, S. Rausch-Schott, W. Retschitzegger, and M. Sakkinen. From rules to rule patterns. Lecture Notes in Computer Science, 1080:99–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kudrass:1996:TBR

[1519] T. Kudrass, M. Lehmbach, and A. Buchmann. Tool-based reengineering of a legacy MIS: An experience report. Lecture Notes in Computer Science, 1080:116–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krone:1996:DIS

[1520] M. Krone, M. Kowsari, P. Hartel, and G. Denker. Developing an information system using TROLL: An application field study. Lecture Notes in Computer Science, 1080:136–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kemper:1996:ARC

[1521] F. Kemper, W. Wilkes, and G. Schlageter. Active relationships for controlling the propagation of information and activities in databases. Lecture Notes in Computer Science, 1080:160–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicolle:1996:MDM

[1522] C. Nicolle, D. Benslimane, and K. Yetongnon. Multi-data models translations in interoperable information systems. Lecture Notes in Computer Science, 1080:176–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weigand:1996:ITB

[1523] H. Weigand, E. Verharen, and F. Dignum. Interoperable transactions in business models: A structured approach. Lecture Notes in Computer Science, 1080:193–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tryfona:1996:IMS

[1524] N. Tryfona and J. Sharma. On information modeling to support interoperable spatial databases. Lecture Notes in Computer Science, 1080:210–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bates:1996:GFF

[1525] B. W. Bates, J.-M. Bruel, R. B. France, and M. M. Larrondo-Petrie. Guidelines for formalizing fusion object-oriented analysis models. Lecture Notes in Computer Science, 1080:222–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castano:1996:TPA

[1526] S. Castano and V. De Antonellis. Techniques for process analysis and unification. Lecture Notes in Computer Science, 1080:234–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costal:1996:HCM

[1527] D. Costal, E. Teniente, T. Urpi, and C. Farre. Handling conceptual model validation by planning. Lecture Notes in Computer Science, 1080:255–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hainaut:1996:DDR

[1528] J.-L. Hainaut, J. Henrard, J.-M. Hick, and D. Roland. Database design recovery. Lecture Notes in Computer Science, 1080:272–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanneghan:1996:WWW

[1529] M. Hanneghan, M. Merabti, and G. Colquhoun. The World-Wide Web as a platform for supporting interactive concurrent engineering. Lecture Notes in Computer Science, 1080:301– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Constantopoulos:1996:CDI

[1530] P. Constantopoulos and Y. Tzitzikas. Context-driven information base update. Lecture Notes in Computer Science, 1080:319–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puustjaervi:1996:MOT

[1531] J. Puustjaervi, H. Tirri, and J. Veijalainen. Managing overlapping transactional work flows. Lecture Notes in Computer Science, 1080:345–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hori:1996:SDM

[1532] M. Hori, Y. Shinoda, and K. Ochimizu. Shared data management mechanism for distributed software development based on a reflective object-oriented model. Lecture Notes in Computer Science, 1080:362–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Norrie:1996:ABT

[1533] M. C. Norrie and M. Wunderli. Agentbased tool integration for distributed information systems. Lecture Notes in Computer Science, 1080:383–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siau:1996:WPN

[1534] K. Siau, Y. Wand, and I. Benbasat.

When parents need not have children: Cognitive biases in information modeling. Lecture Notes in Computer Science, 1080:402–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daneva:1996:RBI

[1535] M. Daneva. The role of benchmarking in information system development. Lecture Notes in Computer Science, 1080:421–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehandjiev:1996:UEO

[1536] N. Mehandjiev and L. Bottaci. Userenhanceability for organizational information systems through visual programming. Lecture Notes in Computer Science, 1080:432–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mirbel:1996:IMD

[1537] I. Mirbel and J.-L Cavarero. An integration method for design schemas. Lecture Notes in Computer Science, 1080:457–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breche:1996:APC

[1538] P. Breche. Advanced primitives for changing schemas of object databases. Lecture Notes in Computer Science, 1080:476–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellahsene:1996:VIS

[1539] Z. Bellahsene, P. Poncelet, and M. Teisseire. Views for information system design without reorganization. Lecture Notes in Computer Science, 1080:496– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Motschnig-Pitrik:1996:SFA

[1540] R. Motschnig-Pitrik and J. Mylopoulos. Semantics, features, and applications of the viewpoint abstraction. Lecture Notes in Computer Science, 1080:514–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lahlou:1996:RIL

[1541] Y. Lahlou and N. Mouaddib. Relaxing the instantiation link: Towards a content-based data model for information retrieval. Lecture Notes in Computer Science, 1080:540–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lammari:1996:DNI

[1542] N. Lammari, R. Laleau, M. Jouve, and X. Castellani. Deriving normalized Isa hierarchies by using applicability constraints. Lecture Notes in Computer Science, 1080:562–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1996:CDI

[1543] J. Anderson and M. Evans. Constraintdirected improvisation for complex domains. Lecture Notes in Computer Science, 1081:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dent:1996:NMH

[1544] M. J. Dent and R. E. Mercer. A new model of hard binary constraint satisfaction problems. Lecture Notes in Computer Science, 1081:14–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wetprasit:1996:RMP

[1545] R. Wetprasit, A. Sattar, and L. Khatib.

Reasoning with multi-point events. Lecture Notes in Computer Science, 1081:26–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allen:1996:SRH

[1546] J. A. Allen and S. Minton. Selecting the right heuristic algorithm: Runtime performance predictors. Lecture Notes in Computer Science, 1081:41–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elkan:1996:RAU

[1547] C. Elkan. Reasoning about unknown, counterfactual, and nondeterministic actions in first-order logic. Lecture Notes in Computer Science, 1081:54– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boutilier:1996:FPB

[1548] C. Boutilier and M. Goldszmidt. The frame problem and Bayesian network action representations. Lecture Notes in Computer Science, 1081:69– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagen:1996:AGC

[1549] E. Hagen and A. Stein. Automatic generation of a complex dialogue history. Lecture Notes in Computer Science, 1081:84–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Popowich:1996:CGS

[1550] F. Popowich. A chart generator for shake and bake machine translation. Lecture Notes in Computer Science, 1081:97–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ardissono:1996:ERU

[1551] L. Ardissono and R. Cohen. Extending the role of user feedback in plan recognition and response generation for advice-giving systems: An initial report. Lecture Notes in Computer Science, 1081:109–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

AltayGuevenir:1996:CBL

[1552] H. Altay Guevenir and A. Tunc. Corpus-based learning of generalized parse tree rules for translation. Lecture Notes in Computer Science, 1081: 121–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahn:1996:PAF

[1553] U. Hahn and M. Strube. PARSETALK about functional anaphora. Lecture Notes in Computer Science, 1081:133– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodner:1996:KBA

[1554] R. C. Bodner and F. Song. Knowledgebased approaches to query expansion in information retrieval. Lecture Notes in Computer Science, 1081:146–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holte:1996:IWU

[1555] R. C. Holte and J. N. Yuen Yan. Inferring what a user is not interested in. Lecture Notes in Computer Science, 1081:159–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kramer:1996:DES

[1556] B. M. Kramer, J. Mylopoulos, M. E.

Benjamin, and Q. B. Chou. Developing an expert system technology for industrial process control: An experience report. Lecture Notes in Computer Science, 1081:172–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charlebois:1996:PLN

[1557] D. Charlebois, D. G. Goodenough, S. Matwin, and A. S. Pal Bhogal. Planning and learning in a natural resource information system. Lecture Notes in Computer Science, 1081:187–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaib-draa:1996:HMA

[1558] B. Chaib-draa. A hierarchical model of agent based on skill, rules, and knowledge. Lecture Notes in Computer Science, 1081:200–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xiang:1996:SMS

[1559] Y. Xiang. Semantics of multiply sectioned Bayesian networks for cooperative multi-agent distributed interpretation. Lecture Notes in Computer Science, 1081:213–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhatia:1996:LSM

[1560] K. Bhatia and C. Elkan. LPMEME: A statistical method for inductive logic programming. Lecture Notes in Computer Science, 1081:227–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rios:1996:EIR

[1561] R. Rios and S. Matwin. Efficient induction of recursive Prolog definitions. Lecture Notes in Computer Science, 1081:240–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1996:CIP

[1562] Y.-J. Hu. Constructive induction: A preprocessor. Lecture Notes in Computer Science, 1081:249–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pendrith:1996:RLR

[1563] M. Pendrith and M. Ryan. Reinforcement learning for real-world control applications. Lecture Notes in Computer Science, 1081:257–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1996:TLA

[1564] W. Lam and S. Mukhopadhyay. A twolevel approach to learning in nonstationary environments. Lecture Notes in Computer Science, 1081:271–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ling:1996:LCM

[1565] C. X. Ling and H. Wang. Learning classifications from multiple sources of unsupervised data. Lecture Notes in Computer Science, 1081:284–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1996:PC

[1566] Z. Lin. Paraconsistent circumscription. Lecture Notes in Computer Science, 1081:296–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Topaloglou:1996:EAQ

[1567] T. Topaloglou. Efficient algorithms for qualitative reasoning about imprecise space. Lecture Notes in Computer Science, 1081:309–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giunchiglia:1996:GPR

[1568] F. Giunchiglia, R. Sebastiani, A. Villafiorita, and T. Walsh. A general purpose reasoner for abstraction. Lecture Notes in Computer Science, 1081:323– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fall:1996:RCI

[1569] A. Fall. Reference constraints and individual level inheritance. Lecture Notes in Computer Science, 1081:336– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koshiba:1996:DTL

[1570] T. Koshiba. Decision tree learning system with switching evaluator. Lecture Notes in Computer Science, 1081:349– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thornton:1996:PPW

[1571] C. Thornton. Parity: The problem that won’t go away. Lecture Notes in Computer Science, 1081:362–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohsawa:1996:PTP

[1572] Y. Ohsawa and M. Ishizuka. A polynomial-time predicate-logic hypothetical reasoning by networked bubble propagation method. Lecture Notes in Computer Science, 1081:375–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wallace:1996:EMS

[1573] R. J. Wallace. Enhancing maximum satisfiability algorithms with pure literal strategies. Lecture Notes in Computer Science, 1081:388–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Culberson:1996:SPD

[1574] J. C. Culberson and J. Schaeffer.

Searching with pattern databases. Lecture Notes in Computer Science, 1081: 402–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noronha:1996:NSL

[1575] S. Noronha and S. Szpakowicz. Negoplan: A system for logic-based decision modelling. Lecture Notes in Computer Science, 1081:417–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barber:1996:ASS

[1576] B. Barber and H. J. Hamilton.

Attribute selection strategies for attribute-oriented generalization. Lecture Notes in Computer Science, 1081: 429–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abu-Hakima:1996:AMA

[1577] S. Abu-Hakima. Automating model acquisition by fault knowledge re-use: Introducing the diagnostic remodeler algorithm. Lecture Notes in Computer Science, 1081:442–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cordova:1996:CNS

[1578] F. Cordova. Communicating NASA’s science to the public. Lecture Notes in Computer Science, 1082:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Billington:1996:LCN

[1579] J. H. Billington. The Library of Congress’s National Digital Library program. Lecture Notes in Computer Science, 1082:17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reddy:1996:ULI

[1580] R. Reddy. The universal library: Intelligent agents and information on demand. Lecture Notes in Computer Science, 1082:27–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smarr:1996:BSA

[1581] L. Smarr. Building a scalable America. Lecture Notes in Computer Science, 1082:35–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolf:1996:VSM

[1582] W. Wolf, B. Liu, A. Wolfe, and M. Yeung. Video as scholarly material in the digital library. Lecture Notes in Computer Science, 1082:45–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shepherd:1996:DLE

[1583] M. A. Shepherd, C. R. Watters, and F. J. Burkowski. Digital libraries for electronic news. Lecture Notes in Computer Science, 1082:55–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desai:1996:WVW

[1584] B. C. Desai and S. Swiercz. Web-Journal: Visualization of a Web journey. Lecture Notes in Computer Science, 1082:63–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nica:1996:USD

[1585] A. Nica and E. A. Rundensteiner.

Uniform structured document handling using a constraint-based object approach. Lecture Notes in Computer Science, 1082:83–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boisvert:1996:DSD

[1586] R. Boisvert, S. Browne, J. Dongarra, and E. Grosse. Digital software and data repositories for support of scientific computing. Lecture Notes in Computer Science, 1082:103–114, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiesener:1996:SHR

[1587] S. Wiesener, W. Kowarschick, P. Vogel, and R. Bayer. Semantic hypermedia retrieval in digital libraries. Lecture Notes in Computer Science, 1082: 115–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myka:1996:FFT

[1588] A. Myka and U. Guentzer. Fuzzy fulltext searches in OCR databases. Lecture Notes in Computer Science, 1082: 131–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouguettaya:1996:DDL

[1589] A. Bouguettaya and S. Milliner. Data discovery in large scale heterogeneous and autonomous databases. Lecture Notes in Computer Science, 1082:149– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rorvig:1996:IAK

[1590] M. E. Rorvig, M. W. Hutchison, R. O. Shelton, and S. L. Smith. An intelligent agent for the K–12 educational community. Lecture Notes in Computer Science, 1082:167–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allen:1996:III

[1591] R. B. Allen. Interface issues for interactive multimedia documents. Lecture Notes in Computer Science, 1082:179– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gershon:1996:SDR

[1592] N. Gershon, W. Ruh, J. Le Vasseur, and J. Winstead. Searching and discovery of resources in digital libraries. Lecture Notes in Computer Science, 1082:191–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choy:1996:ADD

[1593] D. M. Choy, R. Dievendorff, C. Dwork, and J. B. Lotspiech. The Almaden Distributed Digital Library System. Lecture Notes in Computer Science, 1082: 203–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1996:ADL

[1594] C. Fischer, J. Frew, M. Larsgaard, and T. R. Smith. Alexandria Digital Library: Rapid prototype and metadata schema. Lecture Notes in Computer Science, 1082:221–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhao:1996:EEL

[1595] D. G. Zhao and A. Ramsden. The ELINOR electronic library. Lecture Notes in Computer Science, 1082:243– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davis:1996:DBP

[1596] J. R. Davis, C. Lagoze, and D. B. Kraft.

Dienst: Building a production technical report server. Lecture Notes in Computer Science, 1082:259–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barany:1996:CLP

[1597] I. Barany and S. Onn. Colourful linear programming. Lecture Notes in Computer Science, 1084:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1996:TSI

[1598] R. R. Thomas and R. Weismantel. Test sets and inequalities for integer programs. Lecture Notes in Computer Science, 1084:16–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roessner:1996:OSC

[1599] C. Roessner and C. P. Schnorr. An optimal, stable continued fraction algorithm for arbitrary dimension. Lecture Notes in Computer Science, 1084: 31–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chopra:1996:AEF

[1600] S. Chopra, I. Gilboa, and S. T. Sastry. Algorithms and extended formulations for one and two facility network design. Lecture Notes in Computer Science, 1084:44–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barnhart:1996:IMF

[1601] C. Barnhart, C. A. Hane, and P. H. Vance. Integer multicommodity flow problems. Lecture Notes in Computer Science, 1084:58–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caprara:1996:HAS

[1602] A. Caprara, M. Fischetti, and P. Toth. A heuristic algorithm for the set covering problem. Lecture Notes in Computer Science, 1084:72–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tseng:1996:RMG

[1603] P. Tseng and D. P. Bertsekas. An ??- relaxation method for generalized separable convex cost network flow problems. Lecture Notes in Computer Science, 1084:85–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolliopoulos:1996:FRV

[1604] S. G. Kolliopoulos and C. Stein. Finding real-valued single-source shortest paths in o(n03) expected time. Lecture Notes in Computer Science, 1084: 94–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fekete:1996:NFT

[1605] S. P. Fekete, S. Khuller, M. Klemmstein, and B. Raghavachari. A networkflow technique for finding low-weight bounded-degree spanning trees. Lecture Notes in Computer Science, 1084: 105–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halldorsson:1996:ASC

[1606] M. M. Halldorsson. Approximating k-set cover and complementary graph coloring. Lecture Notes in Computer Science, 1084:118–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapoor:1996:MCA

[1607] S. Kapoor. On minimum 3-Cuts and approximating k-Cuts using cut trees. Lecture Notes in Computer Science, 1084:132–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goemans:1996:PDA

[1608] M. X. Goemans and D. P. Williamson.

Primal-dual approximation algorithms for feedback problems in planar graphs. Lecture Notes in Computer Science, 1084:147–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pataki:1996:CLS

[1609] G. Pataki. Cone-LP’s and semidefinite programs: Geometry and a simplex-type method. Lecture Notes in Computer Science, 1084:162–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helmberg:1996:QKR

[1610] C. Helmberg, F. Rendl, and R. Weismantel. Quadratic knapsack relaxations using cutting planes and semidefinite programming. Lecture Notes in Computer Science, 1084:175– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahale:1996:SBM

[1611] N. Kahale. A semidefinite bound for mixing rates of Markov chains. Lecture Notes in Computer Science, 1084: 190–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burkard:1996:QAP

[1612] R. E. Burkard, E. Cela, G. Rote, and G. J. Woeginger. The quadratic assignment problem with a monotone Anti-Monge and a symmetric Toeplitz matrix: Easy and hard cases. Lecture Notes in Computer Science, 1084:204– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1996:OMS

[1613] E. Cohen. On optimizing multiplications of sparse matrices. Lecture Notes in Computer Science, 1084:219– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anstreicher:1996:CRC

[1614] K. M. Anstreicher, M. Fampa, J. Lee, and J. Williams. Continuous relaxations for constrained maximumentropy sampling. Lecture Notes in Computer Science, 1084:234–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartvigsen:1996:SOP

[1615] D. Hartvigsen. A submodular optimization problem with side constraints. Lecture Notes in Computer Science, 1084:249–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murota:1996:CSE

[1616] K. Murota. Convexity and Steinitz’s exchange property. Lecture Notes in Computer Science, 1084:260–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Novick:1996:ICM

[1617] B. Novick and A. Sebo. On ideal clutters, metrics and multiflows. Lecture Notes in Computer Science, 1084:275– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goemans:1996:SRS

[1618] M. X. Goemans. A supermodular relaxation for scheduling with release dates. Lecture Notes in Computer Science, 1084:288–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz:1996:SMT

[1619] A. S. Schulz. Scheduling to minimize total weighted completion time: Performance guarantees of LP-Based heuristics and lower bounds. Lecture Notes in Computer Science, 1084:301– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simonetti:1996:ILT

[1620] N. Simonetti and E. Balas. Implementation of a linear time algorithm for certain generalized traveling salesman problems. Lecture Notes in Computer Science, 1084:316–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertsimas:1996:DRR

[1621] D. Bertsimas, C.-P. Teo, and R. Vohra. On dependent randomized rounding algorithms. Lecture Notes in Computer Science, 1084:330–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:CBH

[1622] H. Chen and A. Frieze. Coloring bipartite hypergraphs. Lecture Notes in Computer Science, 1084:345–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teo:1996:IRA

[1623] C.-P. Teo and D. Bertsimas. Improved randomized approximation algorithms for lot-sizing problems. Lecture Notes in Computer Science, 1084: 359–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoogeveen:1996:MTC

[1624] J. A. Hoogeveen and T. Kawaguchi.

Minimizing total completion time in a two-machine flowshop: Analysis of special cases. Lecture Notes in Computer Science, 1084:374–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1996:NAC

[1625] P. Martin and D. B. Shmoys. A new approach to computing optimal schedules for the job-shop scheduling problem. Lecture Notes in Computer Science, 1084:389–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoogeveen:1996:OLA

[1626] J. A. Hoogeveen and A. P. A. Vestjens.

Optimal on-line algorithms for singlemachine scheduling. Lecture Notes in Computer Science, 1084:404–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goemans:1996:SFA

[1627] M. X. Goemans and L. A. Hall. The strongest facets of the acyclic subgraph polytope are unknown. Lecture Notes in Computer Science, 1084:415– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1996:TP

[1628] R. Mueller and A. S. Schulz. Transitive packing. Lecture Notes in Computer Science, 1084:430–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Funke:1996:PAF

[1629] M. Funke and G. Reinelt. A polyhedral approach to the feedback vertex set problem. Lecture Notes in Computer Science, 1084:445–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cart:1996:SCT

[1630] R. Cart. Separating over classes of TSP inequalities defined by 0 node-lifting in polynomial time. Lecture Notes in Computer Science, 1084:460–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fleischer:1996:SMV

[1631] L. Fleischer and E. Tardos. Separating maximally violated comb inequalities in planar graphs. Lecture Notes in Computer Science, 1084:475–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burkard:1996:TSP

[1632] R. E. Burkard, V. G. Deineko, and G. J. Woeginger. The travelling salesman and the PQ-Tree. Lecture Notes in Computer Science, 1084:490–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDerAcker:1996:ISA

[1633] C. Van Der Acker and J. Vanthienen. Integrating statistical audit evidence with belief function theory. Lecture Notes in Computer Science, 1085:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniou:1996:CSD

[1634] G. Antoniou. A comparative survey of default logic variants. Lecture Notes in Computer Science, 1085:15–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balbiani:1996:MLR

[1635] P. Balbiani. Modal logics with relative accessibility relations. Lecture Notes in Computer Science, 1085:29–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balbiani:1996:GSM

[1636] P. Balbiani, L. F. Del Cerro, T. Tinchev, and D. Vakarelov. Geometrical structures and modal logic. Lecture Notes in Computer Science, 1085: 43–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1996:UFH

[1637] S. Das, J. Fox, and P. Krause. A unified framework for hypothetical and practical reasoning (1) — theoretical foundations. Lecture Notes in Computer Science, 1085:58–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1996:UFH

[1638] J. Fox and S. Das. A unified framework for hypothetical and practical reasoning (2) — lessons from medical applications. Lecture Notes in Computer Science, 1085:73–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doherty:1996:GDC

[1639] P. Doherty, W. Lukaszewicz, and A. Szalas. General domain circumscription and its first-order reduction. Lecture Notes in Computer Science, 1085: 93–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duc:1996:RAR

[1640] Ho Ngoc Duc. Reasoning about rational, but not logically omniscient agents. Lecture Notes in Computer Science, 1085:110–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1996:SNR

[1641] J. Engelfriet and J. Treur. Specification of nonmonotonic reasoning. Lecture Notes in Computer Science, 1085: 111–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Errico:1996:IAS

[1642] B. Errico and L. C. Aiello. Intelligent agents in the situation calculus: An application to user modelling. Lecture Notes in Computer Science, 1085:126–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelCerro:1996:TCW

[1643] L. F. Del Cerro, A. F. Delgado, and A. Herzig. Talkin’bout consistency, or: When logically possible becomes possible. Lecture Notes in Computer Science, 1085:141–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feteris:1996:AEL

[1644] E. T. Feteris. The analysis and evaluation of legal argumentation from a pragma-dialectical perspective. Lecture Notes in Computer Science, 1085:151– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finocchiaro:1996:RAR

[1645] M. A. Finocchiaro. Reasoning about reasoning. Lecture Notes in Computer Science, 1085:167–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisher:1996:RBP

[1646] M. Fisher, M. Wooldridge, and C. Dixon. A resolution-based proof method for temporal logics of knowledge and belief. Lecture Notes in Computer Science, 1085:178–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabbay:1996:MIT

[1647] D. Gabbay and O. Rodrigues. A methodology for iterated theory change. Lecture Notes in Computer Science, 1085:193–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geffner:1996:FFC

[1648] H. Geffner. A formal framework for causal modeling and argumentation. Lecture Notes in Computer Science, 1085:208–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilbert:1996:GA

[1649] M. A. Gilbert. Goals in argumentation. Lecture Notes in Computer Science, 1085:223–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giordano:1996:APP

[1650] L. Giordano, A. Martelli, and M. L. Sapino. An abductive proof procedure for conditional logic programming. Lecture Notes in Computer Science, 1085:231–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Girle:1996:CDL

[1651] R. A. Girle. Commands in dialogue logic. Lecture Notes in Computer Science, 1085:246–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giunchiglia:1996:IRBb

[1652] E. Giunchiglia and F. Giunchiglia. Ideal and real belief about belief. Lecture Notes in Computer Science, 1085: 261–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gotoda:1996:ARO

[1653] H. Gotoda, J. An, and Y. Fujiwara. Analogical reasoning of organic reactions based on the structurized compound-reaction diagram. Lecture Notes in Computer Science, 1085:276– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guarini:1996:MMR

[1654] M. Guarini. Mind, morals, and reasons. Lecture Notes in Computer Science, 1085:305–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hansen:1996:AWT

[1655] H. V. Hansen. Aristotle, Whately, and the taxonomy of fallacies. Lecture Notes in Computer Science, 1085:318– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1996:NRM

[1656] J. Engelfriet, H. Herre, and J. Treur.

Nonmonotonic reasoning with multiple belief sets. Lecture Notes in Computer Science, 1085:331–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grosse:1996:SGV

[1657] G. Grosse, C. S. Herrmann, and E. Sandner. SEdit — graphically validating technical systems. Lecture Notes in Computer Science, 1085:345– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnson:1996:NDT

[1658] R. H. Johnson. The need for a dialectical tier in arguments. Lecture Notes in Computer Science, 1085:349–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerber:1996:TKN

[1659] M. Kerber and E. Melis. Two kinds of non-monotonic analogical inference. Lecture Notes in Computer Science, 1085:361–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloosterhuis:1996:NRA

[1660] H. Kloosterhuis. The normative reconstruction of analogy argumentation in judicial decisions: a pragma-dialectical perspective. Lecture Notes in Computer Science, 1085:375–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kreitz:1996:FRA

[1661] C. Kreitz, K.-K. Lau, and M. Ornaghi. Formal reasoning about modules, reuse, and their correctness. Lecture Notes in Computer Science, 1085: 384–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

May:1996:TCF

[1662] W. May and P. H. Schmitt. A tableau calculus for first-order branching time logic. Lecture Notes in Computer Science, 1085:399–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1996:PWS

[1663] J.-J. C. Meyer and J. C. Van Leeuwen. Possible world semantics for analogous reasoning. Lecture Notes in Computer Science, 1085:414–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montesi:1996:UTI

[1664] D. Montesi, C. Renso, and F. Turini. Using temporary integrity constraints to optimize databases. Lecture Notes in Computer Science, 1085:430–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neumann:1996:GIN

[1665] I. Neumann. Graded inheritance nets for knowledge representation. Lecture Notes in Computer Science, 1085:436– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parsons:1996:DNS

[1666] S. Parsons. Defining normative systems for qualitative argumentation. Lecture Notes in Computer Science, 1085:449– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plug:1996:CAJ

[1667] J. Plug. Complex argumentation in judicial decisions. analysing conflicting arguments. Lecture Notes in Computer Science, 1085:464–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pohl:1996:CPM

[1668] W. Pohl. Combining partitions and modal logic for user modeling. Lecture Notes in Computer Science, 1085: 480–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pollock:1996:RCW

[1669] J. L. Pollock. Reason in a changing world. Lecture Notes in Computer Science, 1085:495–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prakken:1996:SDA

[1670] H. Prakken and G. Sartor. A system for defeasible argumentation, with defeasible priorities. Lecture Notes in Computer Science, 1085:510–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prendinger:1996:MLM

[1671] H. Prendinger. Modal logics for modelling actions and agents. Lecture Notes in Computer Science, 1085:525– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radzikowska:1996:FRA

[1672] A. Radzikowska. Formalization of reasoning about default action (preliminary report). Lecture Notes in Computer Science, 1085:540–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reed:1996:AAD

[1673] C. Reed, D. Long, and M. Fox. An architecture for argumentative dialogue planning. Lecture Notes in Computer Science, 1085:555–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaub:1996:SQA

[1674] T. Schaub and M. Thielscher. Skeptical query-answering in constrained default logic. Lecture Notes in Computer Science, 1085:567–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shiu:1996:TTS

[1675] S. Shiu, Z. Luo, and R. Garigliano.

Type theoretic semantics for Sem-Net. Lecture Notes in Computer Science, 1085:582–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tindale:1996:SAP

[1676] C. W. Tindale. From syllogisms to audiences: The prospects for logic in a rhetorical model of argumentation. Lecture Notes in Computer Science, 1085:596–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogel:1996:HRN

[1677] C. Vogel. Human reasoning with negative defaults. Lecture Notes in Computer Science, 1085:606–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1996:SU

[1678] S. Weber. On the semantics of the unknown. Lecture Notes in Computer Science, 1085:622–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weydert:1996:SJR

[1679] E. Weydert. System J — revision entailment — default reasoning through ranking measure updates. Lecture Notes in Computer Science, 1085:637– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Woods:1996:DDP

[1680] J. Woods. Deep disagreements and public demoralization. Lecture Notes in Computer Science, 1085:650–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wooldridge:1996:PRP

[1681] M. Wooldridge. Practical reasoning with procedural knowledge. Lecture Notes in Computer Science, 1085:663– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yi:1996:TAL

[1682] C.-H. Yi. Towards the assessment of logics for concurrent actions. Lecture Notes in Computer Science, 1085:679– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniou:1996:DRB

[1683] G. Antoniou and M. A. Williams. Default reasoning and belief revision in the CIN project. Lecture Notes in Computer Science, 1085:691–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cimatti:1996:MMA

[1684] A. Cimatti and L. Serafini. Mechanizing multi-agent reasoning with belief contexts. Lecture Notes in Computer Science, 1085:694–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Green:1996:AMM

[1685] D. W. Green. Arguments and mental models: A position paper. Lecture Notes in Computer Science, 1085:697– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parsons:1996:ADM

[1686] S. Parsons and J. Fox. Argumentation and decision making: A position paper. Lecture Notes in Computer Science, 1085:705–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oliveira:1996:IL

[1687] C. M. G. M. Oliveira. The implementation of LENA. Lecture Notes in Computer Science, 1085:710–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oliveira:1996:IC

[1688] C. M. G. M. Oliveira. The implementation of CondLP. Lecture Notes in Computer Science, 1085:713–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Omelyanchyk:1996:HRA

[1689] V. Omelyanchyk. How to reason about akratic action practically? Lecture Notes in Computer Science, 1085:716– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloom:1996:PTA

[1690] C. P. Bloom. Promoting the transfer of advanced training technologies. Lecture Notes in Computer Science, 1086:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boy:1996:LES

[1691] G. Boy. Learning evolution and software agents emergence. Lecture Notes in Computer Science, 1086:10– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collins:1996:WTS

[1692] A. Collins. Whither technology and schools: Collected thoughts on the last and next quarter centuries. Lecture Notes in Computer Science, 1086:26– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merrill:1996:IDT

[1693] M. D. Merrill. Instructional design tools based on knowledge objects. Lecture Notes in Computer Science, 1086: 27–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schank:1996:MMS

[1694] R. C. Schank. Multi-media software development and the tools to support it. Lecture Notes in Computer Science, 1086:28–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanLehn:1996:CML

[1695] K. VanLehn. Conceptual and meta learning during coached problem solving. Lecture Notes in Computer Science, 1086:29–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubourg:1996:ALF

[1696] X. Dubourg. Une architecture logicielle fondée sur le concept d’évenement — logiciel: le cas de REPERES. Lecture Notes in Computer Science, 1086:48– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frasson:1996:ABA

[1697] C. Frasson, T. Mengelle, E. Aimeur, and G. Gouarderes. An actor-based architecture for intelligent tutoring systems. Lecture Notes in Computer Science, 1086:57–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gagne:1996:HFS

[1698] D. Gagne and A. Trudel. A highly flexible student driven architecture for computer based instruction. Lecture Notes in Computer Science, 1086:66– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leman:1996:GAB

[1699] S. Leman, P. Marcenac, and S. Giroux.

A generic architecture for ITS based on a multi-agent approach. Lecture Notes in Computer Science, 1086:75– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1996:LCO

[1700] H. Wang. LearnMedia: A co-operative intelligent tutoring system for learning multimedia. Lecture Notes in Computer Science, 1086:84–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murray:1996:HIA

[1701] T. Murray. Having it all, maybe: Design tradeoffs in ITS authoring tools. Lecture Notes in Computer Science, 1086:93–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoelles:1996:TUE

[1702] M. Schoelles and H. Hamburger. Teacher-usable exercise design tools. Lecture Notes in Computer Science, 1086:102–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charlot:1996:MCP

[1703] J.-M. Charlot, B. Marcos, and J. Lapointe. Modèles cognitifs pour l’apprentissage en thermodynamique. Lecture Notes in Computer Science, 1086:111–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corbett:1996:PSI

[1704] A. T. Corbett and S. Knapp. Plan scaffolding: Impact on the process and product of learning. Lecture Notes in Computer Science, 1086:120–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khuwaja:1996:MTB

[1705] R. Khuwaja and V. Patel. A model of tutoring: Based on the behavior of effective human tutors. Lecture Notes in Computer Science, 1086:130–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsubara:1996:MSH

[1706] Y. Matsubara and M. Nagamachi. Motivation system and human model for intelligent tutoring. Lecture Notes in Computer Science, 1086:139–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayers:1996:MFD

[1707] A. Mayers and B. Lefebvre. La modelisation fine du processus r`’solution de problème dans MIA CE. Lecture Notes in Computer Science, 1086:148– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayala:1996:LMS

[1708] G. Ayala and Y. Yano. Learner models for supporting awareness and collaboration in a CSCL environment. Lecture Notes in Computer Science, 1086: 158–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBarrosCosta:1996:MCI

[1709] E. De Barros Costa and A. Perkusich. Modeling the cooperative interactions in a teaching/learning situation. Lecture Notes in Computer Science, 1086:168–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuji:1996:CBA

[1710] T. Fuji, T. Tanigawa, M. Kozeni, and M. Inui. A case-based approach to collaborative learning for systems analyst education. Lecture Notes in Computer Science, 1086:177–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heh:1996:DDD

[1711] J.-S. Heh, W.-T. Shu, J.-C. J. Jehng, and T.-W. Chan. Design and development of a distributed multi-user visual learning environment. Lecture Notes in Computer Science, 1086:187–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leroux:1996:MPC

[1712] P. Leroux and M. Vivet. De la modelisation d’un processus de cooperation a la conception de systèmes cooperatifs d’apprentissage. Lecture Notes in Computer Science, 1086:197–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lund:1996:MDB

[1713] K. Lund, M. Baker, and M. Baron. Modeling dialogue and beliefs as a basis for generating guidance in a CSCL environment. Lecture Notes in Computer Science, 1086:206–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tadie:1996:MMN

[1714] S. G. Tadie, B. Lefebvre, and C. Frasson. MONACOT: un MOdèle à NAture COoperative pour la représentation de Tâches cooperatives dans un STI. Lecture Notes in Computer Science, 1086: 215–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khuwaja:1996:IGC

[1715] R. Khuwaja, M. Desmarais, and R. Cheng. Intelligent guide: Combining user knowledge assessment with pedagogical guidance. Lecture Notes in Computer Science, 1086:225–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsuda:1996:PCM

[1716] N. Matsuda and T. Okamoto. Parallel computing model for problem solver towards ITSs — epistemological articulation of human problem solving. Lecture Notes in Computer Science, 1086:234–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Millet:1996:AQR

[1717] S. Millet and G. Gouarderes. Approche qualitative de la ré-ingénierie d’un système tuteur intelligent à partir d’une méthodologie d’évaluation. Lecture Notes in Computer Science, 1086: 243–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1996:ICI

[1718] L.-H. Wong, C.-K. Looi, and H.-C.

Quek. Issues in computerizing the inquiry dialogue planning process. Lecture Notes in Computer Science, 1086: 252–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brusilovsky:1996:EAI

[1719] P. Brusilovsky, E. Schwarz, and G. Weber. ELM-ART: An intelligent tutoring system on World Wide Web. Lecture Notes in Computer Science, 1086: 261–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Espinosa:1996:IAB

[1720] E. Espinosa, A. Brito, and F. Ramos. Intelligent agent-based virtual education using the Java technology. Lecture Notes in Computer Science, 1086:270– 278, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kabbaj:1996:USN

[1721] A. Kabbaj, K. Rouane, and C. Frasson. The use of a semantic network activation language in an ITS project. Lecture Notes in Computer Science, 1086:279–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tabak:1996:CGD

[1722] I. Tabak, B. K. Smith, W. A. Sandoval, and B. J. Reiser. Combining general and domain-specific strategic support for biological inquiry. Lecture Notes in Computer Science, 1086:288–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fix:1996:UCP

[1723] V. Fix and S. Wiedenbeck. Using cognition of programming literature in the design of a tool for learning a second programming language. Lecture Notes in Computer Science, 1086:297– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1996:CBA

[1724] S. X. Huang. A content-balanced adaptive testing algorithm for computerbased training systems. Lecture Notes in Computer Science, 1086:306–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teutsch:1996:MSE

[1725] P. Teutsch. Un modèle de situation d’évaluation pour le suivi de formation en langue étrangère. Lecture Notes in Computer Science, 1086:315–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1996:FEL

[1726] A. K. W. Wu. On the formal evaluation of learning systems. Lecture Notes in Computer Science, 1086:324–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouhineau:1996:PLP

[1727] D. Bouhineau and S. Channac. La programmation logique par contraintes pour l’aide à l’enseignant. Lecture Notes in Computer Science, 1086:333– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerri:1996:CMT

[1728] S. A. Cerri. Computational mathematics toolkit: Architectures for modeling dialogues. Lecture Notes in Computer Science, 1086:343–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Folckers:1996:IPS

[1729] J. Folckers, C. Moebus, O. Schroeder, and H.-J. Thole. An intelligent problem solving environment for designing explanation models and for diagnostic reasoning in probabilistic domains. Lecture Notes in Computer Science, 1086:353–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giroux:1996:EPE

[1730] S. Giroux, G. Paquette, F. Pachet, and J. Girard. EpiTalk — A platform for Epiphyte advisor systems dedicated to both individual and collaborative learning. Lecture Notes in Computer Science, 1086:363–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paolucci:1996:AAG

[1731] M. Paolucci, D. Suthers, and A. Weiner. Automated advice-giving strategies for scientific inquiry. Lecture Notes in Computer Science, 1086:372–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paquette:1996:ACE

[1732] G. Paquette and J. Girard. AGD: A course engineering support system. Lecture Notes in Computer Science, 1086:382–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBuen-Rodriguez:1996:KBF

[1733] P. R. De Buen-Rodriguez, E. F. Morales, and S. Vadera. A knowledgebased framework for learning, applying and consulting engineering procedures. Lecture Notes in Computer Science, 1086:392–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elorriaga:1996:HAE

[1734] J. A. Elorriaga and I. Fernandez-Castro. The HSIIP approach. an extension for a teacher’s apprentice. Lecture Notes in Computer Science, 1086: 401–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohnuma:1996:PTT

[1735] H. Ohnuma and T. Watanabe. The program tutoring text model. Lecture Notes in Computer Science, 1086:411– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nkambou:1996:MRD

[1736] R. Nkambou and G. Gauthier. Un module de représentation du curriculum dans un STI. Lecture Notes in Computer Science, 1086:420–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akhras:1996:PSL

[1737] F. N. Akhras and J. A. Self. A processsensitive learning environment architecture. Lecture Notes in Computer Science, 1086:430–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alexe:1996:LES

[1738] C. Alexe and J. Gecsei. A learning environment for the surgical intensive care unit. Lecture Notes in Computer Science, 1086:439–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernat:1996:NWV

[1739] P. Bernat and J. Morinet-Lambert. A new way for visual reasoning in geometry education. Lecture Notes in Computer Science, 1086:448–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burow:1996:EEL

[1740] R. Burow and G. Weber. Example explanation in learning environments. Lecture Notes in Computer Science, 1086:457–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirashima:1996:TLE

[1741] T. Hirashima, A. Kashihara, and J. Toyoda. Toward a learning environment allowing learner-directed problem practice — helping problem-solving by using problem simplification. Lecture Notes in Computer Science, 1086:466– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lester:1996:FPS

[1742] J. C. Lester, B. A. Stone, M. A. O’Leary, and R. B. Stevenson. Focusing problem solving in design-centered learning environments. Lecture Notes in Computer Science, 1086:475–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blanc:1996:UMQ

[1743] F. Blanc. Une utilisation de la modelisation qualitative pour la planification pédagogique. Lecture Notes in Computer Science, 1086:484–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goel:1996:TDL

[1744] A. K. Goel, A. G. De Silva Garza, N. Grue, and J. W. Murdock. Towards design learning environments — I: Exploring how devices work. Lecture Notes in Computer Science, 1086: 493–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leclet:1996:MSB

[1745] D. Leclet and G. Weidenfeld. Un module de simulation basée sur une représentation de type “objets-règles” pour l’enseignement des métiers de ventes. Lecture Notes in Computer Science, 1086:502–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liddle:1996:GAD

[1746] J. Liddle, R. Leitch, and K. Brown.

Generic approaches to developing practical intelligent industrial training systems. Lecture Notes in Computer Science, 1086:512–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Obradovich:1996:TMT

[1747] J. H. Obradovich, P. J. Smith, S. Guerlain, and S. Rudmann. The transfusion medicine tutor: Using expert systems technology to teach domain-specific problem-solving skills. Lecture Notes in Computer Science, 1086:521–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schewe:1996:EKB

[1748] S. Schewe, T. Quak, B. Reinhardt, and F. Puppe. Evaluation of a knowledgebased tutorial program in rheumatology — A part of a mandatory course in internal medicine. Lecture Notes in Computer Science, 1086:531–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eliot:1996:IDV

[1749] C. Eliot and B. P. Woolf. Iterative development and validation of a simulation-based medical tutor. Lecture Notes in Computer Science, 1086: 540–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moebus:1996:WIP

[1750] C. Moebus, O. Schroeder, and H.-J. Thole. WULPUS — an intelligent problem solving environment delivering knowledge based help and explanations in business management simulation. Lecture Notes in Computer Science, 1086:550–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1996:UCB

[1751] D. M. Zhang and L. Alem. Using case-based reasoning for exercice design in simulation-based training. Lecture Notes in Computer Science, 1086:560– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collins:1996:AAU

[1752] J. A. Collins, J. E. Greer, and S. X. Huang. Adaptive assessment using granularity hierarchies and Bayesian nets. Lecture Notes in Computer Science, 1086:569–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dowling:1996:AAT

[1753] C. E. Dowling, C. Hockemeyer, and A. H. Ludwig. Adaptive assessment and training using the neighbourhood of knowledge states. Lecture Notes in Computer Science, 1086:578–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitrovic:1996:SSI

[1754] A. Mitrovic. SINT — A symbolic integration tutor. Lecture Notes in Computer Science, 1086:587–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reye:1996:BNB

[1755] J. Reye. A belief net backbone for student modeling. Lecture Notes in Computer Science, 1086:596–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stern:1996:APP

[1756] M. Stern, J. Beck, and B. P. Woolf.

Adaptation of problem presentation and feedback in an intelligent mathematics tutor. Lecture Notes in Computer Science, 1086:605–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yacef:1996:SEM

[1757] K. Yacef and L. Alem. Student and expert modeling for simulationbased training: A cost effective framework. Lecture Notes in Computer Science, 1086:614–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1996:PLS

[1758] K.-E. Chang, B.-C. Chiao, and R.-S. Hsiao. A programming learning system for beginners — A completion strategy approach. Lecture Notes in Computer Science, 1086:623–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freedman:1996:GRH

[1759] R. Freedman and M. W. Evens. Generating and revising hierarchical multiturn text plans in an ITS. Lecture Notes in Computer Science, 1086:632– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katz:1996:TDM

[1760] S. Katz, A. Lesgold, G. Eggan, and L. Greenberg. Towards the design of more effective advisors for learningby-doing systems. Lecture Notes in Computer Science, 1086:641–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ragnemalm:1996:CDL

[1761] E. L. Ragnemalm. Collaborative dialogue with a learning companion as a source of information on student reasoning. Lecture Notes in Computer Science, 1086:650–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Retalis:1996:ADT

[1762] S. Retalis, H. Pain, and M. Haggith. Arguing with the devil: Teaching in controversial domains. Lecture Notes in Computer Science, 1086:659–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slator:1996:LLR

[1763] B. M. Slator and H. Chaput. Learning by learning roles: A virtual roleplaying environment for tutoring. Lecture Notes in Computer Science, 1086: 668–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srisethanil:1996:EDI

[1764] C. Srisethanil and N. C. Baker. ITS-Engineering: A domain independent ITS for building engineering tutors. Lecture Notes in Computer Science, 1086:677–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1996:DAC

[1765] H. Yang and C. Zhang. Definition and application of a comprehensive framework for distributed problem solving. Lecture Notes in Computer Science, 1087:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cavedon:1996:LFM

[1766] L. Cavedon and G. Tidhar. A logical framework for multi-agent systems and joint attitudes. Lecture Notes in Computer Science, 1087:16–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shehory:1996:KOM

[1767] O. Shehory and S. Kraus. A kerneloriented model for autonomous-agent coalition-formation in general environments. Lecture Notes in Computer Science, 1087:31–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ossowski:1996:KLM

[1768] S. Ossowski and A. Garcia-Serrano. A knowledge-level model of coordination. Lecture Notes in Computer Science, 1087:46–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlsson:1996:EMM

[1769] B. Carlsson. An evolutionary model of multi-agent systems. Lecture Notes in Computer Science, 1087:58–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luck:1996:ECM

[1770] M. Luck and M. D’Inverno. Engagement and cooperation in motivated agent modelling. Lecture Notes in Computer Science, 1087:70–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kendall:1996:MDA

[1771] E. A. Kendall, M. T. Malkoun, and C. H. Jiang. A methodology for developing agent based systems. Lecture Notes in Computer Science, 1087:85– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wobcke:1996:PRI

[1772] W. Wobcke. Plans and the revision of intentions. Lecture Notes in Computer Science, 1087:100–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DInverno:1996:FVS

[1773] M. D’Inverno and M. Luck. A formal view of social dependence networks. Lecture Notes in Computer Science, 1087:115–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sommaruga:1996:PTD

[1774] L. Sommaruga and N. Catenazzi. From practice to theory in designing autonomous agents. Lecture Notes in Computer Science, 1087:130–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munday:1996:MPM

[1775] C. Munday, J. Dangedej, T. Cross, and D. Lukose. Motivation and perception mechanisms in mobil agent for electronic commerce. Lecture Notes in Computer Science, 1087:144–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rao:1996:IAA

[1776] A. S. Rao. Integrated agent architecture: Execution and recognition of mental-states. Lecture Notes in Computer Science, 1087:159–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1996:NNS

[1777] M. Zhang and C. Zhang. Neural network strategies for solving synthesis problems in non-conflict cases in distributed expert systems. Lecture Notes in Computer Science, 1087:174– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davidsson:1996:LQA

[1778] P. Davidsson. A linearly quasianticipatory autonomous agent architecture: Some preliminary experiments. Lecture Notes in Computer Science, 1087:189–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belo:1996:CAO

[1779] O. Belo and J. Neves. Cooperation among opportunistic agents on a distributed environment. Lecture Notes in Computer Science, 1087:204–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moulin:1996:SBD

[1780] B. Moulin and M. Brassard. A scenario-based design and an environment of multi-agent systems. Lecture Notes in Computer Science, 1087:216– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taft:1996:PIA

[1781] S. T. Taft. Programming the Internet in Ada 95. Lecture Notes in Computer Science, 1088:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneidewind:1996:RMS

[1782] N. F. Schneidewind. Reliability modeling for safety critical software. Lecture Notes in Computer Science, 1088: 17–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerraoui:1996:FTR

[1783] R. Guerraoui and A. Schiper. Faulttolerance by replication in distributed systems. Lecture Notes in Computer Science, 1088:38–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burns:1996:AEC

[1784] A. Burns and A. J. Wellings. Ada 95:

An effective concurrent programming language. Lecture Notes in Computer Science, 1088:58–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeLaPuente:1996:MHH

[1785] J. A. De La Puente, A. Alonso, and A. Alvarez. Mapping HRT-HOOD designs to Ada 95 hierarchical libraries. Lecture Notes in Computer Science, 1088:78–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1996:AIS

[1786] H. Kim and C. Boldyreff. An approach to increasing software component reusability in Ada. Lecture Notes in Computer Science, 1088:89– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kruchten:1996:ISD

[1787] P. Kruchten and C. J. Thompson. Iterative software development for large Ada programs. Lecture Notes in Computer Science, 1088:101–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emery:1996:HUD

[1788] D. Emery and J. S. Madhur. HCSD unit development process: Step-wise process improvement. Lecture Notes in Computer Science, 1088:111–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madhav:1996:TAP

[1789] N. Madhav. Testing Ada 95 programs for conformance to rapide architectures. Lecture Notes in Computer Science, 1088:123–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1996:TDA

[1790] J. Cheng and K. Ushijima. Tasking deadlocks in Ada 95 programs and their detection. Lecture Notes in Computer Science, 1088:135–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:SCP

[1791] Z. T. Chen and Y. T. Yu. On some characterisation problems of subdomain testing. Lecture Notes in Computer Science, 1088:147–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fletcher:1996:FTO

[1792] R. Fletcher and A. S. M. Sajeev. A framework for testing object-oriented software using formal specifications. Lecture Notes in Computer Science, 1088:159–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Craigen:1996:ACS

[1793] D. Craigen, M. Saaltink, and S. Michell. Ada 95 and critical systems: An analytical approach. Lecture Notes in Computer Science, 1088:171–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosskopf:1996:USA

[1794] A. Rosskopf. Use of a static analysis tool for safety-critical Ada applications: A critical assessment. Lecture Notes in Computer Science, 1088:183– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kienzle:1996:SCD

[1795] J. Kienzle, T. Wolf, and A. Strohmeier.

Secure communication in distributed ada. Lecture Notes in Computer Science, 1088:198–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lundqvist:1996:UOO

[1796] K. Lundqvist and G. Wall. Using object-oriented methods in Aria 95 to implement Linda. Lecture Notes in Computer Science, 1088:211–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wall:1996:SPT

[1797] G. Wall and K. Lundqvist. Shared packages through Linda. Lecture Notes in Computer Science, 1088:223–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miranda:1996:DAE

[1798] F. J. Miranda, A. Alvarez, S. Arevalo, and F. J. Guerra. Drago: An Ada extension to program fault-tolerant distributed applications. Lecture Notes in Computer Science, 1088:235–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brosgol:1996:DPA

[1799] B. M. Brosgol. The dining philosophers in Ada 95. Lecture Notes in Computer Science, 1088:247–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Real:1996:UAP

[1800] J. Real, A. Espinosa, and A. Crespo. Using Ada 95 for prototyping real-time systems. Lecture Notes in Computer Science, 1088:262–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oh:1996:GIP

[1801] D.-I. Oh, T. P. Baker, and S.-J Moon. The GNARL implementation of POSIX/Ada signal services. Lecture Notes in Computer Science, 1088:275– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mundie:1996:IPT

[1802] D. Mundie, J. Fardo, and E. Kuzemchak. Implementing protected types on embedded targets. Lecture Notes in Computer Science, 1088:287–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rybin:1996:AGP

[1803] S. Rybin, A. Strohmeier, A. Kuchumov, and V. Fofanov. ASIS for GNAT: From the prototype to the full implementation. Lecture Notes in Computer Science, 1088:298–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sotirovski:1996:HIG

[1804] D. Sotirovski and P. Kruchten. Handling irregular grammars in Ada. Lecture Notes in Computer Science, 1088: 312–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birus:1996:ICO

[1805] T. Birus, C. Cipriani, and D. Sutherland. Interprocedural call optimization. Lecture Notes in Computer Science, 1088:319–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blieberger:1996:AAA

[1806] J. Blieberger, R. Lieger, and B. Burgstaller.

Augmenting Ada 95 with additional real-tune features. Lecture Notes in Computer Science, 1088:330–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oudshoorn:1996:BAA

[1807] M. J. Oudshoorn and S. C. Crawley.

Beyond Ada 95: The addition of persistence and its consequences. Lecture Notes in Computer Science, 1088:342– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzmueller:1996:EOO

[1808] B. Holzmueller. Extending the objectorientedness of Ada 95. Lecture Notes in Computer Science, 1088:357–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBondeli:1996:AVS

[1809] P. De Bondeli. An Ada 95 view of some difficult or controversial points in object-oriented programming. Lecture Notes in Computer Science, 1088:370– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gellerich:1996:WDG

[1810] W. Gellerich, M. Kosiol, and E. Ploedereder. Where does GOTO go to? Lecture Notes in Computer Science, 1088: 385–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Millan:1996:AOC

[1811] T. Millan and P. Bazex. Ada/O2 coupling: A solution for an efficient management of persistence in Ada 83. Lecture Notes in Computer Science, 1088:396–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gliss:1996:AHC

[1812] B. Gliss. An Ada 95 harness for converting legacy Fortran applications. Lecture Notes in Computer Science, 1088:413–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bremmon:1996:FMM

[1813] C. Bremmon. The funds management modernization: Experiences with developing an object-oriented, clientserver management information system in Ada 95. Lecture Notes in Computer Science, 1088:427–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Molenmaker:1996:CPT

[1814] K. Molenmaker. Converting the part task nautical simulator to Ada 95. Lecture Notes in Computer Science, 1088: 439–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Celier:1996:VCM

[1815] V. Celier. Visibility control and migration of interfaces in large Ada systems. Lecture Notes in Computer Science, 1088:451–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waroquiers:1996:ATD

[1816] P. Waroquiers. Ada tasking and dynamic memory: To use or not to use, that’s a question! Lecture Notes in Computer Science, 1088:460–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emery:1996:EAP

[1817] D. E. Emery, R. F. Hilliard, and T. B.

Rice. Experiences applying a practical architectural method. Lecture Notes in Computer Science, 1088:471–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smart:1996:DDD

[1818] J. D. Smart. A decade of development and deployment of distributed Ada systems. Lecture Notes in Computer Science, 1088:485–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feith:1996:PTA

[1819] R. Feith and M. Tonndorf. Planning the transition to Ada 95 for a large real-time project. Lecture Notes in Computer Science, 1088:500–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramalingam:1996:BIC

[1820] G. Ramalingam. Bounded incremental computation. Lecture Notes in Computer Science, 1089:xi + 190, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrews:1996:IBL

[1821] M. Andrews, M. X. Goemans, and L. Zhang. Improved bounds for online load balancing. Lecture Notes in Computer Science, 1090:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xue:1996:ATA

[1822] G. Xue and D.-Z. Du. An O(n log n)- average-time algorithm for shortest networks under a given topology. Lecture Notes in Computer Science, 1090: 11–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1996:SPD

[1823] T.-S. Hsu, K.-H. Tsai, D.-W. Wang, and D. T. Lee. Steiner problems on directed acyclic graphs. Lecture Notes in Computer Science, 1090:21–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karaivazoglou:1996:WVD

[1824] E. Karaivazoglou, P. Spirakis, and V. Triantafilou. Wormhole versus deflection routing: A case study on the mesh. Lecture Notes in Computer Science, 1090:31–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lefmann:1996:SPC

[1825] H. Lefmann, P. Pudlak, and P. Savicky. On sparse parity check matrices. Lecture Notes in Computer Science, 1090:41–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:FHC

[1826] Z. Chen, C. Cunha, and S. Homer.

Finding a hidden code by asking questions. Lecture Notes in Computer Science, 1090:50–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dai:1996:ILL

[1827] H. K. Dai and K. E. Flannery. Improved length lower bounds for reflecting sequences. Lecture Notes in Computer Science, 1090:56–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imai:1996:CGA

[1828] H. Imai, S. Iwata, K. Sekine, and K. Yoshida. Combinatorial and geometric approaches to counting problems on linear matroids, graphic arrangements, and partial orders. Lecture Notes in Computer Science, 1090: 68–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiBattista:1996:OSR

[1829] G. Di Battista, R. Tamassia, and L. Vismara. Output-sensitive reporting of disjoint paths. Lecture Notes in Computer Science, 1090:81–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rahman:1996:RGD

[1830] S. M. Rahman, S.-I. Nakano, and T. Nishizeki. Rectangular grid drawings of plane graphs. Lecture Notes in Computer Science, 1090:92–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shin:1996:AEA

[1831] C.-S. Shin, Sung Kwon Kim, and K.- Y. Chwa. Area-efficient algorithms for upward straight-line tree drawings. Lecture Notes in Computer Science, 1090:106–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aichholzer:1996:SSG

[1832] O. Aichholzer and F. Aurenhammer. Straight skeletons for general polygonal figures in the plane. Lecture Notes in Computer Science, 1090:117–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allender:1996:NUC

[1833] E. Allender. A note on uniform circuit lower bounds for the counting hierarchy. Lecture Notes in Computer Science, 1090:127–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmeister:1996:NSE

[1834] T. Hofmeister. A note on the simulation of exponential threshold weights. Lecture Notes in Computer Science, 1090:136–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grolmusz:1996:HAR

[1835] V. Grolmusz. Harmonic analysis, real approximation, and the communication complexity of Boolean functions. Lecture Notes in Computer Science, 1090:142–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calinescu:1996:FLP

[1836] G. Calinescu and C. G. Fernandes.

Finding large planar subgraphs and large subgraphs of a given genus. Lecture Notes in Computer Science, 1090: 152–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shahrokhi:1996:EDA

[1837] F. Shahrokhi and W. Shi. Efficient deterministic algorithms for embedding graphs on books. Lecture Notes in Computer Science, 1090:162–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1996:OBL

[1838] T.-S. Hsu and M.-Y. Kao. Optimal bilevel augmentation for selectivity enhancing graph connectivity with applications. Lecture Notes in Computer Science, 1090:169–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domingo:1996:ELS

[1839] C. Domingo. Exact learning of subclasses of CDNF formulas with membership queries. Lecture Notes in Computer Science, 1090:179–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teng:1996:FSD

[1840] S.-H. Teng. Fast separator decomposition for finite-element meshes. Lecture Notes in Computer Science, 1090: 189–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1996:RAC

[1841] H. L. Bodlaender and B. De Fluiter. Reduction algorithms for constructing solutions in graphs with small treewidth. Lecture Notes in Computer Science, 1090:199–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uehara:1996:FRN

[1842] R. Uehara, Z.-Z. Chen, and X. He. Fast RNC and NC algorithms for finding a maximal set of paths with an application. Lecture Notes in Computer Science, 1090:209–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaerkkaeinen:1996:SST

[1843] J. Kaerkkaeinen and E. Ukkonen. Sparse suffix trees. Lecture Notes in Computer Science, 1090:219–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yeh:1996:DET

[1844] C.-H. Yeh and E. A. Varvarigos. Depthefficient threshold circuits for multiplication and symmetric function computation. Lecture Notes in Computer Science, 1090:231–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1996:NSW

[1845] V. Arvind. A note on the selfwitnessing property of computational problems. Lecture Notes in Computer Science, 1090:241–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kavvadias:1996:ISP

[1846] D. Kavvadias and M. Sideri. The inverse satisfiability problem. Lecture Notes in Computer Science, 1090:250– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemaspaandra:1996:JCL

[1847] L. A. Hemaspaandra, Z. Jiang, J. Rothe, and O. Watanabe. The join can lower complexity. Lecture Notes in Computer Science, 1090:260–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yong:1996:DEG

[1848] X. Yong. On the distribution of eigenvalues of graphs. Lecture Notes in Computer Science, 1090:268–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grigni:1996:DDG

[1849] M. Grigni, V. Mirelli, and C. H. Papadimitriou. On the difficulty of designing good classifiers. Lecture Notes in Computer Science, 1090:273–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kumar:1996:ALS

[1850] S. R. Kumar, A. Russell, and R. Sundaram. Approximating Latin square extensions. Lecture Notes in Computer Science, 1090:280–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akutsu:1996:AMK

[1851] T. Akutsu and F. Bao. Approximating minimum keys and optimal substructure screens. Lecture Notes in Computer Science, 1090:290–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belanger:1996:RCR

[1852] J. Belanger and W. Jie. Reductions and covergence rates of average time. Lecture Notes in Computer Science, 1090:300–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yanbe:1996:CCP

[1853] A. Yanbe and K. Sakurai. On the complexity of computational problems associated with simple stochastic games. Lecture Notes in Computer Science, 1090:310–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibarra:1996:CCA

[1854] O. Ibarra, P. Diniz, and M. Rinard.

On the complexity of commutativity analysis. Lecture Notes in Computer Science, 1090:323–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clementi:1996:INA

[1855] A. E. F. Clementi and L. Trevisan. Improved non-approximability results for vertex cover problems with density constraints. Lecture Notes in Computer Science, 1090:333–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1996:SNN

[1856] M. Li, J. Tromp, and L. Zhang. Some notes on the nearest neighbour interchange distance. Lecture Notes in Computer Science, 1090:343–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shanbhogue:1996:DCA

[1857] V. Shanbhogue and M. Yung. Distributed computing in asynchronous networks with Byzantine edges. Lecture Notes in Computer Science, 1090: 352–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cho:1996:WBL

[1858] S. Cho and S. Sahni. Weight biased leftist trees and modified skip lists. Lecture Notes in Computer Science, 1090:361–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lau:1996:PAL

[1859] Hoong Chuin Lau. Probabilistic analysis of local search and NP-Completeness result for constraint satisfaction. Lecture Notes in Computer Science, 1090:371–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whitesides:1996:RC

[1860] S. Whitesides and N. Pei. On the reconfiguration of chains. Lecture Notes in Computer Science, 1090:381–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tan:1996:TGR

[1861] X. Tan and B. Zhu. Two-guarding a rectilinear polygon. Lecture Notes in Computer Science, 1090:391–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Safavi-Naini:1996:TSS

[1862] R. Safavi-Naini. Three systems for shared generation of authenticators. Lecture Notes in Computer Science, 1090:401–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1996:EGE

[1863] K.-Y. Lam, L. San, and L. C.-K Hui. Efficient generation of elliptic curve cryptosystems. Lecture Notes in Computer Science, 1090:411–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meng:1996:SMM

[1864] J. Meng. Superconnectivity for minimal multi-loop networks. Lecture Notes in Computer Science, 1090:417–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agha:1996:MCS

[1865] G. Agha. Modeling concurrent systems: Actors, nets, and the problem of abstraction and composition. Lecture Notes in Computer Science, 1091:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murata:1996:TUF

[1866] T. Murata. Temporal uncertainty and fuzzy-timing high-level Petri nets. Lecture Notes in Computer Science, 1091: 11–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valmari:1996:CSS

[1867] A. Valmari. Compositionality in state space verification methods. Lecture Notes in Computer Science, 1091:29– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barkaoui:1996:LCS

[1868] K. Barkaoui and J.-F. Pradat-Peyre.

On liveness and controled siphons in Petri nets. Lecture Notes in Computer Science, 1091:57–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benalycherif:1996:BSC

[1869] M.-L. Benalycherif and C. Girault.

Behavioral and structural composition rules preserving liveness by synchronization for colored FIFO nets. Lecture Notes in Computer Science, 1091: 73–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bilinski:1996:HLS

[1870] K. Bilinski and E. L. Dagless. High level synthesis of synchronous parallel controlers. Lecture Notes in Computer Science, 1091:93–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Busi:1996:NSS

[1871] N. Busi and G. M. Pinna. Non sequential semantics for contextual P/ T nets. Lecture Notes in Computer Science, 1091:113–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donatelli:1996:PMI

[1872] S. Donatelli and G. Franceschinis. The PSR methodology: Integrating hardware and software models. Lecture Notes in Computer Science, 1091:133– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Floreani:1996:DVC

[1873] D. J. Floreani, J. Billington, and A. Dadej. Designing and verifying a communications gateway using colored Petri nets and design/CPN. Lecture Notes in Computer Science, 1091:153– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

German:1996:EIR

[1874] R. German, A. P. A. Van Moorsel, M. A. Qureshi, and W. H. Sanders. Expected impulse rewards in Markov regenerative stochastic Petri nets. Lecture Notes in Computer Science, 1091: 172–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haddad:1996:ACH

[1875] S. Haddad and P. Moreaux. Asynchronous composition of high level Petri nets: A quantitative approach. Lecture Notes in Computer Science, 1091:192–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1996:FDH

[1876] X. He. A formal definition of hierarchical predicate transition nets. Lecture Notes in Computer Science, 1091: 212–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hiraishi:1996:RSS

[1877] K. Hiraishi. Reduced state space representation for unbounded vector state spaces. Lecture Notes in Computer Science, 1091:230–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joergensen:1996:MAD

[1878] J. B. Joergensen and K. H. Mortensen.

Modeling and analysis of distributed program execution in BETA using colored Petri nets. Lecture Notes in Computer Science, 1091:249–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kemper:1996:RAB

[1879] P. Kemper. Reachability analysis based on structured representations. Lecture Notes in Computer Science, 1091:269– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kindler:1996:ATP

[1880] E. Kindler and R. Walter. Arc-typed Petri nets. Lecture Notes in Computer Science, 1091:289–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kleinjohann:1996:SLS

[1881] B. Kleinjohann, E. Kleinjohann, and J. Tacken. The SEA language for system engineering and animation. Lecture Notes in Computer Science, 1091: 307–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knaup:1996:NIN

[1882] W. Knaup. A new iterative numerical solution algorithm for a class of stochastic Petri nets. Lecture Notes in Computer Science, 1091:327–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kondratyev:1996:SAA

[1883] A. Kondratyev, M. Kishinevsky, A. Taubin, and S. Ten. A structural approach for the analysis of Petri nets by reduced unfoldings. Lecture Notes in Computer Science, 1091:346–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kovalyov:1996:AVI

[1884] A. Kovalyov. An O(dSd × dT d)- algorithm to verify if a net is regular. Lecture Notes in Computer Science, 1091:366–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakos:1996:CUN

[1885] C. Lakos. The consistent use of names and polymorphism in the definition of object Petri nets. Lecture Notes in Computer Science, 1091:380–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LinnebergRasmussen:1996:DSS

[1886] J. Linneberg Rasmussen and M. Singh.

Designing a security system by means of colored Petri nets. Lecture Notes in Computer Science, 1091:400–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nigro:1996:MAD

[1887] L. Nigro and F. Pupo. Modeling and analysing DART systems through highlevel Petri nets. Lecture Notes in Computer Science, 1091:420–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Recalde:1996:CMH

[1888] L. Recalde, E. Teruel, and M. Silva.

0SC0*ECS: A class of modular and hierarchical cooperating systems. Lecture Notes in Computer Science, 1091:440– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sunesen:1996:BEI

[1889] K. Sunesen and M. Nielsen. Behavioral equivalence for infinite systems — partially decidable! Lecture Notes in Computer Science, 1091:460–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanOosten:1996:TAT

[1890] J. Van Oosten. Topological aspects of traces. Lecture Notes in Computer Science, 1091:480–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varshavsky:1996:ACD

[1891] V. I. Varshavsky and V. B. Marakhovsky. Asynchronous control device design by net model behavior simulation. Lecture Notes in Computer Science, 1091: 497–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vernadat:1996:CSG

[1892] F. Vernadat, P. Azema, and F. Michel. Covering step graph. Lecture Notes in Computer Science, 1091:516–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuan:1996:PAC

[1893] J. Yuan, L. Liu, and Z. Dong. Performance analysis of a connection management scheme in IWU interconnecting LANs across ATM networks. Lecture Notes in Computer Science, 1091: 536–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1996:IFO

[1894] M. Baaz, A. Leitsch, and R. Zach. Incompleteness of a first-order Gödel logic and some temporal logics of programs. Lecture Notes in Computer Science, 1092:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barros:1996:SNT

[1895] J. Barros and J. Goguen. Semantics of non-terminating rewrite systems using minimal coverings. Lecture Notes in Computer Science, 1092:16–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1996:CT

[1896] G. Barthe and H. Geuvers. Congruence types. Lecture Notes in Computer Science, 1092:36–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beckert:1996:DCS

[1897] B. Beckert and R. Haehnle. Deduction by combining semantic tableaux and integer programming. Lecture Notes in Computer Science, 1092:52–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beckert:1996:LLE

[1898] B. Beckert and J. Posegga. leanEA:

A lean evolving algebra compiler. Lecture Notes in Computer Science, 1092: 64–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blackburn:1996:PSF

[1899] P. Blackburn, W. Meyer-Viol, and M. De Rijke. A proof system for finite trees. Lecture Notes in Computer Science, 1092:86–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1996:RUL

[1900] J. Brown and L. A. Wallen. Representing unification in a logical framework. Lecture Notes in Computer Science, 1092:106–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caferra:1996:DPU

[1901] R. Caferra and N. Peltier. Decision procedures using model building techniques. Lecture Notes in Computer Science, 1092:130–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clote:1996:NRB

[1902] P. Clote. A note on the relation between polynomial time functionals and Constable’s class K. Lecture Notes in Computer Science, 1092:145–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dawar:1996:FOL

[1903] A. Dawar, S. Lindell, and S. Weinstein. First order logic, fixed point logic and linear order. Lecture Notes in Computer Science, 1092:161–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degtyarev:1996:SRU

[1904] A. Degtyarev and A. Voronkov. Simultaneous rigid E-unification is undecidable. Lecture Notes in Computer Science, 1092:178–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelCastillo:1996:EAA

[1905] G. Del Castillo, I. Durdanovic, and U. Glaesser. An evolving algebra abstract machine. Lecture Notes in Computer Science, 1092:191–??, 1996.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiCosmo:1996:REP

[1906] R. Di Cosmo and D. Kesner. Rewriting with extensional polymorphic lambdacalculus. Lecture Notes in Computer Science, 1092:215–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Droste:1996:LLD

[1907] M. Droste and D. Kuske. Languages and logical definability in concurrency monoids. Lecture Notes in Computer Science, 1092:233–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grumbach:1996:GID

[1908] S. Grumbach, Z. Lacroix, and S. Lindell. Generalized implicit definitions on finite structures. Lecture Notes in Computer Science, 1092:252–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gurevich:1996:RCP

[1909] Y. Gurevich and J. K. Huggins. The railroad crossing problem: An experiment with instantaneous actions and immediate reactions. Lecture Notes in Computer Science, 1092:266–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasegawa:1996:LAP

[1910] R. Hasegawa. A logical aspect of parametric polymorphism. Lecture Notes in Computer Science, 1092:291–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heuerding:1996:MLK

[1911] A. Heuerding and S. Schwendimann. On the modal logic K plus theories. Lecture Notes in Computer Science, 1092:308–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hudelmaier:1996:IDP

[1912] J. Hudelmaier. Inproved decision procedures for the modal logics K,T, and S4. Lecture Notes in Computer Science, 1092:320–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ingolfsdottir:1996:FAD

[1913] A. Ingolfsdottir and A. Schalk. A fully abstract denotational model for observational precongruence. Lecture Notes in Computer Science, 1092:335– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johannsen:1996:SBL

[1914] J. Johannsen. On sharply bounded length induction. Lecture Notes in Computer Science, 1092:362–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kummer:1996:ESE

[1915] M. Kummer and M. Ott. Effective strategies for Enumeration Games 368. Lecture Notes in Computer Science, 1092:368–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leiss:1996:BFP

[1916] H. Leiss. Bounded fixed-point definability and tabular recognition of languages. Lecture Notes in Computer Science, 1092:388–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mossakowski:1996:EAV

[1917] T. Mossakowski. Equivalences among various logical frameworks of partial algebras. Lecture Notes in Computer Science, 1092:403–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pandya:1996:SEP

[1918] P. K. Pandya. Some extensions to propositional mean-value calculus: Expressiveness and decidability. Lecture Notes in Computer Science, 1092:434– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rubio:1996:TPM

[1919] A. Rubio. Theorem proving modulo associativity. Lecture Notes in Computer Science, 1092:452–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vigneron:1996:PDM

[1920] L. Vigneron. Positive deduction modulo regular theories. Lecture Notes in Computer Science, 1092:468–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crowley:1996:MFN

[1921] J. L. Crowley. Mathematical foundations of navigation and perception for an autonomous mobile robot. Lecture Notes in Computer Science, 1093:9–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voorbraak:1996:RUA

[1922] F. Voorbraak. Reasoning with uncertainty in AI. Lecture Notes in Computer Science, 1093:52–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elfes:1996:RNI

[1923] A. Elfes. Robot navigation: Integrating perception, environmental constraints and task execution within a probabilistic framework. Lecture Notes in Computer Science, 1093:93–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hummel:1996:URO

[1924] R. Hummel. Uncertainty reasoning in object recognition by image processing. Lecture Notes in Computer Science, 1093:131–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaelbling:1996:POM

[1925] L. P. Kaelbling, M. L. Littman, and A. R. Cassandra. Partially observable Markov decision processes for artificial intelligence. Lecture Notes in Computer Science, 1093:146–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pagac:1996:EAP

[1926] D. Pagac, E. M. Nebot, and H. Durrant-Whyte. An evidential approach to probabilistic map-building. Lecture Notes in Computer Science, 1093:164–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paris:1996:BFC

[1927] J. Paris and A. Vencovska. Belief formation by constructing models. Lecture Notes in Computer Science, 1093:171–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shafer:1996:CR

[1928] G. Shafer. Causal relevance. Lecture Notes in Computer Science, 1093:187– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alferov:1996:RCS

[1929] G. V. Alferov and O. A. Malafeyev.

The robot control strategy in a domain with dynamical obstacles. Lecture Notes in Computer Science, 1093: 211–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bacchus:1996:RAN

[1930] F. Bacchus, J. Y. Halpern, and H. J.

Levesque. Reasoning about noisy sensors (and effectors) in the situation calculus. Lecture Notes in Computer Science, 1093:218–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boley:1996:RTL

[1931] D. L. Boley, E. S. Steinmetz, and K. T. Sutherland. Recursive total least squares: An alternative to using the discrete Kalman filter in robot navigation. Lecture Notes in Computer Science, 1093:221–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouilly:1996:SBM

[1932] B. Bouilly and T. Simeon. A sensorbased motion planner for mobile robot navigation with uncertainty. Lecture Notes in Computer Science, 1093:235– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brafman:1996:KCR

[1933] R. I. Brafman and Y. Shoham. Knowledge considerations in robotics. Lecture Notes in Computer Science, 1093: 248–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDam:1996:NNA

[1934] J. W. M. Van Dam, B. J. A. Kroese, and F. C. A. Groen. Neural network applications in sensor fusion for an autonomous mobile robot. Lecture Notes in Computer Science, 1093:263– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Froehlinghaus:1996:SUK

[1935] T. Froehlinghaus. Structuring uncertain knowledge with hierarchical Bayesian networks. Lecture Notes in Computer Science, 1093:279–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez:1996:UTS

[1936] E. Gonzalez, A. Suarez, C. Moreno, and F. Artigue. Uncertainty treatment in a surface filling mobile robot. Lecture Notes in Computer Science, 1093: 294–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hebert:1996:PML

[1937] P. Hebert, S. Betge-Brezetz, and R. Chatila. Probabilistic map learning: Necessity and difficulties. Lecture Notes in Computer Science, 1093:307– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koenig:1996:RNM

[1938] S. Koenig, R. Goodwin, and R. G. Simmons. Robot navigation with Markov models: A framework for path planning and learning with limited computational resources. Lecture Notes in Computer Science, 1093:322–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konolige:1996:RMO

[1939] K. Konolige. A refined method for occupancy grid interpretation. Lecture Notes in Computer Science, 1093:338– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kristensen:1996:SPB

[1940] S. Kristensen. Sensor planning with Bayesian decision theory. Lecture Notes in Computer Science, 1093:353–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saffiotti:1996:PBS

[1941] A. Saffiotti and L. P. Wesley.

Perception-based self-localization using fuzzy locations. Lecture Notes in Computer Science, 1093:368–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellahsene:1996:VMS

[1942] Z. Bellahsene. View mechanism for schema evolution in object-oriented DBMS. Lecture Notes in Computer Science, 1094:18–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dinn:1996:ARL

[1943] A. Dinn, N. W. Paton, M. H. Williams, and A. A. A. Fernandes. An active rule language for ROCK and ROLL. Lecture Notes in Computer Science, 1094:36–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doucet:1996:ICM

[1944] A. Doucet, S. Gancarski, G. Jomier, and S. Monties. Integrity constraints in multiversion databases. Lecture Notes in Computer Science, 1094:56– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibrahim:1996:DSI

[1945] H. Ibrahim, W. A. Gray, and N. J. Fiddian. The development of a semantic integrity constraint subsystem for a distributed database. Lecture Notes in Computer Science, 1094:74–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delannoy:1996:UTB

[1946] X. C. Delannoy. Understanding the tension between transition rules and confidentiality. Lecture Notes in Computer Science, 1094:92–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiseman:1996:SED

[1947] S. Wiseman and B. Pomeroy. SPEAR:

Extending ER for dynamic behaviour and refinement. Lecture Notes in Computer Science, 1094:107–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freitas:1996:SKD

[1948] A. A. Freitas and S. H. Lavington.

Speeding up knowledge discovery in large relational databases by means of a new discretization algorithm. Lecture Notes in Computer Science, 1094: 124–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flach:1996:ILM

[1949] G. Flach and H. Meyer. Integration of load measurement parameters into the cost evaluation of database queries. Lecture Notes in Computer Science, 1094:134–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boncz:1996:HPO

[1950] P. A. Boncz, F. Kwakkel, and M. L.

Kersten. High performance OO traversals in monet. Lecture Notes in Computer Science, 1094:152–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Embury:1996:MCA

[1951] S. M. Embury and P. M. D. Gray.

A modular compiler architecture for a data manipulation language. Lecture Notes in Computer Science, 1094:170– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayres:1996:QGD

[1952] R. Ayres and P. J. H. King. Querying graph databases using a functional language extended with second order facilities. Lecture Notes in Computer Science, 1094:189–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griffiths:1996:SAT

[1953] A. Griffiths and B. Theodoulidis. SQL+i: Adding temporal indeterminacy to the database language SQL. Lecture Notes in Computer Science, 1094:204–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chalmers:1996:PSS

[1954] M. Chalmers. Pearls, swines and sows’ ears: Interface research inside a multinational bank. Lecture Notes in Computer Science, 1094:222–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zdonik:1996:DBI

[1955] S. B. Zdonik. Dissemination-based information systems: Your data may be where you least expect it. Lecture Notes in Computer Science, 1094:230– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stanley:1996:MDT

[1956] N. Stanley. Microsoft database technologies. Lecture Notes in Computer Science, 1094:231–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butterfield:1996:PMQ

[1957] M. Butterfield, N. Caine, and S. Ross-Talbot. Predicate maintained queries: An active OODBMS for financial applications. Lecture Notes in Computer Science, 1094:232–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bailey:1996:UDM

[1958] A. Bailey. Universal data management. Lecture Notes in Computer Science, 1094:233–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCune:1996:ADE

[1959] W. McCune and R. Padmanabhan. Automated deduction in equational logic and cubic curves. Lecture Notes in Computer Science, 1095:ix + 231, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schael:1996:WMS

[1960] T. Schael. Workflow management systems for process organisations. Lecture Notes in Computer Science, 1096:xii + 200, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alon:1996:DSS

[1961] N. Alon. Derandomization via small sample spaces. Lecture Notes in Computer Science, 1097:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodal:1996:RCM

[1962] G. S. Brodal, S. Chaudhuri, and J. Radhakrishnan. The randomized complexity of maintaining the minimum. Lecture Notes in Computer Science, 1097:4–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

RanchHenzinger:1996:FAN

[1963] M. Ranch Henzinger and J. A. Telle. Faster algorithms for the nonemptiness of Streett automata and for communication protocol pruning. Lecture Notes in Computer Science, 1097:16– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marathe:1996:SCN

[1964] M. V. Marathe, R. Ravi, and R. Sundaram. Service-constrained network design problems. Lecture Notes in Computer Science, 1097:28–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelsen:1996:AHC

[1965] P. Kelsen, S. Mahajan, and H. Ramesh. Approximate hypergraph coloring. Lecture Notes in Computer Science, 1097: 41–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chandra:1996:FDR

[1966] B. Chandra and M. M. Halldorsson. Facility dispersion and remote subgraphs. Lecture Notes in Computer Science, 1097:53–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravi:1996:CMS

[1967] R. Ravi and M. X. Goemans. The constrained minimum spanning tree problem. Lecture Notes in Computer Science, 1097:66–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lau:1996:RAC

[1968] H. C. Lau and O. Watanabe. Randomized approximation of the constraint satisfaction problem. Lecture Notes in Computer Science, 1097:76–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klauck:1996:HGL

[1969] H. Klauck. On the hardness of global and local approximation. Lecture Notes in Computer Science, 1097:88– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asano:1996:AAM

[1970] T. Asano, T. Ono, and T. Hirata. Approximation algorithms for the maximum satisfiability problem. Lecture Notes in Computer Science, 1097:100– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirata:1996:HAM

[1971] K. Hirata, S. Shimozono, and A. Shinohara. On the hardness of approximating the minimum consistent OBDD problem. Lecture Notes in Computer Science, 1097:112–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1996:CUM

[1972] T. W. Lam, W. K. Sung, and H. F. Ting. Computing the unrooted maximum agreement subtree in subquadratic time. Lecture Notes in Computer Science, 1097:124–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asahiro:1996:GFD

[1973] Y. Asahiro, K. Iwama, H. Tamaki, and T. Tokuyama. Greedily finding a dense subgraph. Lecture Notes in Computer Science, 1097:136–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez-Baca:1996:USP

[1974] D. Fernandez-Baca, G. Slutzki, and D. Eppstein. Using sparsification for parametric minimum spanning tree problems. Lecture Notes in Computer Science, 1097:149–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1996:VPP

[1975] A. Gupta, D. Kaller, S. Mahajan, and T. Shermer. Vertex partitioning problems on partial k-trees. Lecture Notes in Computer Science, 1097:161– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blair:1996:MAF

[1976] J. R. S. Blair, P. Heggernes, and J. A. Telle. Making an arbitrary filled graph minimal by removing fill edges. Lecture Notes in Computer Science, 1097:173–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersson:1996:SSR

[1977] A. Andersson. Sorting and searching revisited. Lecture Notes in Computer Science, 1097:185–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Husfeldt:1996:LBD

[1978] T. Husfeldt, T. Rauhe, and S. Skyum.

Lower bounds for dynamic transitive closure, planar point location, and parentheses matching. Lecture Notes in Computer Science, 1097:198–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alstrup:1996:OPA

[1979] S. Alstrup and M. Thorup. Optimal pointer algorithms for finding nearest common ancestors in dynamic trees. Lecture Notes in Computer Science, 1097:212–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelsen:1996:NGD

[1980] P. Kelsen. Neighborhood graphs and distributed cap delta + 1-Coloring. Lecture Notes in Computer Science, 1097:223–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gargano:1996:CCG

[1981] L. Gargano, A. A. Rescigno, and U. Vaccaro. Communication complexity of gossiping by packets. Lecture Notes in Computer Science, 1097:234– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Przytycka:1996:OCS

[1982] T. Przytycka and L. Higham. Optimal cost-sensitive distributed minimum spanning tree algorithm. Lecture Notes in Computer Science, 1097:246– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auletta:1996:LTA

[1983] V. Auletta, A. Monti, M. Parente, and P. Persiano. A linear time algorithm for the feasibility of pebble motion on trees. Lecture Notes in Computer Science, 1097:259–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levcopoulos:1996:LTH

[1984] C. Levcopoulos and A. Oestlin. Lineartime heuristics for minimum weight rectangulation. Lecture Notes in Computer Science, 1097:271–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aronov:1996:VMR

[1985] B. Aronov, A. R. Davis, T. K. Dey, and S. P. Pal. Visibility with multiple reflections. Lecture Notes in Computer Science, 1097:284–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levcopoulos:1996:FHA

[1986] C. Levcopoulos and D. Krznaric. A fast heuristic for approximating the minimum weight triangulation. Lecture Notes in Computer Science, 1097:296– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodnik:1996:NG

[1987] A. Brodnik and J. I. Munro. Neighbours on a grid. Lecture Notes in Computer Science, 1097:309–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Azar:1996:TDP

[1988] Y. Azar and L. Epstein. On two dimensional packing. Lecture Notes in Computer Science, 1097:321–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biedl:1996:OOD

[1989] T. C. Biedl. Optimal orthogonal drawings of triconnected plane graphs. Lecture Notes in Computer Science, 1097: 333–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez-Ortiz:1996:WSF

[1990] A. Lopez-Ortiz and S. Schuierer. Walking streets faster. Lecture Notes in Computer Science, 1097:345–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preminger:1996:SET

[1991] S. Preminger and E. Upfal. Safe and efficient traffic laws for mobile robots. Lecture Notes in Computer Science, 1097:357–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paterson:1996:PS

[1992] M. Paterson. Progress in selection. Lecture Notes in Computer Science, 1097:

368–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonnet:1996:PAS

[1993] G. H. Gonnet and S. A. Benner. Probabilistic ancestral sequences and multiple alignments. Lecture Notes in Computer Science, 1097:380–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasieniec:1996:EAL

[1994] L. Gasieniec, M. Karpinski, W. Plandowski, and W. Rytter. Efficient algorithms for Lempel–Ziv encoding. Lecture Notes in Computer Science, 1097:392–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeumker:1996:DCP

[1995] A. Baeumker, W. Dittrich, and A. Pietracaprina. The deterministic complexity of parallel multisearch. Lecture Notes in Computer Science, 1097: 404–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodal:1996:PQP

[1996] G. S. Brodal. Priority queues on parallel machines. Lecture Notes in Computer Science, 1097:416–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fagerberg:1996:BST

[1997] R. Fagerberg. Binary search trees: How low can you go? Lecture Notes in Computer Science, 1097:428–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boros:1996:BAI

[1998] E. Boros, T. Ibaraki, and K. Makino.

Boolean analysis of incomplete examples. Lecture Notes in Computer Science, 1097:440–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldberg:1996:MS

[1999] Adele Goldberg. Measurement strategies. Lecture Notes in Computer Science, 1098:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1098/10980001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1098/10980001.pdf.

Bancilhon:1996:WEE

[2000] François Bancilhon. Will Europe ever produce and sell objects? Lecture Notes in Computer Science, 1098:2– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1098/10980002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1098/10980002.
pdf.

Boyland:1996:TSC

[2001] John Boyland and Giuseppe Castagna.

Type-safe compilation of covariant specialization: A practical case. Lecture Notes in Computer Science, 1098:3– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/1098/10980003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1098/10980003.
pdf.

Gawecki:1996:ISM

[2002] Andreas Gawecki and Florian Matthes. Integrating subtyping, matching and type quantification: A practical perspective. Lecture Notes in Computer Science, 1098:26–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1098/10980026.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1098/10980026.pdf.

Qian:1996:TOO

[2003] Zhenyu Qian and Bernd Krieg-Brückner. Typed object-oriented functional programming with late binding. Lecture Notes in Computer Science, 1098:48–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1098/10980048.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1098/10980048.
pdf.

Baumer:1996:LSO

[2004] Dirk Bäumer, Rolf Knoll, Guido Gryczan, and Heinz Züllighoven. Large scale object-oriented softwaredevelopment in a banking environment — an experience report. Lecture Notes in Computer Science, 1098:73– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1098/10980073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1098/10980073.
pdf.

Banavar:1996:AFM

[2005] Guruduth Banavar and Gary Lindstrom. An application framework for module composition tools. Lecture Notes in Computer Science, 1098:91– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1098/10980091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1098/10980091.
pdf.

Engelson:1996:AGU

[2006] Vadim Engelson, Dag Fritzson, and Peter Fritzson. Automatic generation of user interfaces from data structure specifications and objectoriented application models. Lecture Notes in Computer Science, 1098:114– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1098/10980114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1098/10980114.
pdf.

Aigner:1996:EVF

[2007] Gerald Aigner and Urs Hölzle. Eliminating virtual function calls in C++ programs. Lecture Notes in Computer Science, 1098:142–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1098/10980142.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1098/10980142.
                                        pdf.

Amiel:1996:SED

[2008] Eric Amiel and Eric Dujardin. Supporting explicit disambiguation of multi-methods. Lecture Notes in Computer Science, 1098:167–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1098/10980167.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1098/10980167.
                                        pdf.

Minsky:1996:TAF

[2009] Naftaly H. Minsky. Towards alias-free pointers. Lecture Notes in Computer Science, 1098:189–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1098/10980189.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1098/10980189.pdf.

Jacobs:1996:ICC

[2010] Bart Jacobs. Inheritance and cofree constructions. Lecture Notes in Computer Science, 1098:210–231, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1098/10980210.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1098/10980210.
pdf.

Lechner:1996:OCR

[2011] Ulrike Lechner, Christian Lengauer, Friederike Nickl, and Martin Wirsing. (objects + concurrency) & reusability — A proposal to circumvent the inheritance anomaly. Lecture Notes in Computer Science, 1098:232–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1098/10980232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1098/10980232.
pdf.

Rossie:1996:MSB

[2012] Jonathan G. Rossie Jr., Daniel P.

Friedman, and Mitchell Wand. Modeling subobject-based inheritance. Lecture Notes in Computer Science, 1098: 248–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1098/10980248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1098/10980248.
pdf.

Jezequel:1996:PO

[2013] Jean-Marc Jézéquel and Jean-Lin Pacherie. Parallel operators. Lecture Notes in Computer Science, 1098:275– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1098/10980275.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1098/10980275.
                                        pdf.

Kono:1996:IMM

[2014] Kenji Kono, Kazuhiko Kato, and Takashi Masuda. An implementation method of migratable distributed objects using an RPC technique integrated with virtual memory management. Lecture Notes in Computer Science, 1098:295–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1098/10980295.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1098/10980295.
                                        pdf.

Garbinato:1996:PCD

[2015] Benoît Garbinato, Pascal Felber, and Rachid Guerraoui. Protocol classes for designing reliable distributed environments. Lecture Notes in Computer Science, 1098:316–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1098/10980316.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1098/10980316.
                                        pdf.

Bullat:1996:DCO

[2016] Frédérique Bullat and Michel Schneider. Dynamic clustering in object databases exploiting effective use of relationships between objects. Lecture Notes in Computer Science, 1098:344– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1098/10980344.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1098/10980344.
pdf.

Silva:1996:CDA

[2017] Mauricio J. V. Silva and C. Robert Carlson. Conceptual design of active object-oriented database applications using multi-level diagrams. Lecture Notes in Computer Science, 1098:366– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1098/10980366.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1098/10980366.
pdf.

Hohenstein:1996:BGB

[2018] Uwe Hohenstein. Bridging the gap between C++ and relational databases. Lecture Notes in Computer Science, 1098:398–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1098/10980398.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1098/10980398.
pdf.

Brandt:1996:GBT

[2019] Søren Brandt and Jørgen Lindskov Knudsen. Generalising the BETA type system. Lecture Notes in Computer Science, 1098:421–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1098/10980421.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1098/10980421.
                                        pdf.

Rinat:1996:MPT

[2020] Ran Rinat and Menachem Magidor. Metaphoric polymorphism: Taking code reuse one step further. Lecture Notes in Computer Science, 1098:449– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1098/10980449.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1098/10980449.
                                        pdf.

Kristensen:1996:AAC

[2021] Bent Bruun Kristensen and Daniel C. M. May. Activities: Abstractions for collective behavior. Lecture Notes in Computer Science, 1098:472– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1098/10980472.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1098/10980472.
                                        pdf.

Anonymous:1996:AIb

[2022] Anonymous. Author index. Lecture Notes in Computer Science, 1098:502– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1098/1098ind.pdf.

Ganzinger:1996:SBT

[2023] H. Ganzinger. Saturation-based theorem proving. Lecture Notes in Computer Science, 1099:1–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ranade:1996:BEP

[2024] A. Ranade. Bandwidth efficient parallel computation. Lecture Notes in Computer Science, 1099:4–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruyere:1996:VLM

[2025] V. Bruyere and M. Latteux. Variablelength maximal codes. Lecture Notes in Computer Science, 1099:24–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Razborov:1996:LBP

[2026] A. A. Razborov. Lower bounds for propositional proofs and independence results in bounded arithmetic. Lecture Notes in Computer Science, 1099:48– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:ACD

[2027] X. J. Chen and R. De Nicola. Algebraic characterizations of decorated trace equivalences over tree-like structures. Lecture Notes in Computer Science, 1099:63–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jenner:1996:FAS

[2028] L. Jenner and W. Vogler. Fast asynchronous systems in dense time. Lecture Notes in Computer Science, 1099: 75–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenzi:1996:HTC

[2029] G. Lenzi. A hierarchy theorem for the λ(?)-calculus. Lecture Notes in Computer Science, 1099:87–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bradfield:1996:ETS

[2030] J. Bradfield, J. Esparza, and A. Mader. An effective tableau system for the linear time λ(?)-calculus. Lecture Notes in Computer Science, 1099:98–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joung:1996:CFI

[2031] Y.-J. Joung. Characterizing fairness implementability for multiparty interaction. Lecture Notes in Computer Science, 1099:110–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucas:1996:TCS

[2032] S. Lucas. Termination of contextsensitive rewriting by rewriting. Lecture Notes in Computer Science, 1099: 122–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1996:CGS

[2033] A. Cheng and D. Kozen. A complete Gentzen-style axiomatization for set constraints. Lecture Notes in Computer Science, 1099:134–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Billaud:1996:FEC

[2034] M. Billaud. Fatal errors in conditional expressions. Lecture Notes in Computer Science, 1099:146–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mossakowski:1996:DTA

[2035] T. Mossakowski. Different types of arrow between logical frameworks. Lecture Notes in Computer Science, 1099: 158–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitchell:1996:EMP

[2036] J. Mitchell and R. Viswanathan. Effective models of polymorphism, subtyping and recursion. Lecture Notes in Computer Science, 1099:170–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosscher:1996:RLC

[2037] D. J. B. Bosscher and W. O. D. Griffioen. Regularity for a large class of context-free processes is decidable. Lecture Notes in Computer Science, 1099: 182–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caucal:1996:ITG

[2038] D. Caucal. On infinite transition graphs having a decidable monadic theory. Lecture Notes in Computer Science, 1099:194–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Senizergues:1996:SGA

[2039] G. Senizergues. Semi-groups acting on context-free graphs. Lecture Notes in Computer Science, 1099:206–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lisovik:1996:HSM

[2040] L. P. Lisovik. Hard sets method and semilinear reservoir method with applications. Lecture Notes in Computer Science, 1099:219–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flajolet:1996:RPP

[2041] P. Flajolet, X. Gourdon, and D. Panario. Random polynomials and polynomial factorization. Lecture Notes in Computer Science, 1099:232–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koppenhagen:1996:OGB

[2042] U. Koppenhagen and E. W. Mayr. Optimal Gröbner base algorithms for binomial ideals. Lecture Notes in Computer Science, 1099:244–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloks:1996:MFC

[2043] T. Kloks, D. Kratsch, and C. K. Wong. Minimum fill-in on circle and circulararc graphs. Lecture Notes in Computer Science, 1099:256–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1996:PAS

[2044] Z.-Z Chen. Practical approximation schemes for maximum inducedsubgraph problems on K3,3-free or K5free graphs. Lecture Notes in Computer Science, 1099:268–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koutsoupias:1996:SFG

[2045] E. Koutsoupias, C. Papadimitriou, and M. Yannakakis. Searching a fixed graph. Lecture Notes in Computer Science, 1099:280–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

RauchHenzinger:1996:ISA

[2046] M. Rauch Henzinger and M. Thorup.

Improved sampling with applications to dynamic graph algorithms. Lecture Notes in Computer Science, 1099:290– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pin:1996:EPE

[2047] J.-E. Pin. The expressive power of existential first order sentences of Buechi’s sequential calculus. Lecture Notes in Computer Science, 1099:300–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaivola:1996:FRT

[2048] R. Kaivola. Fixpoints for Rabin tree automata make complementation easy. Lecture Notes in Computer Science, 1099:312–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hashiguchi:1996:NUB

[2049] K. Hashiguchi. New upper bounds to the limitedness of distance automata. Lecture Notes in Computer Science, 1099:324–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raymond:1996:RRE

[2050] P. Raymond. Recognizing regular expressions by means of dataflow networks. Lecture Notes in Computer Science, 1099:336–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ablayev:1996:PRB

[2051] F. Ablayev and M. Karpinski. On the power of randomized branching programs. Lecture Notes in Computer Science, 1099:348–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreev:1996:HSD

[2052] A. E. Andreev, A. E. F. Clementi, and J. D. P. Rolim. Hitting sets derandomize BPP. Lecture Notes in Computer Science, 1099:357–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Book:1996:TPQ

[2053] R. Book, H. Vollmer, and K. Wagner. On type-2 probabilistic quantifiers. Lecture Notes in Computer Science, 1099:369–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiedermann:1996:SST

[2054] J. Wiedermann. Speeding-up singletape nondeterministic computations by single alternation, with separation results. Lecture Notes in Computer Science, 1099:381–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Julia:1996:OGC

[2055] S. Julia. On omega-generators and codes. Lecture Notes in Computer Science, 1099:393–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeLuca:1996:SSM

[2056] A. De Luca. On standard Sturmian morphisms. Lecture Notes in Computer Science, 1099:403–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ateniese:1996:CBV

[2057] G. Ateniese, C. Blundo, A. De Santis, and D. R. Stinson. Constructions and bounds for visual cryptography. Lecture Notes in Computer Science, 1099:416–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Azar:1996:CI

[2058] Y. Azar, Y. Bartal, E. Feuerstein, and A. Fiat. On capital investment. Lecture Notes in Computer Science, 1099:429–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miltersen:1996:LBS

[2059] P. B. Miltersen. Lower bounds for static dictionaries on RAMs with bit operations but no multiplication. Lecture Notes in Computer Science, 1099: 442–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bradford:1996:LBR

[2060] P. G. Bradford and K. Reinert. Lower bounds for row minima searching. Lecture Notes in Computer Science, 1099: 454–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shukla:1996:CRP

[2061] S. K. Shukla, H. B. Hunt, D. J. Rosenkrantz, and R. E. Stearns. On the complexity of relational problems for finite state processes. Lecture Notes in Computer Science, 1099:466–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jancar:1996:DFP

[2062] P. Jancar and J. Esparza. Deciding finiteness of Petri nets up to bisimulation. Lecture Notes in Computer Science, 1099:478–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodei:1996:MPD

[2063] C. Bodei, P. Degano, and C. Priami. Mobile processes with a distributed environment. Lecture Notes in Computer Science, 1099:490–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanGlabbeek:1996:MNP

[2064] R. J. Van Glabbeek. The meaning of negative premises in transition system specifications II. Lecture Notes in Computer Science, 1099:502–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albers:1996:ACA

[2065] S. Albers and M. Mitzenmacher. Average case analyses of list update algorithms, with applications to data compression. Lecture Notes in Computer Science, 1099:514–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz:1996:SOD

[2066] F. Schulz and E. Schoemer. Selforganizing data structures with dependent accesses. Lecture Notes in Computer Science, 1099:526–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1996:LTA

[2067] V. S.-N. Choi and M. Golin. Lopsided trees: Analyses, algorithms, and applications. Lecture Notes in Computer Science, 1099:538–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farach:1996:OLT

[2068] M. Farach and S. Muthukrishnan. Optimal logarithmic time randomized suffix tree construction. Lecture Notes in Computer Science, 1099:550–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alon:1996:IPA

[2069] N. Alon and A. Srinivasan. Improved parallel approximation of a class of integer programming problems. Lecture Notes in Computer Science, 1099:562– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bermond:1996:ECC

[2070] J.-C. Bermond, L. Gargano, S. Perennes, and A. A. Rescigno. Efficient collective communication in optical networks. Lecture Notes in Computer Science, 1099:574–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chlebus:1996:SMS

[2071] B. S. Chlebus, A. Gambin, and P. Indyk. Shared-memory simulations on a faulty-memory DMM. Lecture Notes in Computer Science, 1099:586–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herley:1996:FDB

[2072] K. T. Herley, A. Pietracaprina, and G. Pucci. Fast deterministic backtrack search. Lecture Notes in Computer Science, 1099:598–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1996:ARD

[2073] X. Yu and M. Yung. Agent rendezvous: A dynamic symmetry-breaking problem. Lecture Notes in Computer Science, 1099:610–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aumann:1996:EAC

[2074] Y. Aumann and M. A. Bender. Efficient asynchronous consensus with the value-oblivious adversary scheduler. Lecture Notes in Computer Science, 1099:622–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cowen:1996:FFE

[2075] L. Cowen, J. Feigenbaum, and S. Kannan. A formal framework for evaluating heuristic programs. Lecture Notes in Computer Science, 1099:634–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chakrabarti:1996:ISA

[2076] S. Chakrabarti, C. A. Phillips, A. S.

Schulz, and D. B. Shmoys. Improved scheduling algorithms for minsum criteria. Lecture Notes in Computer Science, 1099:646–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paterson:1996:CSF

[2077] M. Paterson and T. Przytycka. On the complexity of string folding. Lecture Notes in Computer Science, 1099: 658–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez-Baca:1996:PTA

[2078] D. Fernandez-Baca and J. Lagergren.

A polynomial-time algorithm for nearperfect phylogeny. Lecture Notes in Computer Science, 1099:670–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godefroid:1996:POM

[2079] Patrice Godefroid. Partial-order methods for the verification of concurrent systems: an approach to the stateexplosion problem, volume 1032 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-60761-7 (Berlin softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). 142 pp. LCCN QA76.58 .G63 1996.

Hajnicz:1996:TSF

[2080] Elzbieta Hajnicz. Time structures: formal description and algorithmic representation, volume 1047 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-60941-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). ix + 244 pp. LCCN QA76.9.A43 H35 1996.

Haddadi:1996:CCA

[2081] Afsaneh Haddadi. Communication and cooperation in agent systems: a pragmatic theory, volume 1056 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61044-8. ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 148 pp. LCCN QA76.9.D5 H33 1996.

SparckJones:1996:ENL

[2082] Karen Sparck Jones and J. R. (Julia Rose) Galliers. Evaluating natural language processing systems: an analysis and review, volume 1083 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61309-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). various pp. LCCN QA76.9.N38S74 1996.